A Complete Bibliography of Publications in the Journal of Statistical Planning and Inference:

1990–1999

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.01

Title word cross-reference

(0, k) [SJG93]. (4k^{2}, 2k^{2} + k, k^{2} + k) [Kou94]. (M, N) [BM93, BKF98]. (m, s) [OR98]. (n, k, 5) [Far92]. (ν, 4, 3) [Qin96]. (ν, 4, 4) [Kej96]. (p, t) [BTV96]. (r, λ, c) [Mon90]. (v_{k−30}, k, 3^{k−1} − 21; 3) [HH93a]. (X, Y) [Let97]. $135.00 [Smi97a]. 1β [Col95]. 2 [BL96, MPW97, Rus93]. 2 − (27, 9, 4) [LT96, LT00]. 2 − p [CL91]. 2^{m} [GT93, Hyo92, NM97, ST96b, SH92c, YFHY96]. 2^{n−k} [CX98b]. 2^{n−m} [CH98]. 2^{n}u^{1} [YGL96]. 2l + 1 [Hyo92]. 2 × 2 [AG99]. 3 [HSS97, SK92]. 3, 3 [HH96a, HM96]. $39.95 [Law97]. 3^{m} [KI98]. 3v_{0} + v_{2} [HM96]. 3v_{1} + v_{2} [HH96a]. 3v_{1} + v_{3} [HM96]. 3v_{2} + v_{3} [HH96a]. $47.50 [Chi92]. $49.95 [AC97b, Iye92]. 5 [BL96]. $51.95 [Sal92]. $54.95 [Smi97c]. $55.50 [Jon92]. $59.95 [Coa97, Dra97]. $74.95 [Smi97b]. 7 ≤ m ≤ 10 [NM97]. [15, 4, 9; 3] [HH96a]. [k − 1, k + 1] [LY96]. ^{2} [Iwa97]. ^{m} [Shi91]. ^{n−k} [CX98a]. _{0} [ZX98]. _{1} [Det93]. _{2} [BG98]. _{v} [Yin96]. A [CK94, CF96, Far91, Far92, GRP99, Jac92, NM97, SS90a, SS91a, ST96b, TY96]. a + b [BD91]. α [NG91]. a × b [BD91, VH98]. B [Lia90a]. B(4, 2; ν) [Din96]. B_{2} = 0 [HH96a]. C [BS93a]. χ^{2} [Ma97a, MH97]. D [BD96b, Bis93, CM91, CC96, Cha90, CL97, Det93, DN97, KKNK96, Lim90, Mik93, MH95, PWS98, SS90a, SS91a, SO93, SW95a, SS97a, Vil91, Det94]. Δ [Gui96, Kha90a]. E [Bag96, Hei96, Jac90, Jac98, SK92, Udd96]. E(X^{2}|Y) [Let97]. ϵ [HF95, MP93, dlHF94]. F [Ruk93]. F_{2}[{F_{2^{m}}, +}] [Cha96c]. g [HMM97]. H [Kus99, WZ96a]. I [BD95a]. ∞ [Lee95]. J [HP90a]. K [Ber95b, DN94, BMV97, GEGM99, HMM97, Liu97b]. k = 3 [BK96]. k = 4 [KVAN90]. k = n − 1, n [Far92]. K_{4} − e [BGL97]. L [BG98, Cha97b, Mae94, MS90a, MH98, Sho97]. L_{k} [AS96a]. L_{p} [Arc99, CADMM95, ESV97, GS97b]. λ_{1} = 0 [BK96]. LQ [MH98]. M [BMV97, Shi97, Sil92, Wie94, YZ97, Zha95, BL96, CLR91a, CLR91b, HP90a, Hai92]. m(n) [Nie92]. m = 3, 4 [BL96]. N [CM98, Che98b]. n = 15 [CF96]. n = 19 [Far91]. n = 90 [KKNK96]. N ≡ 3 mod 4 [SS91a, CK94]. N ≡ 3 (mod 4) [SS90a, Far92]. n × n [LB92]. ν(m, t) [MY96b]. P [AG98, OD99, AG99, BH98, BTV96, CW98, Moo96, SO93]. P (Y < X) [AFJ97]. p, q [dML93b, dML93a]. q [BSS93, Cha96a, Lim90, Mik93, RS96, dML93b, dML93a]. q + 1 [Kes92]. R [CC90, Ras98, aS90, MM93]. R/S [TTW99]. R^{k} [JF91]. R^{m} [Ken90]. R^{n} [Ses97]. S [MY96a, BA95, OR98]. s^{m} [HK94]. s^{n} [BA95]. s_{1}^{n1} × s_{2}^{n2} [LI98b]. S_{n} [BCFS92]. S_{n} × S_{n} [BCFS92]. T [Iwa97, BA94, HSS96, LP91, LMT97, SS96a, SK99, vTrWM96]. θ\ − gq = n + λΣ [EK97]. θ\ − νgq = n + λΣ [EEK97]. U [BS92, CS90, Chr92, DP92, Ios92, Jan94b, Mae91, Mae97, Mik94, Rag97, RvZ92, Sch98a]. υ = 9 [KVAN90]. V [Mae94, NM97, ST96b, HS92]. {v_{2} + 2v_{3}, v_{1} + 2v_{2}; k − 1, 3} [HH93a]. V_{p,q} [HK94]. φ [DPSvZ96]. x [AG98].

-analogues [BSS93]. -and- [HMM97]. -auswertung [Tho97]. -balanced [TvR97]. -ball [HS92]. -best [BA94, CADMM95]. -BIBDs [Kej96]. -codes [HH93a, HH96a, HW96b, HH96b, WZ96a]. -concurrence [SK92]. -contaminated [dlHF94]. -contamination [MP93]. -correct [Gui96]. -corrected [Kha90a]. -cube [Det94]. -cycle [BL96, BKF98]. -dependent [HP90a, Hai92, Nie92]. -designs [LMT97, vTrWM96]. -differences [Cha96a]. -dimensional [Rus93, CM98]. -distribution [Ruk93]. -distributions [Cha97b]. -efficiency [GRP99, NM97]. -error [BG98]. -estimated [Wie94]. -estimates [YZ97]. -estimation [Zha95]. -estimators [Arc99, CC90, Ras98, aS90]. -factorials [Cha96a]. -factorizations [LY96]. -groups [MPW97]. -hypergeometric [RS96]. -matrix [BS93a]. -means [GEGM99]. -methods [Sil92]. -minihypers [HH93a, HH96a, HM96]. -mixing [DPSvZ96]. -moments [MH98]. -norm [GS97b]. -optimal [Bag96, BD96b, Bis93, CM91, CK94, CC96, CF96, Cha90, CL97, DN97, Far91, Far92, Hei96, Jac90, Jac98, KKNK96, Lim90, MH95, PWS98, SS90a, SS91a, SW95a, SS97a, SK92, TY96, Udd96, Vil91]. -optimality [Det93, OR98, Jac92, ST96b]. -optimum [Mik93, SO93]. -perfect [BL96].

-PMD [Qin96]. -policy [BM93]. -projections [BD95a]. -ranks [Moo96]. -record [Ber95b, DN94]. -RNN [BMV97]. -simplex [Lim90, Mik93]. -smoothing [AS96a]. -stage [Liu97b, BMV97]. -statistic [Iwa97, DP92, Mae97]. -statistics [BS92, CS90, Chr92, Ios92, Jan94b, Mae91, Mae94, MS90a, Mik94, Rag97, RvZ92, Sch98a, Sho97]. -symbol [HSS97]. -symmetric [Kus99]. -test [Ma97a]. -type [Shi97]. -upper [HP90a]. -value [AG99, AG98, OD99]. -values [CW98]. -way [Che98b]. -wheel [CLR91a, CLR91b].

0-12-50846-X [AC97b]. 0-4702-3413-X [Smi97c]. 0-471-55239-9 [Fou97]. 0-471-55779-X [Law97]. 0-471-58100-3 [Dra97]. 0-471-82221-3 [Coa97]. 0-8247-9341-2 [Smi97a]. 0-940600-32-3 [Jen97a].

1 [BM93, SJG93, Wip98]. 1-110 [Ano90g]. 1-111 [Ano97u]. 1-120 [Ano94k]. 1-126 [Ano93i, Ano93j]. 1-129 [Ano95l, Ano96v, Ano96w]. 1-130 [Ano90h]. 1-132 [Ano92k, Ano95m, Ano96x]. 1-134 [Ano92l, Ano94l]. 1-135 [Ano95n]. 1-136 [Ano90i]. 1-138 [Ano91g, Ano93k]. 1-140 [Ano94m]. 1-142 [Ano91h]. 1-145 [Ano92m]. 1-156 [Ano92n]. 1-158 [Ano93l]. 1-160 [Ano94n]. 1-165 [Ano98x]. 1-168 [Ano99o]. 1-178 [Ano99p]. 1-180 [Ano98y]. 1-189 [Ano97v, Ano97w]. 1-190 [Ano98z]. 1-191 [Ano97x, Ano98-28, Ano98-27]. 1-192 [Ano97y]. 1-197 [Ano97z]. 1-200 [Ano99q]. 1-204 [Ano98-29]. 1-216 [Ano97-27]. 1-228 [Ano91i]. 1-264 [Ano99r]. 1-290 [Ano94o]. 1-292 [Ano99s]. 1-298 [Ano99t]. 1-380 [Ano98-30]. 1-402 [Ano99u]. 1-408 [Ano98-31]. 1-420 [Ano98-32]. 1-99 [Ano96y]. 105 [dML93a]. 111-205 [Ano90j]. 121-256 [Ano94p]. 127-270 [Ano93m]. 127-432 [Ano93n]. 131-252 [Ano90k]. 131-259 [Ano95o]. 131-262 [Ano96z]. 131-264 [Ano96-27]. 133-263 [Ano95p]. 133-271 [Ano92o]. 133-283 [Ano96-28]. 135-292 [Ano92p]. 135-352 [Ano94q]. 135-366 [Ano97-28]. 137-263 [Ano95q]. 137-270 [Ano90l]. 139-268 [Ano93o]. 139-270 [Ano91j]. 141-278 [Ano94r]. 143-270 [Ano91k]. 147-280 [Ano92q]. 157-293 [Ano92r]. 159-315 [Ano93p]. 161-392 [Ano94s]. 167-344 [Ano98-33]. 169-356 [Ano99v]. 179-368 [Ano99w]. 181-464 [Ano99x]. 191-382 [Ano98-34]. 191-386 [Ano97-29]. 191-388 [Ano97-30]. 193-384 [Ano97-31]. 193-385 [Ano97-32]. 193-386 [Ano98-36, Ano98-35]. 1983 [Shu95]. 199-390 [Ano97-33]. 1991 [Bas95a].

201-396 [Ano99y]. 203-400 [Ano95r]. 205-378 [Ano98-37]. 207-468 [Ano90m]. 21-30 [Ano92i, Ano92j]. 217-408 [Ano97-34]. 229-368 [Ano91l]. 253-372 [Ano90n]. 257-392 [Ano94t]. 261-392 [Ano95s]. 261-406 [Ano96-29]. 263-396 [Ano96-30]. 265-396 [Ano95t]. 265-400 [Ano95u, Ano96-31]. 269-400 [Ano93q]. 271-406 [Ano90o, Ano91m, Ano91n]. 271-414 [Ano93r]. 273-392 [Ano92s]. 279-412 [Ano94u]. 281-426 [Ano92t]. 285-438 [Ano96-32]. 291-394 [Ano94v]. 293-439 [Ano92u]. 295-428 [Ano92v]. 2nd [Fou97].

3-312 [Ano95v]. 3-486-23146-4 [Tho97]. 3-7908-0930-6 [Ter97]. 305-416 [Ano96-33]. 311-404 [Ano95w]. 313-430 [Ano95x]. 317-452 [Ano93s]. 34 [dML93a]. 353-426 [Ano94w]. 36 [KS11].

425-439 [Ano92w]. 49 [AG99].

56 [LT00]. 58 [Ano98q].

69 [AG99].

72 [Bed00].

A-efficient [MG95]. A-optimal [AM91, KS97e]. A. [Les97, LR97, Nel97, Wir97]. Aalen [PR93]. abelian [MPW97, RCX97, Xia98]. ability [Hut96]. abrupt [Hus99]. absence [CX98a]. Absolute [DF95, Sko98, BS98b, Mus91]. absolutely [HP91, Har92]. absolutes [Bai98]. abstract [vRR96]. Academic [AC97b, FF97]. Accelerated [MS92, CL92, LY95a, Tse94]. accuracy [BS97, cIC95, Koc98]. accurate [Loh93]. Acknowledgement [Gup90, Gup91, Gup92]. action [KC93]. acyclic [AMPT95]. adapted [KBH98]. Adaptive [Lie96, RV95, Sch91, SW96b, Wef97, Ano98q, Büh99, Eis94, Mac97a, MP97, RS97b, Sun97b]. adapts [DS99]. addition [Hoo90, JRM96, Ten91, Ten92]. addition-multiplication [JRM96]. additional [BH92a, GT93]. additive [KR93, Lan96, Wan98c]. additivity [SW98]. adjusted [BvB91, EJW93, GHRG^{+}99]. Adjusting [Kon99]. adjustment [Kak97]. adjustments [RM95]. Admissibility [Ber98a, Eat97, Hei91, Mal95, PL97]. Admissible [BM90, BMR95, Hei92, Zou97, HS92, Hof95, PR91]. admit [GH99]. admitting [Pad94]. affine [LT96, LT00, MT99, MP96, PRS97, Van97]. after [CW97b, Wan96]. against [BB91, BH92b, BH98, Cha97b, El 96, HLM90, SS92]. age [Ahm99, BBTV95, Bos97, Fin97, LP91]. age-dependent [Bos97]. ageing [MB96]. ahead [GM96]. aid [BT97]. AIDS [ED94]. Akaike [CS98]. algebra [Cha96c, Di 99, Tun98, Zhu96]. algebraic [WCP91]. algorithm [Cha95c, Mac97a, McL99, Mey99, Tju93]. algorithms [Böh95, FH91, YH90]. Aliasing [CX98a, CX98b]. Aligned [AA93, Man99, Akr91, Man98]. allocation [BHT91, BENP91, RS97b]. allocations [GM96]. Almost [Mik94, Pur95]. alphabetic [MMC94]. alternate [TM96]. Alternative [MR99, Akr91, Bel95, Cha97a, CMB95, SSE98, TSS96, WS95]. alternatives [BB91, BB99a, Lee98a, Man90a]. Alvin [Smi97b]. American [Jen97a]. among [Lya97]. amount [HB95]. analog [BG91]. Analogs [MH98]. analogues [BSS93, DM93]. analyses [GW94, Lav94]. analysing [PPO95]. Analysis [HK94, Osk95, RC98, Vos99, Ano97j, Ano97-35, AB97b, BABM99, Ber90, BM91, BRM94, BB99b, BA92b, Car99, Cha97a, Cha96d, DEM95, Den96, Den98, GNWK99, Ham97, HT92b, Jen97b, Kad94, Ken97, Khu98, KC93, Kun98, LH97b, LT97, MGdW96, Man90a, MBS99, MST91, MBR99, MB97b, NK94, Nay92, OS96, PMR97, RN95, RS92, Róz92a, Sei92, Sen95, Sen99a, Sil92, Smi90, Smi97c, Spe94, TR90, TG93, Tuk97, VY98, VW99, Wip98, Ye94, dLS99, dlHF94, Smi97b, Jon92]. analytic [Sho97]. analyzing [CS99c, Mau98]. Anant [Smi97a]. ancestors [DY94a]. Ancillarity [AP98, Ram90]. ancillary [Wan98b, vBLO95]. Anderson [Jen97b]. Annette [Iye92]. Announcement [Sri90b, Sri90c, Sri92b, Sri98]. ANOVA [DV97, DG95, KL99, PP94b, Van98]. answering [LR92]. appearance [Jar98]. appendix [EEK97]. application [Bis93, CMS93, DK94, DP91, Ein92, Fro90, GNWK99, JV97, Kad94, KC93, Li98a, Muk98a, Nan95, PPO95, PRW99, Rou90, SS95, Ste91, VM95, Zhe96]. Applications [BBTV95, DS99, MH92b, AP98, AB98, AC94, BBB96, BO99, BN91, BD94c, DM96a, ESV97, FZ98a, GB95, GSGMPS97, GMR96, HM96, HB94, Jan93b, Jan99, JKW92, KP97, KK95a, KS97c, KMMN93, LSP99, LDG98, MNWZ95, MBL97, MM93, Mat94, MKR93, MPSM94, Nie96, PP92a, Pfa94, RL99, Ruy94, SS90c, Sri94, Ste93, TD99, Ull96, Jen97b, Jen97a]. applied [AP99, MBS99, Seb98, SMS96]. Applying [Con96a, LM99]. approach [ASK97, AB97a, AM96, BE91, BD95a, BM93, BFV94, Cha97a, Cri92, FCA94, GP99, GLL96, HS93, HM97, Lav94, LS99b, Liu95a, MG96a, Mol96, MG98a, MW97, NRZ98, Pow95, QP99, RS96, SS90b, SJ93, SJG93, Sho97, dML93b, dML93a, AC97b, Shu95]. Approaches [SW97, Che96a, KP93b, NK95]. Approximate [Bal90, CW97b, Sch93, SJ96b, GH95, MP98, SC97]. approximately [LW97, Wie92, Wie94, WZ96b]. approximating [MG96b]. Approximation [KR95, DF94, Enn95, HYT98, Mae91, MM98a, MI90, MF96, MG98b, Seo94]. Approximations [GH96a, HW96a, CADMM95, CR93, Deh92b, Erk97, GJ98a, KT92, KM96, Mae94, MM93, MH97, RW95, UT97]. April [Ano92k, Ano93k, Ano94l, Ano94q, Ano95p, Ano96y, Ano97-29, Ano97z, Ano98-33]. arbitrary [AB98, BM90, BMR95, BGK95, Hof95, Sar95]. ARCH [BZ96, Haf98]. area [DDB99, SR99]. ARFIMA [Hau99]. ARIMA [KP93b]. arising [BSS93, LW99b, PK99]. arithmetic [GJ98c]. arm [LZ98]. ARMA [AM98b, Bro94, Ger94, Hau99, Mar96, Pol93]. armed [GC99, LZ98]. Arnold [Smi97c, Wes96]. Arnold-Ghosh [Wes96]. arrangement [Hay98]. array [LB92]. arrays [Bie96, FZ98a, FHKM96, HSS96, HSS97, Kag90, Man95, Muk98a, RCZ97, Siu97, SS96b, WC98a]. article [RP97]. artificial [CHF92]. Aspects [GPS98, BDG90b, Mak94, Mül95b, SS93b, Sri90e, Ter97, Wu94]. assays [DBG99]. Assessing [Ten92, JM98]. assessment [PFvR97, WC98b]. asset [O’H94]. assisted [CR97a]. associated [GEGM99]. Association [Jen97a, SS98b, WJ96, Ban96, CC96, JM98, Mei98, RS92, SH90, Tun98, WL96]. assuming [SCM94]. assumption [DMNN99]. astronomy [BF96]. asymmetric [BE91, Hua95, MH92b, Sil92, WC98a]. asymmetrical [Man95]. Asymptotic [Bar97a, BR97, BG91, CC90, CL95, CGZ98, CdlH96, DG91, DP92, Die95, FR99, FW93, Fuj93, Iwa97, JM96, JW92, Lie92, LR95, Lus94, MB95, Man90b, Mei98, MPSM94, MPV95, MC90, Nev95, NZ91, Ral92, Rat96, Ros97a, RFD97, Rou90, Ten98, vRR96, BG98, Ber98b, BS94, Bou90, Cai99, CFM^{+}98, CS99a, Cha90, CL93, CH95, Das93, Das94, Dat96, Fro90, FP98, GMD99, HP92, Hor95, Kre90, Lee98a, LM99, Luo96, Ma97a, MS90a, Mei97, MS90b, NR90, Pfa93, Ral93a, RS93, Tah95]. Asymptotically [Kim97a, MMY94, BH98, BA94, CB94, Fro90]. Asymptotics [Coa97, CL96, GEGM99, SC97, Böh96, DF95, HF95, SE95]. attached [Ban96]. attaining [Kag90]. attenuation [Kon99]. attraction [Alv95]. August [Ano92q, Ano93n, Ano94k, Ano95q, Ano95t, Ano96x, Ano96-28, Ano96-32, Ano97-28, Ano98-30, Ano99s]. August-September [Ano93n]. auswertung [Tho97]. Author [Ano90a, Ano90b, Ano91a, Ano91b, Ano91c, Ano92a, Ano92b, Ano92c, Ano92d, Ano93a, Ano93b, Ano93c, Ano93d, Ano94a, Ano94b, Ano94c, Ano94d, Ano94e, Ano95a, Ano95b, Ano95c, Ano95d, Ano95e, Ano95f, Ano96e, Ano96f, Ano96g, Ano96h, Ano96i, Ano96j, Ano96k, Ano96l, Ano96m, Ano96a, Ano96b, Ano96c, Ano96d, Ano97a, Ano97b, Ano97c, Ano97d, Ano97e, Ano97f, Ano97g, Ano97h, Ano97i, Ano98a, Ano98b, Ano98c, Ano98d]. autocorrelated [DM96c, Kar90, SB94a, WZ96b]. autocorrelation [Fuh92, KS99]. autocorrelogram [GGJ98]. automorphism [Ma96, MP96]. Automorphisms [Ben98]. Autoregression [HP98, DS97, HJPZ99, HTY98, Kou91, Kre90, LLL97, McK97, Qia98, Wef97]. Autoregressive [Bha99, BD94b, BS98c, BH99b, GL98a, GL98b, HJPZ99, HP98, HB93, HB98b, Lee98b, Lee98c, Lüt96, Mar97b, SSR94, SB94b, SB95b, Wan98a]. Auxiliary [PRF94, ASK97, Con96b, Llo98, Zha95]. availability [Ana99]. average [Ahm94, BH96]. averages [KS93]. avoiding [Zei96].

B [Ano97j]. back [RT90]. Bagchi [KS97a]. Balakrishnan [Tik97]. Balance [Cal93, CK92]. Balanced [BKM93, Dav94, KI98, Tan99, AD97, BCM98, Che98b, Dru99, FHKM96, Hyo92, HK94, IS98, Kes92, Kun98, MGdW96, Man98, Mon91, MU97, Muk98a, MB94b, PK99, SS93a, SK92, TvR97, Tun98, Udd92, Udd97, Van98, Ye94]. ball [HS92]. balls [GH95, MI90]. band [Bas95b, Tho97]. bandit [Che96b, GC99]. bandits [LZ98]. bands [CV92, CGZ98, Ein93]. Bandwidth [AL95, Chu95, Pat93, AM98a, BFM97, Fal92, Haz99, Kim97a, Lie92, Siv94, YT98]. Barnard [Goo90a]. barrels [DP91]. Barry [AC97b]. Bartlett [CF98, Kak97, RM95]. Bartlett-type [CF98, Kak97, RM95]. based [AS96b, AP97, Asa99, BA92a, BB99a, BLBL93, BE96, Bel95, BH92b, BR98, CGZ98, CM92, DP98, FS97, Fro90, GH98b, GL98b, GLL96, HJPZ99, Hay98, HR95, JK94b, KZ99, KS11, LK98, Lee98a, Lee98c, LR95, LB91, Mec92, NN97, PP92a, PT97, PP92b, RN95, RW95, Sar95, SS93a, SC97, Str97, TP97, TK91, TAT99, VS97, YT98, YB99, ZWF95]. bases [Mey99]. basic [Fed93, KS97b]. bathtub [HS97]. Bayes [AFJ97, BZ91, CP98, DG91, DG95, DDB99, DS98, Gas96a, GNWK99, Giu95, GSZ92, GW94, GH93, GL94, GLL96, GL99c, Hei93, Hu98, Hua95, JM93, KSDS95, LP91, LM95, LS95d, Lia90b, LH93, Lin97b, LS99b, Mai98, O’H91, PS90a, SJ96b, Sin95, TY96, VR99, ZdBPL90]. Bayes-Hermite [O’H91]. Bayes-type [JM93]. Bayesian [AHJ96, AB97b, BSV93, BE91, BABM99, BF99, Ber90, BM91, BM94a, BM94b, BX91, BH99a, BRM94, BB99b, Bos94, CL92, Cha93, Cha96d, CMB95, Cza94, DV97, DD94, Den96, Den98, DN96, DN97, ESA98, Esc95, FCA94, FR94, GGR99, GC99, GM96, Hay98, JK94b, KS90, Kad94, KT99a, KT99b, Lav94, Lec99, LH97b, LN99, Lya97, ML96, MGdW96, Mak94, MG96a, Mal98, MP93, MC95, MG98a, MBR99, MH95, MB97b, MW97, Nan95, Oga96, OS96, PST98, PP94b, Sal94, Sha99, Siv91, Siv93, STL96, TT92a, TZ93, TN91, TG93, VS97, WLW93, WS94, WC98b, Wip98, Ye94, YC95, dlHF94]. Bayesianism [Lev90]. be [BL96, Ham97, SMS96]. beans [JMVH97]. beat [SH97]. beet [Dou96]. behavior [BR97, HP91, JM96, Lee98a]. behaviour [BB95, CAGM98, CdlH96, GR90, KS92, NZ91]. Behnken [HJ98]. Behrens [MBR99]. beliefs [IM94]. benefit [Spe94]. benefit-cost [Spe94]. Berger [Jos96c]. Berman [Ass96]. Bernardo [Cox97, IS97]. Bernoulli [Csá96, Dut96, GC99, JK94a, LH97a]. Berry [BS92, Che97, Iso94, SH97]. Bessel [Fed93]. Best [BGK95, Slu92, BA94, CH92, CADMM95, DDB99, FCA94, Fro90, GM96, Hsu92, Lie96, MM93, MS92]. better [Ahm94, GH99, LW94]. between [AMPT95, FG96, JM98, LS95a, Pfa93, Ral92, Was96, dlHF94]. Beukenhout [CQ98]. Bhaskar [SS93a, Seb98]. bias [BSV93, CW97a, CS99a, FP94, MST99, MF97, Sha95]. Biased [BSV93, CK92, Eis94, PST97, Sha95, Yan96a]. biasing [FP94, Pak97]. BIB [KVAN90]. BIBDs [Kej96, Mor98, ZC92]. bicolored [TS97]. bideterminant [Pro93]. bidimensional [BM94a, MC95]. bifurcating [BH99b]. bilinear [BH92b, BH98]. binary [AM98a, cIC95, CS92, Cza94, HR97, JK94b, Siu97, TR90, WZ96a]. binomial [BG91, CL95, GM93b, KT99a, LH93, Ma97a, RJ98, ZL95]. binomials [CL95]. bins [HM97]. bioassay [KSDS95]. bioassays [MG95]. bioavailability [CH94]. biographical [Kat97]. biomedical [Li98a]. biometric [TD99]. Biostatistics [Lee99]. Birnbaum [Jos96c]. birth [Phe96]. bitableaux [Gho96]. Bivariate [DT91, BB91, BB99a, Bar97a, LB91, NJ99, SS95, Wan99]. blackjack [YK92]. Block [FZ98a, LMT97, AM91, AiFL99, AC99, BC93, BH92a, BC99, BKM93, Cal93, CK96, Car96, Che98b, Cut93a, Cut93b, Dav94, DBG99, GLMN98, GRP99, Jac92, Jac98, JR99, KM97, Kes92, KV98, KS97e, KS97f, LS99a, ME91, ME93, MEG93, MP96, MU97, OR98, Ras98, RJ99, SS92, SG91, TvR97, Tju93, Udd92, Udd96, VK99]. Block-crossed [FZ98a]. block-transitive [MP96]. blocking [DDN98, SCM94]. blocks [Bai90, FP94, GL99b, Man98, Stu91a, WN98]. Board [Ano90d, Ano90e, Ano90f, Ano91d, Ano91e, Ano91f, Ano92e, Ano92f, Ano92g, Ano92h, Ano93e, Ano93f, Ano93g, Ano93h, Ano94f, Ano94g, Ano94h, Ano94i, Ano94j, Ano95g, Ano95h, Ano95i, Ano95j, Ano95k, Ano96n, Ano96o, Ano96p, Ano96q, Ano96r, Ano96s, Ano96t, Ano96u, Ano97k, Ano97l, Ano97m, Ano97n, Ano97o, Ano97p, Ano97q, Ano97r, Ano97s, Ano98k, Ano98l, Ano98m, Ano98n, Ano98o]. Bock [Tho97]. bond [PPO95]. Bonferroni [CGZ98, HB99, MS93a]. Book [AC97a, AC97b, Ano97j, Ano98e, Ano98f, Ano98g, Ano98h, Ano99a, Ano99b, Ano99c, Ano99d, Ano99e, Bas95a, Chi92, Coa97, Dra97, Fou97, Iye92, Jen97a, Jon92, Law97, Sal92, Shu95, Smi97a, Smi97b, Smi97c, Ter97, Sri90c]. Bootstrap [Bab95, Fal92, FH91, RPK97, AiF97, AiFL99, BB98, BY95, CR93, DK99, Dat96, DS97, FS95b, Kul99, Lah97, LS95e, MM98a, Tan99, VM95]. bootstrapped [Mik94]. Bootstrapping [AF92, Bar97b, Fuh92, KSDS95, Nav95, VV98, ZL95, Jan94b, LR95, Sho97]. bootstraps [Gui99a, Gui99b]. Bose [Shr98, CQ98, Seb98, Tun98]. Bose-Srivastava [Tun98]. both [HR97, Kle91]. bound [BD96c, DL98a, HH93a, Ham98, Iso94, SS96b]. boundaries [Böh96]. Boundary [Erk97]. Boundary-mode [Erk97]. Bounded [LY95b, Zou92, Iso99, SH92a, Siv94, Siv97]. Bounds [Das94, KD99, KT92, Siv94, UT97, AHJ96, CP97a, Che97, DH92, GL98c, Hoo90, Kag90, KV91, MS93a, RS97a]. Box [HJ98]. Boyce [Smi97a]. Bradley [HZzY91]. Bradley-Terry [HZzY91]. branching [BD96a, Bas95a, DY94a, Kim98]. break [NL97]. break-point [NL97]. Breakdown [Mül95a, CRV96, Ruc97, VY98, Zha96]. breaking [GP94]. breast [dSDF95]. breeding [BT97, Dou96, HT96]. brief [Gov93, Kat97]. broad [HP92, Seo94]. Brownian [KW92, MR99]. Bruce [Jen97a]. Bruck [CQ98]. Buffon [RW98b]. Burgers [LW99b]. burn [Spi91]. burn-in [Spi91]. Burr [AFJ97, AHJ96]. Burr-type [AFJ97]. burst [BD96c].

C [Dra97, Shr98, Smi97c, Wip98]. c.d.f [AiF97]. CA [AC97b, Jen97b]. calculations [KT99a]. Calibration [FK98, BMS91, Hag92]. can [BL96, Ham97, SMS96]. cancer [KP97, dSDF95]. canonical [Ruk92, SA97]. capability [HL95]. Capacity [DL90, RP97]. carcinogen [KSDS95]. Carlo [DM96c, Hau99, Koc98, NK95, Pow95, TAT99]. case [AF92, CX98b, CdlH96, Dou96, Fer94, GJ96, GJ98c, HJL95, Hof95, MBL97, Mey99, Pfa93, PR91, SA97, Sen91, Ten92, WC95]. case-control [WC95]. case-deletion [HJL95]. cases [Bab95, NJ99]. Catalan [DM93]. catch [Smi90]. categorical [MPZ98, Nis95, Sim95]. Cauchy [DM96a]. causal [Rub90]. Causality [Dem90, AE93, Sup90]. cause [GUH91, MB97b]. caused [GL99a]. CBMS [Jen97a]. cell [BH99b, Khu90a, WB94]. cells [BA94, Wan98c]. censored [AB98, Cha95c, CP98, GB95, GB98, Gas93, HC92, JN97, KZ99, Mau98, MBS99, Osk95, PZ99, PT92, Ren97, Slu92, TP97, VV98, Yan93, Yan95]. censoring [CW97b, Ein93, SS90c, SL95, Wu94, YC95]. censorship [BCvdM92, Dik98, Hor98]. centered [CFM^{+}98]. Central [GS99a, Kim97b, Nie92, Yan93]. certain [BH92a, Höp97, LB92, SA97, Zin96]. certainty [SS90b]. chain [Wef97]. chains [AF92, BB99b, Eat97, Fuh92, Ios92, Pen91, Sko98]. Change [Fer94, AHP97, BR97, Bou90, GLS96, GH96a, GH98b, HK97, HC93, JMVH97, JM93, KS91, LZ98, Lya97, MK96a]. Change-point [Fer94, BR97, GLS96, GH96a, GH98b, HK97, HC93, JM93]. changeover [DLC99]. changes [Hor98, Hus99, JM91, WC98a]. changing [RN95]. channel [YS92]. chaotic [Far96]. Chapman [Chi92, Iye92, Jon92, Shu95]. Characteristic [Hyo92, GNZ97, Ram97, WJ96, WL96]. Characterization [DPSvZ96, GNZ97, JM99, KL99, Pak97, AS96b, Ass96, Bon97, BS98b, FS97, HH93a, HH96a, HM96, KS97b, KLP98, McK97, NRZ98, RMZ93, Wes96]. Characterizations [Arn97, BDG90a, CP97a, Kem97, AFJ97, Asa99, CK98, DCK95, Kam91, NN97, SK97]. chart [LPS98]. charts [BH96]. Cheng [Fro90]. Chernoff [AS96b, Nik91, Ral93b, RS97a]. Chernoff-Savage [Ral93b]. Chernoff-type [RS97a]. Chessboard [LW99a]. Chi [Law97, KH95]. Chi-squared [Law97, KH95]. choice [Pat93, AC97a]. circuits [Yin96]. circulant [KM96, KKNK96, Str98]. circular [HM92, KN94, Kun98, Pur95, SS94]. Circumstances [SMS96]. class [Ahm94, AH99, ADP91, BvB91, BB91, BN91, Bos94, Dre98, Gup98, HP92, Ios92, KR97, MPW97, MT94, PR91, Ruy94, Siv94, Sun97b, TG93, VP94, Wan98a]. Classes [MC95, MB96, RW95, Vos93, ZX98]. classical [BTV96, CM96, GR97, MM93]. Classification [BM99a, LT96, LT00, FS95a, FK98, JK94b, ON95, Smi97c, vBLO95]. clear [CH98]. clinical [Coa94, CW97b, Eis94, EMS98, ED94, FP94, FS94, LW94, Lee99, Sen94, Sim94, WRKL99, Wu94]. cliques [BEHW96]. clock [RT90]. close [Lia97]. Closed [FG94, GH99, BD96c, CM96, RH95]. closed-loop [BD96c]. closeness [Bos98, Kou95]. closerness [Jos96g]. CLT [CP97a, Seo90, Sho97]. Cluster [DK94, BF99]. clustered [Cro98]. clustering [CK99]. Cochran [TS95, Wan97a, WC99]. codes [Ass96, BD96c, Cha96c, HH93a, HH96a, HM96, HW96b, HH96b, Ham98, HW98, JV97, TW97, WZ96a, Zhu96, Zin96]. coding [HS98b]. Coefficient [Lee98b, Ber95a, Bro93, Fuh92, HP98, HB93, HB98b, KSM92, Oht97]. coefficients [BH99b, Kub90, Wan96, dLS99]. coherent [BENP91, PST97]. coin [BSV93, Eis94]. coinsertion [AJ93]. cointegration [TAT99, VM95]. cokriging [HB98a]. collection [MC95]. Collins [Shr98]. Colorado [Shr98]. colored [Nie93]. colorings [Mah98]. column [Bag96, DD92a, EJW93, Jac90, JPW97, Kun93, MT96, MU93, MU97, SS93b, SW98, Udd97, Wan98c]. columns [SK90, Udd92]. combinations [GP99, GGBT95, Ma98b, NZ91]. Combinatorial [SJ93, Bar98, BM93, BSS93, Zin96, dML93b, dML93a]. Combinatorics [Di 99, Siu97]. combined [Alb92]. Combining [Yan90a, GM93b, KT99b]. Comment [Bar90b, Car97, Wir97, Con96b, DEM95, Goo90a, Goo96c, Jos96e]. Comments [Cox97, Daw97, God90, Lia90a, JK94a, Lin97b, Lin97a]. common [Kub90, PL97, PT97, Sar94, Yan90a, YSS99]. communication [BM91]. compactness [CM95]. comparative [CH94, Pow95]. Comparing [CW98, LD93, Vos93, AM91, CW97b, Cut93a, Cut93b, GLMN98, GG92, GJ98b, GL99c, Liu97a, MT96, Stu91a, TY96]. Comparison [OD99, RM95, BF99, CHF92, FR99, LYC98, MS93a, Rot99, Sha99, Siv91, Yu96, Zou92]. comparisons [BH99a, CH92, DKR97, Gup98, GRP99, HT91, Hsu92, JR93, KH99a, Koc98, LS95a, MS98b, Ras98, Sen99a, Som99, TN91, VH98]. compartment [KM98b]. competing [KD99, WS95]. complement [Hoo90]. complete [Jos96a, KN98, KS92]. Complex [Ram97, CC96, CMB95, LW94]. Complex-valued [Ram97]. complexity [BQA99, Ger94, QGG96, QK98, WF94]. component [BDG90a, DGK97, GUH91, HT92b, LB91, Lyn99, WB94, XB91]. Components [DM91, BGK95, DKR97, EBS96, GL99b, HT92b, KWZ93, Mie90, MS95a, Mol96, SW97, Van98, VS97, Wit98, ZM93]. Composed [Tal98]. Composite [MN97, GAS96b]. compound [SCM94]. comprehensive [FS94]. Compressible [EEK97]. compressing [KN98]. Compromise [DMS92]. Compromising [Lev90]. computational [MM95, Oga96, Ter97]. Computer [CMB95, Sal92, Di 99, Par94]. Computing [BMS91, GH95, Hei96]. concealed [Fin97]. Concentration [FR94]. concept [Kam95a]. concepts [Cal93, Gro98a]. conceptual [KGS96]. Concerning [MY96b, Goo90b, HP90a, JV97, Khu92, Sri90c]. concomitants [KD90]. concurrence [SK92]. condition [Rag97, Vil91]. Conditional [Ruy94, AMPT95, Arn97, BE96, Ber98a, BJ90, Che96a, Che97, HB94, KL96, Kot90, LOS95, Liu94, LLL97, NRZ98, PW99, Pfa93, RMZ93, Yan95]. Conditionality [Jos96f]. conditionally [GL98a, MT94, Mül94]. Conditions [KV91, BRM94, FH96, GTL98, HP90b, LY95b, Ma97b, MS90a, RS93, RCX97, Rou90, VSB98, WM90, Wu98, vR93]. cone [Gly97, Gly98, Iwa98]. cone-restricted [Iwa98]. conference [Ano96e, Jen97a, Shr98]. Confidence [AiF97, CV92, Ein93, Fut99, GGBT95, Hol99, Joh91, KH95, BR97, BS97, CW97b, CGZ98, Fie97, GS99b, GL98c, Mol96, MS98b, PT97, Pfa98, aS90, TN95, Tak98, VW99, Wan98b, YSS99, ZL95]. confidentiality [Ten91, Ten92]. configuration [BS94]. configurations [GF91]. conflict [Was96]. conjecture [GS97a, Qiu96]. Conjectures [Bed98, Bed00]. connected [BH92a, BD91, KS97e]. connection [Eat97]. consensus [SS90b]. consequences [SS96a]. considerations [DL90, RP97]. Consistency [CADMM95, Chi91, Eks98, Hu98, LC97, SW96a]. Consistent [AM98a, GGR99, KK95c, Kot90, Ren97]. constancy [Lee98b]. Constant [GR93, Hei91, Siv94, Zin96]. constants [HP90b, Seo90]. constraint [Wan98b]. constraints [El 96, ER99, McK92, Mey99, MF97, MO96, ST93, Wit98]. construct [Seb98]. constructed [KS99]. constructibility [Far92]. Constructing [HB98a, Mon98, Ros92]. Construction [Bie96, Cha95a, JWT93, Kou94, Man95, MD94, SS90a, WD98, YIHN99, Cha90, GJ98b, HS93, Jac98, JRM96, MY96a, MU93, Muk98a, SS91a, WL96]. Constructions [Far91, SS93a, SK92, Udd92, AH96, Col96, Mor90]. Contagion [CK99]. containing [CH98, HP97, MPW97]. contaminated [LY97, Mül94, dlHF94]. contamination [BR98, MP93, SP91]. contaminations [Bos94, Das94]. Contents [Ano98i, Ano98v, Ano98w]. contiguous [HS98a, Lee98a]. contingency [PW99, RS92, Tom93].

Continuity [MH97, DF95, Sko98]. continuous [AC94, BG91, Bro94, Dat95, HR97, Kim98, MPZ98, RMZ93, TR90]. continuous-time [BG91, Bro94]. contoured [TS95, WC99]. contractions [Gro98b]. contrasts [GS99c, HT91]. contribute [Ham97]. Contribution [Nel97]. control [AM91, BH96, Cut93a, Cut93b, DR96, Gup98, GRP99, HT91, HS98a, HL95, Lia97, Liu97a, LPS98, MT96, PD99, Ras98, SJG93, SW95b, TY96, TN91, WC95]. controlled [MV96, TN98]. Controlling [LB90, YB99]. controls [BL99, GJ98b, KSDS95, TS91]. Convergence [BN99, GT99, LOS95, LC96, Add99, Arc99, Ber98b, Ber98c, BD94b, CS90, CM92, DPSvZ96, ESV97, HP90b, Kou91, KS92, Lia90b, Mik94, Mus91, ND98, PRW99, Rag97, Tho93, Tom95]. converges [Cha90]. convex [BRZ93, BD97, ER99, Ma98b, ST93]. convolution [SS90c, Pfa91]. correct [Gui96, GL98c, Liu94, Wan98c]. corrected [Kha90a, MH97]. correcting [BD96c, HM96, Pel96]. Correction [Ano90c, CF98, FS96]. Correlated [ALMS97, AGLS98, Bis93, CP95, Cut93a, Cut93b, KR93, Osk95, QP99, SB94b, UM91, YB99]. Correlation [BC99, BH96, CS99a, Fou99, LB92, SA97, Shu95]. correlations [MST91, Sen91]. corresponding [BDG90a]. Corrigendum [LT00]. Cost [Spi94, GHRG^{+}99, Mal95, Spe94]. cost-effectiveness [GHRG^{+}99]. Cost-efficient [Spi94]. costs [Ham91]. count [Dea93, DEM95, MG98a]. Countable [Kul99, AF92]. Counting [Ges93, CH90, Nie96, YK92]. courses [Wei99]. covariables [Man90b]. covariance [CL96, DEM95, Haw92, HOGF99, JSH99, NK94, Oma97, PFvR97, Smi97c, ST93, Ste97, Str97, WL93, Yok97]. covariances [Giu95]. Covariate [GHRG^{+}99, BA92b, Buz98, CM91, RN95, Spi94]. Covariate-adjusted [GHRG^{+}99]. covariates [CC99, HN99, HRR97, LY95a]. Cox [AS96b, BY95, CH90, Kon99]. Craig [Pro96]. credible [Siv97]. criteria [JD91, Lya97, SA97, SMS96]. criterion [Bos98, CS98, Chi91, Det93, Mal95, Nay90, NP93, Pol93, Tak98, TK92]. Critical [CMM94, Som99, FR99, FFR97, TTW99]. Cross [YT98, JWJB99, MG98b, OS96, Stu92]. cross-over [JWJB99]. cross-sectional [OS96]. cross-validation [Stu92]. Cross-validatory [YT98]. crossed [FZ98a]. crossing [Lab93]. crossover [BS98a, Kun98, Mat94]. Crowder [Jon92]. cryptology [Zhe96]. crystals [BDG90b]. cube [Det94]. cubes [GH95]. cubic [GH95, GH98a, HS92, Lim90, Mik93]. cuboid [Nan95]. cum [TSS96]. cumulative [BH96, Stu94]. curious [SC99a]. current [SK98]. curve [ALMS97, AGLS98, CW98, CGZ98, Fuj93, KSS97, KSM92, LN99, Mal98, PFvR97, Sil92, Smi90, Tru91, Wu98, Yok97]. curves [BFV94, CW97b, CGZ98, Fut99, KP97, KS97d, LH97b, Smi97a]. cusp [Arc99]. CW [Str98]. cycle [BL96, BKF98]. Cyclic [FHKM96, NG91, Ban96, CC96, GS97a, JWT93, Vos93].

D [Dra97, Jon92, Tho97]. Data [Ken97, LW94, Ano97-35, BRW99, BABM99, BD95b, BQA99, Cai99, Car99, CS98, Cha97a, Cha95c, CC99, CS99c, CC95, Con96b, Dea93, DEM95, Gas96a, GNWK99, GJ96, GL98b, GP94, GH96b, Ham91, Ham97, JN97, JK94b, KZ99, KH99a, LT97, Lie92, MNWZ95, Mau98, MBS99, MST91, MG98a, MB97b, Mus91, Nis95, ON95, OS96, PZ99, PT92, PRF94, PW99, Pur97, Ren97, Ros92, SS92, SS90c, SSR94, Sha95, She98, Sho97, Shu95, Sim95, Sin91, Slu92, Spe94, Ten91, TP97, Tho93, Tuk97, VV98, VY98, WS95, Wan97b, Wei97, Wu97, Yan93, YT98, dR97]. data-dependent [Lie92]. data-oriented [BRW99]. database [DKS98]. David [Ano97j]. Deal [PL97]. death [Phe96]. DeBruijn [Sta97]. decay [Phe96]. December [Ano92v, Ano93r, Ano94v, Ano95s, Ano97v, Ano97-30, Ano99r]. decision [Den96, Hay94b, HM97, HT92b, RV95, vDMS97]. decisions [Sal94]. Decomposability [BS93a]. Decomposable [BS93c]. decomposition [Tom93]. decreasing [Ebr97]. Default [MBR99]. defects [Mac97a, MR98]. Deficiencies [Alb92]. deficiency [Zou92]. defined [BL96]. Defining [Mah98, CK99]. definition [Jos96g]. Definitions [DCK95, Gui96, Jos96e]. deformed [GGJ98]. degree [DP99, KP93b, Mae91, NG91, TS97]. Dekker [Smi97a]. Delaunay [Ken90]. deletion [HJL95]. delivered [STL96]. densities [AC94, Dat95]. Density [BT99, RW95, Tra94, Arc99, BD96c, CP97b, CdlH96, Fal92, GMD99, Haz99, Iso94, Iso99, Kar91, Kim97a, LCKP95, PR96, Ren97, Siv94, Siv97, Stu92, SW96b, Ten98, Wer92, Wu97, Yan93, ZK98, dR97]. density-bounded [Siv97]. dependence [AARM99, AHP97, BH92b, BH98, BCFS97, GL98b, HLM90, Hay98, HK97, IT90, KN94, KD90, LA99, LH97b, ME91, Rou90, Tra90, Wu97]. dependent [Add99, AM98b, Bos97, BD94c, DT91, GA99, GLS96, HP90a, Hai92, HMT99, JSH99, Kim97a, Lie92, LY95a, Man90b, ME93, MEJ98, MJE98, Nie92, PRW99, She98, Sta92, Tho93, WC95, YT98, dLS99, dR97]. derivation [Akr91]. derivative [MMY94]. derivatives [Iso94, Iso99, LCKP95, dR97]. Design [DKS98, FMN99, JWJB99, MP98, Wu94, AIV95, AC97a, BS93a, BMS91, Ben96, CL92, Cha93, Cha90, CL98, CN97, Eis94, FZ98a, FMN98, HM95, Kot90, Li98a, Lim90, Mat94, McL99, Mik93, Muk98a, Mül96, OR96, Pat94, PP92b, PWS98, RS97b, Sch91, Smi90, Sri90e, Sri94, SC99b, STL96, TN91, TG93, VV98, Vil91, VH98, YIHN99]. designed [Dra97, Mül95a]. Designing [BHT91, McL99, ED94]. Designs [DDN98, HP97, ALMS97, AGLS98, AM91, AH96, AC99, AD97, BvB91, BC93, Bag96, BH92a, BD96b, BSV93, BK96, Bar92, BC99, BCM98, BS98a, BKM93, BGL97, BD91, Bis93, BD94c, Cal93, CK96, CK92, CM91, CK94, CC96, CF96, Cha95a, CH92, CH96, CL97, CM93a, CM93b, CL91, CH94, CH98, Che98b, CLR91a, CLR91b, Col96, Cut93a, Cut93b, DD92a, DMS92, Dav94, DLC99, DL98a, Det91, Det94, DN96, DN97, DBG99, Din96, DL90, DP99, Dru99, EJW93, ESA98, Far91, Far92, FH96, GH95, GH98a, GLMN98, GAS96b, GL99b, GJ98b, Gup98, GRP99, HS93, HS98a, Hei91, Hei92, HS92, Hei96, Hei98, HB95, HJ98, HH93b, HM92, Hyo92, HK94, IS98, IJ94, Jac90, Jac92, Jac98, JPW97, Jeo95, JWT93, JJMY95, JR99]. designs [JMY90, JKW92, KM97, KM98a, Kar90, Kes92, Key96, KN94, KVAN90, KN98, KW98, KSWyX97, Kou98, KS99, KKNK96, KS97e, KS97f, Kun93, Kun98, Kus99, KI98, LT96, LT00, LW99a, LI98b, LMT97, LS99a, LW97, LYC98, MFS98, MT96, Mal95, MH92a, Mar97a, ME91, ME93, MEG93, MEJ98, MJE98, MP96, MS97, MMY94, MT94, Mon90, MF97, Mor90, MU93, MU97, Mor97, MM95, MD94, MG95, MH95, MV96, Mul98, MG96b, MG98b, Mül93, Mül94, Mül95b, MMC94, NM97, OR98, PT94, PS98a, Par94, RP97, RN95, Ras98, RCZ97, RFD97, RJ99, SP97, SS93a, SS90a, SS91a, SS92, SO93, SW95a, Sch98b, SS97a, SS98a, Seb98, Sen94, SS93b, SH92b, ST96b, SG91, SK90, SK92, Spi94, Spr90, SM98, SH92c, SL96]. designs [SW98, Str92, SS94, Stu91a, Stu91b, SSSD99, Tal98, TW94, TvR97, TY96, Tju93, Udd92, Udd96, Udd97, VSB98, Vos93, Vos99, VW99, Wal96, Wan98c, WM90, Wie92, Wie94, WZ96b, Won94, WD98, YFHY96, Zin96, vTrWM96]. desirable [HM95]. Desmond [Car97, Les97, LR97, Nel97, Wir97]. Detecting [GL99a, Haw92, NK95, GH96b, LR92]. detection [DM91, Sza92, Sal92]. determination [AHP97, GJ98b, HT91, Jac98, Liu97a, RJ98]. determine [PS98a]. determined [AMPT95]. Determining [KZ99, AG98, AG99, HT92b, Khu96, KP93b, MW97]. deterministic [MMY94]. development [vDMS97]. developments [Hey92]. deviation [Bai98, BS98b, CS96, KS95, Mus91]. deviations [Deh92b, Ral92]. DeWet [Yan95]. diagnostics [Siv93]. diagonal [BH98, Du96]. diagrams [Fed93]. dialogue [IS97]. dichotomous [KH99a]. Diego [AC97b, FF97]. difference [And93, CC96, DJ97, EEK97, GS97a, HSS96, JV97, KP97, KSWyX97, Ma97b, MPW97, PRS97, RCX97, Ruy94, Van97, yXL96, Xia98]. differenced [SB95b]. differences [BR98, Cha96a, Seb98]. differencing [KP93b, NL97, TAT99]. Different [Gui96, SMS96]. dilation [WS94]. diluted [HZzY91]. dimension [LM99]. dimensional [AHT98, Bar90a, CM98, CMS93, Jos91, Mar97b, Mor90, ON95, Rus93, SS90c, SB95b, TN98]. dimensions [CW93, Mey99]. Dimitri [Sal92]. direct [Col96]. directed [AMPT95, BL96, Yan96b]. direction [CM98]. Directional [Iwa98]. directions [DY94b, Ham97]. Dirichlet [SE92]. discontinuous [Pfa98]. discount [LH97a]. discovery [BL99, YB99]. discrepancy [BRZ93]. Discrete [LB98, AC94, Csà93, DM97, GB95, GGT97, Joh91, Kem97, Kim98, KH99b, KLP98, MBL97, Mau98, MPV95, Rot99, Ten98]. discrete-continuous [AC94]. Discriminant [LT97, VY98]. discriminating [Lya97]. discrimination [BMV97, OR96]. Discussion [Les97, LR97, Gho97, Goo90c, Ham97, Nel97, Sha95]. disease [CK99, GNWK99, Mai98]. disjoint [Chu96, GGBT95]. disorders [Fer94, JM98]. disparity [BSV97]. Dispersion [GL90, CXL93, GG99]. distance [HHH94, JJMY95, JMY90, Mag92, Ull96, WWE95]. distances [MPZ98]. distinct [Udd96]. distributed [YC95]. Distribution [GB95, HLM90, LC97, AHJ96, AK93, AL95, Alv95, Bal90, BBB96, BES90, Bar90a, Bas95b, Ber98c, CM98, Che97, Che98a, CL95, Chr94, CG98b, CK98, Das93, Ebr97, EM99, Fin97, FS96, GSGMPS97, Gas96a, GS97b, HS97, Iwa97, Jan93a, KL99, Kem97, KH95, KS96, Kou95, KSDS95, KLP98, Lad95, Let97, LR95, Luo96, Lyn99, Man98, MM98a, McK97, MT94, NJ99, Oga98, PP92b, Pen91, PS90a, Ruk92, Ruk93, Sar93, Sul93, Sun97b, SWA97, Tah95, TN95, Ten98, TP97, Wei97, Wes96, YC95, dSDF95]. Distribution-free [HLM90, LC97, MT94, Sun97b]. Distributional [vdL96]. Distributions [JM93, SS99, AB98, Ahm94, AH99, ADP91, Arn90, ALW98, BJL98, BN91, BB97, BS97, BJ90, CP97a, CL93, Cha97b, CK99, DCK95, DH92, GB98, GR90, GM93a, GP98, GNZ97, HW96a, HMM97, JM96, JM99, Kam91, KT92, KH99a, KK95b, KL96, Lec99, LG91, Lia90b, Lin98, LB98, Ma97a, Ma98b, ML96, MB94a, MB96, Mon98, MPV95, NG91, Nev95, PZ99, RT90, RMZ93, SR90, Sar95, SA97, SS91b, Sun97a, UT97, ZX98, vBLO96]. distributive [AMPT95]. disturbances [DM96c, Kna99, LK98]. divergence [MPSM94, MPV95, Ull96]. divisible [AH96, BK96, KS97f, Mon90, SM98, Stu91a]. DK [Yin96]. do [Daw97, Gho97, IS97, Lin97a]. Dobson [Iye92]. does [SH97]. domain [Alv95]. Dominance [DL98b, MKR93]. domination [LW99a, Sar94]. dose [BDG90b, MW97]. double [MC90, VM95]. doubly [Eis94, Ren97, WZ96a]. down [BL99, FR99, GP98, HT91, Som99]. downdating [Cha95b]. Draper [RP97]. drift [KW92, Phe96]. driven [RFD97]. Dropping [TS96]. Duadic [Zhu96]. dual [Cha96c, GLMN98, MKR93, Mey99, MB94a, WZ96a]. Duality [FG96, BD95a]. dummy [NK95]. Duncan [Sha99]. Dunnett [Liu97a]. duration [Ahm99, Ame96]. dwell [YS92]. Dyck [Lab93]. Dynamic [LL95b, Hen96, MYTH96, PP96]. dynamical [MB95]. dynamics [Vin96].

E-optimal [CH96, Hei98]. each [Oht97]. early [PT94]. Econometric [DM96b, Far96, Ull96]. econometrics [Cha96b, Con96a, God96, Zel96]. economic [Cha96b, GJ96]. edge [FS96, JS95]. Edgeworth [AHT98, GSGMPS97, KL96, Mae97, SH97]. edition [Fou97, FF97]. Editorial [Ano98j, Ano99f, Ano99g, Ano99h, Ano99i, Ano90d, Ano90e, Ano90f, Ano91d, Ano91e, Ano91f, Ano92e, Ano92f, Ano92g, Ano92h, Ano93e, Ano93f, Ano93g, Ano93h, Ano94f, Ano94g, Ano94h, Ano94i, Ano94j, Ano95g, Ano95h, Ano95i, Ano95j, Ano95k, Ano96n, Ano96o, Ano96p, Ano96q, Ano96r, Ano96s, Ano96t, Ano96u, Ano97k, Ano97l, Ano97m, Ano97n, Ano97o, Ano97p, Ano97q, Ano97r, Ano97s, Ano98k, Ano98l, Ano98m, Ano98n, Ano98o]. eds [Jen97b, Ter97, Tho97]. Effect [AHP97, BA92b, BH96, Cha98, CS92, DK99, EL93, FP94, GT93, HK97, Khu96, Shi91, Ye94]. effective [Mül93]. effectiveness [GHRG^{+}99, JW92, Ten92]. Effects [WL93, AC99, AD97, BM99b, CM93a, CM93b, Das93, GL90, HP97, JKW92, Khu92, KV98, LN99, NK94, RH97, Rub90, SH92c, SC99b, SMS96, VK99, Yok97]. efficiencies [Das94, Jos91, Nik91]. Efficiency [JR93, She98, Ste97, WB94, BS94, BS95, Cal93, GRP99, LB92, Llo98, NM97, PR93, SCM94, SSE98, vdL96]. efficiency-robust [BS95]. Efficient [Büh99, Cut93a, Cut93b, GLMN98, GAS96b, Gui99a, HRR97, IJ94, JPW97, Pen91, Sch97, Bai90, MG95, Mül95b, RM95, Spi94, Stu91b]. EGD [SH90]. eigensystem [WL93]. eigenvalue [MH92b]. eight [MS97]. Einstein [Ben98]. Elation [BTV96]. electron [BDG90b]. element [Liu94]. elementary [Seo90]. elements [CM92]. eliminating [CS99a]. elimination [BS93a]. elliptical [Iwa97]. elliptically [TS95, WC99]. embeddable [MFS98]. embedded [TV98]. Embedding [Mon90]. embeddings [Chu96]. Empirical [AFJ97, DDB99, El 96, Hua95, RvZ92, Sin95, CGZ98, DEM95, DPSvZ96, DM95, GS99a, GH93, GL94, GLL96, GL99c, KSDS95, KS93, LP91, Lee98a, Lia90b, LH93, PZ99, PRW99, PS90a, Rag97, Ral92, Ral93a, Ruy94, Yan95]. Empirically [Vin96]. employing [MC90]. endpoint [EMS98]. endpoints [AiF97, KJS98, Sen94, WRKL99, ZK98]. enhancement [Koc98]. entropies [NG91]. Entropy [Ull96, CB94, GLL96, LR95, PN96, ZWF95]. Enumeration [CMS93, Fed93, TS97, HSS97, Nie93, WL96]. envelope [MST97]. environment [LG91]. environments [AM96]. epidemic [TX99, YC98]. epidemiology [DK94]. Epstein [FF97]. equal [IJ94]. equality [CM98, KS97d]. equation [DM96c, EEK97, EK97, GNZ97, SS96a]. equationally [BL96]. equations [Cha97a, Fei94, MBL97, MY97, QP99, Sch93, ST96a]. equineighboured [Str92]. equireplicated [CM91]. Equivalence [JD91, GF91, Pfa93, Wan99, Zin96]. equivalent [KKNK96]. equivariant [BD97, PC94]. Er/M/1 [Wip98]. Er/M/c [Wip98]. Erdos [Kat97, Sta97]. ergodic [Pen91]. Erlang [MB97a]. Errata [UT97]. Erratum [Ano98q, Ano98p, Bed00, KS11]. error [BG98, BS96, Buz98, DG95, HM96, Lie92, LTT93, LB90, MB95, Muk94, Oht97, Pel96, SSR94, Seo94, Spi94, SR99, Ten98, Wan96, Yan93]. error-correcting [HM96, Pel96]. error-in-variables [Muk94]. errors [AA93, BD96c, BZS96, CK92, Chu95, Cut93a, Cut93b, GA99, HB93, Kim97b, KS98b, LN99, MG96a, Muk98b, Sch97, SB94a, Sil92, Tru91, UM91, WZ96b, Won94]. errors-in-variables [MG96a]. Esseen [BS92, Che97, Iso94, SH97]. Esseen-type [Che97, Iso94]. estimate [CS99a, DS99, LOS95, Rou90, SMS96]. Estimated [Wan98b, Con96a, Fut99, Wie94]. estimates [Das94, Fal92, FS95b, Fre91, Fuj93, Hu98, Hut96, KK95c, Lie92, MPV95, Sch97, Sch99, Siv91, Str97, Van98, WB94, Yan90a, Yan96a, YZ97]. Estimating [AF97, BD96a, Bel95, BL92, Chr94, FZ98b, GUH91, Haf98, JS95, KP97, Kot90, Kou95, KM98b, MG98a, PR94, Sch98a, VM95, Wan96, BB98, Cha97a, CG98a, CM93a, CM93b, CH95, Des97a, Dur92, Fei94, FS96, GP98, GH93, JD91, LCKP95, Les97, LR97, LH97c, MH92a, MBL97, Mül95b, Nel97, QP99, SK98, SH92c, Ste93, SNN90, Tah95, TH99, Wir97]. Estimation [AK93, Alv95, Ana99, Bai98, BD95b, BS98c, cIC95, Cro98, Dut96, ER99, EA96, GMD99, HC92, Kub92, KMMN93, KS91, LK98, MYTH96, Mag92, MB97a, Nay90, PN96, PS90a, RC95, RS97b, Ruk93, SSS94, TA91, TK92, WR96, YS92, ALMS97, AGLS98, AC94, AFJ97, Ahm99, AL95, Ano98q, BDG90a, BRM92, BP93, BMR95, BGK95, Bal90, BA92a, BE91, BSV97, BBTV95, BG98, Ben96, BB97, BT99, Bro93, BS96, CW97a, CH90, CS99b, CR97b, Che98a, CC95, Coa94, CL96, DDB99, Dea93, Ein92, ESV97, FRS98, FP94, GSGMPS97, Gas96a, GL90, GSZ92, GC95, GP94, GM93b, HT92a, Haz99, He93, Hei93, Hey96, HRR97, Höp97, HC93, HB93, HB98b, IK98, Iso99, IT90, JF91, Joh91, Jos96b, Kar91, KW92, Kim98, Kna99, Kub90, KJS98, Lan96]. estimation [LA99, LS95d, LM99, LTT93, LN95, LY95b, LY97, Liu97b, LYC98, Loh93, Mad95, ML96, Mai98, Mal98, MP97, MY97, Mec92, MS98a, MST91, MO96, MS95b, Muk98b, MS93b, Mül94, Niy97, ND98, OR96, Oga98, Oht97, PP92a, PK99, Pat93, Pat97, Pen91, Phe96, Pol93, PWS98, PC94, RW97, RCGMC98, Róz92b, SSR94, SB94b, SB95b, SW96a, Sri90a, SS96c, SK99, SW97, ST93, Ste91, SR99, Stu92, SMS96, Tah94, Tak98, TP97, TT92a, TS92, TZ93, TG93, Tra94, Tru91, Vel96, VR99, Wan97b, WF94, Wer92, WD99, Wit98, WWE95, Wu97, Yan95, YT98, ZdBPL90, Zel90, Zha95, ZK98, Zou97, dHR93, dR97, Sal92]. estimator [BB95, BD97, Ber98b, Ber98c, BD94b, Che97, Das93, Dat95, GG96, GTL98, HF95, Kim97a, KR93, Kot90, KSDS95, Mae97, May90, PL97, PR93, Rat96, SSE98, SO92, TSS96]. estimators [Arc99, BM90, BA92a, BLBL93, BR98, BG91, BFM97, Büh99, Car96, CC90, CP97b, CRV96, DG91, DG95, DF95, DM96c, Dre98, Fer94, Fro90, GH98b, GL98b, Hau99, HHH94, Hof95, HK97, Iso94, Jos91, KHAH90, KT92, Kim97b, KS98a, KWZ93, KS96, KS93, KS11, LP91, LH93, Man90b, MST99, MC90, Pad94, PP92b, Pfa93, PR96, PR91, Qia98, RC98, Ras98, Ren97, RFD97, Ruk92, Sar94, Sch93, ST96a, SF97, aS90, SC97, Stu94, Ten98, UT97, VS93, Wef97, WN98, Wu98, Yan93, Yoh97, vRR96, vR93]. Ethics [FS94]. Eulerian [Jan93a]. Evaluating [SCM94, Lav94]. Evaluation [Oga96]. evaluations [SH97]. even [Van97, WZ96a, Yin96]. event [AF97, Pur97, dLS99]. events [Ber95c, CP95]. everywhere [Wan98b]. everywhere-valid [Wan98b]. evidence [KS98d, NK95, SS90b, Ano97j]. EWMA [SSK99]. Exact [BS98a, CK94, CW98, Kar90, Khu90a, KH99a, RJ98, Sei92, SL95, AM98b, Ber98a, DH92, KS96, MH97, TK91, WCP91, YSS99]. examination [FS94]. example [Cha90, LM99]. exceedance [Lea95]. exceedances [KR95]. excess [HKL98, Kou94]. exchange [BO99, Haf98]. exchangeable [CdlH96, LY97, Mon98]. exclusion [Hoo90]. excursions [Csá96]. exist [Daw97, Gho97, IS97, Lin97a, SO93]. Existence [Muk98a, Wu98, YGL96, ZZ96, GS97a, Kej96, Liu95b, Pel96, Qin96, SSSD99]. exogeneity [OS96]. exogenous [BS98c]. Expansion [VR99, FP98]. expansions [AHT98, BJ90, Fuj93, GSGMPS97, KL96, Mae97, SH97]. expectation [Vin96]. expectations [Asa99, Erk97, Kha97, LDG98, NRZ98, RMZ93, Siv94]. expected [Ahm99, GT97, GGBT95, LB98]. experiment [BMS91, Kad94, Koc98, SH92c, STL96]. experimental [BH92a, CH94, GG92, HM95, Hei91, Hei92, Mar97a, MMY94, PP92b, Sri90e, Sri94, TN91, TG93, AC97a]. experimentation [Llo98]. experiments [BT97, BHT91, CK96, Den98, FMN99, GL90, GT93, GG92, GT97, HP97, HB95, Jeo95, JJMY95, JR99, KT99a, LI98b, MM95, Muk98a, Mül95a, MP98, Par94, RJ98, RW98b, SP97, SW95a, Won94, YS92, Dra97]. explanatory [TR90]. explicit [Vin96]. exploration [LR96a]. Exploratory [MM95]. exponent [LA99]. Exponential [KS94, Ana99, BSV97, Bon97, CFM^{+}98, CW97b, DN97, DCK95, DKR97, Ham91, HC92, HB94, KHAH90, KS96, Kou95, KS98c, KK95c, LB91, ML96, MS92, MH95, PR91, Ses97, Sin95, SS95, Tah94, TP97, VS93, YC95]. exponentiality [BB91, Cha97b, GR97]. exponentially [Pen91]. expressions [Wan97a, WCP91]. Extendable [Key96]. extended [FS97, Wu98]. extension [Mey99]. Extensions [HR95, BS95]. extraneous [SO92]. extrapolation [Spr90]. extremal [Alv95, BN99, FH96, WN98]. Extreme [NR96, Bar90a, Ber95b, Chr94, Dre98, Hüs95, Rév95, SE95, Wei97]. extreme-value [Ber95b]. Extremes [Fer95, HKP95, BES90, Bar97a, Hai92].

F [LR97, Nel97, NL97, Smi97c, Wir97]. F-tests [NL97]. F25 [Jos96e]. F28 [Jos96c]. F42 [Jos96a]. F43 [Jos96b]. F44 [Goo96a, Jos96c]. F44A [Goo96a, Jos96d]. F44B [Jos96d]. F45 [Jos96e]. F46 [Goo96b]. F47 [Jos96f]. F48 [Jos96g]. F49 [Goo96c, Hut96]. F49A [Goo96c]. facility [DGK97]. factor [CM93a, CM93b, CH98, Gup98, HP97, LS99b, MD94, Ruc97, SP97, SC99b, Won94]. factorial [BD94c, DL90, GL90, GT93, HP97, Hyo92, HK94, JJMY95, JR99, KI98, Li98a, LI98b, MEJ98, MJE98, Mon91, NM97, RP97, SH92b, ST96b, SH92c, Vos93, YFHY96]. factorials [Cha96a, CM93a, CM93b, Shi91]. factorizations [LY96]. factors [Bar92, CX98b, DS98, GL99a, Giu95, HSS97, Jeo95, Lin97b, Man95, Mül93, SW95a, WD98, dP97]. Failure [RS98, AP98, AP99, Ahm94, AM96, Ana99, BB91, BABM99, Bos97, Cai99, DP91, GUH91, GGT97, KS91, MB97b, SB95a, SK98, SS91b]. Failure-time [RS98]. failures [Lyn99]. fall [KV91]. fallacy [Goo96b, Jos96c]. false [BL99, YB99]. families [Bon97, Cza97, GC95, GP99, Hua95, HB94, Kes92, KK95b, KS94, KS98c, KLP98, RH95, Ses97, Stu91b, Wet97]. family [Arn90, MFS98, MST91, MB94a, PZ99, PS90a, RT90, Sin95, Tah94, yXL96, Yok97]. Fang [Jen97b]. fatigue [DP91]. favor [PT94]. favorable [BS94, CB94, GF91]. Feasible [ST96a, Vin96]. February [Ano90l, Ano91k, Ano92p, Ano93p, Ano94r, Ano95w, Ano96-33, Ano99t, Ano99o]. Ferrers [Fed93]. fertility [ZW99]. few [CF98]. Feynman [Csà93]. Fibonacci [Zei96]. fiducial [MN97]. field [Róz92b, WJ96, WL96]. fields [AARM99, BT99, EA96, Hey96, KM96, LW99b, MG98b, PR96, Róz92a, SJ96b]. filling [KSS97]. filtering [BM98]. finding [LS99a, WM90]. Finite [CH90, Zhe96, AMPT95, Ben98, BB99b, BZ91, FS95b, FL91, Fuh92, HS98b, Kar90, Key96, MMM95, MS95b, Muk94, SSS94, TV98, WJ96, WL96, WF94, ZdBPL90, Zou97]. finite-intersection [MMM95]. finite-state [WF94]. First [SS98a, BH92b, CK94, CF98, DP99, Kub92, SH92a]. first- [DP99].

First-order [SS98a, BH92b, CK94]. Fisher [MBR99]. fit [BY95, Har99, Jan94a, Kha90a, KLP98, Sha95, SM96, Wan98a]. fits [ON95, QP99]. Fitting [BJ90, Wei97, HB98a, KS98b, Lad95, Mül96]. five [KM97]. Fixed [KW92, Mec92, TN95, BS97, CMS93, Col95, Das93, Hol99, Kla90, Mal95, MS98b, VV98]. fixed-cost [Mal95]. fixed-proportional [BS97]. Fixed-width [Mec92, TN95, Col95, MS98b]. flat [PR96]. flat-top [PR96]. flats [BA95, Li98a, LI98b, SL96]. flexible [Arn90, DDN98]. flows [Phe96, Rao94]. fn1 [GG99, KT99b]. fold [RH97, SR99]. forecasting [RL99, TT92b]. forecasts [BO99]. foreign [BO99, Haf98]. Foreword [PS92, PS95a, PS95b, Sri90d, Sri91, Sri92a, Sri92c, Sri93, Sri97]. form [GMR96]. Formal [Rub90, McK92]. forms [GT99, MR99, MS93b]. formula [Cha95b, Csà93]. formulas [PS90b, RS96]. formulation [Gui96]. Fort [Shr98]. forum [Goo90c]. foundations [Ham97, Tuk97]. four [BKM93, CW93, GL99b, KM98a, Kou98, KS97f]. Fourth [TK91]. fractal [AHT99]. fractional [AARM99, DL90, Hyo92, HK94, KI98, MR99, NM97, RP97, ST96b, TAT99, YFHY96]. fractionally [SB95b]. fractions [BA95, CX98a, CX98b]. frailty [dSDF95]. free [Cha95a, GP99, HLM90, HJ98, LS99a, LC97, MT94, Sun97b]. free-combinations [GP99]. frequencies [Khu90a]. Frequency [Sta92, GP98, Haf98]. Frequency-dependent [Sta92]. full [DM96c, Mik93]. Fuller [Fou97]. function [AK93, AL95, BRZ93, BD97, BMR95, BLBL93, Bar90a, BE91, BBTV95, BM98, Bou90, CC90, Die95, Dur92, Ein93, Gas96a, GH96a, GLL96, Hua95, IT90, KS99, Lin97c, Mec92, MS93b, ON95, PN96, PR94, PRW99, Pol93, Róz92b, Sch99, SO92, Ste97, Stu94, TZ93, Wat93, Wef97, Zel90, Zou97]. Functional [Iml95, ND98, Deh92a, GNZ97, Man90b]. functionals [Jan99, Pfa98]. functions [BD96a, BM90, BP93, BS92, CG98a, CH95, CP97b, Des97a, Dut96, Fed93, Fin97, FR94, Gho96, GR93, GP94, GNZ97, Gut93, Haw92, Iwa98, JD91, Kra93, LK98, Les97, LR97, LH97c, MM93, MMY94, Nel97, NZ91, Ome95, RS96, Ram97, Ron97, Ros92, Sar93, Sat93, SS99, Ses97, SNN90, TH99, Wir97]. Further [HM92, JK94a, KS97b, gL96, SS91a, KP93a]. future [Goo90b, Ham97, McL99, Zel96]. Fuzziness [GG92]. fuzzy [LDG98].

G [Jen97a, Ter97, Tho97, Won94]. G-optimal [Won94]. G. [Goo90a]. gambling [FF97]. gamma [LS95b, LS95c, MB97a, TN95]. Gauss [BM90, BP92, BRM92, BP93, BMR95]. Gaussian [And93, DM91, GA95, Hüs95, KM96, LA99, LPS98]. Gaussianity [RW98a]. General [AS96b, BS98b, AH99, AP97, BM90, BP92, BP93, BMR95, BM99b, BD95a, Böh96, Bos97, CM93a, CM93b, Chu95, DS99, Dre98, Fin97, GF91, GP98, GSZ92, GC95, GP99, JM96, LS95e, Loh93, LPS98, ME91, NR90, PR91, Pur97, SO92, Wan97a, XB91, dP97]. general-scale [GC95]. generalised [HMM97, MM98b, WD99]. Generalization [Qia94, Bon97, SC99a]. generalizations [CP97a, Pro96]. Generalized [AJ93, SR90, SJ96a, Sen91, TV98, WZ96a, AP98, Bal90, BRM94, BTV96, Cha97a, Cha96a, CR97b, CC99, CW98, CS90, Cza97, DM96a, DGP97, GMR96, GPS98, HP98, HB98b, IK98, Iye92, Jan93a, JK94a, Kam95a, KT92, Lab93, LL95b, Man90b, McL99, MC90, NR96, Nay90, ON95, Pay98, RCGMC98, RS92, SS93a, SM98, UT97, Vos93]. generated [AM96]. Generating [HM95, Kra93, Sat93, BLBL93]. Geometric [GJ98c, CK98, CAGM98, GB98, MY96a, Wes96]. geometries [Bar98, Ben98, Key96, Ma96]. geometry [CG98a, Jen97a, PST98, WL96]. Ghosh [Wes96]. Gibbs [SJ96b]. Gibbs-Markov [SJ96b]. Giuseppe [Bar98]. given [MMY94, Nie93]. GLMs [GNWK99]. global [Chi91, DF94, Fut99, KH99b, PST98, Yoh97]. GLR [Bou90]. GM [SC97]. golf [CN97]. Good [Spr90, FG94, Lin90]. Goodness [BY95, Sha95, Har99, Jan94a, Kha90a, KLP98, SM96, Wan98a]. Goodness-of-fit [BY95, Sha95, Har99, Jan94a, KLP98, SM96, Wan98a]. gradients [KD99]. Gradual [Hus99]. graph [BEHW96, Mah98, Wal96]. Graphical [BT97, Nis95, Ros92, JV97]. graphics [SC97]. graphs [AMPT95, LW99a, LY96, TS97]. Graybill [PL97]. great [BBTV95]. Greenwood [Law97]. grids [RT97]. Griesmer [HH93a, Ham98]. Group [KC93, PT94, AH96, BK96, Ban96, Cha96c, CX98a, Cri92, KS97f, MP96, SM98, Stu91a, VP94, YH90, ZZ96, Zhu96]. group-divisible [KS97f]. grouped [BD95b]. groups [LD93, Ma96, MPW97]. Growth [BFV94, ALMS97, AGLS98, BG91, CW98, Fuj93, Joh91, KSM92, LH97b, LN99, MH95, PFvR97, Sil92, Wu98, Yok97, Smi97a]. guarded [KS98d]. Guiard [Tho97]. guide [Law97]. Gumbel [LB91]. gun [DP91]. Guttorp [Bas95a].

H [AC97a, Dra97, Shu95, Smi97c]. Hadamard [Bed00, Bed98, DJ97, GS97a, Kou94, RCX97, yXL96, Xia98, Yam96]. half [Che95]. half-space [Che95]. Hall [Chi92, Iye92, Jon92, Shu95]. Hand [Jon92]. Handbücher [Tho97]. Härdle [Ter97]. harmonic [EBS96]. Harmonizable [Rao94]. having [Jac90, Man95]. Hayward [Jen97b]. hazard [GP94, IT90, KD99, Pat93, Pat97, PR94, Rou90, SS99, Stu94, dSDF95]. hazardous [SW98]. hazards [BCvdM92, Buz98, Kon99]. health [JM98]. heavy [ALW98, MS98a]. Heidelberg [AC97a, Ter97]. height [KP93a]. Hellinger [WWE95]. Hermite [O’H91]. Hermitian [GR93, Moo96]. Herrendörfer [Tho97]. heterogeneity [BS93a, TN91]. heterogeneous [DKR97, HMT99, SF97]. heteroscedastic [AA93, DMS92, GL98a, Khu92, KW98, LLL97, Sch97, Sch99, Won94]. heteroscedasticity [GMR96]. hidden [EA96, Hey96]. hidden-Markov [Hey96]. Hierarchical [DG95, GNWK99, Mai98, DG91, DS98, Esc95, MK96b, MP93, RH95, Sch98b]. High [vTrWM96, AHT98, Haf98, Lea95, NM97, ON95, SH92b, dHR93]. high-dimensional [ON95]. high-frequency [Haf98]. Higher [Kak97, JW92, SC99b]. Higher-order [Kak97]. Hilbert [Gho96, SNN90]. histogram [HM97]. histograms [CT96, Tra94]. historical [KSDS95, TS91]. history [CS99c]. HIV [TX99]. Hodges [Nik91, KS95]. Hodges-Lehmann [Nik91]. holes [YGL96]. homogeneity [CK92, Hay94a, JSH99, JK94a, JK97, MMM95, QGG96]. Homogeneous [Chu96, DMNN99, DP98, GLL96, PR96]. homoscedasticity [GG96]. homosexual [TX99]. honest [Tuk97]. Horvitz [Ber98b, Ber98c]. Hotelling [Iwa97]. hybrid [dKgL96]. Hyperbolic [MG98b]. hyperboloid [CM98]. hypercube [Par94]. hypercubes [Mul98]. hypercubic [SH90]. hypergeometric [GR93, RS96]. hyperplanes [KK95a, Sat93]. Hypotheses [HMT99, ZM93, BL99, HW96a, HB99, Iwa98, LS95b, LS95c, MPSM94, RH95, Sei93]. Hypothesis [Kin96, PT92, AP97, BM94b, Cai99, PT94, PPO95, RCGMC98, Seo90, Sin95].

i.i.d [LS95e]. ideals [Cha96c, Ste93]. idempotent [AZZ96]. identical [BABM99]. Identifiability [FL91, CL98, Pfa94, YS92]. Identification [EM99, SWA97, LW99b, Wan98c]. identifies [Mac97a]. identifying [SH92c, SC99b]. identities [GR93]. identity [BBB96, CL91, Pro93]. IFRA*t [ZX98]. II [AGLS98, Eub99, PT92, Ram97, WC99]. III [CH98]. ill [Hag92]. ill-posed [Hag92]. Image [CM97]. images [Ken97]. imaging [BDG90b]. imbalance [Khu96]. Implied [BP92]. importance [BENP91, XB91]. important [GH90]. imposed [KK95c]. imprecise [IM94]. Improper [Wet97, Was96]. Improved [AS96a, GC95, Llo98, Hoo90, IK98, SW95b]. Improvement [Tak98]. Improving [KSM92, Sch99, TK91]. impulse [Lüt96]. in-control [BH96]. Inaccuracy [KS95]. inactive [DGK97]. inadequacy [GPS98]. Inadmissibility [KR93, KS93, KS11]. inclusion [Hoo90]. inclusion-exclusion [Hoo90]. Incomplete [CC99, DBG99, ME93, Sen94, AM91, BC99, BQA99, CS98, Che98b, Col96, GJ96, Ham91, Kes92, Man98, ME91, MG98a, MU97, MB97b, OR98, Ras98, TvR97, Udd92, Udd96]. Incorporating [SSR94, TS91]. incorrect [Ste97]. increasing [Ahm94, BB91, DY94a, Hay98, SS90b, Stu94, SW96b]. increments [SE95]. indecomposable [Din96]. independence [AMPT95, BL99, BM94a, HJPZ99, Mei97, PW99, Ram90]. independent [BES90, GH90, LY95b]. indeterminacy [Lev90]. Index [Ano91c, Ano98r, Ano98s, Ano98t, Ano98u, Ano99j, Ano99k, Ano99l, Ano99m, Ano99n, Ano90a, Ano90b, Ano91a, Ano91b, Ano92a, Ano92b, Ano92c, Ano92d, Ano92i, Ano92j, Ano93a, Ano93b, Ano93c, Ano93d, Ano94a, Ano94b, Ano94c, Ano94d, Ano94e, Ano95a, Ano95b, Ano95c, Ano95d, Ano95e, Ano95f, Ano96e, Ano96f, Ano96g, Ano96h, Ano96i, Ano96j, Ano96k, Ano96l, Ano96m, Ano96a, Ano96b, Ano96c, Ano96d, Ano97a, Ano97b, Ano97c, Ano97d, Ano97e, Ano97f, Ano97g, Ano97h, Ano97i, Ano98a, Ano98b, Ano98c, Ano98d, BB95, BB98, Ber95b, Dre98, HSS97, HL95, WN98]. indexed [DM95, KS92]. indices [BP95]. indifference [Gui96]. indirect [Wan97b]. indistinguishable [MI90]. individual [Oht97]. induce [KP93b]. induced [GP98, HOGF99]. Inequalities [Kam91, CS90, NR90]. inequality [AS96b, El 96, RS97a, Wan96]. inequivalent [HKL98]. infection [TX99]. Inference [AG99, Ano98q, ALW98, Bed00, BH99b, Fei94, KS11, LT00, LB91, SB95a, Sin91, dML93a, vBLO96, AS96a, Adk94, AB98, Bas95a, BN91, CMB95, CS92, Cza94, DY94a, GGR99, Hey92, HB94, JR93, Jos96f, Lea95, LY95a, LS95e, Lus94, MB95, MH97, Nav95, PS90b, Pur97, Rao94, Ron97, Rub90, SB94a, SS95, WS94, YC98, vDMS97]. inferences [KS95, TK91, VS97]. inferential [vBLO96]. infinite [Bag96, Cha90, Kre90, MFS98, SS90c]. Infinitesimal [RW95, BJL98]. infinity [Arc99]. influence [BRM92, HJL95, PFvR97, Ron97, WC98b]. influential [GH96b]. Information [Iro93, JKW92, Lan96, NP93, RW98b, ASK97, Bha91, BA95, CK96, CS98, CHF92, DM96c, EMS98, GG92, GJ96, GT97, GUH91, Gut93, Hyo92, LB90, Mar97a, MMY94, MO96, MP98, Sha92, SO92, ZC92, Zha95]. Informational [MPZ98]. informative [Daw97, Gho97, IS97, Lin97a, Wu94]. inherent [Nis95]. initial [SS99]. injection [Phe96]. innovations [BL92]. inputs [CG98b]. inspection [GL94]. Institute [Jen97a, Jen97b]. Instrumental [BS96, Car96]. instrumental-variables [Car96]. insufficiency [Jos96e]. integer [JWT93, Nie93]. integrated [LCKP95, Lie92, MYTH96, SS97b, Ten98, Yan93]. integration [LDG98, RS96]. intensity [DMNN99, Róz92b, SK98]. intensive [CMB95]. inter [CK96]. inter-block [CK96]. interacting [SW95a]. interaction [Sch98b, VH98, ZC92]. Interactions [dP97, CM93a, CM93b, CH98, CX98a, HP97, SH92b, SC99b]. interactive [Dav94]. interarrival [MB97a]. interblock [EMS98, Sha92]. intercept [CH96]. interim [Sen99a]. Intermediate [CM95, Ein92, GT95]. Interpolating [Gas93]. interpretation [DS99]. interrelations [Coa97]. Intersection [BHW97, BGL97, HS95, MMM95, vTrWM96]. Intersection-sets [BHW97]. Interval [TP97, Col95, IK98, KW92, Mec92, TN95]. intervals [AiF97, BS97, CW97b, Fie97, GS99b, GGBT95, KH95, Mol96, MS98b, Nan95, VW99, YSS99, ZL95]. Intrinsic [LS99b]. introduction [Iye92, Fou97]. introductory [Wei99]. Invariance [GR97, Ban96, DD92b, HP90a, Hai92, Har92, KR97, Pak97, Ste91]. Invariant [HS92, DR96, GH95, GH98a, GM93a, GSZ92, GC95, JM99, Let97, Mad95, Ram90, Wer92]. inverse [CLH98, CHH99, Dut96, Fou99, GA95, Hag92, LN95]. investigation [SS94]. investigations [Tun98]. involving [Arn97, LC97, Shi97]. Irrelevance [Jos96b]. ISBN [Coa97, Dra97, Fou97, Jen97a, Law97, Smi97a, Smi97b, Smi97c, Ter97, Tho97]. isomorphic [BK96]. Isotonic [TW94, FRS98, LM95, PS90b, Qia94]. issue [Ano96e]. issues [ED94, IM94, KGS96]. iterated [Deh92a, Iml95, LDG98]. IV [Bar92, CH98, KI98].

J [Ano98q, Bed00, Chi92, Iye92, Jon92, Smi97c, Tho97, dML93a]. J. [Cox97]. jackknife [Mae97]. Jana [Coa97]. January [Ano90i, Ano91h, Ano92l, Ano93l, Ano94m, Ano98-28, Ano99x]. Jeffreys [CB94]. jeopardy [TM96]. job [Bel95]. John [Smi97b]. Joint [Che98a, Sar95, AC94, BBB96, Lyn99, MM98a, Oga98]. Jones [Shu95]. José [IS97]. Joseph [Ano97j]. Journal [KS11, LT00, AG99]. July [Ano90m, Ano91m, Ano92m, Ano93i, Ano94s, Ano95n, Ano96-30, Ano98z, Ano98-34, Ano99w]. June [Ano90j, Ano91j, Ano92s, Ano93q, Ano94n, Ano95x, Ano96v, Ano96z, Ano97-32, Ano98-27, Ano98-36, Ano99p]. Jurecková [Coa97].

K-stationarity [CT98]. Kac [Csà93, Hor95]. Kadane [Ano97j]. Kaplan [LOS95, Stu94]. Kariya [UT97]. Kazakos [Sal92]. Kelley [GC97]. Kernel [GGJ98, IT90, Wu97, AL95, AM98a, Che97, DM96a, Fal92, Haz99, Kar91, Kim97a, KJS98, Lie92, LC97, PR96, Ren97, Rou90, Ten98, Yan93, ZK98]. kernels [JW92]. Kiefer [DP99, Mec92]. Kirkman [gL96]. Knockout [HZzY91]. known [BHT91, SS99]. Kolmogorov [DM91, Kha90a, MS90b]. Koul [Yan95]. Koul-DeWet-type [Yan95]. Kreweras [Sul93]. Krzanowski [Smi97c]. Kshirsagar [Smi97a]. Kumar [Coa97]. Kurt [AC97b]. kurtosis [BGK95].

labels [Jos96b]. labor [Bel95]. lacuna [Jos96g]. lag [Luo96]. Laha [FS97]. language [Zei96]. Large [Deh92b, HB94, dK97, SB94a, SF97, CS96, Chr92, Jeo95, dKgL96, KS95, Man95, Ruy94, Yan96b]. largest [BHT91, GM93b, HL95, SH92a]. last [Khu90a]. latin [Bai90, CMM94, FFR97, MT94, AZZ96, Du96, Par94, UM91]. Latin-hypercube [Par94]. Lattice [PW99, SJG93, Böh96, BSS93, CMS93, Ges93, Gho96, KP93a, KC93, Kra96, Mar96, Mar97b, MKR93, Nie96, Sat93, SB95b]. lattices [AMPT95, Gov93]. law [Chr92, Deh92a, LS99b, Ral93a, SK98]. laws [DN94, Iml95]. layout [PT92, Sch91]. layouts [BDG90a, BGK95, BS93c, ME93]. leading [MB96]. learning [YK92]. Least [BS94, Bai98, BD94b, CS99a, CB94, GF91, GTL98, Jan93b, KT92, May90, Qia94, Sch99, UT97]. least-square [Qia94]. least-squares [BD94b, Sch99]. lecture [Jen97b]. Leffler [Lin98]. left [TZ93]. Lehmann [KS95, Nik91]. Lehr [Tho97]. Lehr- [Tho97]. lemma [Fro90, Kha97]. length [BH96, Pak97, SSK99, Sha95]. length-biased [Sha95]. length-biasing [Pak97]. Level [WC98a, DL90, HP97, JJMY95, Lea95, MEJ98, MJE98, RP97, YIHN99]. levels [Man95, SS96b]. Lévy [KS94]. lie [Mac97b]. life [Ahm94, AK93, BE91, BBTV95, CL92, CS99b, CS99c, LP91, LY95a, LB91, MB94a, MB96, Mus91, RT90, Tse94, Zah91, ZX98]. Lifetime [LG91, Ebr97]. lifetimes [CV92]. light [YFHY96]. Likelihood [BM98, CFM^{+}98, Ma98a, PT97, AM98b, Bal90, Böh95, CH90, CL95, Dea93, Des97a, El 96, FS95a, FP98, GL98b, Hau99, Hey96, Iwa98, JSH99, KM96, KBH98, KS96, LK98, Les97, LR97, LM99, Lin97c, Liu95b, Ma97a, MST99, MBL97, Mau98, MST91, Nel97, PZ99, Pfa93, Qia98, Rat96, RFD97, Ros92, SB95b, WD99, Wir97, vR93]. Likelihood-based [PT97, GL98b]. likely [AF97]. LIL [Sho97]. Limit [Bar90a, DN94, Ein92, HS95, BES90, GS99a, Kim97b, KS92, Nie92, Yan93]. limitations [Erd98]. Limited [MO96, BM91, DM96c]. Limiting [HP91, Man98, BB95, Lad95]. limits [PT97]. Lin [RP97]. Lindsay [Jen97a]. Line [CM96, SW95b, Sza92]. Line-closed [CM96]. lineage [BH99b]. Linear [Hei93, HJ98, Ruk92, ALMS97, AA93, BRZ93, BM90, BDG90a, BP92, BMR95, BA92a, BM99b, Ben96, BZ96, Bis93, BL92, CM96, CT96, CC96, Cha95a, Cha96d, CS91, CC99, CU92, Col95, Cza97, DMS92, DDB99, Dav94, DP92, DGP97, DH92, ESV97, FW93, FZ98b, Fre91, GS99a, GJ98b, GSZ92, GW94, Gro98b, Gro98a, GGBT95, GH90, GS99c, HJL95, Ham98, HP90b, HP97, Hei91, Hen96, Hof95, HH93b, HL94, Hua96, Iye92, Jac98, JM91, JR93, Jar98, Kna99, KS96, KM98b, Lee98a, LL95b, LTT93, Loh93, MNWZ95, Man99, Mar97a, MEG93, MM98b, McL99, Mei98, MO96, Muk98b, Mül94, Mül95b, NR90, Nik91, NZ91, ON95, Oht97, QK98, Rei99, RCGMC98, SO93, Sch97, Sch99, SS98a, Sei93, SB95a, SM92, SB94a, Sin95, SS96b, SS96c]. linear [ST93, Ste97, SM96, TA91, Tho93, TS96, VSB98, VS97, Wie92, Wie94, WZ96b, vRR96, vR93, AC97b]. linear-optimal [HH93b]. Linearity [HYT98, Har99, RW98a, Rio98]. Linearization [WLW93]. LINEX [GG96, Hua95, Zou97]. link [CS92, Cza94, Cza97, ON95]. List [Ano97t]. lists [EEK97]. load [Lyn99]. load-sharing [Lyn99]. Local [Kre90, PFvR97, Vil91, Yoh97, Bou90, DM95, Haz99, HYT98, Mül96, PST98, Ral93a]. Locally [BH98, SP91, DR96, FMN99, Hüs95, YZ97]. location [Ano98q, CFM^{+}98, CXL93, CAGM98, GGR99, Haw92, HMW91, KP97, KM99, Kou95, Lee95, MP97, Nay90, Oga98, PWS98, PC94, RW97, Ruk92, Shi92, TS92]. location-scale [PC94]. locations [Das94]. log [CU92, ER99]. log-convex [ER99]. log-linear [CU92]. logarithm [Deh92a, Iml95]. logic [FF97]. logically [HB99]. logistic [Bal90, FL91, KL96, MBL97, MMC94, RC98, RC95, SO93, SS97a, WC95]. logit [CU92]. loglinear [MK96b]. lognormal [PR94]. logrank [SL95]. London [Chi92, Iye92, Jon92, Shu95, Smi97c]. Long [BCvdM92, Add99, AARM99, DT91, GA99, GS99a, HK97, KS98b, LA99, LR96a, Muk98b, RL99, RZ98, Wan97b]. long-memory [Muk98b, RL99, Wan97b]. long-range [AARM99, DT91, GA99, HK97, LA99]. long-range-dependent [Add99]. longitudinal [AE93, Cha97a, SSR94, Shu95]. look [GM96, Sta97, TTW99]. loop [BD96c]. Lorenz [CGZ98]. Loss [Bha91, BD97, BE91, GG92, GG96, Hua95, KM99, Kub90, Mak94, PN96, Sin95, Zou97]. loss-robustness [Mak94]. low [BD96c, BDG90b]. low-density [BD96c]. low-dose [BDG90b]. lower [Bar90a, GL98c]. lower-upper [Bar90a]. Lukacs [FS97]. LV [MEG93].

M [BM93, Cox97, IS97, Jon92, SJG93, Smi97a, Ter97, Wip98, BM93, SJG93]. M/M/1 [BM93, SJG93]. Machine [YK92]. Magic [Eva96, JRM96]. Main [GT93, Shi91, Cha98, CM93a, CM93b, HP97, Nie92]. main-effect [Cha98]. maintenance [Sin91]. Majorization [KK95a]. makes [SW98]. making [GRP99, RV95, vDMS97]. Mallows [Chi91]. management [O’H94]. Mann [Di 99]. MANOCOVA [Sen99c]. MANOVA [BRZ93, aS90]. Many [TvR97, KH99a]. many-to-one [KH99a]. map [CQ98]. mapping [BN91, GNWK99, Mai98]. Marcel [Smi97a]. March [Ano90o, Ano91n, Ano92u, Ano93s, Ano94u, Ano95m, Ano97-27, Ano97-34, Ano97w, Ano98-35, Ano98x, Ano99v]. marginal [LB90, Pfa93, WC98b]. Marginally [HH93b]. market [Bel95]. Markov [AF92, BM90, BP92, BRM92, BP93, BMR95, BB99b, Eat97, EA96, Fuh92, HKP95, Hey96, Ios92, KS98c, Kul99, MS95a, Pen91, SJ96a, SJ96b, Sko98, SK97, Wef97]. Markov-Polya [SJ96a]. Markovian [SJ93]. Marriott [Smi97c]. martingale [Fei94]. masked [BABM99]. Master [Ano92i, Ano92j]. matched [Che96a]. matching [LD93]. Mathematical [Jen97a, Jen97b, JM98, Sri94]. Matrices [Bed00, Bed98, DEM95, DGP93, GR93, HKL98, Hyo92, JSH99, JKW92, Kou94, KS97c, KS99, Mar97a, MP98, Str98, SH90, WJ96, Wan97a, yXL96, Yam96]. matrix [BS93a, BA95, Fou99, KSM92, WL93]. Matveychuk [JK94a]. maxima [CM95]. Maximal [BEHW96, BE96, CS90, HSS97, Kou94]. maximally [Ma97a]. Maximin [BS95, Mül95b, JMY90]. maximizers [Fut99]. Maximum [Hau99, KBH98, MST91, SB95b, AiFL99, Bal90, Böh95, CH90, Eks98, FS95a, KS96, LB98, Ma97a, MST99, MM98a, MW97, Pfa93, Qia98, Rat96, RFD97, vR93]. May [Ano90g, Ano91g, Ano92o, Ano93o, Ano94w, Ano96-29, Ano97-33, Ano97y, Ano99u, Ham97]. MDS [MM93]. mean [AC97b, BHT91, BBTV95, BS98b, BB97, BS97, CW97a, CS99b, EBS96, GGBT95, Gui99a, HMW91, HC92, KV91, LP91, Lee95, LTT93, Liu97b, Oht97, PL97, PRF94, SH92a, Sar94, Sch99, Tah94, Tah95, TN95, YSS99, Zah91]. means [ESV97, GEGM99, He93, Hol99, HMT99, Iwa98, LY95b, Mik94, SSS94, VS93, WB94, Zou92]. measure [Kou95, XB91]. measurement [AIV95, BZS96, BS96, Buz98, KGS96, SSR94, SH92b, Spi94]. measurements [IJ94, JKW92, Kus99, LL95b, Stu91b, Smi97c]. measures [Car99, CAGM98, DF95, GW94, HJL95, LN95, MT99, MP98, Pat94, PRW99, RN95, Ras98, Róz92a, SWA97, Ull96, Wu94, Jon92]. mechanics [Sup90]. median [CW97a, Che95, Mus91, PD99, Pur95]. medical [JM98, Kad94, Mat94]. meeting [HH93a, Ham98]. Meier [LOS95, Stu94]. Mejzler [Wei97]. memorial [Shr98]. memory [GS99a, KS98b, LR96a, Muk98b, RL99, RZ98, Sta92, Wan97b]. Mendelsohn [Chu96]. Menon [MPW97]. method [BA92a, BY95, CS91, Chr94, Dut96, Eks98, Erd98, FS96, GMR96, Hoo90, Khu96, KT99b, LM95, Mag92, MS98a, MF96, Nav95, PRF94, SS90a, Seb98, Ten92, TS92, Yan96a]. methodologies [MS98b]. methodology [DM96b, Dra97]. Methods [Car99, CW97a, CMB95, DK99, Esc95, GP98, Ham97, Har99, KK95a, KSI99, Oga96, Pat97, PP94b, PWS98, Ros92, SS91a, Sil92, WH92, Smi97b]. Metrics [BCFS97, Cri92, CM92]. Metz [CQ98]. Midzuno [BA99]. MIFRA [SS91b]. Mikhail [Law97]. Miller [Chi92]. minihypers [HH93a, HH96a, HM96]. Minimal [Cha98, BD96b, Bar92, DDN98, KS97e, OR98, Vil91, Jos96a]. Minimax [JJMY95, JMY90, KM99, Kub90, MH92a, SS96c, ST93, Wan97b, Wie92, WZ96b, BA94, BZ91, Dat95, Hei93, KW98, May90, Sch93, ST96a, vRR96]. minimizer [Chi91]. minimizing [CRV96, Fro90]. Minimum [BH92a, BSV97, GH98a, HHH94, LTT93, MF97, Oht97, WWE95, DF94, EM99, FFR97, Mag92, Mec92, Shi92, Wu98]. MISE [Iso99]. Mises [Ios92]. missing [Car99, CC95, Gas93, KK95c, PW99, SS90c]. mission [vDMS97]. Misspecification [God96, CS92, PP96, SS96a, Spe94]. misspecified [GL98b, GTL98]. Mittag [Lin98]. Mittag-Leffler [Lin98]. Mixed [VS93, AB97a, BDG90a, BGK95, DK99, JKW92, Khu92, LI98b, Mey99, MPZ98, Oma97, PK99, Sei92, Sei93, SS96b, TR90, VSB98, Van98, VH98, ZM93, Sen99c]. mixed-effect [DK99]. mixed-MANOCOVA [Sen99c]. mixing [DPSvZ96, HP90b, LOS95, Rag97]. Mixture [DP98, Böh95, Che96a, CL95, Det93, DP99, DR96, GL99b, HB95, LL95a, WWE95, ZW99, Jen97a]. mixtures [Cha95c, Das94, FL91, Gut93, Pfa94, SP91, TG93]. ML [LN99]. MLE [KSM92, Luo96]. MML [BA92a]. mode [Erk97, Rub90]. Model [BRW99, CR97a, ALMS97, AGLS98, AFJ97, AA93, AC97a, AC97b, BRZ93, BM90, BP92, BRM92, BP93, BMR95, BA92a, BM99b, Ben96, BZ96, Bha99, BG91, Bis93, Bos97, BS98c, BQA99, Büh99, BH99b, BY95, BS96, CFM^{+}98, CK96, CS98, CM91, CH90, Cha96d, Che96a, CS99c, Dur92, DR96, DP91, FS95a, FMN98, Fuj93, GA95, GPS98, HS97, HB93, JM98, JMVH97, JK94a, Jos91, Khu92, KZ99, Kna99, KS96, Kon99, KS98b, KSM92, KM98b, Lee98b, LN99, LY95a, Loh93, LB91, MNWZ95, Mag92, MGdW96, MEG93, McL99, MI90, Mol96, MH95, MMC94, NK94, OR96, ON95, Oht97, Oma97, PFvR97, PK99, PP94a, Pur97, QK98, RH97, RZ98, RCGMC98, RS92, RC95, SO93, Sch93, ST96a, SS97a, SB95a, SB94a, SB95b, SJ96b, ST99, SSE98]. model [SS96a, SS96c, SW97, Sta92, Ste91, TX99, Tom93, TS95, TM96, TT92b, VV98, WS95, WR96, WF94, WC99, Ye94, YC98, Yok97, Zah91, ZW99, dLS99, vR93]. modeling [Lea95, Spi94]. Modelling [Ber95c, Mat94, SK97, BX91, Des97a, Nel97, O’H94, dSDF95, Les97, LR97, Wir97]. models [AP98, AP99, ALMS97, AGLS98, AB97a, AM96, Ame96, AC97b, AMPT95, AHT99, AM98b, BDG90a, BGK95, Ban96, BSV97, BABM99, BCvdM92, Bel95, BM94a, BM98, Bis93, Böh95, BZS96, BD94b, Car96, cIC95, CS91, CC99, CW98, CU92, CMB95, CP95, Cza94, Cza97, DK99, DMS92, DG95, Dav94, DS98, DN96, DN97, DGP97, Dik98, DP98, DP99, DD92b, Esc95, Far96, FL91, GG99, GL90, GSZ92, GH96a, GL98a, GL98b, GH90, HT92a, Haf98, Hau99, HP97, Hei93, Hen96, HB98a, Höp97, Hu98, HC93, HL94, Hua96, HOGF99, JM91, Jen97a, Joh91, JKW92, JM99, Khu90a, KV98, KW98, KR93, KL96, KP93b, Kub92, Kub90, KS91, Lec99, LL95a, LL95b, LN95, LS95e, Lus94, Lya97, MG96a, Man99, MM96, Mar97a, MY97, MM98b, MK96b, Mic90, MB96]. models [MP93, MN97, Muk98b, Muk94, Mül94, Mül95b, OR96, Oga96, OS96, PPO95, PW99, PP94b, PP96, PC94, Ram90, RL99, Ruc97, SS90c, SSR94, SS98a, Sei93, SJ93, Sen95, SJ96a, Sen99c, SS91b, SS95, SR99, SM96, TA91, TG93, VK99, VSB98, Van98, Vin96, VH98, VS97, Wan98a, WZ96b, Wu98, Yam96, ZM93, Iye92]. modes [HW96a]. modification [Mol96]. modifications [JSH99]. Modified [BA99, Stu92, TR90, BN91, DS97, EBS96, Lad95, LK98, TTW99]. Modular [Ban96, Eva96]. modulus [Ral93a]. Moment [BLBL93, DGP93, GMR96, TK91, Yan90a]. Moments [SA97, BRM94, Chr94, CF98, Kam91, Kam95b, MH98]. monitoring [LW94]. monograph [Jen97b]. monotone [CADMM95, Dat95, Lia90b, LH93, Lyn99, MM93, PW99, Spe94]. Monotonic [Kha97, GGT97, MB96]. Monotonicity [CP98, Iwa98, SSK99]. Monte [DM96c, Hau99, Koc98, NK95, Pow95, TAT99]. Montgomery [Dra97]. Morris [DL98a]. mortality [Mai98]. Moser [AC97b]. most [AF97, DR96, Mie90, MS95a, SK92, Tse94]. motion [KW92, MR99]. motivated [DKS98]. moving [AiFL99]. MRBP [TK91]. Muller [Ass96]. multi [BM99a, BM92, Gup98, MEJ98, Muk98a, Won94, dLS99]. multi-arrays [Muk98a]. multi-event [dLS99]. multi-factor [Gup98, Won94]. multi-level [MEJ98]. multi-stage [BM99a, BM92]. Multidimensional [CS96, CP98, CAGM98, GH95, Har92, MS95b, RT97, Tun98]. Multigraph [MK96b]. multinomial [AF97, AG98, AG99, BA94, BS94, ER99, GS99b, GH93, GLL96, SE92, TS91]. multinormal [Hol99]. multiparameter [DD94, RFD97]. multipath [WCP91]. Multiple [CLH98, Ras98, Rot99, Sen99a, VH98, BM99a, Bai98, BL99, BH99a, Bis93, CH92, CHH99, FG96, FR99, GH98a, GP99, Hay94a, HT91, Hei92, HR95, Hsu92, Mal95, NP93, RH95, Ros92, Sen99b, SB94b, Sha99, Som99, TD99, WRKL99, YB99]. multiple-objective [Mal95]. multiple-regime [Bai98]. multiplication [JRM96]. multiplicity [Sim94]. multiplier [PRS97, Qiu96]. multipliers [KSWyX97]. Multiresponse [Khu90b, KV98, VK99, Sen94]. Multistage [Che96b, MS98b, SC99b, HRR97]. multistate [CS99c]. multistep [Bha99]. multitableaux [AJ93]. multitype [YC98]. multivariable [HB98a]. Multivariate [AM96, CP95, ASK97, Arn90, BRZ93, Cai99, DDB99, Det94, GNZ97, HJL95, Hai92, Haw92, Hay98, HL94, JM96, JSH99, JK94b, KR95, PD99, PWS98, Rag97, RW98a, Rei99, RW97, RMZ93, Sar94, SA97, SP91, SB94a, Smi97b, Ten91, Ten92, Ten98, TA91, TS95, WC99, Yok97, ZM93, ZWF95, vR93, Jen97b, Smi97c]. multiway [PPO95]. München [Tho97]. mutually [AZZ96]. Myers [Dra97].

N [Tho97, BHW97]. N. [Tik97]. natural [BRM92, Ses97]. NBU*t [ZX98]. NBUE [DGK97]. Near [Bar92, KR93, KJS98, UM91, ZK98].

Near-minimal [Bar92]. nearest [BM99a, FS96]. nearly [BD96b, Cha95a, DDN98, KS97e]. nearness [Nay90, TK92]. Necessary [GTL98, MS90a, RS93, KP93b, Ma97b, RCX97, Vil91]. NEF [BB97, BS97]. NEF-PVF [BB97, BS97]. negative [BSV97, KS96]. neglected [AD97]. negligible [SH92b, SH92c, SC99b]. Neighbor [CLR91a, CLR91b, BM99a, FS96, Mor90]. neighbour [AD97, Dru99]. Nelson [PR93]. Nelson-Aalen [PR93]. Nested [DJ97, DLC99, KM97, KM98a, Bag96, CRV96, KN94, Khu92, MU93, MU97, RH97, SK90, SR99, Udd92, Udd97, YH90]. nets [CMM94]. network [FS95a]. networks [CHF92, WCP91]. neural [CHF92, FS95a, WH92]. neuronal [HT92a]. neurons [McK92]. Neyman [Kha97, Tak98]. Nikulin [Law97]. nine [MS97]. NN [SM96]. No [Coa97, Law97, Ter97, TAT99, EL93, SP91]. No-cointegration [TAT99]. noise [BL92, DM91, Ten91, Ten92]. Non [BK96, Gho97, IS97, LA99, BABM99, CK92, CF96, DD92a, Daw97, DMNN99, Dur92, GGBT95, JN97, KKNK96, Kun93, Lab93, Lin97a, LS95e, MFS98, Mac97a, Man95, MPW97, MB96, PW99, SW95a, SS98a, SS93b, Sin95, SH92c, SW98, SSSD99, SW96b, TA91, Wan98c]. non-abelian [MPW97]. non-adaptive [Mac97a]. non-additive [Wan98c]. non-additivity [SW98]. non-crossing [Lab93]. non-disjoint [GGBT95]. non-embeddable [MFS98]. non-equivalent [KKNK96]. non-existence [SSSD99]. non-homogeneity [CK92]. non-homogeneous [DMNN99]. non-i.i.d [LS95e]. non-identical [BABM99]. non-increasing [SW96b]. Non-informative [Gho97, IS97, Daw97, Lin97a]. non-interacting [SW95a]. Non-isomorphic [BK96]. non-linear [SS98a, TA91]. non-monotone [PW99]. non-monotonic [MB96]. non-negligible [SH92c]. non-optimality [DD92a]. non-orthogonal [Dur92, Kun93, SS93b]. Non-parametric [LA99]. non-predictable [JN97]. non-prime [Man95]. non-regular [Sin95]. non-saturated [CF96]. nonadaptive [Mac97b, VP94]. nonadditivity [CU92]. noncentral [KH95]. noncentrality [Ruk93]. noncyclic [NG91]. nonexistence [HW96b, HH96b, Ham98, HW98, Pfa98]. nonexponential [Hua95, PS90a]. nonhomogeneity [PMR97]. noninformative [Sun97a]. nonintersecting [Kra96]. nonisotropic [BT99]. Nonlinear [DD92b, Kna99, PP92b, RZ98, Cha93, Das93, DK99, DN96, FZ98b, Hu98, HB94, LC97, MM96, Mül95b, OR96, PR91, Sch97]. nonlinearities [Cha96b]. nonnegative [Pad94]. nonnormal [DM96c, Kna99]. nonnull [Iwa97]. Nonparametric [AC94, Ahm99, CC95, EMS98, Esc95, Gas96a, HJPZ99, HTY98, HC93, KJS98, LP91, Pat97, SW96a, TZ93, Tru91, dR97, dSDF95, AS96a, BB99a, BFM97, BFV94, Chi91, Chu95, Die95, EL93, GSGMPS97, GP94, Haf98, HP92, HMW91, Jan99, JK94a, Kim97b, KS97d, LM95, LC97, MB94a, MW97, Pat93, Rio98, RCGMC98, Sch98a, TT92a, Wan98a, Zel90]. nonparametrically [Fut99]. nonparametrics [Sen95, Sen99c]. nonsequential [OR96]. Nonstandard [DM95, Bab95]. nonstationary [HP91, Har92, Nev95]. Nonunique [Hut96]. nonzero [Lüt96]. norm [GS97b]. normal [BHT91, Ber98c, DM96c, DR96, EM99, FCA94, Fro90, GL99c, HMT99, Iwa98, KL99, KT92, Kem97, KSS97, Lad95, Lec99, Lia97, Liu97b, Mad95, Mae91, Mae94, McK97, PL97, SH92a, Sar94, SSK99, SP91, Seo94, Sri90a, Tah95, Ten92, TG93, UT97, Wan97a, Wan99, YSS99, Zou92]. normality [BG98, CC90, DP92, FW93, Kre90, Lie92, Man90b, MS90a, Mei97, Mei98, NR90, RS93, RFD97, Rou90, ZWF95]. normalized [GT99, KK95b]. normals [JSH99]. Note [RP97, Asa99, Böh96, Cha93, CK98, CF98, DL98a, Det91, DY94b, EEK97, EK97, Gro98a, Hüs95, Kun93, Liu94, PST98, PL97, Ral93b, RJ99, SC99a, SS97a, Str92, Sun97a, Tik97, Vel96]. notes [Jen97b]. November [Ano90n, Ano91l, Ano92r, Ano93m, Ano94o, Ano95l, Ano95o, Ano96-31, Ano97-31, Ano98-37, Ano98y, Ano99y]. NPMLE [SW96b]. NSF [Jen97a]. NSF-CBMS [Jen97a]. nuisance [BM94b, Bha91, Kin96, KS95]. null [Iwa97, PT94]. Number [Smi97c, BH92a, BD96b, BES90, DDN98, DY94a, Enn95, Fer95, GAS96b, GT97, HW96a, HM97, HSS97, HF95, HT92b, Jeo95, KS97e, Man95, Nie93, SH92a, Sar95, Sat93, Ste93]. numbers [BSS93, CMS93, Chr92, Zei96, dML93b, dML93a, vTrWM96]. Numerical [SH97].

objective [Mal95]. objects [LD93]. observables [AH99]. observations [BM99a, BD91, Bis93, BL92, BD94c, CK94, CF96, CP98, DDN98, DT91, GLS96, GT97, Kim97a, KS97e, KK95c, Ma97a, ME93, MEJ98, MJE98, Osk95, SB94b, Sza92, WL93, vBLO95]. observed [MB95]. obtained [ON95]. obtaining [GM93a]. occurrence [AF97]. October [Ano90k, Ano91i, Ano92n, Ano93j, Ano94t, Ano95r, Ano96w, Ano96-27, Ano97x, Ano98-29, Ano99q]. odd [CM91]. odd-equireplicated [CM91]. odds [LC96, PT97]. off [HM95, dlHF94]. often [GH99]. old [Ahm99]. Oldenbourg [Tho97]. Olkin [Jen97b]. on-line [SW95b, Sza92]. Once [DG98]. one [Bos94, DG95, DN96, DGK97, GM96, HP91, KL99, KM99, KH99a, LZ98, Lya97, Mac97b, Mol96, Mül93, PT94, PZ99, PT92, PP94a, PR91, RN95, Sch91, aS90, Shi91, Som99, Ste91, Ye94]. one-parameter [PZ99, PR91]. one-sided [PT94]. one-stage [GM96]. one-way [DG95, KL99, Mol96, PT92, RN95, Sch91, aS90, Ste91, Ye94]. only [BL96]. onset [Neu99]. open [YS92]. optima [Ros92]. Optimal [ALMS97, AGLS98, Ano98q, AC99, BD97, BHT91, BH92b, BD91, CH92, CR97b, Des97a, Dur92, ESA98, FMN98, GL99b, HT92a, HB95, Jeo95, KN94, Kim98, KW98, KSWyX97, LS95d, MP97, ME91, MST97, MS95b, MG96b, Mül94, Mül96, Nel97, Oga98, Par94, Sch98b, SM92, Siv97, SNN90, Sza92, Ten91, UM91, Udd97, Yin96, YZ97, AR97, AM91, BC93, Bag96, BD96b, BLBL93, BMS91, BH98, Bis93, CM91, CK94, CC96, CF96, Cha93, Cha90, CH96, CL97, CBM94, Dat95, DN96, DN97, Fal92, Far91, Far92, GH95, GJ98b, HS93, Haz99, HS92, Hei96, Hei98, HH93b, HM92, Jac90, Jac98, Kim97a, KKNK96, KS97e, KS98d, Kun93, Kus99, Lim90, MH95, MMC94, PS98a, PWS98, SS90a, SS91a, SW95a, SS97a, SS98a, SP91]. optimal [SK92, Stu91b, TY96, TN98, TN91, Udd96, Vil91, WM90, Won94, YK92, YH90, Les97, LR97, Wir97]. Optimality [AD97, Dru99, MU93, SS93b, SM98, BvB91, Cha95a, CH90, CH95, DD92a, DG91, Det93, Jac92, JKW92, JD91, KR93, Mar97a, Mat94, MC90, OR98, RJ99, ST96b, Vil91]. optimization [Dra97, Smi90, Tju93]. Optimum [Adk94, AIV95, CK92, CH94, Mie90, MS95a, Kar90, MH92a, Mik93, MMY94, SO93]. order [AB98, BEHW96, BBB96, BH92b, Ber98a, BTV96, CK94, CN97, FP98, GB95, GT95, Gov93, HP90a, HP91, HF95, HKL98, JRM96, JW92, Kak97, Kam91, Kam95a, KD90, Kou98, Kre90, Kub92, LDG98, Ma98a, Mal98, NN97, NR96, NZ91, PL97, PK99, PS90b, Qi97, Ros97a, SR90, Sar95, SS98a, SS98b, SH92b, Ste93, Str98, TW97, Van97, vTrWM96]. ordered [LS95a, LS95b, LS95c, Man90a, MMM95, RCZ97, VS93]. ordering [DP99, Ma98a]. orderings [BCFS92, DM97]. orders [AZZ96, Ste93]. ordinal [Ber95c, RS92]. oriented [BRW99, dK97]. Orthant [DM97]. Orthogonal [CG98a, Du96, HT91, Li98a, LI98b, LY96, MU97, Mul98, RCZ97, SL96, Tom93, AZZ96, BvB91, Bie96, BJ90, Cha98, Dur92, FHKM96, GL99b, HSS96, HSS97, JS95, Kag90, Kou98, KS99, Kun93, Man95, Muk98a, Pat97, SS93b, SS96b, VSB98, Vos99, VW99, WC98a]. orthogonality [Cal93, EJW93]. Oscillating [Kra96]. oscillation [Ral93a]. Other [Ano98v, Ano98w, Sar95, Sin91]. outcome [Con96b, HRR97, PRF94]. outcomes [GG92, QP99]. outlier [GR90]. outliers [AHJ96, NK95]. overall [CG98b]. Overdispersed [DGP97]. overview [WRKL99]. oviposition [JMVH97].

P [Sal92]. paced [Enn95]. packing [HS98b, Yin96]. Pages [Ano90l, Ano90i, Ano90j, Ano90m, Ano90o, Ano90g, Ano90n, Ano90k, Ano90h, Ano91k, Ano91h, Ano91j, Ano91m, Ano91n, Ano91g, Ano91l, Ano91i, Ano92w, Ano92k, Ano92q, Ano92v, Ano92p, Ano92l, Ano92s, Ano92m, Ano92u, Ano92o, Ano92r, Ano92n, Ano92t, Ano93k, Ano93n, Ano93r, Ano93p, Ano93l, Ano93q, Ano93i, Ano93s, Ano93o, Ano93m, Ano93j, Ano94l, Ano94q, Ano94k, Ano94v, Ano94r, Ano94m, Ano94n, Ano94s, Ano94u, Ano94w, Ano94o, Ano94t, Ano94p, Ano95p, Ano95q, Ano95t, Ano95s, Ano95w, Ano95x, Ano95n, Ano95m, Ano95l, Ano95o, Ano95r, Ano95u, Ano95v, Ano96y, Ano96x, Ano96-28, Ano96-32, Ano96-33, Ano96v, Ano96z, Ano96-30, Ano96-29, Ano96-31, Ano96w, Ano96-27, Ano97-29, Ano97z, Ano97-28, Ano97v, Ano97-30, Ano97-32, Ano97-27, Ano97-34, Ano97w, Ano97-33]. Pages [Ano97y, Ano97-31, Ano97x, Ano97u, Ano98-33, Ano98-30, Ano98-28, Ano98-27, Ano98-36, Ano98z, Ano98-34, Ano98-35, Ano98x, Ano98-37, Ano98y, Ano98-29, Ano98-31, Ano98-32, Ano99s, Ano99r, Ano99t, Ano99o, Ano99x, Ano99p, Ano99w, Ano99v, Ano99u, Ano99y, Ano99q]. pairs [Che96a, Lab93, Mor97, Pel96]. pairwise [BCM98, IS98, Koc98, LD93, MS98b]. Paley [BEHW96]. Papantoni [Sal92]. Papantoni-Kazakos [Sal92]. paper [Car97, Cox97, Les97, LR97, Nel97, Tik97, Wir97]. parallel [Ana99, BA95, DKR97, Li98a, LI98b, Sat93, SL96]. parallel-flats [LI98b]. Parameter [HB93, HB98b, LW99b, SB94b, ALMS97, AGLS98, Alv95, Bou90, CS99a, DN96, Dut96, GGR99, GF91, GM93b, Hei93, Hof95, Höp97, JM91, KM99, KHAH90, Kou95, LH93, Lus94, OR96, PZ99, PN96, PR91, Ruk92, Ruk93, Sar93, TP97, TK92, Yan90a]. parameters [BM94b, Bha91, BL92, CM98, CXL93, Che98a, Chr94, CMB95, CL96, Cro98, EM99, Haw92, Hay94a, HRR97, KP97, Kin96, KS95, Kub92, KM98b, LW99a, LS95b, LS95c, Ma97b, May90, Mon90, Nay90, Oga98, PP92a, Pfa94, Shi92, Sri90a, SNN90, TS92, Vel96, VM95, WR96, Wef97, Wie94]. Parametric [HOGF99, Phe96, BM90, BP93, BSV97, CW97a, Cza94, Cza97, KS95, LA99, MBS99, WS95, Wan98a, ZW99]. Parametrics [Sen95]. parents [GB95]. Pareto [Arn90, BB95, BB98]. Pareto-index [BB95, BB98]. parities [TS97]. part [Nie92, Smi97c, ALMS97, WC99]. Partial [Ruc97, Bha91, BCFS92, CH90, Cha98, Giu95, GUH91, Jan93b, KS92, Ma96, Ma97b, ST96a, Ste93, Fou99]. partially [CS91, Hei93, Tun98, Udd97]. participants [Ano97t]. partition [Liu94]. partitioning [BG98]. Partitions [MS97, Kra93, Pro93]. parts [Ano96e]. Past [Zel96]. path [MG96b, Nie96, SJG93]. paths [Böh96, BSS93, CMS93, Ges93, Gho96, KP93a, Kra96, Lab93, MKR93, Sat93]. patterns [LM99, PW99]. Paul [Kat97]. PBIB [SK90]. peakedness [Ma98b]. Pearson [Kha97, MH97]. penalized [SW96b]. penalty [BRW99]. percentile [CV92]. perfect [BL96]. performance [GC99]. period [IJ94]. periodic [Fer95, Ste97]. periodogram [GMD99]. permutation [Cri92]. permutations [BCFS97, PST97]. permuted [DM93, FP94]. persistence [MT99]. Personal [Mic90]. perspective [KS90, MN97]. perspectives [BH99a]. perturbed [Ral92, RS93]. Peter [Bas95a]. Petunin [JK94a]. PG [BHW97]. pharmacokinetic [JWJB99]. phase [ESA98, SSS94]. phenomenon [WS94]. philosophy [DM96b, Lin90, Sri90e]. phylogenetic [Nav95]. Physica [AC97a, Ter97]. Physica-Verlag [Ter97]. Physiological [McK92]. piecewise [CS99c]. Pitman [Bos98, JF91, Jos96g, Kou95, Nay90, TK92]. Pitman-closerness [Jos96g]. plane [Kra93, Pro93]. planes [Gly97]. Plann [Ano98q, Bed00, dML93a]. planned [Sin91]. Planning [AG99, KS11, LT00, Tse94]. plans [Cha98, GT93, HM95, Kla90, MC90, SM92, Shi91, YC95]. plant [BT97]. play [BB99a]. play-the-winner [BB99a]. plea [Lev90]. plot [NK95]. plots [CW93]. plotting [Kle91]. plug [BFM97]. plug-in [BFM97]. PMD [Qin96]. Poincaré’s [Bon97]. point [BF96, BDG90a, BR97, DV97, DN96, Fer94, GLS96, GH96a, GH98b, Höp97, HK97, HC93, JMVH97, JM93, KS91, LZ98, Lya97, MH92a, Mec92, NL97, Rat96, Tom93, VY98]. point-symmetry [Tom93]. points [Cha90, DG98, DF94, LN99, Lus94, MS97, Mül95a, Zha96]. Poisson [Deh92b, DMNN99, JS95, LY95b, MI90, Rév95, Róz92b, Ruk92]. Polars [Gut93]. policies [Mie90, MS95a]. policy [BM93, SJG93, SMS96]. Polya [SJ96a]. Polygonal [SSSD99]. polyhedral [ST93]. polynomial [CH96, CL97, Det91, Det93, Det94, Hei92, Hei96, Hei98, HYT98, JM93, Lim90, LW97, Mal98, Spr90]. polynomials [And93, DP98, Hyo92, KR97]. polytomous [RC95]. pool [SK99]. pooling [OS96]. population [BZS96, FCA94, GM96, HW96a, Lie96, LYC98, MS95b, MS92, Muk94, MB94b, Nay92, SSS94, SSE98, Yat95, ZdBPL90, Zou97]. populations [BM92, BZ91, DR96, FG94, GLL96, GL99c, Kar90, Lia97, PL97, PT92, QGG96, SH92a, Sar94, TX99, YSS99, Zou92]. posed [Hag92]. Positive [CRV96, BCFS97, Hay98, KWZ93, Shi92]. Positive-breakdown [CRV96]. Possible [AARM99, Lya97]. post [Büh99, CLH98, CHH99]. post-model-selection [Büh99]. post-stratification [CLH98, CHH99]. Posterior [LY97, PP94a, Bas95b, BJL98, Bos98, CG98b, DM94, Erk97, Lad95, LC96, OD99, RC98, Siv94]. posteriors [GG99]. potential [NK95]. Poupard [Sul93]. poverty [KGS96]. power [BN91, Bou90, GH96a, HMW91, Iwa98, KL99, LS99b, Man95, MS90b, SK98]. power-law [SK98]. powerful [DR96, Liu95b]. powers [Kou95, Ram97, SL95]. pp [AC97a, AC97b, Chi92, Coa97, Dra97, Fou97, Iye92, Jen97a, Jon92, Law97, Sal92, Smi97a, Smi97b, Smi97c, Ter97, Tho97]. practical [Kle91]. Pranab [Coa97]. Pre [Ben96, TT92b, GG96, Wan96]. Pre-test [Ben96, TT92b, GG96, Wan96]. precedence [Slu92]. Precision [KSI99, dlHF94]. predetermined [SJ96a]. predictable [JN97]. Predicting [AH99, MB94b, BZS96]. Prediction [Ber95a, LN95, Muk94, TH99, ALMS97, AGLS98, AR97, AHJ96, Bha99, BZ91, DDB99, HB98a, Kha90b, KW98, Kub98, MMY94, Nan95, Rao94, RS98, TR90]. Predictive [WF94, AE93, CdlH96, Ger94, Lec99]. predictor [CR97b]. predictors [DG91, KT92, ON95, Ste97, UT97]. Preface [AS97, BS93b, BP94, Bas96, Bas91, CFT97, Dwi95, HWD92, Hai95, HR92, Hal98, Hin90, JL95, Kor92, MR96, Moh93, Moh96, Sha97, Sin96, Sri96, Tei97]. preference [HZzY91]. preferences [IM94]. Preliminary [BP93, Muk98b, Sri94]. Preorderings [MM93]. prescribed [cIC95]. presence [AHJ96, AC99, BM94b, Bha91, CS98, GJ98b, Jac98, Kin96, SM92, TN91, Wu94, Zha95]. present [Jan94a, LB92, Zel96]. preserving [Ten91, Ten92]. Press [AC97b, FF97, Sal92]. prevalence [Sha95]. Price [Coa97, Law97, Smi97a, Smi97c]. prices [LR96a]. primal [Mey99]. primal-dual [Mey99]. prime [BTV96, BA95, Man95, Str98]. principal [Cha96c, HT92b]. principle [HP90a, Hai92]. Prior [BM94a, Ber90, CB94, GM93a, LB90, Mon98, PS90a, Siv91, Ye94]. priors [Daw97, Gho97, IS97, Kub98, Lin97a, LY97, MC95, RW95, Siv97, Sun97a, Was96, Wet97, dlHF94]. Priscilla [Law97]. Probabilistic [Sup90, Ano97j, DP91]. probabilities [ER99, GH93, GL98c, Hay94b, HS98a, Yu96]. Probability [Wan98c, AC94, AF97, Bas95b, BD94a, DCK95, Erd98, GR90, Jen97a, Lin90, Liu94, NG91, Niy97, OD99, Ros97b, SWA97]. probing [SH92c]. probit [BS96]. problem [BR97, BGL97, BA92b, DM94, Fin97, GLS96, Hag92, HS98b, HKL98, JV97, Liu95a, LY97, Mac97b, MH92b, MBR99, Pay98, Qi97, Rus93, RT97]. Problems [Sim94, AHT98, BRM92, Cha93, Che96b, DV97, DD94, DKS98, Den96, DM91, Eat97, GJ96, HMW91, Kle91, NRZ98, Nis95, WLW93, YS92]. procedure [BB98, BL99, BA94, EBS96, HT92b, KBH98, Liu97a, Liu97b, Mec92, MS92, NP93, OR96, RH95, SH92a, Sha99, SW95b, Tah95]. Procedures [WRKL99, BM92, BS94, Coa97, FG94, FR99, GH99, GP99, GL99c, HT91, Hay94b, HR95, HB99, HL94, Hua95, Hua96, Kha90b, LR92, LB92, Lec99, LR96b, LB91, PS98a, Pan96, Pur97, Ras96, Rei99, Rot99, SB95a, SJ96b, Som99, Spi91, SC99b, Sun97b, TD99, TT92b, YB99]. Process [Dra97, AiFL99, BZ96, BL92, GS99a, Höp97, JS95, Kul99, Lee95, Lee98a, Lee98c, LS99b, LPS98, MYTH96, Mar97b, Mau98, MBS99, MG96b, Mus91, Ral93a, SK98, SB94b, Yan95]. processes [Add99, Adk94, AR97, AHT99, BF96, BD96a, Bar97b, BE96, BD94a, Bro94, CT96, CH90, Dat96, DPSvZ96, Deh92a, DM95, DMNN99, DY94a, Ein92, Far96, Fei94, FW93, FZ98b, Ger94, GGJ98, HP91, Har92, HP98, Hey92, HL95, HOGF99, Hüs95, HB98b, Kim98, KS94, KS98c, LA99, Lüt96, Mar96, MB97a, Mon98, NR90, Rag97, Ral92, Rao94, Rat96, Rév95, RV95, RvZ92, Ruy94, SW96a, SK97, Sta92, Ste97, SE95, TT92a, WF94, Bas95a]. processing [CHF92]. product [BN99, Dra97, PS98a, Pro93, SSS94, TSS96, Tse94, Wat93]. profiles [BT97]. programming [JWT93, SS96b]. programs [Spe94]. progress [Qiu96]. progressive [AB98]. projection [PS98b, ZWF95]. projections [BD95a, CG98a, PRS97]. projective [HS98b, TV98]. promised [STL96]. proof [Jos96c, Pro93, Seo90]. proofs [CP97a]. Properties [AP99, GP98, Mar97a, Mor98, AP98, AB98, BH92a, BLBL93, Bar97a, BG91, Car96, CR93, Dat96, DCK95, GR97, GGT97, Iwa98, Jan93a, Kha97, MPSM94, ND98, Rat96, RW95, SF97, WB94]. property [Ban96, Dea93, DGK97, FS97, GG99, GT93, Hay98, HM92, Jan93b, KR97, RT90, Ses97, SH90]. Proportional [Zah91, BCvdM92, BS97, Buz98, DMNN99, Kon99, Ros97b]. proportions [Den98, GS99b, TS91, WWE95]. provision [PT94]. proxy [TS96]. pseudo [Dea93, WL96]. pseudo-likelihood [Dea93]. pseudo-symplectic [WL96]. pseudofactors [Vos93]. psychological [Pfa94]. published [Fou97]. pure [Qin96]. pursuit [PS98b, ZWF95]. PVF [BB97, BS97].

quadrangles [BTV96, Pay98, TV98]. Quadratic [Ste91, AGLS98, BMR95, Dut96, GT99, GSZ92, Gro98a, Kim97b, Kub90, LR95, MH92a, MS93b, Ses97, TR90, Wan97a]. quadratically [Eat97]. quadrature [O’H91]. qualitative [WD98]. quality [GL98b, HL95]. quantal [STL96]. quantification [Kar91]. quantified [Nis95]. Quantile [HN99, AK93, CP97b, DPSvZ96, Deh92a, Ein92, Ein93, Lee98c, ON95, Ral92, Ruy94, Zel90, Zha95]. quantiles [Che97, Fro90, GP98, HP98, KS92, Lan96, Oga98, RS93, VV98, dHR93]. quantitative [WD98]. quantities [BJL98]. quantum [Sup90]. Quasi [FH96, AM98b, CS99a, Des97a, Hey96, Les97, LR97, LM99, MFS98, Mon90, Nel97, Van98, Vil91, Wer92, WD99, Wir97]. quasi-balanced [Van98]. quasi-invariant [Wer92]. quasi-least [CS99a]. quasi-likelihood [AM98b, Des97a, Hey96, Les97, LR97, LM99, Nel97, WD99, Wir97]. quasi-necessary [Vil91]. quasi-residual [MFS98, Mon90]. quaternary [Ham98]. question [Goo96a]. questions [TM96]. queueing [AB97b, SJ93]. queues [BM93, Wip98]. Quintessential [BCM98]. Quota [SE92]. quotient [BN99]. quotients [SWA97].

R [Dra97, Shu95, Tho97]. R. [Shr98]. Random [Ken90, Kla90, LSP99, Róz92a, Adk94, AM96, AARM99, AC99, BES90, Bar97a, BN99, BF99, BCvdM92, Ber95a, BD96c, BN91, BT99, BH99b, CM92, Dik98, DY94a, DM97, Ein93, EA96, Fer95, FW93, GJ98c, HP98, He93, Hey96, Hor98, HMT99, HB93, HB98b, JMVH97, KR97, KM96, Khu90a, Khu92, KP96, LH97a, Lee98b, LN99, LW99b, LM99, Liu95a, LDG98, Mal98, Mol96, MPSM94, MG98b, NK94, Pak97, PR96, RH97, Róz92b, SH92a, SR90, SS90c, SJ96b, Ste91, SL95, TZ93, VK99, Ye94, YC95, Yok97]. random-effects [Yok97]. random-size [Liu95a]. Randomised [BD94a]. Randomization [KS90, CK96, DK94, GL99a, LB92, Pat94, SP97]. randomized [BB99a, CR97a, LR92, TM96]. randomly [Adk94, KS92, Yan93]. Range [CXL93, Add99, AARM99, DT91, FMN98, GA99, GF91, HK97, LA99, Seo94]. Ranges [Bas95b]. Rank [BD94c, GH98b, Man90a, PP92a, Tra90, Akr91, AA93, AB97a, Alb92, BM99a, BH92b, Cri92, DP92, DH92, Gov93, HJPZ99, HP90b, HP91, Har92, HP98, JN97, Jan99, Man98, Man99, MM93, Mei97, Mei98, MT94, NR90, Nik91, RN95, Rei99, SR90, Sar95, Seo90, Seo94, Tho93]. Rank-based [PP92a, BH92b, RN95]. Ranked [PST97, KS93, KS11, LSP99, LD93]. ranking [BM99b, GH93, HMM97]. rankings [AP97, PST97]. ranks [HLM90, Moo96]. rao [Kag90, SS93a, Seb98]. rare [Yat95]. Rasch [Tho97]. Rate [Ber98b, Ber98c, Ahm94, BB91, BL99, GP94, Haf98, IT90, KS91, Pat93, Pat97, Rou90, SS99, SB95a, YB99]. Rates [Arc99, Tho93, Tom95, BO99, Bos97, GGT97, KR95, KZ99, KS95, Lia90b, LB90, Mai98, Phe96, vRR96]. ratio [BD94a, CL95, DBG99, El 96, FP98, Iwa98, JSH99, Liu95b, Ma97a, Ma98a, Mad95, Mau98, PP92a, PZ99, PT97, SSS94, SSE98, Sri90a, TSS96, Ye94]. ratio-cum-product [TSS96]. rational [KSS97, Vin96]. rationality [Mic90]. ratios [Den98, GHRG^{+}99, GGBT95]. real [CHF92, MST97, Ram97]. real-valued [MST97]. recapture [LYC98, Llo98]. reciprocal [KW92]. reciprocals [ESV97]. record [Ber95b, BP95, DN94, Enn95, GP94, Kam95b, NN97]. record-breaking [GP94]. records [Iml95, Nev95, Nie93]. Recovering [GJ96]. Recovery [Sha92, ZC92, CK96, EMS98]. rectangles [Eva96, LMT97]. Recurrence [Kam95b, Eat97, dSDF95]. recurrent [Ber95c]. Recurring [Pur97]. Recursive [MM98b, Róz92b, Iso94, PS90b]. reduced [Rei99]. reduced-rank [Rei99]. Reducing [Fer97]. reduction [CW97a, vBLO95]. Redundancy [BENP91]. Reed [Ass96]. referee [Gup90, Gup91, Gup92]. Reference [Kub98, GF91, Kha90b, Ye94, vBLO98]. refinement [FK98]. refinements [DL98b]. regime [Bai98]. Regional [Jen97a]. regions [BR97, Fut99, aS90]. Regression [KS98b, SS97b, AS96a, Akr91, AM98a, AC99, BA92a, BG98, BZ96, Ber95a, BR98, BFM97, BZS96, Buz98, Cha95b, cIC95, CH96, CL97, Cha96d, CR97b, CC90, Chi91, Chu95, CC95, Col95, CRV96, CS92, Cza94, Das93, Det91, Det93, Det94, DN97, Die95, ESV97, EL93, FMN99, FRS98, Fie97, FS95b, FL91, FCA94, FS97, Fou99, Fre91, GH95, GH98a, GA99, GG96, GSZ92, Goo96b, Gro98b, GTL98, GH90, GH96b, GPS98, HJL95, HP90b, Hei91, Hei92, HS92, Hei96, Hei98, Hof95, HRR97, Hu98, HL94, HR97, HB93, JM91, JM93, Jan93b, KZ99, Kna99, KS96, KL96, Kon99, Kot90, KS97d, Kub92, Kub90, KJS98, LK98, Lav94, LM95, LB94, Lie92, Lim90, LTT93, LW97, LC97, Lya97, MYTH96, MM96, MBL97, Mec92, MC90, MMC94, OR96, Oht97]. regression [PS98b, PP96, PS90b, QP99, Qia94, QK98, RC98, Rei99, RCGMC98, RC95, Sch97, Sch98a, Sch99, SS97a, Seo90, Shi97, ST99, SS96c, ST93, SR99, TT92b, TS96, VV98, Vil91, WR96, Wie92, Wie94, WZ96b, Yan95, YT98, Yoh97, YZ97, Chi92]. regressions [Bai98, GMR96]. regressors [HR97, SS97b]. regret [Tah95]. Regular [Ma96, Wal96, Eat97, HP91, Har92, MD94, Sin95]. regularity [DM96a]. regularization [Yan96a]. regularly [Ome95]. Rejoinder [Ber97, Des97b, Yan90b]. related [HP98, HB99, LW99a, Moo96, MPZ98, Mul98, Mus91, Pur97, SS95, SK97, Tun98]. relation [AMPT95]. relations [Kam95b]. relationship [Tan99]. relationships [CL91]. Relative [Jos91, LB92, Das94, Seo94, Spi94]. reliabilities [GUH91]. reliability [BE91, BX91, BG91, GA95, Joh91, Kau96, Sen95, WCP91, ZB92]. reliable [Böh95, Mie90, MS95a, Tse94]. remaining [Zah91]. remark [Oma97]. Remarks [JK97, GS99c, Sen99b]. remedians [CL93]. REML [CL96]. removal [LYC98]. removing [GG92]. Rencher [Smi97b]. renewal [Bar97b, Enn95, MB97a, SW96a, SE95]. repair [Ana99, DGK97, Fin97]. Repeated [SH92b, Car99, CP95, GS99c, IJ94, Jon92, JKW92, Kus99, LL95b, LN95, Pat94, RN95, Ras98, Smi97c, Stu91b, Tan99, Wu94]. repeated-measures [LN95]. replacement [Bos97, Kla90, Mie90, MS95a, WC98a]. replicate [JPW97, MD94, Vos93]. replicated [GL90, Jac90, MT94]. replication [Cha98, Tan99]. replications [MP98, Vil91]. reply [Goo90a, Jos96c, Jos96d]. representation [AS96b, GMD99, HJL95, Hor95, McK92, Nis95, Pur95]. representations [FZ98b, Gly97, Gly98, MK96b, Stu94]. required [KV91]. Resampling [YB99, KS98a, RPK97, Str97]. Resampling-based [YB99]. rescaling [DD92b]. research [FS94, Li98a]. Residual [CW93, AK93, BBTV95, CS99b, CV92, LP91, MFS98, Mon90, Mus91, SC97]. residual-based [SC97]. residuals [AP98, AP99, BR98, Jan93b, Lee98c, MM98b]. resighting [LYC98]. Resistance [SG91, WC98a]. Resistant [QP99]. resolution [Bar92, CH98, Hyo92, HK94, KI98, NM97, ST96b]. resolutions [LMT97]. Resolvable [IS98, JR99, JPW97, LT96, LT00, RJ99]. respect [BD97, BMR95]. respondent [TM96]. Response [Niy97, Bis93, CR97a, KV91, KR93, LR92, MH92a, Pow95, RFD97, Sch99, Sri90e, TM96, WC95, WM90, WD98, Dra97]. response-dependent [WC95]. response-driven [RFD97]. responses [Lüt96, STL96]. restoration [CM97]. Restricted [WS95, BB99a, CMS93, GTL98, Hei93, HH93b, Iwa98, KM99, KK95b, PK99, SP97, Sat93]. restriction [BRM92, Wan96]. restrictions [BP92, GL99a, PS90b, SO93, ST96a, TT92b]. result [FS97, Kou91, MKR93, SSK99, TK91]. results [Cai99, CS90, Csá96, GB98, GF91, GL98a, HP92, HM92, KP93a, KK95b, gL96, Mar96, MEJ98, MJE98, NJ99, PR93, SS91a, Sri94, SSSD99, TAT99, vdL96]. retrieval [Let97]. retrospective [Sin91, Sta97]. Return [KP96]. reversible [Xia98]. Reversing [LDG98]. Review [AC97a, AC97b, Ano97j, Bas95a, Chi92, Coa97, Dra97, Fou97, Iye92, Jen97a, Jen97b, Jon92, Law97, Sal92, Shu95, Smi97a, Smi97b, Smi97c, Ter97, Ano98e, Ano98f, Ano98g, Ano98h, Ano99a, Ano99b, Ano99c, Ano99d, Ano99e, Böh95, CM92, Gov93, KS99, LL95a, Pan96]. reviews [Sri90c]. revised [FF97]. revisited [vBLO98]. Reweighting [SC97]. Rezension [Tho97]. Richard [FF97]. ridge [Fou99]. Risk [GG96, BMR95, CB94, LY95b, Spi94, WS95, Wu98]. risks [KD99]. Rissanen [Ger94]. RNN [BMV97]. Robust [Ano97-35, Ber90, BM94b, BRM94, BFM97, Cha96d, Det94, Fie97, HR97, LW97, MM96, MY97, MEG93, MMM95, O’H94, PP94b, RW97, Ron97, Sen99c, Siv93, Str97, TG93, VY98, Wie94, BA92a, BR98, BS95, CC90, Das94, Dea93, Det91, FZ98a, Ham97, Jos91, Kad94, Kou91, LS95d, MH92b, MS98a, Mül94, PWS98, QK98, SS92, Siv97, SW97, WC95, WS94, YZ97, Coa97]. Robustness [Che95, DR96, GW94, HMM97, IM94, LR96b, MNWZ95, ADP91, BJL98, BM94a, Bos94, CR93, DM94, DD94, FR94, LY97, Mak94, MP93, MC95, PST98, Pan96, PP94a, She98, WLW93, Was96, Yoh97, dlHF94]. role [Far96, Gov93]. room [LMT97]. root [Cha96b]. roots [MK96a]. Rotatability [DGP93, DY94b, Khu90b]. rotation [Let97]. Row [MT96, SW98, Bag96, DD92a, EJW93, Jac90, JPW97, Kun93, MU93, SS93b, Udd97, Wan98c]. Row-column [MT96, SW98, Bag96, DD92a, EJW93, Jac90, JPW97, SS93b, Udd97, Wan98c]. row-column-structure [Kun93]. rows [MU97, SK90, Udd92]. rule [BMV97, BM99a, BB99a]. rules [GH93, GL94, HMM97, Iro93, KM99, Kim97a, Lin97c, VR99, vBLO96]. run [BCvdM92, BH96, SSK99]. runs [GAS96b, Sar95, WN98].

S [Law97]. Sacco [Ano97j]. same [KZ99]. Sample [CBM94, DV97, HT91, Lan96, ASK97, BHT91, BD95b, Car96, CM98, CH90, DM94, DP92, DMNN99, DL98b, ESV97, Fie97, FS95b, Fro90, Gas93, GJ98a, GC99, GH93, HP91, He93, HB94, Jan94a, JN97, KT99a, Liu97a, Mal95, Mei98, Nay92, Oga98, PR93, Pur95, RJ98, RS93, Sar95, SB94a, SF97, SO92, Tan99, TSS96, WL93, Zha96, ZWF95]. Sample-size-optimal [CBM94]. sampled [Wu97]. samples [BF99, FH91, JM96, MB94b, SF97]. Sampling [HS98a, vBLO98, BA99, CLH98, CHH99, CR97a, Dut96, GB95, GB98, GL94, GM96, HM95, HC92, Iro93, Kar90, Kau96, KSI99, Kim98, KBH98, Kla90, KS93, KS11, LSP99, Liu97b, MPSM94, MC90, MV96, PST97, PP94a, Ros97a, Ros97b, SM92, SSS94, SE92, SO92, YC95]. San [AC97b, FF97]. saturated [CK94, CF96, Vos99, VW99]. Savage [Ral93b]. SBIBD [Kou94]. scale [CRV96, Das94, GG96, GC95, KHAH90, Kou95, LS95b, LS95c, Nay90, Oga98, PP92a, PN96, PC94, TP97, TS92, Vel96]. scaling [Pak97]. scatter [PWS98]. Scheffé [MS93a]. scheme [BA99, SSK99, Tun98]. schemes [Ban96, HSS96, HZzY91, KBH98, Nev95, SS98b, SH90, WJ96, WL96]. Schimek [Ter97]. Schröder [BSS93]. Schum [Ano97j]. Schur [Wat93]. Science [Sal92, DM96b, Lin90, Zei96]. score [BH96, Con96a, HP98, PRF94, RM95]. scores [Buz98, HJPZ99, HP90b, HP91, Har92, Seo94]. screening [KP97, PS98a]. Search [CM93a, CM93b, Bel95, DL98a, GT93, GM93b, Mac97a, Rus93, RT97, Van97]. searching [CM93a, CM93b, MR98]. seasonal [NL97]. seasonality [SMD95]. Second [HF95, DP99, Kub92, PL97]. second-degree [DP99]. Second-order [HF95, PL97]. sectional [OS96]. select [LB90]. selected [HP97, Ma97a, Oga98, Vel96]. Selecting [FCA94, GH90, HM97, HL95, Mal95, BHT91, BA94, Cza97, FG94, FG96, GH96b, Lia97, MS92, SH92a, Tse94]. Selection [HT96, Hua96, PS98a, AL95, AM98a, BRW99, BM92, BS94, Bha99, BQA99, Büh99, CS98, Chi92, Chu95, Dou96, Fal92, FG94, FP94, Gui96, GM93b, GL94, GM96, GLL96, GL99c, HMM97, Hay94b, HS98a, HL94, Kha90b, Kim97a, LSP99, Lie96, LR96b, Liu94, Liu95a, Pan96, QK98, Ras96, SP97, Sar93, SJ96b, TN98, vdL96]. selections [GL98c, ST99, YT98]. selector [Haz99]. selects [SH92a]. Self [Cha96c, MT99, Zei96, MKR93, Ren97, WZ96a]. Self-affine [MT99]. Self-avoiding [Zei96]. self-consistent [Ren97]. Self-dual [Cha96c, MKR93, WZ96a]. semi [Bai90, Mon90, Sen95, Sha99, ZW99]. semi-Bayesian [Sha99]. semi-latin [Bai90]. semi-parametric [ZW99]. semi-parametrics [Sen95]. semi-symmetric [Mon90]. SEMIFAR [BO99]. semigroup [Fei94]. semimartingale [HT92a, Lus94]. Semiparametric [GA99, LR96a, LY95a, MG96a, PMR97, ST99, Böh95, BS98c, Dik98, GGR99, LL95a, Mag92, WR96]. semisymmetric [FH96]. Sen [Coa97]. Sensitivity [Siv91, Spe94, Ber90, CG98b, Lav94, Sha95]. separation [CX98b]. September [Ano90h, Ano91i, Ano92t, Ano93n, Ano94p, Ano97u, Ano98-31, Ano98-32]. September-October [Ano91i]. sequence [GS99a, KS92, MM98a]. sequences [Fer95, HP90a, Hai92, HKP95, KS99, Mon91, Nie93, Nie96, Nie92, Yu96]. Sequential [BB99a, BB97, BS97, Coa94, Iso99, Spi91, Sri90a, SH92c, Tah94, vDMS97, BD94a, BA92b, CW97b, CP98, CBM94, Eis94, ESV97, GT97, Liu97b, Mau98, MBS99, McL99, MF96, MS92, PT94, RPK97, Róz92a, Tah95, TN95, TW94, ZdBPL90]. serial [AIV95, BH96, HLM90, HP90b, NR90, Shu95, Tra90]. Serially [Mon91]. series [Bag96, BN91, GMD99, HJPZ99, Har99, HB94, JS95, Jen97a, Jen97b, KC93, LC97, MT99, MST97, Mor90, NK95, ND98, OS96, RW98a, RL99, Rio98, RZ98, Ruy94, Shi97, SK90, TA91, Tom95, Tra94, Tru91, ZB92, Fou97]. service [AB97b]. services [Gup90, Gup91, Gup92]. set [Cha90, El 96, Fut99, GJ98b, GC95, KD99, KS93, KS11, LSP99, MC95, PST97, Tak98, Zhe96, vBLO98]. sets [BHW97, CC96, CMM94, DM93, DJ97, DM95, FFR97, GS97a, GGBT95, Hol99, JV97, dKgL96, dK97, KSWyX97, Ma97b, Mah98, MS97, MPW97, MU97, NG91, Pfa98, PRS97, RCX97, Ros92, Róz92a, Siv97, Stu94, UM91, VP94, Van97, yXL96, Xia98, Yan96b]. Setting [RT90]. seven [MS97]. several [CM93a, CM93b, CXL93, HN99, HL95, Kub90, Liu97a, PL97, QGG96, Sar94, Shi92, YSS99]. Shanghai [Ano96e]. shape [RW97]. shapes [Ken97]. sharing [Lyn99]. Sharpening [KS98a]. Sheffer [Nie96]. shift [Sch98a]. shifted [Kra93, ML96]. shifts [Haw92]. Shock [MB96, SS91b, HOGF99]. shock-induced [HOGF99]. short [Sta92]. short-term [Sta92]. shortcut [GH99]. shortness [Tak98]. Shrinkage [Sar94, Muk98b, Rei99, SK99]. Shrinking [LB94]. shrunken [KHAH90]. sibling [MST91]. side [BA92b]. sided [DM91, MS90b, PT94]. Sign [DM96a]. signal [DM91]. signals [KK95c]. signed [DH92, HLM90, Seo90, Seo94]. significance [GS99c, Sen99b]. Simes [HR95, KH99b, SC99a, Sen99b]. Simes-type [Sen99b]. similar [DCK95]. Simple [DH92, BB95, CP97a, Col95, Eub99, Fre91, FP98, Kej96, KP96, McL99, MS98a, PP94b, Ses97, SMD95, Yu96]. simplest [SW97]. simplex [DP99, Lim90, Mik93]. simplexes [Ken90]. Simplified [MR98]. Simulation [CW97a, McK97, VS97, AB98]. Simulation-based [VS97]. simulations [RL99]. Simultaneous [FRS98, GS99b, GLL96, GL98c, PC94, VW99, DM96c, GL99c, LY95b, LY97, MY97, MS98b, Sch93, ST96a]. Simultaneously [Lia97]. Single [GH93, BA92a, CL95, GLMN98, MD94, Vos93, vDMS97]. single-replicate [MD94]. Single-sample [GH93]. singular [BRM92, DF95, LW99b, Lus94, Sen91, ST93]. sites [LB90]. situations [Jos91]. size [BC93, Bai90, BHT91, BKM93, BD95b, CBM94, DM94, DV97, FFR97, HT91, Hol99, KM97, KM98a, KT99a, Kes92, KS97f, KH99b, Liu95a, Liu97a, LYC98, PST97, RJ98, Ros97b, Stu91a, ZdBPL90]. size-biased [PST97]. sizes [AM91, He93, IJ94, KVAN90, RJ99, Sar95, Udd96]. sketch [Kat97]. skew [Fed93, Wat93]. slippage [HMW91]. slope [ALMS97, DBG99, DY94b, Fre91]. Small [GJ98a, GC99, KS97f, PR93, Car96, DDB99, Du96, Fer94, Fie97, FH91, GAS96b, SS98b, SR99]. small-area [SR99]. Small-sample [GC99, Car96]. Smirnov [DL98b, DM91, Kha90a, MS90b]. Smooth [GP94, Zel90, CS99b, CP97b, ON95, Sar93]. smoothed [DPSvZ96]. Smoothing [Sar93, Sim95, AS96a, Ano98q, BM98, CS91, Eub99, Fer97, GPS98, MP97, SW96b, TW94, Ter97]. smoothness [Kar91]. Software [Ras96, BX91, Kau96]. SOLS [YGL96]. solution [BM93, SJG93]. solutions [BK96, Vin96]. Some [AB98, BC93, BCFS92, BSS93, BDG90b, Cai99, Car96, CM91, Col96, DK99, EJW93, FS95b, GB98, GF91, GL98a, GS99c, HP92, HL94, Jac90, Jan93a, KP93a, KS97a, Kha90b, KK95b, Kou98, KGS96, Lin97b, Lin97a, Ma97b, Mak94, Mar96, MEJ98, MJE98, Moo96, Mor90, MMC94, NJ99, NK95, NR90, RL99, SS92, Sen99b, SS96a, Sri90e, Stu91b, TAT99, Tun98, WM90, WL96, WCP91, BLBL93, BS92, BS95, cIC95, Csá96, DD92a, DM96c, GR97, HJL95, HH93a, HH96a, HM96, Ham98, Ham97, HKL98, HMW91, Hua95, Hyo92, JW92, Jos96e, Kha97, KVAN90, Kou94, KK95c, KLP98, Lee98a, Ram97, SR90, Sar95, Sar94, SS90a, SS91a, Sen95, SF97, SK92, Siv91, SK97, Ste93, SSSD99, Udd92, ZZ96, ZM93]. Sometimes [SK99]. Space [KSS97, BM98, Che95, Hei93, Shu95, Sri94, SNN90, TX99, TV98]. Space-filling [KSS97]. spaces [CM96, GR93, HS98b, LH97c, PRW99]. spacing [Eks98, GJ98a, Oga98]. spacings [Bar97b, KK95b]. spares [BENP91]. Spatial [BF96, CL96, GNWK99, GL98b, HB98a, LB92, LM99, ME93, Oga96, SM92]. spatio [Rat96]. spatio-temporal [Rat96]. special [GF91, Ios92, Kha90b, Kub92]. species [Nay92, Yat95]. Specification [BZ96, Arn97]. specifications [Bel95]. specificity [Sha95]. specified [KV91, MC95]. Spectral [KM96, GMD99, LCKP95, MST97, PR96]. spectrum [BA95, GR98]. Speculations [Goo90b]. Speed [BD94b]. spherical [GS97b]. spin [Ban96, Yam96]. spline [CS91, Eub99, Hei98, ON95, RC98]. splines [AS96a, HN99, Shi97, TW94]. spring [CK92]. SPRT [Kha97]. square [Bai90, BEHW96, JRM96, LTT93, MT94, Qia94, Ten98, Yan93]. squared [KH95, Law97, LCKP95, Lie92, Oht97]. squares [AZZ96, BD94b, CS99a, CMM94, Du96, FFR97, GTL98, GGBT95, Jan93b, KT92, Khu98, May90, Sch99, UM91, UT97]. Srivastava [Tun98]. Stability [BJL98, BR98, Fou99, Sal94, Wes96, GT95, Qi97, Yan96a]. stabilizing [Lah97]. stable [BL92, SMD95]. stage [BMV97, BM99a, BM92, BFV94, CS91, Coa94, GMR96, GM96, Hay94b, Khu90a, Liu97b, Loh93, SH92a, ZB92]. stages [Tah94]. standard [AJ93, BRZ93, GL99c, KS11, SP91, Siv91, Stu91a]. standardized [Fuj93]. standby [DGK97]. state [BM98, Kim98, Kul99, Shu95, TX99, WF94]. state-space [BM98, Shu95, TX99]. statics [Jen97a]. stationarity [CT98, DS97, KP93b]. stationary [AiFL99, GMD99, Ger94, HP90a, Hai92, Hüs95, KM96, KN94, Lee95, MM98a, McK97, Pen91, RW98a, Rév95, SB94b, SK97]. Statist [Ano98q, Bed00, dML93a]. statistic [DP92, GF91, HP90b, Har92, Iwa97, Jos96a, Mae97, RM95, TTW99, Wan98b]. Statistical [AG99, BN91, DY94a, ED94, Goo90c, GT97, Jen97a, KS11, LT00, Sen95, YC98, Adk94, BMV97, BDG90b, Coa97, Den96, Des97a, FS95a, Fou97, FF97, GL98a, Jan99, KT99b, KGS96, Les97, LR97, LDG98, MPSM94, Nay92, Nel97, PS90b, Rub90, Sri90e, Sri94, SNN90, Wir97, Bas95a, Ter97].

Statistics [Jen97a, Jen97b, AP99, AB98, Ahm94, Ano97-35, AHP97, BBB96, BS92, BE96, BSS93, BD94c, CS90, Chr92, CF98, Cri92, DT91, Dem90, DP92, DH92, FS94, GB95, GT95, GJ98a, Goo90b, Gov93, HP90a, HW96a, HP92, HP91, HS98b, HMW91, Ios92, JM93, Jan94b, Kam91, Kam95a, KD90, KP96, KH99b, Lan96, Lin90, LR95, Ma98a, Mae91, Mae94, MS90a, Mei98, Mik94, MPSM94, MPZ98, NN97, NR96, NR90, Nik91, NZ91, Qi97, Rag97, Rot99, RvZ92, SR90, Sar95, Sch98a, Seo90, Seo94, Sho97, Tho93, Tra90, Wei99, YB99]. Statistik [Tho97]. Stein [DS99, Mol96]. Steiner [CM96, DG98, Key96, KS97a, MY96a, TW97]. step [BL99, FR99, HT91, Som99]. step-down [BL99, FR99, HT91]. step-up [FR99]. stepup [GP99]. Stepwise [Hsu92, TD99]. Stirling [dML93a, dML93b]. Stochastic [AHT99, BQA99, DF94, DKR97, WH92, AR97, Ber98a, BD94a, Cha95c, CM95, Fei94, Ger94, Hey92, Hu98, HOGF99, JMVH97, KS94, MBS99, MG96b, Mus91, NR90, QGG96, QK98, Rao94, RZ98, RV95, TT92a, WF94]. stock [LR96a]. stopped [Adk94]. Stopping [Lin97c, PT94, Róz92a]. storage [CHF92]. strata [BF99, Yan90a]. Strategies [Pad94, Wei99, CR97a, FP94, GC99, SJ96a, YK92, ZZ96]. stratification [CLH98, CHH99]. stratified [MPSM94]. stratum [He93]. strength [DP91, HSS96, HSS97, Nan95]. stress [DP91]. stress-strength [DP91]. Strictly [KWZ93]. string [HM92]. Strong [Enn95, Hu98, Rév95, Stu94, CS96, Chr92, CM92, HP90a, Hai92, LOS95, LC97, MT99]. strongly [Str92]. Structural [Ame96, MK96a, DCK95, Jos96f, Pfa94, SS96a]. structure [DP92, GL98b, KD90, Kub92, Kun93, Oma97, SS93b, SM98, SH92c, ST93, SS94, VP94]. structured [Den96, Den98]. structures [BNJ97, ME91, Sch98b, Smi97c, Yok97]. studentization [Lah97]. studentized [Mae97, Fuj93]. Studies [SO92, ADP91, AE93, BH99b, CH94, GA95, HRR97, JWJB99, Llo98, Mat94, WC95, ZW99]. study [BRM92, DM96c, Dou96, Hau99, JMVH97, Sha99, Spi94]. sub [PZ99]. sub-family [PZ99]. subblock [KM97]. subclass [Ome95]. subgradients [Gut93]. subject [MO96]. submodels [LB94]. subsampling [PRW99, She98]. Subset [Hoo90, SP97, Dou96, FG94, Hof95, LR96b, Liu95a, vdL96, Chi92]. subsets [CM96, KSS97, May90]. subsquares [Du96]. subuniform [DM94]. Successes [Erd98]. Successive [Kau96, LS95a, And93]. Sufficiency [Ram90, Bha91, Gro98a, Jos96e, Jos96a, Lan96]. Sufficient [VSB98, GTL98, Jos96a, MS90a, RS93, Vil91]. sugar [Dou96]. sum [LH97a, MM98a]. summation [RS96]. sums [CM95, CM92, FW93, Jan93b, Khu98, KS92]. super [Kej96, KK95c, SSE98]. super-simple [Kej96]. superblock [KM98a]. superdiagonal [BH92b]. superimposed [Mor98]. superiority [GTL98]. superpopulation [Muk94]. Supersaturated [Mül93, CL98, YIHN99]. supplement [Pfa91]. supplementary [CC96, KSWyX97, yXL96]. support [GH98a, KVAN90]. supported [DGK97]. Supports [Din96]. sure [Mik94, Pur95]. surface [Dra97, KV98, MH92a, Pow95, Sri90e, VK99, WM90, WD98]. surfaces [DGP93]. surrogate [Con96b, EMS98, Sen94]. Survey [Smi90, CR97a, Gly98, LT97, Mah98]. surveys [ASK97, Tan99, TSS96]. survival [CW97b, CP95, Hay98, IT90, KP97, KZ99, Mau98, MBS99, SF97, TZ93, dLS99]. suspected [MEG93]. Switching [BM98, WCP91]. symbol [HSS97]. Symmetric [Nie96, DT91, Eat97, Kus99, Mon90, Oga98, SP91, Sul93, WJ96]. symmetries [Pro93]. symmetry [NG91, SCM94, Seo90, Tom93]. symplectic [BNJ97, WL96]. system [Ana99, AB97b, BABM99, DGK97, GUH91, LB91, MY96a]. systematic [KSI99]. systems [BENP91, BL96, BKF98, CM96, Chu96, CHF92, CLR91a, CLR91b, DG98, DKR97, Hen96, dKgL96, dK97, KS97a, gL96, Lyn99, MB95, TW97, WH92, Yan96b, ZB92, vDMS97].

T [Jen97b, GR98]. T. [Jen97b]. tableaux [Kra96]. tables [PPO95, PW99, RS92, Tom93]. tactics [JMVH97]. Taguchi [SW95b]. tail [Alv95, BB95, BB98, BBTV95, Bro93, Deh92a, Ein92, Höp97, Lea95, Tom95]. tailed [ALW98]. tails [MS98a]. Tallini [Bar98]. target [HF95]. technique [HM95]. techniques [MH92b]. technologies [Wei99]. temporal [Hey92, JM98, Rat96]. tensile [Nan95]. tensile-strength [Nan95]. term [GR93, Hei91, Sta92]. ternary [BKM93, HW98, SS93a]. Terry [HZzY91]. Test [Hay94a, QGG96, AP99, AM91, BP93, BE96, Ben96, BZ96, CM98, CL95, CF98, Cut93a, Cut93b, DL98b, Di 99, Die95, DM91, DD92b, DR96, EBS96, El 96, FP98, GG96, GJ98b, GP99, Gup98, HT91, HR95, HS95, HMW91, JSH99, Kha90a, KH99b, KLP98, LR92, Lee98b, Lia90b, LS99b, Liu95b, LB91, Ma97a, MS90b, MPSM94, Muk98b, PD99, Rio98, Rot99, SA97, Sha95, Stu91a, SL95, SMD95, TD99, TY96, TT92b, TAT99, WS95, Wan96, Wan98a, YB99, ZWF95]. Testing [Cha97b, CU92, DMNN99, Ebr97, EL93, Har99, Jan99, Jar98, Koc98, KS97d, Lüt96, NK94, PPO95, RCGMC98, Sei93, Ahm94, AP97, Ana99, BE91, BL99, BM94b, Cai99, CL92, CBM94, FG96, GF91, HW96a, HYT98, HR95, HMT99, Hua95, Kin96, LK98, MH92b, MO96, Neu99, NP93, Oma97, PT94, PK99, RH95, Sin95, SS96a, Som99, TS91, VP94, WRKL99, YH90, ZZ96, Law97]. Tests [Hor98, JM91, Khu92, KS98d, LS95b, LS95c, PS98b, RW98a, Yok97, Akr91, AA93, AB97a, Alb92, BRZ93, BB91, BB99a, BH92b, BH98, Ber98a, BD94a, BS98a, BS95, Bou90, BY95, CXL93, CW98, Con96a, CP98, God96, GR97, GS99c, HLM90, HJPZ99, Iwa98, Jan94a, JN97, Jan99, JK94a, JK97, KN94, Khu90a, KT99b, Liu95b, MGdW96, Man90a, Man98, Man99, MM96, Mei97, MT94, MMM95, NL97, PT92, Pfa94, Sei92, Sen99b, SP91, Slu92, SM96, Tse94, Zha96, ZB92, ZM93]. their [God96, GNZ97, KK95a, KS97c, Nie96, TW97]. them [SW98]. theorem [BS92, GC97, GS99a, Jos96c, PRS97, Pro96, Ral93b, SC99a, SS90c, Sta97, Wan97a, Yan93, Pfa91]. theorems [AS96b, Bar90a, BS98b, CS96, Ein92, HS95, Kim97b, Nie92, Rév95, TS95, WL96, WC99, ZZ96]. theoretic [HM97]. theoretical [Kle91]. Theory [Jen97a, AP97, Bel95, BN91, DKS98, Die95, DK94, FS95b, FF97, HS98b, NR96, Ros97a, Sri90e, Ter97, Zhe96]. thickness [MS98a]. Thompson [Ber98b, Ber98c]. Three [AZZ96, Loh93, Bar92, BS93c, CMS93, Coa94, DM93, dK97, Kou98, Mac97a, Mar97b, RH97, Tah94, YIHN99]. three-dimensional [CMS93, Mar97b]. three-fold [RH97]. three-level [YIHN99]. Three-stage [Loh93, Coa94]. three-way [BS93c]. threshold [Bro94, LLL97, Luo96, Qia98]. ties [Jan94a]. tightening [PPO95]. Time [dLS99, AP98, AP99, AK93, AM98b, BG91, Bro94, Cai99, CK99, Fou97, GMD99, GGJ98, GH96a, HJPZ99, Har99, HB94, KC93, LY95a, LC97, MT99, Mau98, MST97, MB97b, NK95, ND98, OS96, RW98a, RL99, RS98, RZ98, Ruy94, SS91b, Shi97, TA91, TP97, Tra94, Tru91, YS92]. time-censored [TP97]. time-deformed [GGJ98]. Time-dependent [dLS99, AM98b, LY95a]. time-series [OS96, RL99]. times [Ana99, BP95, DN94, Enn95, Hay98, HF95, JM91, MB97a, Neu99, Rio98]. tolerance [GEGM99]. tolerated [MW97]. tool [Kle91]. top [PR96]. total [BZS96, MS95b, MB94b]. totals [LB90]. tournaments [HZzY91]. Toutenburg [AC97a]. Toyooka [UT97]. tracers [Phe96]. trade [HM95, dlHF94]. trade-off [HM95, dlHF94]. trades [GR98, KS97b]. traffic [Add99]. Training [VY98]. transform [Man98, Man99]. transformation [Cza97, HS97, MN97, WR96]. transformations [Lah97, MST97]. Transient [Phe96, BM93, SJG93]. transition [CS99c]. transitions [Bel95]. transitive [KS97a, KSI99, gL96, LMT97, MP96]. transitivity [Bos98]. translation [Gly97, Wet97]. transversal [Col96]. treatment [BM99b, DR96, GRP99, LW94, Luo96, Stu91a, TN91]. treatment-control [GRP99, TN91]. treatment-treatment [GRP99]. treatments [AM91, Bai90, CW98, Cut93a, Cut93b, ED94, GLMN98, GJ98b, JM98, Jac90, KZ99, LS95a, Liu97a, MT96, SH92b, Stu91a, TY96]. tree [Den96, Den98, FP98, SH92c]. tree-structured [Den96, Den98]. trees [AIV95, FK98, Mon98]. trend [Cha95a, GJ98b, HJ98, Jar98, LS99a, SM92, TS91, WC98a]. trend-free [Cha95a, HJ98, LS99a]. trended [GJ98c]. trends [Jac98, PPO95]. trial [CW97b]. trials [AG98, AG99, Coa94, DK94, Eis94, EMS98, ED94, FP94, JWJB99, LW94, Lee99, Sen94, Sim94, WRKL99, Wu94]. triangular [Car96, SK90]. Trimmed [Fre91, GEGM99, Lee95, MS90a, Sho97]. triple [CM96, Chu96, DG98, dKgL96, dK97, KS97a, gL96, TW97, Yan96b]. trivariate [EM99]. Truncated [CHH99, Asa99, Gas96a]. truncation [TZ93]. TTT [HS97, Kle91]. TTT-plotting [Kle91]. tumor [LB90, Neu99]. turbulence [LW99b]. turns [CMS93]. twelve [Bai90]. Two [Bag96, BM92, GMR96, Jan94a, JN97, LZ98, Lec99, Oga96, ZB92, BC93, Bai90, BDG90a, BS93a, BGK95, Bar90a, BFV94, CM93a, CM93b, CS91, CH98, CL95, Coa94, CW97b, DP92, DMNN99, DL98b, DL90, DM91, ESA98, GL99b, GC99, HJPZ99, Hay94b, HP97, JPW97, JJMY95, Jos91, KP97, KM97, KZ99, KS99, KS97d, LD93, Liu94, MR98, Mae91, MGdW96, MJE98, Mei98, MS90b, Mor90, MD94, Mül93, PT94, RP97, Ruc97, SH92a, SP97, Sar95, Sat93, SB95b, Shi91, SSS94, SC99b, SR99, TP97, TN98, TN91, Tom93, TM96, Udd96, VS93, VS97, WJ96, Wan99, WL96, Yu96, Zou92]. Two- [BM92]. two-armed [GC99]. two-dimensional [Bar90a, Jos91, Mor90, SB95b, TN98]. two-element [Liu94]. two-factor [CM93a, CM93b, CH98, HP97, MD94, Ruc97, SP97, SC99b]. two-fold [SR99]. two-level [DL90, HP97, JJMY95, MJE98, RP97]. two-parameter [TP97]. two-phase [ESA98, SSS94]. two-replicate [JPW97]. Two-sample [Jan94a, JN97, DL98b, Mei98, Sar95]. two-sided [DM91, MS90b, PT94]. Two-stage [GMR96, BFV94, CS91, Coa94, Hay94b, SH92a]. two-variance [VS97]. two-way [BDG90a, BS93a, BGK95, MGdW96, TN91, Tom93]. type [AFJ97, AHJ96, BRZ93, BE96, Che97, CF98, DKS98, GB98, HC92, Iso94, JM93, Kak97, Li98a, PT92, RM95, RS97a, Sen99b, Shi97, SL96, YGL96, Yan95]. types [dK97]. Typologies [Hen96].

Ulam [Mac97b]. unavailability [MNWZ95, SS92]. unbalanced [Khu98, LN99, Sei92, Ste91, VH98]. Unbiased [Buz98, DDB99, Kna99, Pad94, Rao94, Ste91, SNN90, Wu98]. unbounded [HP90b, HP91, Har92]. uncertainty [Ebr97, Siv91]. unconditional [KH99a]. underdetermined [GJ96]. understanding [BT97]. unequal [AM91, BHT91, IJ94, Khu90a, RJ99]. unequally [Jac90]. Unified [CP97b, AB97a, AM96, NRZ98, RS96, dML93b, dML93a]. Uniform [DM94, ND98, Sho97, AB98, BC99, Lia90b]. uniformly [Liu95b, Wu98, YC95]. unifying [HJL95]. unilateral [Mar96, Mar97b]. union [HS95]. union-intersection [HS95]. Uniqueness [vR93, HH96b, Mah98]. Unit [Cha96b, MK96a]. unitals [CQ98]. units [BH92a, BD96b, HS98a]. univariate [HJL95, NJ99]. Universal [DD92a, Mar97a]. unknown [GMR96, JM91, Kou95, LG91, Nay92]. unlabelled [TS97]. unlimited [AB97b]. unobservable [MB97a]. unsigned [HLM90]. unstructured [PFvR97]. untruthful [LR92]. unweighted [Khu98]. up-and-down [GP98]. update [Sha92]. upper [Bar90a, HP90a, Hoo90]. urn [MI90, SJ96a]. USA [Jen97b]. Use [BA92b, ASK97, Con96b, DEM95, Hey96, SW98, TS96]. used [Ahm94, HW96a]. useful [Ahm94, BCFS97, FP98, Kou91, Lec99, PS90b, VP94]. uses [God96]. Using [LS95e, PZ99, AS96a, BM99a, BE91, BABM99, BJ90, CW98, Dra97, Ham91, HMM97, HYT98, Hua95, JWT93, KS98a, KS99, LB90, Mal95, Mau98, Mie90, MS95a, Mol96, NRZ98, Nay90, Ras98, RCGMC98, SP97, Ste97, TG93, TK92, Wei99, Yan95]. usual [Sar94]. utilizing [SO92].

V [Tho97, GR98]. vaccines [ED94]. valid [Wan98b]. validation [HRR97, Stu92]. validatory [YT98]. validity [FMN98]. value [AG98, AG99, Ber95b, Chr94, Dre98, FCA94, HF95, LB98, NR96, OD99, SE95, Wei97]. valued [MST97, Ram97, SNN90]. values [Bar90a, Ber95b, Car99, CW98, FR99, Gas93, Hüs95, Kam95b, NN97, Rév95, Som99]. Vanzetti [Ano97j]. variable [BS96, HL94, Kou98, NK95, TX99, YC95]. variables [BES90, BS98c, Car96, Fer95, GJ98c, GH90, GH96b, HMT99, LH97a, LDG98, MG96a, MPZ98, Muk94, SSK99, TR90, TS96]. Variance [Lah97, BDG90a, BGK95, Ber98b, CP97a, CR97b, DG95, DGP93, EBS96, Fer97, Fro90, Gui99b, HK94, IK98, Khu98, KR93, KWZ93, Kot90, KMMN93, Mad95, Mae97, MGdW96, Man90a, Mol96, MS93b, Pad94, RS97a, Sch99, Sei92, Ses97, SW97, TG93, Van98, VS97, Wan96, WB94, Wit98, ZM93]. variances [BHT91, CK92, LLL97, MMM95, Wan99, Ye94]. variation [Sul93]. varieties [Moo96, Goo96b]. various [JD91]. varying [Ome95]. vector [BM90, BP93, CC96, EM99, HTY98, He93, Hof95, Lüt96, MY96b, Wan99, WC99]. vector-elliptically [WC99]. vectors [DM97]. Verfahrensbibliothek [Tho97]. Verlag [AC97a, Ter97, Tho97]. version [BMV97, Mac97b, SW95b, Ten98]. Versions [Wan97a, AS96b, CFM^{+}98]. Versuchsplanung [Tho97]. versus [GJ98c, Gup98, Hus99, STL96, TS96, dSDF95]. very [AHT98]. very-high-dimensional [AHT98]. via [AiF97, BS98b, BY95, Cha97a, Cri92, GMD99, Har99, QK98, Sch98a, SE92, SK99, SWA97, TH99, Wan97b]. Victor [Tho97]. view [BDG90a, DV97, Lee99, LN99]. viewpoint [Hor95]. violation [SA97]. vol [Jen97b]. volatility [Haf98, RZ98]. Volume [Ano90a, Ano90b, Ano90p, Ano91c, Ano91a, Ano91b, Ano92a, Ano92b, Ano92c, Ano92d, Ano93a, Ano93b, Ano93c, Ano93d, Ano94a, Ano94b, Ano94c, Ano94d, Ano94e, Ano95a, Ano95b, Ano95c, Ano95d, Ano95e, Ano95f, Ano96f, Ano96g, Ano96h, Ano96i, Ano96j, Ano96k, Ano96l, Ano96m, Ano97a, Ano97b, Ano97c, Ano97d, Ano97e, Ano97f, Ano97g, Ano97h, Ano97i, Ano98a, Ano98b, Ano98c, Ano98d, Jen97a]. volumes [Ano92i, Ano92j].

W [Jen97b, Ter97]. W. [Smi97c]. walk [GJ98c, SR90]. walks [KP96, Zei96]. warranties [Sin91]. Wasserstein [CM92]. wavelet [Pat97]. wavelets [CT98, Wan97b]. way [BDG90a, BS93a, BGK95, BS93c, Che98b, DG95, KL99, MGdW96, Mol96, PT92, RN95, Sch91, aS90, Ste91, TN91, Tom93, Ye94]. Wayne [Fou97]. Weak [BB95, HP90b, Har92, PRW99, Rag97, Add99, DPSvZ96, IT90, Kou91, MT99, Mus91, Stu94, Wu97]. weakly [TV98]. weevils [JMVH97]. Weibull [BABM99, BM92, Bro93, Che98a, PT92, Sun97a]. Weighing [KS97c, CK92, CF96, Far91, Far92, KS99, SS90a, SS91a, Str98]. weight [SO92, Zin96]. Weighted [Gui99b, GRP99, Jan94b, MBL97, ST96b, CL97, CM92, ER99, FMN99, Gui99a, GA95, Lie92, NJ99, Rag97, Sch99]. weights [JN97, KS98d, WC95]. wheel [CLR91a, CLR91b]. whether [Ebr97]. which [Cha90, Cha96c, LW94, Ros92, SMS96]. white [DM91]. Whitney [Di 99]. whose [SH92a]. width [Col95, KW92, Mec92, MS98b, TN95]. Wien [Tho97]. Wilcoxon [Di 99, SL95]. Wilcoxon-Mann-Whitney [Di 99]. Wiley [Ano97j, Bas95a, Coa97, Dra97, Fou97, Law97]. William [Smi97a]. Williams [SS94]. winner [BB99a]. within [CX98a, KV91]. without [CH96, DS97, MP98, SCM94, Ste91, Vos93]. WLLN [Sho97]. Wolfowitz [Mec92]. work [Bar98, Lin90].

X [AC97b, Law97, Smi97c, Tho97, AFJ97]. XII [AHJ96]. xiv [Coa97]. xvi [Smi97b]. xxi [Sal92].

Yokes [BNJ97]. York [Ano97j, Bas95a, Coa97, Dra97, Fou97, Law97, Sal92, Smi97a, Smi97b]. Youden [Mor97]. Young [Ges93, Gho96].

zero [Giu95, KM99, KS99, RT90]. zero-one [KM99]. zone [Gui96]. zones [GEGM99].

References

Akritas:1993:ART

[AA93] Michael G. Akritas and Willem Albers. Aligned rank tests for the linear model with heteroscedastic errors. Journal of Statistical Planning and Inference, 37(1):23–41, October 1993. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

            URL http://www.sciencedirect.com/science/article/pii/
            037837589390077J.

Anh:1999:PLR

[AARM99] V. V. Anh, J. M. Angulo, and M. D. Ruiz-Medina. Possible longrange dependence in fractional random fields. Journal of Statistical Planning and Inference, 80(1–2):95–110, August 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

            URL http://www.sciencedirect.com/science/article/pii/
            S0378375898002444.

Akritas:1997:UAR

[AB97a] Michael G. Akritas and Edgar Brunner. A unified approach to rank tests for mixed models. Journal of Statistical Planning and Inference, 61(2):249–277, June 16, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

            URL http://www.sciencedirect.com/science/article/pii/
            S0378375896001772.

Armero:1997:BAQ

[AB97b] C. Armero and M. J. Bayarri. A Bayesian analysis of a queueing system with unlimited service. Journal of Statistical Planning and Inference, 58(2):241–261, March 15, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896000869.

Aggarwala:1998:SPP

[AB98] Rita Aggarwala and N. Balakrishnan. Some properties of progressive censored order statistics from arbitrary and uniform distributions with applications to inference and simulation. Journal of Statistical Planning and Inference, 70(1):35–49, July 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897001730.

Ahmad:1994:NEJ

[AC94] Ibrahim A. Ahmad and Patricia B. Cerrito. Nonparametric estimation of joint discrete-continuous probability densities with applications. Journal of Statistical Planning and Inference, 41(3):349–364, October 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375894900280.

Anderson-Cook:1997:BRE

[AC97a] C. M. Anderson-Cook. Book review: Experimental design and model choice: H. Toutenburg, (Physica Verlag, Heidelberg, 1995, pp. 463). Journal of Statistical Planning and Inference, 59(1):187–189, March 30, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896001000.

Anderson-Cook:1997:BRL

[AC97b] C. M. Anderson-Cook. Book review: Linear models: a mean model approach: Barry Kurt Moser, (Academic Press, San Diego, CA, 1996, $49.95, 0-12-50846-X, 228 pp.). Journal of Statistical Planning and Inference, 64(1):153–155, October 30, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897865202.

Atkins:1999:ORD

[AC99] Joel E. Atkins and Ching-Shui Cheng. Optimal regression designs in the presence of random block effects. Journal of Statistical Planning and Inference, 77(2):321–335, March 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837589800189X.

Azais:1997:ONB

[AD97] Jean-Marc Azaïs and Pierre Druilhet. Optimality of neighbour balanced designs when neighbour effects are neglected. Journal of Statistical Planning and Inference, 64(2):353–367, November 15, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375896002133.

Addie:1999:WCL

[Add99] R. G. Addie. On weak convergence of long-range-dependent traffic processes. Journal of Statistical Planning and Inference, 80(1–2):155–171, August 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375898002481.

Adke:1994:OSI

[Adk94] S. R. Adke. Optimum statistical inference from randomly stopped random processes. Journal of Statistical Planning and Inference, 39(2):255–271, April 15, 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894902089.

Albert:1991:CDR

[ADP91] James Albert, Mohan Delampady, and Wolfgang Polasek. A class of distributions for robustness studies. Journal of Statistical Planning and Inference, 28(3):291–304, July 1991. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589190068P.

Arjas:1993:PCL

[AE93] Elja Arjas and Mervi Eerola. On predictive causality in longitudinal studies. Journal of Statistical Planning and Inference, 34(3): 361–386, March 1993. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589390146W.

Athreya:1992:BMC

[AF92] K. B. Athreya and C. D. Fuh. Bootstrapping Markov chains: countable case. Journal of Statistical Planning and Inference, 33(3):311–331, December 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290002A.

Alam:1997:EPO

[AF97] Khursheed Alam and Zhuojen Feng. Estimating probability of occurrence of the most likely multinomial event. Journal of Statistical Planning and Inference, 59(2):257–277, April 15, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896001127.

Ahmad:1997:EBE

[AFJ97] K. E. Ahmad, M. E. Fakhry, and Z. F. Jaheen. Empirical Bayes estimation of P (Y < X) and characterizations of Burr-type X model. Journal of Statistical Planning and Inference, 64(2):297–308, November 15, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897000384.

Andres:1998:DVM

[AG98] A. Martín Andrés and J. M. Tapia García. On determining the P -value in 2 x 2 multinomial trials. Journal of Statistical Planning and Inference, 69(1):33–49, June 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897001535.

Andres:1999:DVM

[AG99] A. Martín Andrés and J. M. Tapia García. On determining the p-value in 2 × 2 multinomial trials: Journal of Statistical Planning and Inference 69 (1998) 33–49. Journal of Statistical Planning and Inference, 79(2):365, July 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898002651.

Abt:1998:ODG

[AGLS98] Markus Abt, Norbert Gaffke, Erkki P. Liski, and Bikas K. Sinha. Optimal designs in growth curve models — II correlated model for quadratic growth: optimal designs for parameter estimation and growth prediction. Journal of Statistical Planning and Inference, 67(2):287–296, April 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897001304.

Arasu:1996:NCG

[AH96] K. T. Arasu and Susan Harris. New constructions of group divisible designs. Journal of Statistical Planning and Inference, 52(2):241–253, June 15, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895001115.

Al-Hussaini:1999:POG

[AH99] Essam K. Al-Hussaini. Predicting observables from a general class of distributions. Journal of Statistical Planning and Inference, 79 (1):79–91, June 1999. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375898002286.

Al-Hussaini:1996:BPB

[AHJ96] Essam K. Al-Hussaini and Zeinhum F. Jaheen. Bayesian prediction bounds for the Burr type XII distribution in the presence of outliers. Journal of Statistical Planning and Inference, 55(1):23– 37, October 15, 1996. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375895001840.

Ahmad:1994:CSU

[Ahm94] Ibrahim A. Ahmad. A class of statistics useful in testing increasing failure rate average and new better than used life distributions. Journal of Statistical Planning and Inference, 41(2):141– 149, September 1994. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375894901597.

Ahmad:1999:NEE

[Ahm99] Ibrahim A. Ahmad. Nonparametric estimation of the expected duration of old age. Journal of Statistical Planning and Inference, 81(2):223–227, November 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375899000208.

Antoch:1997:EDS

[AHP97] J. Antoch, M. Huskova, and Z. Prásková. Effect of dependence on statistics for determination of change. Journal of Statistical Planning and Inference, 60(2):291–310, May 15, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896001383.

Anderson:1998:EEV

[AHT98] Niall H. Anderson, Peter Hall, and D. M. Titterington. Edgeworth expansions in very-high-dimensional problems. Journal of Statistical Planning and Inference, 70(1):1–18, July 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897001717.

Anh:1999:SMF

[AHT99] V. V. Anh, C. C. Heyde, and Q. Tieng. Stochastic models for fractal processes. Journal of Statistical Planning and Inference, 80(1–2):123–135, August 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898002468.

Athreya:1997:CIE

[AiF97] Krishna B. Athreya and Jun ichiro Fukuchi. Confidence intervals for endpoints of a c.d.f. via bootstrap. Journal of Statistical Planning and Inference, 58(2):299–320, March 15, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896000870.

Athreya:1999:BMB

[AiFL99] Krishna B. Athreya, Jun ichiro Fukuchi, and Soumendra N. Lahiri. On the bootstrap and the moving block bootstrap for the maximum of a stationary process. Journal of Statistical Planning and Inference, 76(1–2):1–17, February 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898001402.

Ahlbrandt:1995:ODS

[AIV95] Robert Ahlbrandt, Hari K. Iyer, and Dominic F. Vecchia. Optimum design of serial measurement trees. Journal of Statistical Planning and Inference, 48(3):379–390, December 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589500012X.

Abhyankar:1993:GCS

[AJ93] Shreeram S. Abhyankar and Sanjeevani B. Joshi. Generalized coinsertion and standard multitableaux. Journal of Statistical Planning and Inference, 34(1):5–18, January 1993. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375893900296.

Alam:1993:EQF

[AK93] Khursheed Alam and K. B. Kulasekera. Estimation of the quantile function of residual life time distribution. Journal of Statistical Planning and Inference, 37(3):327–337, December 1993. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390111I.

Akritas:1991:ADA

[Akr91] Michael G. Akritas. An alternative derivation of aligned rank tests for regression. Journal of Statistical Planning and Inference, 27(2):171–186, February 1991. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375891900146.

Altman:1995:BSK

[AL95] Naomi Altman and Christian Léger. Bandwidth selection for kernel distribution function estimation. Journal of Statistical Planning and Inference, 46(2):195–214, August 1, 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894001022.

Albers:1992:DCR

[Alb92] Willem Albers. Deficiencies of combined rank tests. Journal of Statistical Planning and Inference, 32(1):5–24, July 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290149M.

Abt:1997:ODG

[ALMS97] Markus Abt, Erkki P. Liski, Nripes K. Mandal, and Bikas K. Sinha. Optimal designs in growth curve models: Part I: Correlated model for linear growth: Optimal designs for slope parameter estimation and growth prediction. Journal of Statistical Planning and Inference, 64(1):141–150, October 30, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896002121.

Alves:1995:ETP

[Alv95] M. I. Fraga Alves. Estimation of the tail parameter in the domain of attraction of an extremal distribution. Journal of Statistical Planning and Inference, 45(1–2):143–173, ???? 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894000689.

Athreya:1998:IHT

[ALW98] K. B. Athreya, S. N. Lahiri, and Wei Wu. Inference for heavy tailed distributions. Journal of Statistical Planning and Inference, 66(1):61–75, January 5, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897000773.

Angelis:1991:OIB

[AM91] Lefteris Angelis and Chronis Moyssiadis. A-optimal incomplete block designs with unequal block sizes for comparing test treatments with a control. Journal of Statistical Planning and Inference, 28(3):353–368, July 1991. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589190072M.

Al-Mutairi:1996:MFM

[AM96] Dhaifalla K. Al-Mutairi. Multivariate failure models generated by random environments: a unified approach. Journal of Statistical Planning and Inference, 55(2):249–253, October 30, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375896000535.

Altman:1998:CBS

[AM98a] Naomi Altman and Brenda MacGibbon. Consistent bandwidth selection for kernel binary regression. Journal of Statistical Planning and Inference, 70(1):121–137, July 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897001766.

Azrak:1998:EQL

[AM98b] Rajae Azrak and Guy Mélard. The exact quasi-likelihood of time-dependent ARMA models. Journal of Statistical Planning and Inference, 68(1):31–45, May 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897001341.

Amemiya:1996:SDM

[Ame96] Takeshi Amemiya. Structural duration models. Journal of Statistical Planning and Inference, 49(1):39–52, January 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895000291.

Andersson:1995:RBC

[AMPT95] Steen A. Andersson, David Madigan, Michael D. Perlman, and Christopher M. Triggs. On the relation between conditional independence models determined by finite distributive lattices and by directed acyclic graphs. Journal of Statistical Planning and Inference, 48(1):25–46, November 1, 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589400150T.

Ananda:1999:ETA

[Ana99] Malwane M. A. Ananda. Estimation and testing of availability of a parallel system with exponential failure and repair times. Journal of Statistical Planning and Inference, 77(2):237–246, March 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375898001840.

Andrews:1993:DSG

[And93] George E. Andrews. On the difference of successive Gaussian polynomials. Journal of Statistical Planning and Inference, 34(1): 19–22, January 1993. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589390030A.

Anonymous:1990:AIVa

[Ano90a] Anonymous. Author index: Volume 24 (1990). Journal of Statistical Planning and Inference, 24(3):405–406, March 1990. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375890900608.

Anonymous:1990:AIVb

[Ano90b] Anonymous. Author index: Volume 26 (1990). Journal of Statistical Planning and Inference, 26(3):371–372, November 1990. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589090139L.

Anonymous:1990:C

[Ano90c] Anonymous. Correction. Journal of Statistical Planning and Inference, 25(2):205, June 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837589090070B.

Anonymous:1990:EBa

[Ano90d] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 24(1):ifc, January 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837589090011I.

Anonymous:1990:EBb

[Ano90e] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 25(1):ii, May 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837589090001B.

Anonymous:1990:EBc

[Ano90f] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 26(1):ii, September 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589090090H.

Anonymous:1990:PMb

[Ano90g] Anonymous. Pages 1-110 (May 1990). Journal of Statistical Planning and Inference, 25(1):??, May 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1990:PS

[Ano90h] Anonymous. Pages 1-130 (September 1990). Journal of Statistical Planning and Inference, 26(1):??, September 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1990:PJa

[Ano90i] Anonymous. Pages 1-136 (January 1990). Journal of Statistical Planning and Inference, 24(1):??, January 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1990:PJb

[Ano90j] Anonymous. Pages 111-205 (June 1990). Journal of Statistical Planning and Inference, 25(2):??, June 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1990:PO

[Ano90k] Anonymous. Pages 131-252 (October 1990). Journal of Statistical Planning and Inference, 26(2):??, October 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1990:PF

[Ano90l] Anonymous. Pages 137-270 (February 1990). Journal of Statistical Planning and Inference, 24(2):??, February 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1990:PJc

[Ano90m] Anonymous. Pages 207-468 (July 1990). Journal of Statistical Planning and Inference, 25(3):??, July 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1990:PN

[Ano90n] Anonymous. Pages 253-372 (November 1990). Journal of Statistical Planning and Inference, 26(3):??, November 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1990:PMa

[Ano90o] Anonymous. Pages 271-406 (March 1990). Journal of Statistical Planning and Inference, 24(3):??, March 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1990:V

[Ano90p] Anonymous. Volume 25 (1990). Journal of Statistical Planning and Inference, 25(3):467–468, July 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589090089D.

Anonymous:1991:AIVa

[Ano91a] Anonymous. Author index volume 27 (1991). Journal of Statistical Planning and Inference, 27(3):405–406, March 1991. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589190053H.

Anonymous:1991:AIVb

[Ano91b] Anonymous. Author index volume 28 (1991). Journal of Statistical Planning and Inference, 28(3):405–406, July 1991. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589190076Q.

Anonymous:1991:AIVc

[Ano91c] Anonymous. Author index volume 29 (1991). Journal of Statistical Planning and Inference, 29(3):367–368, November 1991. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375891900113.

Anonymous:1991:EBa

[Ano91d] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 27(1):ii, January 1991. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837589190077R.

Anonymous:1991:EBb

[Ano91e] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 28(1):ifc, May 1991. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837589190054I.

Anonymous:1991:EBc

[Ano91f] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 29(1–2):ii, September/October 1991. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375892901148.

Anonymous:1991:PMb

[Ano91g] Anonymous. Pages 1-138 (May 1991). Journal of Statistical Planning and Inference, 28(1):??, May 1991. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1991:PJa

[Ano91h] Anonymous. Pages 1-142 (January 1991). Journal of Statistical Planning and Inference, 27(1):??, January 1991. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1991:PSO

[Ano91i] Anonymous. Pages 1-228 (September-October 1991). Journal of Statistical Planning and Inference, 29(1–2):??, September/ October 1991. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1991:PJb

[Ano91j] Anonymous. Pages 139-270 (June 1991). Journal of Statistical Planning and Inference, 28(2):??, June 1991. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1991:PF

[Ano91k] Anonymous. Pages 143-270 (February 1991). Journal of Statistical Planning and Inference, 27(2):??, February 1991. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1991:PN

[Ano91l] Anonymous. Pages 229-368 (November 1991). Journal of Statistical Planning and Inference, 29(3):??, November 1991. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1991:PJc

[Ano91m] Anonymous. Pages 271-406 (July 1991). Journal of Statistical Planning and Inference, 28(3):??, July 1991. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1991:PMa

[Ano91n] Anonymous. Pages 271-406 (March 1991). Journal of Statistical Planning and Inference, 27(3):??, March 1991. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1992:AIVa

[Ano92a] Anonymous. Author index volume 30 (1992). Journal of Statistical Planning and Inference, 30(3):423–424, March 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290168R.

Anonymous:1992:AIVb

[Ano92b] Anonymous. Author index volume 31 (1992). Journal of Statistical Planning and Inference, 31(3):391–392, June 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290145I.

Anonymous:1992:AIVc

[Ano92c] Anonymous. Author index volume 32 (1992). Journal of Statistical Planning and Inference, 32(3):425–426, September 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290024M.

Anonymous:1992:AIVd

[Ano92d] Anonymous. Author index volume 33 (1992). Journal of Statistical Planning and Inference, 33(3):427–428, December 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290010P.

Anonymous:1992:EBa

[Ano92e] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 30(1):ii, January 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375892901007.

Anonymous:1992:EBb

[Ano92f] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 31(1):ii, April 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837589290036R.

Anonymous:1992:EBc

[Ano92g] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 32(1):ii, July 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837589290146J.

Anonymous:1992:EBd

[Ano92h] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 33(1):ii, October 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837589290089B.

Anonymous:1992:MIVa

[Ano92i] Anonymous. Master index volumes 21-30. Journal of Statistical Planning and Inference, 30(??):425–439, ???? 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290136G.

Anonymous:1992:MIVb

[Ano92j] Anonymous. Master index volumes 21-30. Journal of Statistical Planning and Inference, 30(3):425–439, March 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290169S.

Anonymous:1992:PAa

[Ano92k] Anonymous. Pages 1-132 (April 1992). Journal of Statistical Planning and Inference, 31(1):??, April 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1992:PJa

[Ano92l] Anonymous. Pages 1-134 (January 1992). Journal of Statistical Planning and Inference, 30(1):??, January 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1992:PJc

[Ano92m] Anonymous. Pages 1-145 (July 1992). Journal of Statistical Planning and Inference, 32(1):??, July 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1992:PO

[Ano92n] Anonymous. Pages 1-156 (October 1992). Journal of Statistical Planning and Inference, 33(1):??, October 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1992:PMb

[Ano92o] Anonymous. Pages 133-271 (May 1992). Journal of Statistical Planning and Inference, 31(2):??, May 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1992:PF

[Ano92p] Anonymous. Pages 135-292 (February 1992). Journal of Statistical Planning and Inference, 30(2):??, February 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1992:PAb

[Ano92q] Anonymous. Pages 147-280 (August 1992). Journal of Statistical Planning and Inference, 32(2):??, August 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1992:PN

[Ano92r] Anonymous. Pages 157-293 (November 1992). Journal of Statistical Planning and Inference, 33(2):??, November 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1992:PJb

[Ano92s] Anonymous. Pages 273-392 (June 1992). Journal of Statistical Planning and Inference, 31(3):??, June 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1992:PS

[Ano92t] Anonymous. Pages 281-426 (September 1992). Journal of Statistical Planning and Inference, 32(3):??, September 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1992:PMa

[Ano92u] Anonymous. Pages 293-439 (March 1992). Journal of Statistical Planning and Inference, 30(3):??, March 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1992:PD

[Ano92v] Anonymous. Pages 295-428 (December 1992). Journal of Statistical Planning and Inference, 33(3):??, December 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1992:P

[Ano92w] Anonymous. Pages 425-439 (1992). Journal of Statistical Planning and Inference, 30(??):425–439, ???? 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1993:AIVa

[Ano93a] Anonymous. Author index volume 34 (1993). Journal of Statistical Planning and Inference, 34(3):451–452, March 1993. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390151U.

Anonymous:1993:AIVb

[Ano93b] Anonymous. Author index volume 35 (1993). Journal of Statistical Planning and Inference, 35(3):399–400, June 1993. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375893900252.

Anonymous:1993:AIVc

[Ano93c] Anonymous. Author index volume 36 (1993). Journal of Statistical Planning and Inference, 36(2–3):431–432, August/September 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589390142S.

Anonymous:1993:AIVd

[Ano93d] Anonymous. Author index volume 37 (1993). Journal of Statistical Planning and Inference, 37(3):413–414, December 1993. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390116N.

Anonymous:1993:EBa

[Ano93e] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 34(1):ifc, January 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375893900263.

Anonymous:1993:EBb

[Ano93f] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 35(1):ii, April 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837589390062B.

Anonymous:1993:EBc

[Ano93g] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 36(1):ifc, July 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837589390096O.

Anonymous:1993:EBd

[Ano93h] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 37(1):ii, October 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837589390074G.

Anonymous:1993:PJc

[Ano93i] Anonymous. Pages 1-126 (July 1993). Journal of Statistical Planning and Inference, 36(1):??, July 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1993:PO

[Ano93j] Anonymous. Pages 1-126 (October 1993). Journal of Statistical Planning and Inference, 37(1):??, October 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1993:PA

[Ano93k] Anonymous. Pages 1-138 (April 1993). Journal of Statistical Planning and Inference, 35(1):??, April 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1993:PJa

[Ano93l] Anonymous. Pages 1-158 (January 1993). Journal of Statistical Planning and Inference, 34(1):??, January 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1993:PN

[Ano93m] Anonymous. Pages 127-270 (November 1993). Journal of Statistical Planning and Inference, 37(2):??, November 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1993:PAS

[Ano93n] Anonymous. Pages 127-432 (August-September 1993). Journal of Statistical Planning and Inference, 36(2–3):??, August/ September 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1993:PMb

[Ano93o] Anonymous. Pages 139-268 (May 1993). Journal of Statistical Planning and Inference, 35(2):??, May 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1993:PF

[Ano93p] Anonymous. Pages 159-315 (February 1993). Journal of Statistical Planning and Inference, 34(2):??, February 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1993:PJb

[Ano93q] Anonymous. Pages 269-400 (June 1993). Journal of Statistical Planning and Inference, 35(3):??, June 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1993:PD

[Ano93r] Anonymous. Pages 271-414 (December 1993). Journal of Statistical Planning and Inference, 37(3):??, December 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1993:PMa

[Ano93s] Anonymous. Pages 317-452 (March 1993). Journal of Statistical Planning and Inference, 34(3):??, March 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1994:AIVa

[Ano94a] Anonymous. Author index volume 38 (1994). Journal of Statistical Planning and Inference, 38(3):411–412, March 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894900213.

Anonymous:1994:AIVb

[Ano94b] Anonymous. Author index volume 39 (1994). Journal of Statistical Planning and Inference, 39(3):425–426, May 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894900965.

Anonymous:1994:AIVc

[Ano94c] Anonymous. Author index volume 40 (1994). Journal of Statistical Planning and Inference, 40(2–3):391–392, July 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894901341.

Anonymous:1994:AIVd

[Ano94d] Anonymous. Author index volume 41 (1994). Journal of Statistical Planning and Inference, 41(3):391–392, October 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894900310.

Anonymous:1994:AIVe

[Ano94e] Anonymous. Author index volume 42 (1994). Journal of Statistical Planning and Inference, 42(3):393–394, December 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589490149X.

Anonymous:1994:EBa

[Ano94f] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 38(1):ii, January 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837589490135X.

Anonymous:1994:EBb

[Ano94g] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 39(1):ii, April 1, 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375894901074.

Anonymous:1994:EBc

[Ano94h] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 40(1):ii, June 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375894901376.

Anonymous:1994:EBd

[Ano94i] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 41(1):ii, August 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375894901503.

Anonymous:1994:EBe

[Ano94j] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 42(1–2):ii, November 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894901856.

Anonymous:1994:PAc

[Ano94k] Anonymous. Pages 1-120 (August 1994). Journal of Statistical Planning and Inference, 41(1):??, August 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1994:PAa

[Ano94l] Anonymous. Pages 1-134 (1 April 1994). Journal of Statistical Planning and Inference, 39(1):??, April 1, 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1994:PJa

[Ano94m] Anonymous. Pages 1-140 (January 1994). Journal of Statistical Planning and Inference, 38(1):??, January 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1994:PJb

[Ano94n] Anonymous. Pages 1-160 (June 1994). Journal of Statistical Planning and Inference, 40(1):??, June 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1994:PN

[Ano94o] Anonymous. Pages 1-290 (November 1994). Journal of Statistical Planning and Inference, 42(1–2):??, November 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1994:PS

[Ano94p] Anonymous. Pages 121-256 (September 1994). Journal of Statistical Planning and Inference, 41(2):??, September 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1994:PAb

[Ano94q] Anonymous. Pages 135-352 (15 April 1994). Journal of Statistical Planning and Inference, 39(2):??, April 15, 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1994:PF

[Ano94r] Anonymous. Pages 141-278 (February 1994). Journal of Statistical Planning and Inference, 38(2):??, February 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1994:PJc

[Ano94s] Anonymous. Pages 161-392 (July 1994). Journal of Statistical Planning and Inference, 40(2–3):??, July 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1994:PO

[Ano94t] Anonymous. Pages 257-392 (October 1994). Journal of Statistical Planning and Inference, 41(3):??, October 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1994:PMa

[Ano94u] Anonymous. Pages 279-412 (March 1994). Journal of Statistical Planning and Inference, 38(3):??, March 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1994:PD

[Ano94v] Anonymous. Pages 291-394 (December 1994). Journal of Statistical Planning and Inference, 42(3):??, December 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1994:PMb

[Ano94w] Anonymous. Pages 353-426 (May 1994). Journal of Statistical Planning and Inference, 39(3):??, May 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1995:AIVa

[Ano95a] Anonymous. Author index volume 43 (1995). Journal of Statistical Planning and Inference, 43(3):403–404, February 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895900330.

Anonymous:1995:AIVb

[Ano95b] Anonymous. Author index volume 44 (1995). Journal of Statistical Planning and Inference, 44(3):399–400, ???? 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895900004.

Anonymous:1995:AIVc

[Ano95c] Anonymous. Author index volume 45 (1995). Journal of Statistical Planning and Inference, 45(3):429–430, June 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895900152.

Anonymous:1995:AIVd

[Ano95d] Anonymous. Author index volume 46 (1995). Journal of Statistical Planning and Inference, 46(3):395–396, August 15, 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895900357.

Anonymous:1995:AIVe

[Ano95e] Anonymous. Author index volume 47 (1995). Journal of Statistical Planning and Inference, 47(3):399–400, October 15, 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895900314.

Anonymous:1995:AIVf

[Ano95f] Anonymous. Author index volume 48 (1995). Journal of Statistical Planning and Inference, 48(3):391–392, December 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895900233.

Anonymous:1995:EBa

[Ano95g] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 43(1–2):ii, January 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895900195.

Anonymous:1995:EBb

[Ano95h] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 44(1):iv, March 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375895900136.

Anonymous:1995:EBc

[Ano95i] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 46(1):iv, July 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375895900071.

Anonymous:1995:EBd

[Ano95j] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 47(1–2):ii, October 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895900012.

Anonymous:1995:EBe

[Ano95k] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 48(1):ii, November 1, 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589590011X.

Anonymous:1995:PNa

[Ano95l] Anonymous. Pages 1-129 (1 November 1995). Journal of Statistical Planning and Inference, 48(1):??, November 1, 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1995:PM

[Ano95m] Anonymous. Pages 1-132 (March 1995). Journal of Statistical Planning and Inference, 44(1):??, March 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1995:PJb

[Ano95n] Anonymous. Pages 1-135 (July 1995). Journal of Statistical Planning and Inference, 46(1):??, July 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1995:PNb

[Ano95o] Anonymous. Pages 131-259 (15 November 1995). Journal of Statistical Planning and Inference, 48(2):??, November 15, 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1995:PAa

[Ano95p] Anonymous. Pages 133-263 (1 April 1995). Journal of Statistical Planning and Inference, 44(2):??, April 1, 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1995:PAb

[Ano95q] Anonymous. Pages 137-263 (1 August 1995). Journal of Statistical Planning and Inference, 46(2):??, August 1, 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1995:PO

[Ano95r] Anonymous. Pages 203-400 (15 October 1995). Journal of Statistical Planning and Inference, 47(3):??, October 15, 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1995:PD

[Ano95s] Anonymous. Pages 261-392 (December 1995). Journal of Statistical Planning and Inference, 48(3):??, December 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1995:PAc

[Ano95t] Anonymous. Pages 265-396 (15 August 1995). Journal of Statistical Planning and Inference, 46(3):??, August 15, 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1995:Pb

[Ano95u] Anonymous. Pages 265-400 (1995). Journal of Statistical Planning and Inference, 44(3):??, ???? 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1995:Pc

[Ano95v] Anonymous. Pages 3-312 (1995). Journal of Statistical Planning and Inference, 45(1–2):??, ???? 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1995:PF

[Ano95w] Anonymous. Pages 311-404 (February 1995). Journal of Statistical Planning and Inference, 43(3):??, February 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1995:PJa

[Ano95x] Anonymous. Pages 313-430 (June 1995). Journal of Statistical Planning and Inference, 45(3):??, June 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1996:AIa

[Ano96a] Anonymous. Author index. Journal of Statistical Planning and Inference, 50(2):i, March 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375896900439.

Anonymous:1996:AIb

[Ano96b] Anonymous. Author index. Journal of Statistical Planning and Inference, 51(1):99, April 1, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375896900610.

Anonymous:1996:AIc

[Ano96c] Anonymous. Author index. Journal of Statistical Planning and Inference, 54(1):139, September 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896900099.

Anonymous:1996:AId

[Ano96d] Anonymous. Author index. Journal of Statistical Planning and Inference, 54(3):359, September 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837589690021X.

Anonymous:1996:AIS

[Ano96e] Anonymous. Author index Shanghai conference issue, parts 1 and 2. Journal of Statistical Planning and Inference, 51(3):403– 404, May 1, 1996. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837589690049X.

Anonymous:1996:AIVa

[Ano96f] Anonymous. Author index volume 49 (1996). Journal of Statistical Planning and Inference, 49(3):415–416, February 1, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896900531.

Anonymous:1996:AIVb

[Ano96g] Anonymous. Author index volume 50 (1996). Journal of Statistical Planning and Inference, 50(3):413–414, March 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896900191.

Anonymous:1996:AIVc

[Ano96h] Anonymous. Author index volume 51 (1996). Journal of Statistical Planning and Inference, 51(3):405–406, May 1, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896900518.

Anonymous:1996:AIVd

[Ano96i] Anonymous. Author index volume 52 (1996). Journal of Statistical Planning and Inference, 52(3):395–396, July 1, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896900336.

Anonymous:1996:AIVe

[Ano96j] Anonymous. Author index volume 53 (1996). Journal of Statistical Planning and Inference, 53(3):437–438, August 30, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896900397.

Anonymous:1996:AIVf

[Ano96k] Anonymous. Author index volume 54 (1996). Journal of Statistical Planning and Inference, 54(3):361, September 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896900233.

Anonymous:1996:AIVg

[Ano96l] Anonymous. Author index volume 55 (1996). Journal of Statistical Planning and Inference, 55(3):399–400, November 8, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375896900476.

Anonymous:1996:AIVh

[Ano96m] Anonymous. Author index volume 56 (1996). Journal of Statistical Planning and Inference, 56(2):325–326, December 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896900452.

Anonymous:1996:EBa

[Ano96n] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 49(1):ii, January 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375896900270.

Anonymous:1996:EBb

[Ano96o] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 50(1):ii, February 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375896900257.

Anonymous:1996:EBc

[Ano96p] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 51(1):ii, April 1, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375896900592.

Anonymous:1996:EBd

[Ano96q] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 52(1):ii, June 1, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375896900014.

Anonymous:1996:EBe

[Ano96r] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 53(1):ii, August 1, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896900373.

Anonymous:1996:EBf

[Ano96s] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 54(1):ii, September 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896900051.

Anonymous:1996:EBg

[Ano96t] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 55(1):ii, October 15, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896900415.

Anonymous:1996:EBh

[Ano96u] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 56(1):ii, December 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375896900130.

Anonymous:1996:PJa

[Ano96v] Anonymous. Pages 1-129 (1 June 1996). Journal of Statistical Planning and Inference, 52(1):??, June 1, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1996:POa

[Ano96w] Anonymous. Pages 1-129 (15 October 1996). Journal of Statistical Planning and Inference, 55(1):??, October 15, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1996:PAb

[Ano96x] Anonymous. Pages 1-132 (1 August 1996). Journal of Statistical Planning and Inference, 53(1):??, August 1, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1996:PAa

[Ano96y] Anonymous. Pages 1-99 (1 April 1996). Journal of Statistical Planning and Inference, 51(1):??, April 1, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1996:PJb

[Ano96z] Anonymous. Pages 131-262 (15 June 1996). Journal of Statistical Planning and Inference, 52(2):??, June 15, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1996:POb

[Ano96-27] Anonymous. Pages 131-264 (30 October 1996). Journal of Statistical Planning and Inference, 55(2):??, October 30, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1996:PAc

[Ano96-28] Anonymous. Pages 133-283 (15 August 1996). Journal of Statistical Planning and Inference, 53(2):??, August 15, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1996:PM

[Ano96-29] Anonymous. Pages 261-406 (1 May 1996). Journal of Statistical Planning and Inference, 51(3):??, May 1, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1996:PJc

[Ano96-30] Anonymous. Pages 263-396 (1 July 1996). Journal of Statistical Planning and Inference, 52(3):??, July 1, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1996:PN

[Ano96-31] Anonymous. Pages 265-400 (8 November 1996). Journal of Statistical Planning and Inference, 55(3):??, November 8, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1996:PAd

[Ano96-32] Anonymous. Pages 285-438 (30 August 1996). Journal of Statistical Planning and Inference, 53(3):??, August 30, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1996:PF

[Ano96-33] Anonymous. Pages 305-416 (1 February 1996). Journal of Statistical Planning and Inference, 49(3):??, February 1, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1997:AIVa

[Ano97a] Anonymous. Author index volume 57 (1997). Journal of Statistical Planning and Inference, 57(2):335–336, February 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897848380.

Anonymous:1997:AIVb

[Ano97b] Anonymous. Author index volume 58 (1997). Journal of Statistical Planning and Inference, 58(2):407–408, March 15, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897900173.

Anonymous:1997:AIVc

[Ano97c] Anonymous. Author index volume 59 (1997). Journal of Statistical Planning and Inference, 59(2):385–386, April 15, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897836415.

Anonymous:1997:AIVd

[Ano97d] Anonymous. Author index volume 60 (1997). Journal of Statistical Planning and Inference, 60(2):389–390, May 15, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897897133.

Anonymous:1997:AIVe

[Ano97e] Anonymous. Author index volume 61 (1997). Journal of Statistical Planning and Inference, 61(2):385, June 16, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897894372.

Anonymous:1997:AIVf

[Ano97f] Anonymous. Author index volume 62 (1997). Journal of Statistical Planning and Inference, 62(2):365–366, August 15, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897842073.

Anonymous:1997:AIVg

[Ano97g] Anonymous. Author index volume 63 (1997). Journal of Statistical Planning and Inference, 63(2):387–388, October 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897844527.

Anonymous:1997:AIVh

[Ano97h] Anonymous. Author index volume 64 (1997). Journal of Statistical Planning and Inference, 64(2):383–384, November 15, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375897802285.

Anonymous:1997:AIVi

[Ano97i] Anonymous. Author index volume 65 (1997). Journal of Statistical Planning and Inference, 65(2):387–388, December 15, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375897817508.

Anonymous:1997:BRP

[Ano97j] Anonymous. Book review: A probabilistic analysis of the Sacco and Vanzetti evidence: Joseph B. Kadane and David A. Schum, (Wiley, New York, 1996). Journal of Statistical Planning and Inference, 64(1):171–191, October 30, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897001055.

Anonymous:1997:EBa

[Ano97k] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 57(1):ii, January 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375897900252.

Anonymous:1997:EBb

[Ano97l] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 58(1):ii, March 1, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897900513.

Anonymous:1997:EBc

[Ano97m] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 59(1):ii, March 30, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837589790029X.

Anonymous:1997:EBd

[Ano97n] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 60(1):ii, April 30, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897900094.

Anonymous:1997:EBe

[Ano97o] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 61(1):ii, May 30, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375897900197.

Anonymous:1997:EBf

[Ano97p] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 62(1):ii, July 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375897900318.

Anonymous:1997:EBg

[Ano97q] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 63(1):ii, September 30, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897900239.

Anonymous:1997:EBh

[Ano97r] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 64(1):ii, October 30, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897900215.

Anonymous:1997:EBi

[Ano97s] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 65(1):ii, December 1, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897900653.

Anonymous:1997:LP

[Ano97t] Anonymous. List of participants. Journal of Statistical Planning and Inference, 57(1):5–6, January 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896000304.

Anonymous:1997:PS

[Ano97u] Anonymous. Pages 1-111 (30 September 1997). Journal of Statistical Planning and Inference, 63(1):??, September 30, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1997:PDa

[Ano97v] Anonymous. Pages 1-189 (1 December 1997). Journal of Statistical Planning and Inference, 65(1):??, December 1, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1997:PMc

[Ano97w] Anonymous. Pages 1-189 (30 March 1997). Journal of Statistical Planning and Inference, 59(1):??, March 30, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1997:PO

[Ano97x] Anonymous. Pages 1-191 (30 October 1997). Journal of Statistical Planning and Inference, 64(1):??, October 30, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1997:PMe

[Ano97y] Anonymous. Pages 1-192 (30 May 1997). Journal of Statistical Planning and Inference, 61(1):??, May 30, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1997:PAb

[Ano97z] Anonymous. Pages 1-197 (30 April 1997). Journal of Statistical Planning and Inference, 60(1):??, April 30, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1997:PMa

[Ano97-27] Anonymous. Pages 1-216 (1 March 1997). Journal of Statistical Planning and Inference, 58(1):??, March 1, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1997:PAc

[Ano97-28] Anonymous. Pages 135-366 (15 August 1997). Journal of Statistical Planning and Inference, 62(2):??, August 15, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1997:PAa

[Ano97-29] Anonymous. Pages 191-386 (15 April 1997). Journal of Statistical Planning and Inference, 59(2):??, April 15, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1997:PDb

[Ano97-30] Anonymous. Pages 191-388 (15 December 1997). Journal of Statistical Planning and Inference, 65(2):??, December 15, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1997:PN

[Ano97-31] Anonymous. Pages 193-384 (15 November 1997). Journal of Statistical Planning and Inference, 64(2):??, November 15, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1997:PJ

[Ano97-32] Anonymous. Pages 193-385 (16 June 1997). Journal of Statistical Planning and Inference, 61(2):??, June 16, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1997:PMd

[Ano97-33] Anonymous. Pages 199-390 (15 May 1997). Journal of Statistical Planning and Inference, 60(2):??, May 15, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1997:PMb

[Ano97-34] Anonymous. Pages 217-408 (15 March 1997). Journal of Statistical Planning and Inference, 58(2):??, March 15, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1997:RSD

[Ano97-35] Anonymous. Robust statistics and data analysis. Journal of Statistical Planning and Inference, 57(1):1, January 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897900276.

Anonymous:1998:AIVa

[Ano98a] Anonymous. Author index volume 66 (1998). Journal of Statistical Planning and Inference, 66(2):385–386, March 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897847805.

Anonymous:1998:AIVb

[Ano98b] Anonymous. Author index volume 67 (1998). Journal of Statistical Planning and Inference, 67(2):343–344, April 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897865792.

Anonymous:1998:AIVc

[Ano98c] Anonymous. Author index volume 68 (1998). Journal of Statistical Planning and Inference, 68(2):417–418, May 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898001050.

Anonymous:1998:AIVd

[Ano98d] Anonymous. Author index volume 69 (1998). Journal of Statistical Planning and Inference, 69(2):385–386, June 15, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898001116.

Anonymous:1998:BRa

[Ano98e] Anonymous. Book review. Journal of Statistical Planning and Inference, 71(1–2):377–378, August 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898001189.

Anonymous:1998:BRb

[Ano98f] Anonymous. Book review. Journal of Statistical Planning and Inference, 74(2):371–376, November 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837589800127X.

Anonymous:1998:BRc

[Ano98g] Anonymous. Book review. Journal of Statistical Planning and Inference, 75(1):175–178, November 15, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898001281.

Anonymous:1998:BRd

[Ano98h] Anonymous. Book review. Journal of Statistical Planning and Inference, 75(1):179–180, November 15, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898001293.

Anonymous:1998:C

[Ano98i] Anonymous. Contents. Journal of Statistical Planning and Inference, 68(1):1–2, May 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375897001523.

Anonymous:1998:Ea

[Ano98j] Anonymous. Editorial. Journal of Statistical Planning and Inference, 72(1–2):vii–ix, September 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898000184.

Anonymous:1998:EBa

[Ano98k] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 66(1):ii, January 5, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898900039.

Anonymous:1998:EBb

[Ano98l] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 67(1):ii, March 16, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898900076.

Anonymous:1998:EBc

[Ano98m] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 68(1):ii, May 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375898900003.

Anonymous:1998:EBd

[Ano98n] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 69(1):ii, June 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375898900131.

Anonymous:1998:EBe

[Ano98o] Anonymous. Editorial Board. Journal of Statistical Planning and Inference, 70(1):ii, July 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375898900015.

Anonymous:1998:Ec

[Ano98p] Anonymous. Erratum. Journal of Statistical Planning and Inference, 75(1):173, November 15, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898001268.

Anonymous:1998:EOS

[Ano98q] Anonymous. Erratum to “Optimal smoothing in adaptive location estimation” [J. Statist. Plann. Inference 58 (1997) 333–348]. Journal of Statistical Planning and Inference, 67(1):165, March 16, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375897001146. See [MP97].

Anonymous:1998:Ia

[Ano98r] Anonymous. Index. Journal of Statistical Planning and Inference, 71(1–2):379–380, August 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375898001633.

Anonymous:1998:Ib

[Ano98s] Anonymous. Index. Journal of Statistical Planning and Inference, 72(1–2):407, September 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375898001657.

Anonymous:1998:Ic

[Ano98t] Anonymous. Index. Journal of Statistical Planning and Inference, 73(1–2):417–419, September 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898001669.

Anonymous:1998:Id

[Ano98u] Anonymous. Index. Journal of Statistical Planning and Inference, 74(2):377–378, November 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375898002006.

Anonymous:1998:OCa

[Ano98v] Anonymous. Other contents. Journal of Statistical Planning and Inference, 71(1–2):v, August 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898001645.

Anonymous:1998:OCb

[Ano98w] Anonymous. Other contents. Journal of Statistical Planning and Inference, 72(1–2):iii–V, September 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898001670.

Anonymous:1998:PMb

[Ano98x] Anonymous. Pages 1-165 (16 March 1998). Journal of Statistical Planning and Inference, 67(1):??, March 16, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1998:PNb

[Ano98y] Anonymous. Pages 1-180 (15 November 1998). Journal of Statistical Planning and Inference, 75(1):??, November 15, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1998:PJd

[Ano98z] Anonymous. Pages 1-190 (1 July 1998). Journal of Statistical Planning and Inference, 70(1):??, July 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1998:PJb

[Ano98-27] Anonymous. Pages 1-191 (1 June 1998). Journal of Statistical Planning and Inference, 69(1):??, June 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1998:PJa

[Ano98-28] Anonymous. Pages 1-191 (5 January 1998). Journal of Statistical Planning and Inference, 66(1):??, January 5, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1998:PO

[Ano98-29] Anonymous. Pages 1-204 (1 October 1998). Journal of Statistical Planning and Inference, 74(1):??, October 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1998:PAb

[Ano98-30] Anonymous. Pages 1-380 (1 August 1998). Journal of Statistical Planning and Inference, 71(1–2):??, August 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1998:PSa

[Ano98-31] Anonymous. Pages 1-408 (1 September 1998). Journal of Statistical Planning and Inference, 72(1–2):??, September 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1998:PSb

[Ano98-32] Anonymous. Pages 1-420 (1 September 1998). Journal of Statistical Planning and Inference, 73(1–2):??, September 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1998:PAa

[Ano98-33] Anonymous. Pages 167-344 (1 April 1998). Journal of Statistical Planning and Inference, 67(2):??, April 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1998:PJe

[Ano98-34] Anonymous. Pages 191-382 (15 July 1998). Journal of Statistical Planning and Inference, 70(2):??, July 15, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1998:PMa

[Ano98-35] Anonymous. Pages 193-386 (1 March 1998). Journal of Statistical Planning and Inference, 66(2):??, March 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1998:PJc

[Ano98-36] Anonymous. Pages 193-386 (15 June 1998). Journal of Statistical Planning and Inference, 69(2):??, June 15, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1998:PNa

[Ano98-37] Anonymous. Pages 205-378 (1 November 1998). Journal of Statistical Planning and Inference, 74(2):??, November 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1999:BRa

[Ano99a] Anonymous. Book review. Journal of Statistical Planning and Inference, 76(1–2):295–296, February 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837589800130X.

Anonymous:1999:BRb

[Ano99b] Anonymous. Book review. Journal of Statistical Planning and Inference, 77(1):167–168, February 15, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898001839.

Anonymous:1999:BRc

[Ano99c] Anonymous. Book review. Journal of Statistical Planning and Inference, 77(2):353–354, March 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898001827.

Anonymous:1999:BRd

[Ano99d] Anonymous. Book review. Journal of Statistical Planning and Inference, 79(1):175–177, June 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898002626.

Anonymous:1999:BRe

[Ano99e] Anonymous. Book review. Journal of Statistical Planning and Inference, 81(2):393, November 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375899000567.

Anonymous:1999:Ea

[Ano99f] Anonymous. Editorial. Journal of Statistical Planning and Inference, 75(2):vii, January 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375898001414.

Anonymous:1999:Eb

[Ano99g] Anonymous. Editorial. Journal of Statistical Planning and Inference, 78(1–2):v, May 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375898002018.

Anonymous:1999:Ec

[Ano99h] Anonymous. Editorial. Journal of Statistical Planning and Inference, 80(1–2):v–vi, August 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898002389.

Anonymous:1999:Ed

[Ano99i] Anonymous. Editorial. Journal of Statistical Planning and Inference, 82(1–2):1–4, December 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375899000488.

Anonymous:1999:Ia

[Ano99j] Anonymous. Index. Journal of Statistical Planning and Inference, 76(1–2):297–298, February 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837589800264X.

Anonymous:1999:Ib

[Ano99k] Anonymous. Index. Journal of Statistical Planning and Inference, 77(2):355–356, March 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375899000154.

Anonymous:1999:Ic

[Ano99l] Anonymous. Index. Journal of Statistical Planning and Inference, 78(1–2):401–402, May 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S037837589900083X.

Anonymous:1999:Id

[Ano99m] Anonymous. Index. Journal of Statistical Planning and Inference, 81(2):395–396, November 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375899001561.

Anonymous:1999:Ie

[Ano99n] Anonymous. Index. Journal of Statistical Planning and Inference, 82(1–2):263–264, December 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375899001652.

Anonymous:1999:PFb

[Ano99o] Anonymous. Pages 1-168 (15 February 1999). Journal of Statistical Planning and Inference, 77(1):??, February 15, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1999:PJb

[Ano99p] Anonymous. Pages 1-178 (June 1999). Journal of Statistical Planning and Inference, 79(1):??, June 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1999:PO

[Ano99q] Anonymous. Pages 1-200 (1 October 1999). Journal of Statistical Planning and Inference, 81(1):??, October 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1999:PD

[Ano99r] Anonymous. Pages 1-264 (1 December 1999). Journal of Statistical Planning and Inference, 82(1–2):??, December 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1999:PA

[Ano99s] Anonymous. Pages 1-292 (1 August 1999). Journal of Statistical Planning and Inference, 80(1–2):??, August 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1999:PFa

[Ano99t] Anonymous. Pages 1-298 (1 February 1999). Journal of Statistical Planning and Inference, 76(1–2):??, February 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1999:PMb

[Ano99u] Anonymous. Pages 1-402 (May 1999). Journal of Statistical Planning and Inference, 78(1–2):??, May 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1999:PMa

[Ano99v] Anonymous. Pages 169-356 (1 March 1999). Journal of Statistical Planning and Inference, 77(2):??, March 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1999:PJc

[Ano99w] Anonymous. Pages 179-368 (1 July 1999). Journal of Statistical Planning and Inference, 79(2):??, July 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1999:PJa

[Ano99x] Anonymous. Pages 181-464 (1 January 1999). Journal of Statistical Planning and Inference, 75(2):??, January 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Anonymous:1999:PN

[Ano99y] Anonymous. Pages 201-396 (1 November 1999). Journal of Statistical Planning and Inference, 81(2):??, November 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

Alvo:1997:GTH

[AP97] Mayer Alvo and Jianhong Pan. A general theory of hypothesis testing based on rankings. Journal of Statistical Planning and Inference, 61(2):219–248, June 16, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897894360.

Aban:1998:APG

[AP98] Inmaculada B. Aban and Edsel A. Peña. Ancillarity properties of generalized residuals with applications in failure time models. Journal of Statistical Planning and Inference, 74(1):31–49, October 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375898000949.

Aban:1999:PTS

[AP99] Inmaculada B. Aban and Edsel A. Peña. Properties of test statistics applied to residuals in failure time models. Journal of Statistical Planning and Inference, 75(2):181–209, January 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898001426.

Adke:1997:OPS

[AR97] S. R. Adke and T. V. Ramanathan. On optimal prediction for stochastic processes. Journal of Statistical Planning and Inference, 63(1):1–7, September 30, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896002017.

Arcones:1999:RCE

[Arc99] Miguel A. Arcones. Rates of convergence of L_{p}-estimators for a density with an infinity cusp. Journal of Statistical Planning and Inference, 81(1):33–50, October 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375899000178.

Arnold:1990:FFM

[Arn90] Barry C. Arnold. A flexible family of multivariate Pareto distributions. Journal of Statistical Planning and Inference, 24(2):249– 258, February 1990. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589090045V.

Arnold:1997:CIC

[Arn97] Barry C. Arnold. Characterizations involving conditional specification. Journal of Statistical Planning and Inference, 63(2):117– 131, October 1997. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375897000050.

Shiraishi:1990:ECR

[aS90] Taka aki Shiraishi. R-estimators and confidence regions in one-way MANOVA. Journal of Statistical Planning and Inference, 24(2):203–214, February 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837589090042S.

Abramovich:1996:IIN

[AS96a] Felix Abramovich and David M. Steinberg. Improved inference in nonparametric regression using L_{k}-smoothing splines. Journal of Statistical Planning and Inference, 49(3):327–341, February 1, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375895000216.

Alharbi:1996:GCT

[AS96b] Abdulghani A. Alharbi and D. N. Shanbhag. General characterization theorems based on versions of the Chernoff inequality and the Cox representation. Journal of Statistical Planning and Inference, 55(2):139–150, October 30, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895001913.

Arasu:1997:P

[AS97] K. T. Arasu and N. M. Singhi. Preface. Journal of Statistical Planning and Inference, 62(1):1, July 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896001607.

Asadi:1999:NCB

[Asa99] M. Asadi. A note on characterizations based on truncated expectations. Journal of Statistical Planning and Inference, 81(2):201–207, November 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375898002584.

Agarwal:1997:NAU

[ASK97] S. K. Agarwal, U. K. Sharma, and Sharmishtha Kashyap. A new approach to use multivariate auxiliary information in sample surveys. Journal of Statistical Planning and Inference, 60(2):261– 267, May 15, 1997. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837589789711X.

Assmus:1996:BCR

[Ass96] E. F. Assmus. On Berman’s characterization of the Reed–Muller codes. Journal of Statistical Planning and Inference, 56(1):17– 21, December 1996. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375896000043.

Abel:1996:TMO

[AZZ96] R. J. R. Abel, Xiafu Zhang, and Hangfu Zhang. Three mutually orthogonal idempotent Latin squares of orders 22 and 26. Journal of Statistical Planning and Inference, 51(2):101–106, April 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375895000739.

Balakrishnan:1992:RME

[BA92a] N. Balakrishnan and R. S. Ambagaspitiya. A robust method of estimation based on the MML estimators for a single linear regression model. Journal of Statistical Planning and Inference, 30(2):267–279, February 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375892900879.

Bilotti-Aliaga:1992:USE

[BA92b] Martha Bilotti-Aliaga. Use of a side effect as a covariate in a problem of sequential analysis. Journal of Statistical Planning and Inference, 32(3):311–323, September 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290014J.

Bhandari:1994:AMP

[BA94] S. K. Bhandari and M. Masoom Ali. An asymptotically minimax procedure for selecting the t-best multinomial cells. Journal of Statistical Planning and Inference, 38(1):65–74, January 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375894901368.

BuHamra:1995:SIM

[BA95] Sana BuHamra and Donald A. Anderson. The spectrum of the information matrix for s^{n} parallel flats fractions when s is prime. Journal of Statistical Planning and Inference, 47(3):333–345, October 15, 1995. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589400139M.

Bedi:1999:MMS

[BA99] P. K. Bedi and S. K. Agarwal. Modified Midzuno scheme of sampling. Journal of Statistical Planning and Inference, 76(1–2):203– 214, February 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375898001153.

Babu:1995:BNC

[Bab95] Gutti Jogesh Babu. Bootstrap for nonstandard cases. Journal of Statistical Planning and Inference, 43(1–2):197–203, January 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589400019R.

Basu:1999:BAM

[BABM99] Sanjib Basu, Asit, P. Basu, and Chiranjit Mukhopadhyay. Bayesian analysis for masked system failure data using nonidentical Weibull models. Journal of Statistical Planning and Inference, 78(1–2):255–275, May 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375898002183.

Bagchi:1996:TIS

[Bag96] Sunanda Bagchi. Two infinite series of E-optimal nested rowcolumn designs. Journal of Statistical Planning and Inference, 52(3):353–357, July 1, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375895001239.

Bailey:1990:ESL

[Bai90] R. A. Bailey. An efficient semi-latin square for twelve treatments in blocks of size two. Journal of Statistical Planning and Inference, 26(3):263–266, November 1990. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589090130M.

Bai:1998:EMR

[Bai98] Jushan Bai. Estimation of multiple-regime regressions with least absolutes deviation. Journal of Statistical Planning and Inference, 74(1):103–134, October 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898000822.

Balakrishnan:1990:AML

[Bal90] N. Balakrishnan. Approximate maximum likelihood estimation for a generalized logistic distribution. Journal of Statistical Planning and Inference, 26(2):221–236, October 1990. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589090127G.

Bannai:1996:MIP

[Ban96] Etsuko Bannai. Modular invariance property and spin models attached to cyclic group association schemes. Journal of Statistical Planning and Inference, 51(2):107–124, April 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895000747.

Barakat:1990:LTL

[Bar90a] H. M. Barakat. Limit theorems for lower-upper extreme values from two-dimensional distribution function. Journal of Statistical Planning and Inference, 24(1):69–79, January 1990. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589090017O.

Barnard:1990:C

[Bar90b] George A. Barnard. Comment. Journal of Statistical Planning and Inference, 25(3):437–439, July 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

037837589090086A. See reply [Goo90a].

Barry:1992:NMR

[Bar92] Ronald Paul Barry. Near-minimal resolution IV designs with three factors. Journal of Statistical Planning and Inference, 33(3):413–425, December 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837589290009H.

Barakat:1997:APB

[Bar97a] H. M. Barakat. Asymptotic properties of bivariate random extremes. Journal of Statistical Planning and Inference, 61(2):203– 217, June 16, 1997. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375896001577.

Barbe:1997:BRS

[Bar97b] Philippe Barbe. Bootstrapping the renewal spacings processes. Journal of Statistical Planning and Inference, 60(2):199–214, May 15, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375896001346.

Barlotti:1998:GTH

[Bar98] A. Barlotti. Giuseppe Tallini and his work in combinatorial geometries. Journal of Statistical Planning and Inference, 72(1–2):1–6, September 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375898000196.

Basu:1991:P

[Bas91] A. P. Basu. Preface. Journal of Statistical Planning and Inference, 29(1–2):3, September/October 1991. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290116A.

Basawa:1995:BRS

[Bas95a] Ishwar V. Basawa. Book review: Statistical inference for branching processes: By Peter Guttorp, Wiley, New York, 1991. Journal of Statistical Planning and Inference, 47(3):393–394, October 15, 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375895900276.

Basu:1995:RPP

[Bas95b] Sanjib Basu. Ranges of posterior probability over a distribution band. Journal of Statistical Planning and Inference, 44(2):149– 166, April 1, 1995. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589400048Z.

Basawa:1996:P

[Bas96] Ishwar Basawa. Preface. Journal of Statistical Planning and Inference, 50(3):309, March 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375896900178.

Bandyopadhyay:1991:CTB

[BB91] Dipankar Bandyopadhyay and Asit P. Basu. A class of tests for bivariate exponentiality against bivariate increasing failure rate alternatives. Journal of Statistical Planning and Inference, 29(3):337–349, November 1991. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375891900083.

Bacro:1995:WLB

[BB95] J. N. Bacro and M. Brito. Weak limiting behaviour of a simple tail Pareto-index estimator. Journal of Statistical Planning and Inference, 45(1–2):7–19, ???? 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894000603.

Bose:1997:SEM

[BB97] Arup Bose and Benzion Boukai. Sequential estimation of the mean of NEF-PVF distributions. Journal of Statistical Planning and Inference, 63(1):55–70, September 30, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896001978.

Bacro:1998:TBP

[BB98] J. N. Bacro and M. Brito. A tail bootstrap procedure for estimating the tail Pareto-index. Journal of Statistical Planning and Inference, 71(1–2):245–260, August 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898000111.

Bandyopadhyay:1999:SNT

[BB99a] Uttam Bandyopadhyay and Atanu Biswas. Sequential nonparametric tests based on randomized play-the-winner rule for restricted bivariate alternatives. Journal of Statistical Planning and Inference, 79(1):45–57, June 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898000792.

Bhat:1999:BAF

[BB99b] B. R. Bhat and M. N. Badade. Bayesian analysis for finite Markov chains. Journal of Statistical Planning and Inference, 81(2):335–348, November 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375898001049.

Balasubramanian:1996:IJD

[BBB96] K. Balasubramanian, M. I. Beg, and R. B. Bapat. An identity for the joint distribution of order statistics and its applications. Journal of Statistical Planning and Inference, 55(1):13–21, October 15, 1996. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375895001387.

Beirlant:1995:MRL

[BBTV95] Jan Beirlant, Michel Broniatowski, Jozef L. Teugels, and Petra Vynckier. The mean residual life function at great age: Applications to tail estimation. Journal of Statistical Planning and Inference, 45(1–2):21–48, ???? 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894000611.

Bagchi:1993:SOD

[BC93] Sunanda Bagchi and Ching-Shui Cheng. Some optimal designs of block size two. Journal of Statistical Planning and Inference, 37(2):245–253, November 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390093L.

Benchekroun:1999:CUI

[BC99] K. Benchekroun and I. M. Chakravarti. Correlation uniform incomplete block designs. Journal of Statistical Planning and Inference, 76(1–2):263–272, February 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898001396.

Block:1992:SNP

[BCFS92] H. W. Block, D. Chhetry, Z. Fang, and A. R. Sampson. Some new partial orderings on S_{n} and S_{n} × S_{n}. Journal of Statistical Planning and Inference, 30(3):351–357, March 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290161K.

Block:1997:MPU

[BCFS97] H. W. Block, D. Chhetry, Z. Fang, and Allan R. Sampson. Metrics on permutations useful for positive dependence. Journal of Statistical Planning and Inference, 62(2):219–234, August 15, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375896001905.

Bennett:1998:QPB

[BCM98] Frank E. Bennett, Charles J. Colbourn, and Ronald C. Mullin. Quintessential pairwise balanced designs. Journal of Statistical Planning and Inference, 72(1–2):15–66, September 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898000214.

Beirlant:1992:LRP

[BCvdM92] Jan Beirlant, An Carbonez, and Edward van der Meulen. Long run proportional hazards models of random censorship. Journal of Statistical Planning and Inference, 32(1):25–44, July 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290150Q.

Birkes:1991:OCD

[BD91] David Birkes and Yadolah Dodge. Optimal a × b connected designs with a + b observations. Journal of Statistical Planning and Inference, 28(1):49–59, May 1991. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589190058M.

Bhat:1994:RSP

[BD94a] B. R. Bhat and Indira A. Devi. Randomised sequential probability ratio tests for stochastic processes. Journal of Statistical Planning and Inference, 39(2):273–289, April 15, 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894902097.

Bose:1994:SCL

[BD94b] Arup Bose and Ratan Dasgupta. Speed of convergence of the least-squares estimator in autoregressive models. Journal of Statistical Planning and Inference, 38(3):371–380, March 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894900175.

Brunner:1994:RSU

[BD94c] Edgar Brunner and Manfred Denker. Rank statistics under dependent observations and applications to factorial designs. Journal of Statistical Planning and Inference, 42(3):353–378, December 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375894000026.

Bhattacharya:1995:GDA

[BD95a] Bhaskar Bhattacharya and Richard Dykstra. A general duality approach to I-projections. Journal of Statistical Planning and Inference, 47(3):203–216, October 15, 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589400142I.

Blumenthal:1995:ESS

[BD95b] Saul Blumenthal and Ram C. Dahiya. Estimation of sample size with grouped data. Journal of Statistical Planning and Inference, 44(1):95–115, March 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589400041S.

Bai:1996:EFB

[BD96a] S. Kalpana Bai and T. M. Durairajan. Estimating functions for branching processes. Journal of Statistical Planning and Inference, 53(1):21–32, August 1, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895001492.

Balasubramanian:1996:ODM

[BD96b] K. Balasubramanian and Aloke Dey. D-optimal designs with minimal and nearly minimal number of units. Journal of Statistical Planning and Inference, 52(2):255–262, June 15, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895001174.

Berardi:1996:BCC

[BD96c] Luigia Berardi and Bal Kishan Dass. A bound for codes correcting random and low-density closed-loop burst errors. Journal of Statistical Planning and Inference, 51(2):125–128, April 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895000755.

Bai:1997:OEE

[BD97] S. Kalpana Bai and T. M. Durairajan. Optimal equivariant estimator with respect to convex loss function. Journal of Statistical Planning and Inference, 64(2):283–295, November 15, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897000360.

Baksalary:1990:CTW

[BDG90a] Jerzy K. Baksalary, Anita Dobek, and Stanisław Gnot. Characterizations of two-way layouts from the point of view of variance component estimation in the corresponding mixed linear models. Journal of Statistical Planning and Inference, 26(1):35– 45, September 1990. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589090093A.

Brillinger:1990:SSA

[BDG90b] David R. Brillinger, Kenneth H. Downing, and Robert M. Glaeser. Some statistical aspects of low-dose electron imaging of crystals. Journal of Statistical Planning and Inference, 25(3): 235–259, July 1990. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375890900756.

Basu:1991:BAL

[BE91] A. P. Basu and Nader Ebrahimi. Bayesian approach to life testing and reliability estimation using asymmetric loss function. Journal of Statistical Planning and Inference, 29(1–2):21–31, September/October 1991. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589290118C.

Beirlant:1996:MTT

[BE96] Jan Beirlant and John H. J. Einmahl. Maximal type test statistics based on conditional processes. Journal of Statistical Planning and Inference, 53(1):1–19, August 1, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375896001358.

Beder:1998:CAH

[Bed98] Jay H. Beder. Conjectures about Hadamard matrices. Journal of Statistical Planning and Inference, 72(1–2):7–14, September 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375898000202. See erratum [Bed00].

Beder:2000:ECA

[Bed00] Jay H. Beder. Erratum to “Conjectures About Hadamard Matrices” [J Statist. Plann. Inference 72 (1998) 7–14]. Journal of Statistical Planning and Inference, 84(1–2):343, March 2000. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375899001482. See [Bed98].

Baker:1996:MCP

[BEHW96] R. D. Baker, G. L. Ebert, J. Hemmeter, and A. Woldar. Maximal cliques in the Paley graph of square order. Journal of Statistical Planning and Inference, 56(1):33–38, December 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896000067.

Belzil:1995:EML

[Bel95] Christian Belzil. Estimating models of labor market transitions based on alternative specifications of job search theory. Journal of Statistical Planning and Inference, 43(1–2):243–255, January 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589400024P.

Benda:1996:PTE

[Ben96] Norbert Benda. Pre-test estimation and design in the linear model. Journal of Statistical Planning and Inference, 52(2):225– 240, June 15, 1996. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375896001152.

Benz:1998:AFE

[Ben98] Walter Benz. Automorphisms of finite Einstein geometries. Journal of Statistical Planning and Inference, 72(1–2):67–77, September 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375898000226.

Boland:1991:RIA

[BENP91] Philip J. Boland, Emad El-Neweihi, and Frank Proschan. Redundancy importance and allocation of spares in coherent systems. Journal of Statistical Planning and Inference, 29(1–2):55–65, September/October 1991. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375892901218.

Berger:1990:RBA

[Ber90] James O. Berger. Robust Bayesian analysis: sensitivity to the prior. Journal of Statistical Planning and Inference, 25(3):303– 328, July 1990. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589090079A.

Beran:1995:PRC

[Ber95a] Rudolf Beran. Prediction in random coefficient regression. Journal of Statistical Planning and Inference, 43(1–2):205–213, January 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589400020V.

Berred:1995:RVE

[Ber95b] Mohamed Berred. K-record values and the extreme-value index. Journal of Statistical Planning and Inference, 45(1–2):49–63, ???? 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589400062X.

Berridge:1995:MOR

[Ber95c] Damon M. Berridge. Modelling ordinal recurrent events. Journal of Statistical Planning and Inference, 47(1–2):71–78, October 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589400122C.

Bernardo:1997:R

[Ber97] JoséM. Bernardo. Rejoinder. Journal of Statistical Planning and Inference, 65(1):185–189, December 1, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897900756.

Berger:1998:AEC

[Ber98a] Vance W. Berger. Admissibility of exact conditional tests of stochastic order. Journal of Statistical Planning and Inference, 66(1):39–50, January 5, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375897000785.

Berger:1998:RCA

[Ber98b] Yves G. Berger. Rate of convergence for asymptotic variance of the Horvitz–Thompson estimator. Journal of Statistical Planning and Inference, 74(1):149–168, October 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898001074.

Berger:1998:RCN

[Ber98c] Yves G. Berger. Rate of convergence to normal distribution for the Horvitz–Thompson estimator. Journal of Statistical Planning and Inference, 67(2):209–226, April 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897001079.

Barakat:1990:LDE

[BES90] H. M. Barakat and M. A. El-Shandidy. On the limit distribution of the extremes of a random number of independent random variables. Journal of Statistical Planning and Inference, 26(3):353– 361, November 1990. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589090137J.

Babu:1996:SPP

[BF96] Gutti Jogesh Babu and Eric D. Feigelson. Spatial point processes in astronomy. Journal of Statistical Planning and Inference, 50(3):311–326, March 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375895000607.

Bayarri:1999:BCC

[BF99] M. J. Bayarri and B. Font. A Bayesian comparison of cluster, strata, and random samples. Journal of Statistical Planning and Inference, 78(1–2):277–294, May 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898002195.

Boente:1997:RPB

[BFM97] Graciela Boente, Ricardo Fraiman, and Jean Meloche. Robust plug-in bandwidth estimators in nonparametric regression. Journal of Statistical Planning and Inference, 57(1):109–142, January 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375896000390.

Boularan:1994:GCT

[BFV94] Joël Boularan, Louis Ferré, and Philippe Vieu. Growth curves: a two-stage nonparametric approach. Journal of Statistical Planning and Inference, 38(3):327–350, March 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894900140.

Bhattacharyya:1991:APE

[BG91] Gouri K. Bhattacharyya and Jayanta K. Ghosh. Asymptotic properties of estimators in a binomial reliability growth model and its continuous-time analog. Journal of Statistical Planning and Inference, 29(1–2):43–53, September/October 1991. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290120H.

Beirlant:1998:ANE

[BG98] Jan Beirlant and László Györfi. On the asymptotic normality of the L _{2}-error in partitioning regression estimation. Journal of Statistical Planning and Inference, 71(1–2):93–107, August 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375898000081.

Baksalary:1995:BEV

[BGK95] J. K. Baksalary, S. Gnot, and S. Kageyama. Best estimation of variance components with arbitrary kurtosis in two-way layouts mixed models. Journal of Statistical Planning and Inference, 44(1):65–75, March 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589400030Y.

Billington:1997:IPD

[BGL97] Elizabeth J. Billington, M. Gionfriddo, and C. C. Lindner. The intersection problem for K_{4} − e designs. Journal of Statistical Planning and Inference, 58(1):5–27, March 1, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896000560.

Baksalary:1992:MNE

[BH92a] Jerzy K. Baksalary and Jan Hauke. Minimum number of experimental units in connected block designs with certain additional properties. Journal of Statistical Planning and Inference, 30(2):173–183, February 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375892900798.

Benghabrit:1992:ORB

[BH92b] Youssef Benghabrit and Marc Hallin. Optimal rank-based tests against first-order superdiagonal bilinear dependence. Journal of Statistical Planning and Inference, 32(1):45–61, July 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290151H.

Bohm:1996:ESC

[BH96] Walter Böhm and Peter Hackl. The effect of serial correlation on the in-control average run length of cumulative score charts. Journal of Statistical Planning and Inference, 54(1):15–30, September 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375895001530.

Benghabrit:1998:LAO

[BH98] Youssef Benghabrit and Marc Hallin. Locally asymptotically optimal tests for AR(p) against diagonal bilinear dependence. Journal of Statistical Planning and Inference, 68(1):47–63, May 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375897001353.

Berry:1999:BPM

[BH99a] Donald A. Berry and Yosef Hochberg. Bayesian perspectives on multiple comparisons. Journal of Statistical Planning and Inference, 82(1–2):215–227, December 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375899000440.

Bui:1999:IRC

[BH99b] Q. M. Bui and R. M. Huggins. Inference for the random coefficients bifurcating autoregressive model for cell lineage studies. Journal of Statistical Planning and Inference, 81(2):253–262, November 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837589900049X.

Bhapkar:1991:LIP

[Bha91] Vasant P. Bhapkar. Loss of information in the presence of nuisance parameters and partial sufficiency. Journal of Statistical Planning and Inference, 28(2):185–203, June 1991. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589190025A.

Bhansali:1999:AMS

[Bha99] R. J. Bhansali. Autoregressive model selection for multistep prediction. Journal of Statistical Planning and Inference, 78(1–2): 295–305, May 1999. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375898002201.

Bechhofer:1991:DES

[BHT91] Robert E. Bechhofer, Anthony J. Hayter, and Ajit C. Tamhane. Designing experiments for selecting the largest normal mean when the variances are known and unequal: Optimal sample size allocation. Journal of Statistical Planning and Inference, 28(3): 271–289, July 1991. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589190067O.

Bruen:1997:ISP

[BHW97] Aiden Bruen, Lucien Haddad, and David Wehlau. Intersectionsets in PG(n,2). Journal of Statistical Planning and Inference, 62 (1):3–11, July 1997. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375896001619.

Bierbrauer:1996:COA

[Bie96] Jürgen Bierbrauer. Construction of orthogonal arrays. Journal of Statistical Planning and Inference, 56(1):39–47, December 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375896000079.

Bischoff:1993:ODL

[Bis93] Wolfgang Bischoff. On D-optimal designs for linear models under correlated observations with an application to a linear model with multiple response. Journal of Statistical Planning and Inference, 37(1):69–80, October 1993. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390080P.

Brunk:1990:FCD

[BJ90] H. D. Brunk and P. W. Jones. Fitting conditional distributions using orthogonal expansions. Journal of Statistical Planning and Inference, 26(3):325–337, November 1990. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589090135H.

Basu:1998:SIR

[BJL98] Sanjib Basu, Sreenivasa Rao Jammalamadaka, and Wei Liu. Stability and infinitesimal robustness of posterior distributions and posterior quantities. Journal of Statistical Planning and Inference, 71(1–2):151–162, August 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898000901.

Banerjee:1996:NIS

[BK96] Shakti Banerjee and Sanpei Kageyama. Non-isomorphic solutions of group divisible designs with k = 3 and λ_{1} = 0. Journal of Statistical Planning and Inference, 56(1):23–32, December 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896000055.

Bryant:1998:CS

[BKF98] Darryn E. Bryant, A. Khodkar, and Hung-Lin Fu. (m, n)- cycle systems. Journal of Statistical Planning and Inference, 74(2):365–370, November 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375898000846.

Billington:1993:BTD

[BKM93] Elizabeth J. Billington, A. Khodkar, and E. S. Mahmoodian. Balanced ternary designs with block size four. Journal of Statistical Planning and Inference, 37(1):95–126, October 1993. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390082H.

Brockwell:1992:ENP

[BL92] Peter J. Brockwell and Jian Liu. Estimating the noise parameters from observations of a linear process with stable innovations. Journal of Statistical Planning and Inference, 33(2):175– 186, November 1992. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589290065Z.

Bryant:1996:PDC

[BL96] Darryn E. Bryant and C. C. Lindner. 2-perfect directed mcycle systems can be equationally defined for m = 3, 4, and 5 only. Journal of Statistical Planning and Inference, 56(1):57– 63, December 1996. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375896000092.

Benjamini:1999:SMH

[BL99] Yoav Benjamini and Wei Liu. A step-down multiple hypotheses testing procedure that controls the false discovery rate under independence. Journal of Statistical Planning and Inference, 82(1–2):163–170, December 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375899000403.

Bar-Lev:1993:MGF

[BLBL93] Shaul K. Bar-Lev, Nitza Barkan, and Naftali A. Langberg. Moment generating function based estimators with some optimal properties. Journal of Statistical Planning and Inference, 35(3): 279–291, June 1993. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589390016Y.

Baksalary:1990:ALE

[BM90] Jerzy K. Baksalary and Augustyn Markiewicz. Admissible linear estimators of an arbitrary vector of parametric functions in the general Gauss–Markov model. Journal of Statistical Planning and Inference, 26(2):161–171, October 1990. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589090124D.

Berger:1991:BAL

[BM91] James O. Berger and Julia Mortera. Bayesian analysis with limited communication. Journal of Statistical Planning and Inference, 28(1):1–24, May 1991. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837589190055J.

Bhandari:1992:TMS

[BM92] Subir Kumar Bhandari and Rahul Mukerjee. Two- and multistage selection procedures for Weibull populations. Journal of Statistical Planning and Inference, 30(1):73–81, January 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375892901085.

Bohm:1993:TSM

[BM93] W. Böhm and S. G. Mohanty. The transient solution of M/M/1 queues under (M, N)-policy. A combinatorial approach. Journal of Statistical Planning and Inference, 34(1):23–33, January 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589390031Z.

Berger:1994:BRB

[BM94a] James Berger and Elías Moreno. Bayesian robustness in bidimensional models: Prior independence. Journal of Statistical Planning and Inference, 40(2–3):161–176, July 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589490118X.

Berger:1994:RBH

[BM94b] James Berger and Julia Mortera. Robust Bayesian hypothesis testing in the presence of nuisance parameters. Journal of Statistical Planning and Inference, 40(2–3):357–373, July 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894901317.

Billio:1998:SSS

[BM98] Monica Billio and Alain Monfort. Switching state-space models: Likelihood function, filtering and smoothing. Journal of Statistical Planning and Inference, 68(1):65–103, May 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897001365.

Bagui:1999:CMO

[BM99a] S. C. Bagui and K. L. Mehra. Classification of multiple observations using multi-stage rank nearest neighbor rule. Journal of Statistical Planning and Inference, 76(1–2):163–183, February 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375898001372.

Bansal:1999:RTE

[BM99b] Naveen K. Bansal and Neeraj Misra. On ranking treatment effects in a general linear model. Journal of Statistical Planning and Inference, 78(1–2):1–11, May 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837589800202X.

Baksalary:1995:ALE

[BMR95] Jerzy K. Baksalary, Augustyn Markiewicz, and C. Radhakrishna Rao. Admissible linear estimation in the general Gauss–Markov model with respect to an arbitrary quadratic risk function. Journal of Statistical Planning and Inference, 44(3):341–347, ???? 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375894000816.

Barlow:1991:COD

[BMS91] Richard E. Barlow, R. W. Mensing, and N. G. Smiriga. Computing the optimal design for a calibration experiment. Journal of Statistical Planning and Inference, 29(1–2):5–19, September/October 1991. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589290117B.

Bagui:1997:SVR

[BMV97] Subhash C. Bagui, K. L. Mehra, and Brandon K. Vaughn. An M-stage version of the k-RNN rule in statistical discrimination. Journal of Statistical Planning and Inference, 65(2):323–333, December 15, 1997. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375897000554.

Berg:1991:SIC

[BN91] Sven Berg and Krzysztof Nowicki. Statistical inference for a class of modified power series distributions with applications to random mapping theory. Journal of Statistical Planning and Inference, 28(2):247–261, June 1991. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589190030I.

Barakat:1999:CRE

[BN99] H. M. Barakat and E. M. Nigm. Convergence of random extremal quotient and product. Journal of Statistical Planning and Inference, 81(2):209–221, November 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375899000129.

Barndorff-Nielsen:1997:YSS

[BNJ97] O. E. Barndorff-Nielsen and P. E. Jupp. Yokes and symplectic structures. Journal of Statistical Planning and Inference, 63(2):133–146, October 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375897000062.

Beran:1999:SFA

[BO99] Jan Beran and Dirk Ocker. SEMIFAR forecasts, with applications to foreign exchange rates. Journal of Statistical Planning and Inference, 80(1–2):137–153, August 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837589800247X.

Bohning:1995:RRM

[Böh95] Dankmar Böhning. A review of reliable maximum likelihood algorithms for semiparametric mixture models. Journal of Statistical Planning and Inference, 47(1–2):5–28, October 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589400119G.

Bohm:1996:NAL

[Böh96] W. Böhm. A note on the asymptotics of lattice paths with general boundaries. Journal of Statistical Planning and Inference, 54(1):3–13, September 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375895001522.

Bondesson:1997:GPC

[Bon97] Lennart Bondesson. A generalization of Poincaré’s characterization of exponential families. Journal of Statistical Planning and Inference, 63(2):147–155, October 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375895000070.

Bose:1994:BRM

[Bos94] Sudip Bose. Bayesian robustness with more than one class of contaminations. Journal of Statistical Planning and Inference, 40(2– 3):177–187, July 1994. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375894901198.

Boshuizen:1997:RMG

[Bos97] Frans A. Boshuizen. A replacement model with general age-dependent failure rates. Journal of Statistical Planning and Inference, 59(2):213–228, April 15, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897836397.

Bose:1998:TPP

[Bos98]  Sudip Bose.  On the transitivity of the posterior Pit-

man closeness criterion. Journal of Statistical Planning and Inference, 69(2):263–274, June 15, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897001559.

Boukai:1990:APF

[Bou90] Benzion Boukai. The asymptotic power function of GLR tests for local change of parameter. Journal of Statistical Planning and Inference, 26(3):291–303, November 1990. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589090133F.

Baksalary:1992:ILR

[BP92] Jerzy K. Baksalary and Pawel R. Pordzik. Implied linear restrictions in the general Gauss–Markov model. Journal of Statistical Planning and Inference, 30(2):237–248, February 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375892900846.

Baksalary:1993:PTE

[BP93] Jerzy K. Baksalary and Pawel R. Pordzik. Preliminary test estimation of a vector of parametric functions in the general Gauss–Markov model. Journal of Statistical Planning and Inference, 36(2–3):227–239, August/September 1993. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390126Q.

Basawa:1994:P

[BP94] I. V. Basawa and N. U. Prabhu. Preface. Journal of Statistical Planning and Inference, 39(2):135, April 15, 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589490202X.

Borovkov:1995:RIR

[BP95] Konstantin Borovkov and Dietmar Pfeifer. On record indices and record times. Journal of Statistical Planning and Inference, 45(1– 2):65–79, ???? 1995. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375894000638.

Bueso:1999:SCM

[BQA99] M. C. Bueso, G. Qian, and J. M. Angulo. Stochastic complexity and model selection from incomplete data. Journal of Statistical Planning and Inference, 76(1–2):273–284, February 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898001128.

Baron:1997:ABC

[BR97] Michael Baron and Andrew L. Rukhin. Asymptotic behavior of confidence regions in the change-point problem. Journal of Statistical Planning and Inference, 58(2):263–282, March 15, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896000766.

Berrendero:1998:SUC

[BR98] JoséR. Berrendero and Juan Romo. Stability under contamination of robust regression estimators based on differences of residuals. Journal of Statistical Planning and Inference, 70(1):149– 165, July 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375897001699.

Baksalary:1992:SIN

[BRM92] Jerzy K. Baksalary, C. Radhakrishna Rao, and Augustyn Markiewicz. A study of the influence of the ‘natural restrictions’ on estimation problems in the singular Gauss–Markov model. Journal of Statistical Planning and Inference, 31(3):335– 351, June 1992. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589290141E.

Betro:1994:RBA

[BRM94] Bruno Betrò, Fabrizio Ruggeri, and Marek Meczarski. Robust Bayesian analysis under generalized moments conditions. Journal of Statistical Planning and Inference, 41(3):257–266, October 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375894900221.

Broniatowski:1993:EWT

[Bro93] Michel Broniatowski. On the estimation of the Weibull tail coefficient. Journal of Statistical Planning and Inference, 35(3): 349–365, June 1993. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589390022X.

Brockwell:1994:CTT

[Bro94] P. J. Brockwell. On continuous-time threshold ARMA processes. Journal of Statistical Planning and Inference, 39(2):291– 303, April 15, 1994. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375894902100.

Bai:1999:MSD

[BRW99] Z. D. Bai, C. R. Rao, and Y. Wu. Model selection with data-oriented penalty. Journal of Statistical Planning and Inference, 77(1):103–117, February 15, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898001682.

Bai:1993:MTT

[BRZ93] Z. D. Bai, C. Radhakrishna Rao, and L. C. Zhao. MANOVA type tests under a convex discrepancy function for the standard multivariate linear model. Journal of Statistical Planning and Inference, 36(1):77–90, July 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390103D.

Bartmanska:1992:BET

[BS92] Barbara Bartmańska and Dominik Szynal. On Berry–Esseen theorem for some functions of U-statistics. Journal of Statistical Planning and Inference, 30(3):385–393, March 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290164N.

Baksalary:1993:DMT

[BS93a] Jerzy K. Baksalary and Idzi Siatkowski. Decomposability of the C-matrix of a two-way elimination of heterogeneity design. Journal of Statistical Planning and Inference, 36(2–3):301–309, August/September 1993. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589390132P.

Baksalary:1993:P

[BS93b] Jerzy K. Baksalary and George P. H. Styan. Preface. Journal of Statistical Planning and Inference, 36(2–3):127, August/ September 1993. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589390117O.

Berube:1993:DTW

[BS93c] Julie Bérubé and George P. H. Styan. Decomposable threeway layouts. Journal of Statistical Planning and Inference, 36(2–3):311–322, August/September 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390133Q.

Bhandari:1994:LFC

[BS94] Subir Kumar Bhandari and Debasis Sengupta. Least favorable configuration and asymptotic efficiency of multinomial selection procedures. Journal of Statistical Planning and Inference, 39(1): 71–83, April 1, 1994. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375894901139.

Bose:1995:MER

[BS95] Sudip Bose and Eric Slud. Maximin efficiency-robust tests and some extensions. Journal of Statistical Planning and Inference, 46(1):105–121, July 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375894E00918.

Buzas:1996:IVE

[BS96] J. S. Buzas and L. A. Stefanski. Instrumental variable estimation in a probit measurement error model. Journal of Statistical Planning and Inference, 55(1):47–62, October 15, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895001808.

Boukai:1997:SCI

[BS97] Benzion Boukai and Jyotirmoy Sarkar. Sequential confidence intervals with fixed-proportional accuracy for the mean of NEF-PVF distributions. Journal of Statistical Planning and Inference, 59(2):241–256, April 15, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896001085.

Bhaumik:1998:ETC

[BS98a] Dulal K. Bhaumik and Arvind K. Shah. Exact tests in crossover designs. Journal of Statistical Planning and Inference, 72 (1–2):79–88, September 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375898000238.

Borzadaran:1998:GCT

[BS98b] G. R. Mohtashami Borzadaran and D. N. Shanbhag. General characterization theorems via the mean absolute deviation. Journal of Statistical Planning and Inference, 74(2):205–214, November 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375898000809.

Bosq:1998:EAS

[BS98c] Denis Bosq and Jia Shen. Estimation of an autoregressive semiparametric model with exogenous variables. Journal of Statistical Planning and Inference, 68(1):105–127, May 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897001377.

Bonin:1993:SAS

[BSS93] Joseph Bonin, Louis Shapiro, and Rodica Simion. Some qanalogues of the Schröder numbers arising from combinatorial statistics on lattice paths. Journal of Statistical Planning and Inference, 34(1):35–55, January 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375893900322.

Ball:1993:BCD

[BSV93] F. G. Ball, A. F. M. Smith, and I. Verdinelli. Biased coin designs with a Bayesian bias. Journal of Statistical Planning and Inference, 34(3):403–421, March 1993. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390148Y.

Basu:1997:MNE

[BSV97] Ayanendranath Basu, Sahadeb Sarkar, and A. N. Vidyashankar. Minimum negative exponential disparity estimation in parametric models. Journal of Statistical Planning and Inference, 58(2):349– 370, March 15, 1997. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837589600078X.

Basford:1997:GPA

[BT97] Kaye E. Basford and John W. Tukey. Graphical profiles as an aid to understanding plant breeding experiments. Journal of Statistical Planning and Inference, 57(1):93–107, January 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896000389.

Bradley:1999:DEN

[BT99] Richard C. Bradley and Lanh T. Tran. Density estimation for nonisotropic random fields. Journal of Statistical Planning and Inference, 81(1):51–70, October 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375899000117.

Bloemen:1996:EGQ

[BTV96] I. Bloemen, J. A. Thas, and H. Van Maldeghem. Elation generalized quadrangles of order (p, t), p prime, are classical. Journal of Statistical Planning and Inference, 56(1):49–55, December 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375897830595.

Buhlmann:1999:EAP

[Büh99] Peter Bühlmann. Efficient and adaptive post-model-selection estimators. Journal of Statistical Planning and Inference, 79(1):

1–9, June 1999. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375898002365.

Buzas:1998:USP

[Buz98] Jeffrey S. Buzas. Unbiased scores in proportional hazards regression with covariate measurement error. Journal of Statistical Planning and Inference, 67(2):247–257, April 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897001080.

Bagchi:1991:ONC

[BvB91] S. Bagchi and E. E. M. van Berkum. On the optimality of a new class of adjusted orthogonal designs. Journal of Statistical Planning and Inference, 28(1):61–65, May 1991. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589190059N.

Bergman:1991:BSR

[BX91] Bo Bergman and Min Xie. On Bayesian software reliability modelling. Journal of Statistical Planning and Inference, 29(1–2):33–41, September/October 1991. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290119D.

Burke:1995:GFT

[BY95] Murray D. Burke and Kam C. Yuen. Goodness-of-fit tests for the Cox model via bootstrap method. Journal of Statistical Planning and Inference, 47(3):237–256, October 15, 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589500141U.

Bolfarine:1991:BMP

[BZ91] H. Bolfarine and S. Zacks. Bayes and minimax prediction in finite populations. Journal of Statistical Planning and Inference, 28(2): 139–151, June 1991. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375891900227.

Bera:1996:STL

[BZ96] Anil K. Bera and Xiao-Lei Zuo. Specification test for a linear regression model with ARCH process. Journal of Statistical Planning and Inference, 50(2):283–308, March 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895000593.

Bolfarine:1996:PPT

[BZS96] Heleno Bolfarine, Shelemyahu Zacks, and Mônica C. Sandoval. On predicting the population total under regression models with measurement errors. Journal of Statistical Planning and Inference, 55(1):63–76, October 15, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895001816.

Cuesta-Albertos:1995:CBM

[CADMM95] J. A. Cuesta-Albertos, J. S. Domínguez-Menchero, and C. Matrán. Consistency of L_{p}-best monotone approximations. Journal of Statistical Planning and Inference, 47(3):295–318, October 15, 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589400144K.

Cuesta-Albertos:1998:GBM

[CAGM98] J. A. Cuesta-Albertos, A. Gordaliza, and C. Matrán. On the geometric behaviour of multidimensional location measures. Journal of Statistical Planning and Inference, 67(2):191–208, April 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375897001092.

Cai:1999:SAR

[Cai99] Jianwen Cai. Some asymptotic results for hypothesis testing in multivariate failure time data. Journal of Statistical Planning and Inference, 75(2):211–221, January 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898001438.

Calinski:1993:BEO

[Cal93] Tadeusz Caliński. Balance, efficiency and orthogonality concepts in block designs. Journal of Statistical Planning and Inference, 36(2–3):283–300, August/September 1993. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390131O.

Carter:1996:SSS

[Car96] R. A. L. Carter. Some small-sample properties of instrumentalvariables estimators of block triangular models. Journal of Statistical Planning and Inference, 50(2):207–221, March 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895000542.

Carroll:1997:CPD

[Car97] R. J. Carroll. Comment on the paper by Desmond. Journal of Statistical Planning and Inference, 60(1):104–106, April 30, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375897900100. See [Des97a].

Carriere:1999:MRM

[Car99] K. C. Carriere. Methods for repeated measures data analysis with missing values. Journal of Statistical Planning and Inference, 77(2):221–236, March 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898001980.

Clarke:1994:JPA

[CB94] Bertrand S. Clarke and Andrew R. Barron. Jeffreys’ prior is asymptotically least favorable under entropy risk. Journal of Statistical Planning and Inference, 41(1):37–60, August 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894901538.

Cressie:1994:SSO

[CBM94] Noel Cressie, Jonathan Biele, and Peter B. Morgan. Samplesize-optimal sequential testing. Journal of Statistical Planning and Inference, 39(2):305–327, April 15, 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894902119.

Cheng:1990:ANR

[CC90] P. E. Cheng and K. F. Cheng. Asymptotic normality for robust R-estimators of regression function. Journal of Statistical Planning and Inference, 24(2):137–149, February 1990. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589090037U.

Chu:1995:NRE

[CC95] C. K. Chu and P. E. Cheng. Nonparametric regression estimation with missing data. Journal of Statistical Planning and Inference, 48(1):85–99, November 1, 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589400151K.

Chadjiconstantinidis:1996:OCC

[CC96] Stathis Chadjiconstantinidis and Theodore Chadjipadelis. Doptimal cyclic complex linear designs and supplementary difference sets with association vector. Journal of Statistical Planning and Inference, 53(1):93–115, August 1, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895001379.

Chen:1999:ICD

[CC99] Yi-Hau Chen and Hung Chen. Incomplete covariates data in generalized linear models. Journal of Statistical Planning and Inference, 79(2):247–258, July 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898002559.

Cuevas:1996:ABP

[CdlH96] Antonio Cuevas and Julián de la Horra. Asymptotic behaviour of the predictive density in the exchangeable case. Journal of Statistical Planning and Inference, 55(1):1–12, October 15, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895001794.

Chadjiconstantinidis:1996:ONS

[CF96] Stathis Chadjiconstantinidis and Nicolas Farmakis. The Aoptimal non-saturated weighing designs for n = 15 observations. Journal of Statistical Planning and Inference, 49(3):401– 413, February 1, 1996. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375895000224.

Cordeiro:1998:NBT

[CF98] Gauss M. Cordeiro and Silvia L. P. Ferrari. A note on Bartletttype correction for the first few moments of test statistics. Journal of Statistical Planning and Inference, 71(1–2):261–269, August 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375898000056.

Cakmak:1998:LCA

[CFM^{+}98] S. Cakmak, D. A. S. Fraser, P. McDunnough, N. Reid, and X. Yuan. Likelihood centered asymptotic model exponential and location model versions. Journal of Statistical Planning and Inference, 66(2):211–222, March 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897000852.

Carroll:1997:P

[CFT97] Raymond Carroll, Luisa Fernholz, and David Tyler. Preface. Journal of Statistical Planning and Inference, 57(1):3–4, January 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375896000298.

Chan:1998:OPG

[CG98a] Schultz Chan and Malay Ghosh. Orthogonal projections and the geometry of estimating functions. Journal of Statistical Planning and Inference, 67(2):227–245, April 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897001183.

Clarke:1998:OSP

[CG98b] Bertrand Clarke and Paul Gustafson. On the overall sensitivity of the posterior distribution to its inputs. Journal of Statistical Planning and Inference, 71(1–2):137–150, August 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898000147.

Csorgo:1998:ACB

[CGZ98] Miklós Csörgő, Joseph L. Gastwirth, and Ricardas Zitikis. Asymptotic confidence bands for the Lorenz and Bonferroni curves based on the empirical Lorenz curve. Journal of Statistical Planning and Inference, 74(1):65–91, October 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898001037.

Chang:1990:FSO

[CH90] I. S. Chang and C. A. Hsiung. Finite sample optimality of maximum partial likelihood estimation in Cox’s model for counting processes. Journal of Statistical Planning and Inference, 25(1): 35–42, May 1990. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589090005F.

Chang:1992:ODM

[CH92] Jane Y. Chang and Jason C. Hsu. Optimal designs for multiple comparisons with the best. Journal of Statistical Planning and Inference, 30(1):45–62, January 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375892901063.

Chen:1994:OED

[CH94] Hegang Chen and A. S. Hedayat. Optimum experimental designs for comparative bioavailability studies. Journal of Statistical Planning and Inference, 42(1–2):271–289, November 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894902011.

Chen:1995:AOE

[CH95] K. Chen and C. C. Heyde. On asymptotic optimality of estimating functions. Journal of Statistical Planning and Inference, 48(1):101–112, November 1, 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589400133G.

Chang:1996:ODP

[CH96] Fu-Chuen Chang and Berthold Heiligers. E-optimal designs for polynomial regression without intercept. Journal of Statistical Planning and Inference, 55(3):371–387, November 8, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375895001964.

Chen:1998:DRI

[CH98] Hegang Chen and A. S. Hedayat. 2^{n−m} designs with resolution III or IV containing clear two-factor interactions. Journal of Statistical Planning and Inference, 75(1):147–158, November 15, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375898001220.

Chan:1990:ECA

[Cha90] Ling-Yau Chan. An example of construction of asymptotic Doptimal design which converges on an infinite set of points. Journal of Statistical Planning and Inference, 25(1):29–34, May 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589090004E.

Chaloner:1993:NOB

[Cha93] Kathryn Chaloner. A note on optimal Bayesian design for nonlinear problems. Journal of Statistical Planning and Inference, 37(2):229–235, November 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390091J.

Chai:1995:CON

[Cha95a] Feng-Shun Chai. Construction and optimality of nearly linear trend-free designs. Journal of Statistical Planning and Inference, 48(1):113–129, November 1, 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589400146M.

Chan:1995:DFR

[Cha95b] N. N. Chan. On a downdating formula for regression. Journal of Statistical Planning and Inference, 46(3):347–350, August 15, 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589400115C.

Chauveau:1995:SEA

[Cha95c] Didier Chauveau. A stochastic EM algorithm for mixtures with censored data. Journal of Statistical Planning and Inference, 46 (1):1–25, July 1995. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589400097F.

Charalambides:1996:DGF

[Cha96a] Ch. A. Charalambides. On the q-differences of the generalized q-factorials. Journal of Statistical Planning and Inference, 54(1):31–43, September 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375895001549.

Charemaz:1996:URE

[Cha96b] Wojciech W. Charemaz. Unit root econometrics and economic nonlinearities. Journal of Statistical Planning and Inference, 49(1):105–126, January 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375895000321.

Charpin:1996:SDC

[Cha96c] Pascale Charpin. Self-dual codes which are principal ideals of the group algebra F_{2}[{F_{2^{m}}, +}]. Journal of Statistical Planning and Inference, 56(1):79–92, December 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895000119.

Chaturvedi:1996:RBA

[Cha96d] Anoop Chaturvedi. Robust Bayesian analysis of the linear regression model. Journal of Statistical Planning and Inference, 50(2): 175–186, March 1996. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375895000526.

Chaganty:1997:AAA

[Cha97a] N. Rao Chaganty. An alternative approach to the analysis of longitudinal data via generalized estimating equations. Journal of Statistical Planning and Inference, 63(1):39–54, September 30, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375896002030.

Chaudhuri:1997:TEA

[Cha97b] Gopal Chaudhuri. Testing exponentiality against L-distributions. Journal of Statistical Planning and Inference, 64(2):249–255, November 15, 1997. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375897000475.

Chang:1998:MOM

[Cha98] Chwen-Ming Chang. Minimal orthogonal main-effect plans and partial replication. Journal of Statistical Planning and Inference, 70(1):167–179, July 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375897001754.

Chen:1995:RHS

[Che95] Zhiqiang Chen. Robustness of the half-space median. Journal of Statistical Planning and Inference, 46(2):175–181, August 1, 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375894001055.

Chen:1996:CMM

[Che96a] Jiahua Chen. On the conditional and mixture model approaches for matched pairs. Journal of Statistical Planning and Inference, 55(3):319–329, November 8, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896000808.

Cheng:1996:MBP

[Che96b] Y. Cheng. Multistage bandit problems. Journal of Statistical Planning and Inference, 53(2):153–170, August 15, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895001298.

Chen:1997:BET

[Che97] Gang Chen. Berry–Esseen-type bounds for the kernel estimator of conditional distribution and conditional quantiles. Journal of Statistical Planning and Inference, 60(2):311–330, May 15, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375896001401.

Chen:1998:JEP

[Che98a] Zhenmin Chen. Joint estimation for the parameters of Weibull distribution. Journal of Statistical Planning and Inference, 66(1):113–120, January 5, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375897000803.

Cheng:1998:WBI

[Che98b] Ching-Shui Cheng. N-way balanced incomplete block designs. Journal of Statistical Planning and Inference, 72(1–2):109–119, September 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375898000263.

Clark:1992:HNN

[CHF92] Gregory A. Clark, Robert D. Hawkins, and William N. Frost. How neural are neural networks? a comparison of information processing and storage in artificial and real neural systems. Journal of Statistical Planning and Inference, 33(1):27–65, October 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375892900938.

Chang:1999:TMI

[CHH99] Kuang-Chao Chang, Chien-Pai Han, and Doyle L. Hawkins. Truncated multiple inverse sampling in post-stratification. Journal of Statistical Planning and Inference, 76(1–2):215–234, February 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375898001359.

Chiu:1991:CGM

[Chi91] Shean-Tsong Chiu. On the consistency of the global minimizer of Mallows’ criterion for nonparametric regression. Journal of Statistical Planning and Inference, 27(2):223–230, February 1991. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375891900179.

Chiu:1992:BRS

[Chi92] Shean Chiu. Book review: Subset selection in regression: A. J. Miller Chapman and Hall, London, 1990 229 pp., $47.50. Journal of Statistical Planning and Inference, 32(3):420–422, September 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589290022K.

Christofides:1992:SLL

[Chr92] Tasos C. Christofides. A strong law of large numbers for Ustatistics. Journal of Statistical Planning and Inference, 31(2): 133–145, May 1992. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589290025N.

Christopeit:1994:EPE

[Chr94] N. Christopeit. Estimating parameters of an extreme value distribution by the method of moments. Journal of Statistical Planning and Inference, 41(2):173–186, September 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894901627.

Chu:1995:BSN

[Chu95]  C. K. Chu.  Bandwidth selection in nonparametric re-

gression with general errors. Journal of Statistical Planning and Inference, 44(3):265–275, ???? 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589400096E.

Chu:1996:HED

[Chu96] Wensong Chu. Homogeneous embeddings of disjoint Mendelsohn triple systems. Journal of Statistical Planning and Inference, 51(2):143–155, April 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837589500078X.

Chang:1995:ESB

[cIC95] Yuan chin Ivan Chang. Estimation in some binary regression models with prescribed accuracy. Journal of Statistical Planning and Inference, 44(3):313–325, ???? 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589400054Y.

Ceranka:1992:OBS

[CK92] Bronislaw Ceranka and Krystyna Katulska. Optimum biased spring balance weighing designs with non-homogeneity of the variances of errors. Journal of Statistical Planning and Inference, 30(2):185–193, February 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290080C.

Chadjiconstantinidis:1994:EOF

[CK94] Stathis Chadjiconstantinidis and Stratis Kounias. Exact Aoptimal first-order saturated designs with n ≡ 3 mod 4 observations. Journal of Statistical Planning and Inference, 42(3):379– 391, December 1994. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375894000042.

Calinski:1996:RME

[CK96] Tadeusz Caliński and Sanpei Kageyama. The randomization model for experiments in block designs and the recovery of inter-block information. Journal of Statistical Planning and Inference, 52(3):359–374, July 1, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895001247.

Cobbers:1998:NCG

[CK98] Henrik Cobbers and Udo Kamps. A note on characterizations of the geometric distribution. Journal of Statistical Planning and Inference, 67(2):187–190, April 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897865780.

Cline:1999:CDD

[CK99] Gary A. Cline and Richard J. Kryscio. Contagion distributions for defining disease clustering in time. Journal of Statistical Planning and Inference, 78(1–2):325–347, May 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898002237.

Chen:1991:IRD

[CL91] Jiahua Chen and Dennis K. J. Lin. On the identity relationships of 2−p designs. Journal of Statistical Planning and Inference, 28 (1):95–98, May 1991. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589190062J.

Chaloner:1992:BDA

[CL92]  Kathryn Chaloner and Kinley Larntz.  Bayesian design

for accelerated life testing. Journal of Statistical Planning and Inference, 33(2):245–259, November 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290071Y.

Chao:1993:ADR

[CL93] M. T. Chao and G. D. Lin. The asymptotic distributions of the remedians. Journal of Statistical Planning and Inference, 37(1): 1–11, October 1993. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589390075H.

Chernoff:1995:ADL

[CL95] Herman Chernoff and Eric Lander. Asymptotic distribution of the likelihood ratio test that a mixture of two binomials is a single binomial. Journal of Statistical Planning and Inference, 43(1–2): 19–40, January 1995. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589400006H.

Cressie:1996:ARE

[CL96] Noel Cressie and Soumendra Nath Lahiri. Asymptotics for REML estimation of spatial covariance parameters. Journal of Statistical Planning and Inference, 50(3):327–341, March 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895000615.

Chang:1997:ODW

[CL97] Fu-Chuen Chang and Ge-Chen Lin. D-optimal designs for weighted polynomial regression. Journal of Statistical Planning and Inference, 62(2):317–331, August 15, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896001838.

Chen:1998:ISD

[CL98] Jiahua Chen and Dennis K. J. Lin. On the identifiability of a supersaturated design. Journal of Statistical Planning and Inference, 72(1–2):99–107, September 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898000251.

Chang:1998:MIS

[CLH98] Kuang-Chao Chang, Jeng-Fu Liu, and Chien-Pai Han. Multiple inverse sampling in post-stratification. Journal of Statistical Planning and Inference, 69(2):209–227, June 15, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897001572.

Colbourn:1991:NDWa

[CLR91a] Charles J. Colbourn, C. C. Lindner, and C. A. Rodger. Neighbor designs and m-wheel systems. Journal of Statistical Planning and Inference, 27(3):335–340, March 1991. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589190046H.

Colbourn:1991:NDWb

[CLR91b] Charles J. Colbourn, C. C. Lindner, and C. A. Rodger. Neighbor designs and m-wheel systems. Journal of Statistical Planning and Inference, 29(3):365, November 1991. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589190010C.

Chadjiconstantinidis:1991:SOO

[CM91] Stathis Chadjiconstantinidis and Chronis Moyssiadis. Some Doptimal odd-equireplicated designs for a covariate model. Journal of Statistical Planning and Inference, 28(1):83–93, May 1991. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589190061I.

Cuesta:1992:RSC

[CM92] Juan A. Cuesta and Carlos Matrán. A review on strong convergence of weighted sums of random elements based on Wasserstein metrics. Journal of Statistical Planning and Inference, 30(3):359– 370, March 1992. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589290162L.

Chatterjee:1993:SDEa

[CM93a] Kashinath Chatterjee and Rahul Mukerjee. Search designs for estimating main effects and searching several two-factor interactions in general factorials. Journal of Statistical Planning and Inference, 35(1):131–138, April 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390073F.

Chatterjee:1993:SDEb

[CM93b] Kashinath Chatterjee and Rahul Mukerjee. Search designs for estimating main effects and searching several two-factor interactions in general factorials. Journal of Statistical Planning and Inference, 37(3):385–392, December 1993. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390114L.

Csorgo:1995:ISS

[CM95] Sándor Csörgö and David M. Mason. Intermediate sums and stochastic compactness of maxima. Journal of Statistical Planning and Inference, 45(1–2):81–90, ???? 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894000646.

Camina:1996:LCS

[CM96] Alan R. Camina and Alice Ann Miller. Line-closed subsets of Steiner triple systems and classical linear spaces. Journal of Statistical Planning and Inference, 56(1):65–77, December 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896000109.

Chan:1997:IR

[CM97] A. Chan and J. Meloche. Image restoration. Journal of Statistical Planning and Inference, 65(2):233–254, December 15, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897000669.

Casalis:1998:TED

[CM98] Muriel Casalis and Hélène Massam. A test for the equality of the direction parameters in a sample from the Ndimensional hyperboloid distribution. Journal of Statistical Planning and Inference, 74(2):229–251, November 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898000974.

Cohen:1995:CIM

[CMB95] Ayala Cohen, Udi Makov, and Ruth Boneh. Computer intensive methods for inference on parameters of complex models — a Bayesian alternative. Journal of Statistical Planning and Inference, 47(1–2):79–94, October 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589400123D.

Cooper:1994:CSN

[CMM94] J. A. Cooper, T. P. McDonough, and V. C. Mavron. Critical sets in nets and latin squares. Journal of Statistical Planning and Inference, 41(2):241–256, September 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894901678.

Chorneyko:1993:ERT

[CMS93] I. Z. Chorneyko, S. G. Mohanty, and Kanwar Sen. Enumeration of restricted three-dimensional lattice paths with fixed numbers of turns and an application. Journal of Statistical Planning and Inference, 34(1):57–62, January 1993. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375893900333.

Colbourn:1997:GDO

[CN97] Charles J. Colbourn and Gillian Nonay. A golf design of order 11. Journal of Statistical Planning and Inference, 58(1):29–31, March 1, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375896000572.

Coad:1994:SET

[Coa94] D. S. Coad. Sequential estimation for two-stage and threestage clinical trials. Journal of Statistical Planning and Inference, 42(3):343–351, December 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375894000336.

Coakley:1997:BRR

[Coa97] Clint W. Coakley. Book review: Robust statistical procedures: Asymptotics and interrelations: Jana Jurecková and Pranab Kumar Sen, (Wiley, New York, 1996, Price: $59.95, ISBN No.: 0-471-82221-3, xiv + 466 pp.). Journal of Statistical Planning and Inference, 64(1):159–161, October 30, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837589700102X.

Coleman:1995:FWI

[Col95] Daniel A. Coleman. A fixed-width interval for 1β in simple linear regression. Journal of Statistical Planning and Inference, 44(3): 291–312, ???? 1995. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589400055Z.

Colbourn:1996:SDC

[Col96] Charles J. Colbourn. Some direct constructions for incomplete transversal designs. Journal of Statistical Planning and Inference, 56(1):93–104, December 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896000122.

Conniffe:1996:AES

[Con96a] Denis Conniffe. Applying estimated score tests in econometrics. Journal of Statistical Planning and Inference, 49(2):261–278, January 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375895000402.

Conniffe:1996:CUA

[Con96b] Denis Conniffe. A comment on the use of auxiliary or surrogate outcome data. Journal of Statistical Planning and Inference, 55(3):353–359, November 8, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837589600081X.

Cox:1997:CPJ

[Cox97] D. R. Cox. Comments on paper by J. M. Bernardo. Journal of Statistical Planning and Inference, 65(1):177–178, December 1, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375897900677.

Crouchley:1995:MSM

[CP95] Robert Crouchley and Andrew R. Pickles. Multivariate survival models for repeated and correlated events. Journal of Statistical Planning and Inference, 47(1–2):95–110, October 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589400124E.

Cacoullos:1997:CDG

[CP97a] T. Cacoullos and V. Papathanasiou. Characterizations of distributions by generalizations of variance bounds and simple proofs of the CLT. Journal of Statistical Planning and Inference, 63(2):157–171, October 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375897000086.

Cheng:1997:UES

[CP97b] Cheng Cheng and Emanuel Parzen. Unified estimators of smooth quantile and quantile density functions. Journal of Statistical Planning and Inference, 59(2):291–307, April 15, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896001103.

Costantini:1998:MBS

[CP98] C. Costantini and D. Pasqualucci. Monotonicity of Bayes sequential tests for multidimensional and censored observations. Journal of Statistical Planning and Inference, 75(1):117–131, November 15, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375898000925.

Casse:1998:BBM

[CQ98] Rey Casse and Catherine T. Quinn. The Bruck–Bose map and Beukenhout–Metz unitals. Journal of Statistical Planning and Inference, 72(1–2):89–98, September 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837589800024X.

Cuevas:1993:RPB

[CR93] Antonio Cuevas and Juan Romo. On robustness properties of bootstrap approximations. Journal of Statistical Planning and Inference, 37(2):181–191, November 1993. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390087M.

Chaudhuri:1997:MAS

[CR97a] Arijit Chaudhuri and Debesh Roy. Model assisted survey sampling strategies with randomized response. Journal of Statistical Planning and Inference, 60(1):61–68, April 30, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837589600119X.

Chaudhuri:1997:OVE

[CR97b] Arijit Chaudhuri and Debesh Roy. Optimal variance estimation for generalized regression predictor. Journal of Statistical Planning and Inference, 60(1):139–151, April 30, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896001267.

Critchlow:1992:RSA

[Cri92] Douglas E. Critchlow. On rank statistics: an approach via metrics on the permutation group. Journal of Statistical Planning and Inference, 32(3):325–346, September 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290015K.

Crowley:1998:ECP

[Cro98] Evelyn M. Crowley. Estimation of clustered parameters. Journal of Statistical Planning and Inference, 74(2):273–291, November 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375898001104.

Croux:1996:PBR

[CRV96] Christophe Croux, Peter J. Rousseeuw, and Annemie Van Bael. Positive-breakdown regression by minimizing nested scale estimators. Journal of Statistical Planning and Inference, 53(2):197– 235, August 15, 1996. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589500128X.

Christofides:1990:MIC

[CS90] Tasos C. Christofides and Robert J. Serfling. Maximal inequalities and convergence results for generalized U-statistics. Journal of Statistical Planning and Inference, 24(3):271–286, March 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589090048Y.

Chen:1991:TSS

[CS91] Hung Chen and Jyh-Jen Horng Shiau. A two-stage spline smoothing method for partially linear models. Journal of Statistical Planning and Inference, 27(2):187–201, February 1991. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375891900157.

Czado:1992:ELM

[CS92] Claudia Czado and Thomas J. Santner. The effect of link misspecification on binary regression inference. Journal of Statistical Planning and Inference, 33(2):213–231, November 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375892900695.

Chaganty:1996:MSL

[CS96] Narasinga R. Chaganty and Jayaram Sethuraman. Multidimensional strong large deviation theorems. Journal of Statistical Planning and Inference, 55(3):265–280, November 8, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896000833.

Cavanaugh:1998:AIC

[CS98] Joseph E. Cavanaugh and Robert H. Shumway. An Akaike information criterion for model selection in the presence of incomplete data. Journal of Statistical Planning and Inference, 67(1):45– 65, March 16, 1998. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375897001158.

Chaganty:1999:EAB

[CS99a] N. Rao Chaganty and Justine Shults. On eliminating the asymptotic bias in the quasi-least squares estimate of the correlation parameter. Journal of Statistical Planning and Inference, 76(1–2):145–161, February 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898001803.

Chaubey:1999:SEM

[CS99b] Yogendra P. Chaubey and Pranab K. Sen. On smooth estimation of mean residual life. Journal of Statistical Planning and Inference, 75(2):223–236, January 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837589800144X.

Chen:1999:PTM

[CS99c] Pai-Lien Chen and Pranab Kumar Sen. A piecewise transition model for analyzing multistate life history data. Journal of Statistical Planning and Inference, 78(1–2):385–400, May 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898002262.

Csaki:1993:DFK

[Csà93] E. Csàki. A discrete Feynman–Kac formula. Journal of Statistical Planning and Inference, 34(1):63–73, January 1993. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375893900344.

Csaki:1996:SRB

[Csá96] Endre Csáki. On some results for Bernoulli excursions. Journal of Statistical Planning and Inference, 54(1):45–54, September 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375895001557.

Carbon:1996:HLP

[CT96] Michel Carbon and Lanh Tat Tran. On histograms for linear processes. Journal of Statistical Planning and Inference, 53(3):403– 419, August 30, 1996. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375895001417.

Cheng:1998:KSW

[CT98] Bing Cheng and Howell Tong. K-stationarity and wavelets. Journal of Statistical Planning and Inference, 68(1):129–144, May 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375897001389.

Christensen:1992:TNL

[CU92] Ronald Christensen and Jessica Utts. Testing for nonadditivity in log-linear and logit models. Journal of Statistical Planning and Inference, 33(3):333–343, December 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290003B.

Cutler:1993:EBDa

[Cut93a] D. Richard Cutler. Efficient block designs for comparing test treatments to a control when the errors are correlated. Journal of Statistical Planning and Inference, 36(1):107–125, July 1993. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390106G.

Cutler:1993:EBDb

[Cut93b] D. Richard Cutler. Efficient block designs for comparing test treatments to a control when the errors are correlated. Journal of Statistical Planning and Inference, 37(3):393–412, December 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589390115M.

Csorgo:1992:CBP

[CV92] Sándor Csörgő and László Viharos. Confidence bands for percentile residual lifetimes. Journal of Statistical Planning and Inference, 30(3):327–337, March 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290159P.

Cook:1993:RPF

[CW93] R. Dennis Cook and Sanford Weisberg. Residual plots in up to four dimensions. Journal of Statistical Planning and Inference, 36(2–3):141–149, August/September 1993. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390119Q.

Cabrera:1997:SMM

[CW97a] J. Cabrera and G. S. Watson. Simulation methods for mean and median bias reduction in parametric estimation. Journal of Statistical Planning and Inference, 57(1):143–152, January 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375897811500.

Coad:1997:ACI

[CW97b] D. S. Coad and M. B. Woodroofe. Approximate confidence intervals after a sequential clinical trial comparing two exponential survival curves with censoring. Journal of Statistical Planning and Inference, 63(1):79–96, September 30, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896002042.

Chi:1998:CTU

[CW98] E. M. Chi and S. Weerahandi. Comparing treatments under growth curve models: Exact tests using generalized p-values. Journal of Statistical Planning and Inference, 71(1–2):179–189, August 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375898000810.

Constantine:1998:AFA

[CX98a] G. Constantine and L. Xue. Aliasing in 2^{n−k} fractions in the absence of within group interactions. Journal of Statistical Planning and Inference, 66(2):345–361, March 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897000840.

Constantine:1998:AFC

[CX98b] Gregory Constantine and Lianyong Xue. Aliasing in 2^{n−k} fractions in the case of a separation of factors. Journal of Statistical Planning and Inference, 72(1–2):121–132, September 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375898000275.

Chen:1993:RTD

[CXL93] H. J. Chen, M. Xiong, and K. Lam. Range tests for the dispersion of several location parameters. Journal of Statistical Planning and Inference, 36(1):15–25, July 1993. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390098Q.

Czado:1994:BIB

[Cza94] Claudia Czado. Bayesian inference of binary regression models with parametric link. Journal of Statistical Planning and Inference, 41(2):121–140, September 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894901589.

Czado:1997:SPL

[Cza97] Claudia Czado. On selecting parametric link transformation families in generalized linear models. Journal of Statistical Planning and Inference, 61(1):125–139, May 30, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896001504.

Das:1993:ADF

[Das93] Sanghamitra Das. The asymptotic distribution of the fixed effects estimator for nonlinear regression. Journal of Statistical Planning and Inference, 35(3):269–277, June 1993. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390015X.

DasGupta:1994:BAR

[Das94] Anirban DasGupta. Bounds on asymptotic relative efficiencies of robust estimates of locations for contaminations by scale mixtures. Journal of Statistical Planning and Inference, 41(1):73– 93, August 1994. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375894901554.

Datta:1995:MOE

[Dat95] Somnath Datta. A minimax optimal estimator for continuous monotone densities. Journal of Statistical Planning and Inference, 46(2):183–193, August 1, 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589400100A.

Datta:1996:APB

[Dat96] Somnath Datta. On asymptotic properties of bootstrap for AR(1) processes. Journal of Statistical Planning and Inference, 53(3):361–374, August 30, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375895001476.

David:1994:BBD

[Dav94] O. David. Balanced block designs under interactive linear models. Journal of Statistical Planning and Inference, 39(1):33–41, April 1, 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375894901104.

Dawid:1997:CNI

[Daw97] A. P. Dawid. Comments on “non-informative priors do not exist”. Journal of Statistical Planning and Inference, 65(1):178–180, December 1, 1997. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375897900690.

Dey:1999:IBD

[DBG99] Aloke Dey, K. Balasubramanian, and Sudhir Gupta. Incomplete block designs for slope ratio assays. Journal of Statistical Planning and Inference, 78(1–2):369–383, May 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898002250.

Dimitrov:1995:DCS

[DCK95] B. Dimitrov, S. Chukova, and Z. Khalil. Definitions, characterizations and structural properties of probability distributions similar to the exponential. Journal of Statistical Planning and Inference, 43(1–2):271–287, January 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589400026R.

Das:1992:UON

[DD92a] Ashish Das and A. Dey. Universal optimality and nonoptimality of some row-column designs. Journal of Statistical Planning and Inference, 31(2):263–271, May 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290035Q.

Dufour:1992:NMR

[DD92b] Jean-Marie Dufour and Marcel G. Dagenais. Nonlinear models, rescaling and test invariance. Journal of Statistical Planning and Inference, 32(1):111–135, July 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290155L.

Delampady:1994:BRM

[DD94] Mohan Delampady and Dipak K. Dey. Bayesian robustness for multiparameter problems. Journal of Statistical Planning and Inference, 40(2–3):375–382, July 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894901325.

Datta:1999:EBL

[DDB99] Gauri Sankar Datta, Bannmo Day, and Ishwar Basawa. Empirical best linear unbiased and empirical Bayes prediction in multivariate small area estimation. Journal of Statistical Planning and Inference, 75(2):269–279, January 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898001475.

Das:1998:DNM

[DDN98] A. Das, A. M. Dean, and W. I. Notz. Designs with nearly minimal number of observations and flexible blocking. Journal of Statistical Planning and Inference, 72(1–2):133–147, September 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375898000287.

Dean:1993:RPP

[Dea93] C. B. Dean. A robust property of pseudo-likelihood estimation for count data. Journal of Statistical Planning and Inference, 35(3): 309–317, June 1993. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375893900193.

Deheuvels:1992:FLI

[Deh92a] Paul Deheuvels. A functional law of the iterated logarithm for tail quantile processes. Journal of Statistical Planning and Inference, 32(1):63–74, July 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290152I.

Deheuvels:1992:LDP

[Deh92b] Paul Deheuvels. Large deviations by Poisson approximations. Journal of Statistical Planning and Inference, 32(1):75–88, July 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589290153J.

Dempster:1990:CS

[Dem90] A. P. Dempster. Causality and statistics. Journal of Statistical Planning and Inference, 25(3):261–278, July 1990. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375890900767.

Dean:1995:CUE

[DEM95] C. B. Dean, D. M. Eaves, and C. J. Martinez. A comment on the use of empirical covariance matrices in the analysis of count data. Journal of Statistical Planning and Inference, 48(2):197–205, November 15, 1995. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589400155O.

Dennis:1996:BAT

[Den96] Samuel Y. Dennis. A Bayesian analysis of tree-structured statistical decision problems. Journal of Statistical Planning and Inference, 53(3):323–344, August 30, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895001123.

Dennis:1998:BAR

[Den98] Samuel Y. Dennis III. A Bayesian analysis of ratios of proportions in tree-structured experiments. Journal of Statistical Planning and Inference, 74(2):293–322, November 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898000986.

Desmond:1997:OEF

[Des97a] A. F. Desmond. Optimal estimating functions, quasi-likelihood and statistical modelling. Journal of Statistical Planning and Inference, 60(1):77–104, April 30, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375896001231.

See discussion [Les97, Car97, Nel97, Wir97, LR97] and rejoinder [Des97b].

Desmond:1997:R

[Des97b] A. F. Desmond. Rejoinder. Journal of Statistical Planning and Inference, 60(1):116–121, April 30, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897900161. See [Des97a].

Dette:1991:NRD

[Det91] Holger Dette. A note on robust designs for polynomial regression. Journal of Statistical Planning and Inference, 28(2):223– 232, June 1991. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589190027C.

Dette:1993:MOC

[Det93] Holger Dette. On a mixture of the D— and D _{1}-optimality criterion in polynomial regression. Journal of Statistical Planning and Inference, 35(2):233–249, May 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390047A.

Dette:1994:RDM

[Det94] Holger Dette. Robust designs for multivariate polynomial regression on the d-cube. Journal of Statistical Planning and Inference, 38(1):105–124, January 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837589200153U.

Dippon:1994:SAG

[DF94] Jürgen Dippon and Václav Fabian. Stochastic approximation of global minimum points. Journal of Statistical Planning and Inference, 41(3):327–347, October 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894900272.

Dion:1995:ACS

[DF95] Jean-Pierre Dion and René Ferland. Absolute continuity, singular measures and asymptotics for estimators. Journal of Statistical Planning and Inference, 43(1–2):235–242, January 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589400023O.

Datta:1991:AOH

[DG91] Gauri Sankar Datta and Malay Glosh. Asymptotic optimality of hierarchical Bayes estimators and predictors. Journal of Statistical Planning and Inference, 29(3):229–243, November 1991. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589190001U.

Datta:1995:HBE

[DG95] Gauri Sankar Datta and Malay Ghosh. Hierarchical Bayes estimators of the error variance in one-way ANOVA models. Journal of Statistical Planning and Inference, 45(3):399–411, June 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589300083T.

Deza:1998:OMA

[DG98] Michel Deza and Viatcheslav Grishukhin. Once more about 80 Steiner triple systems on 15 points. Journal of Statistical Planning and Inference, 72(1–2):165–184, September 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898000305.

Dharmadhikari:1997:NPO

[DGK97] A. D. Dharmadhikari, J. K. Ghosh, and M. M. Kuber. On NBUE property of one component system supported by an inactive standby and a repair facility. Journal of Statistical Planning and Inference, 64(1):1–8, October 30, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896002169.

Draper:1993:RVS

[DGP93] Norman R. Draper, Norbert Gaffke, and Friedrich Pukelsheim. Rotatability of variance surfaces and moment matrices. Journal of Statistical Planning and Inference, 36(2–3):347–355, August/September 1993. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589390136T.

Dey:1997:OGL

[DGP97] Dipak K. Dey, Alan E. Gelfand, and Fengchun Peng. Overdispersed generalized linear models. Journal of Statistical Planning and Inference, 64(1):93–107, October 30, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896002078.

Dufour:1992:SEB

[DH92] Jean-Marie Dufour and Marc Hallin. Simple exact bounds for distributions of linear signed rank statistics. Journal of Statistical Planning and Inference, 31(3):311–333, June 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290140N.

deHaan:1993:EHQ

[dHR93] Laurens de Haan and Holger Rootzén. On the estimation of high quantiles. Journal of Statistical Planning and Inference, 35(1): 1–13, April 1993. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589390063C.

DiBucchianico:1999:CCA

[Di 99] A. Di Bucchianico. Combinatorics, computer algebra and the Wilcoxon-Mann-Whitney test. Journal of Statistical Planning and Inference, 79(2):349–364, July 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898002614.

Diebolt:1995:NTR

[Die95] Jean Diebolt. A nonparametric test for the regression function: Asymptotic theory. Journal of Statistical Planning and Inference, 44(1):1–17, March 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837589400045W.

Dikta:1998:SRC

[Dik98] Gerhard Dikta. On semiparametric random censorship models. Journal of Statistical Planning and Inference, 66(2):253– 279, March 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375897000918.

Ding:1996:SID

[Din96] Changmin Ding. Supports of indecomposable B(4, 2; ν) designs. Journal of Statistical Planning and Inference, 51(2):157–161, April 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375895000798.

Davis:1997:NHD

[DJ97] James A. Davis and Jonathan Jedwab. Nested Hadamard difference sets. Journal of Statistical Planning and Inference, 62(1): 13–20, July 1997. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375896001620.

Donner:1994:CRT

[DK94] Allan Donner and Neil Klar. Cluster randomization trials in epidemiology: theory and application. Journal of Statistical Planning and Inference, 42(1–2):37–56, November 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894901880.

Kang:1997:LST

[dK97] Qing de Kang. Large sets of three types of oriented triple systems. Journal of Statistical Planning and Inference, 58(1):151– 176, March 1, 1997. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375896000675.

Das:1999:SBM

[DK99] Shubhabrata Das and Alok Krishen. Some bootstrap methods in nonlinear mixed-effect models. Journal of Statistical Planning and Inference, 75(2):237–245, January 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898001451.

Kang:1996:LSH

[dKgL96] Qing de Kang and Jian guo Lei. On large sets of hybrid triple systems. Journal of Statistical Planning and Inference, 51(2):

181–188, April 1996. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375895000828.

Dykstra:1997:SCP

[DKR97] Richard Dykstra, Subhash Kochar, and Javier Rojo. Stochastic comparisons of parallel systems of heterogeneous exponential components. Journal of Statistical Planning and Inference, 65(2):203–211, December 15, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837589700058X.

Demetrovics:1998:DTP

[DKS98] János Demetrovics, Gyula O. H. Katona, and Attila Sali. Design type problems motivated by database theory. Journal of Statistical Planning and Inference, 72(1–2):149–164, September 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375898000299.

Draper:1990:CCT

[DL90] Norman R. Draper and Dennis K. J. Lin. Capacity considerations for two-level fractional factorial designs. Journal of Statistical Planning and Inference, 24(1):25–35, January 1990. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589090014L. See note [RP97].

Deng:1998:NMB

[DL98a] Lih-Yuan Deng and Dennis K. J. Lin. A note on Morris’ bound for search designs. Journal of Statistical Planning and Inference, 74(1):203–204, October 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898000998.

DiBucchianico:1998:DRS

[DL98b] A. Di Bucchianico and D. E. Loeb. Dominance refinements of the Smirnov two-sample test. Journal of Statistical Planning and Inference, 66(1):51–60, January 5, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897000815.

Dean:1999:NCD

[DLC99] Angela M. Dean, Susan M. Lewis, and Jane Y. Chang. Nested changeover designs. Journal of Statistical Planning and Inference, 77(2):337–351, March 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898001979.

delaHorra:1994:BAU

[dlHF94] Julian de la Horra and Carmen Fernandez. Bayesian analysis under ϵ-contaminated priors: a trade-off between robustness and precision. Journal of Statistical Planning and Inference, 38(1): 13–30, January 1994. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589200152T.

deLima:1999:TDC

[dLS99] Antonio C. Pedroso de Lima and Pranab K. Sen. Time-dependent coefficients in a multi-event model for survival analysis. Journal of Statistical Planning and Inference, 75(2):393–414, January 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375898001578.

Drees:1991:CTS

[DM91] Holger Drees and Hartmut Milbrodt. Components of the two-sided Kolmogorov–Smirnov test in signal detection problems with Gaussian white noise. Journal of Statistical Planning and Inference, 29(3):325–335, November 1991. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375891900072.

Dale:1993:PAT

[DM93] M. R. T. Dale and J. W. Moon. The permuted analogues of three Catalan sets. Journal of Statistical Planning and Inference, 34(1): 75–87, January 1993. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375893900355.

DasGupta:1994:USP

[DM94] Anirban DasGupta and Saurabh Mukhopadhyay. Uniform and subuniform posterior robustness: the sample size problem. Journal of Statistical Planning and Inference, 40(2–3):189–204, July 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375894901201.

Deheuvels:1995:NLE

[DM95] Paul Deheuvels and David M. Mason. Nonstandard local empirical processes indexed by sets. Journal of Statistical Planning and Inference, 45(1–2):91–112, ???? 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894000654.

Dette:1996:SRG

[DM96a] Holger Dette and Axel Munk. Sign regularity of a generalized Cauchy kernel with applications. Journal of Statistical Planning and Inference, 52(2):131–142, June 15, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895001085.

Dharmapala:1996:EMP

[DM96b] Dhammika Dharmapala and Michael McAleer. Econometric methodology and the philosophy of science. Journal of Statistical Planning and Inference, 49(1):9–37, January 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895000283.

Donatos:1996:MCS

[DM96c] George S. Donatos and George C. Michailidis. A Monte Carlo study of some limited and full information simultaneous equation estimators with normal and nonnormal autocorrelated disturbances. Journal of Statistical Planning and Inference, 50(2): 273–282, March 1996. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375895000585.

Dyckerhoff:1997:OOD

[DM97] Rainer Dyckerhoff and Karl Mosler. Orthant orderings of discrete random vectors. Journal of Statistical Planning and Inference, 62(2):193–205, August 15, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837589600184X.

deMedicis:1993:UCAb

[dML93a] Anne de Médicis and Pierre Leroux. A unified combinatorial approach for q- (and p, q-) Stirling numbers, J. Statist. Plann. Inference 34 (1993) 89–105. Journal of Statistical Planning and Inference, 35(2):267, May 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390049C.

deMedicis:1993:UCAa

[dML93b] Anne de Médicis and Pierre Leroux. A unified combinatorial approach for q- (and p, q-) Stirling numbers. Journal of Statistical Planning and Inference, 34(1):89–105, January 1993. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375893900366.

Deshpande:1999:TTS

[DMNN99] J. V. Deshpande, M. Mukhopadhyay, and U. V. Naik-Nimbalkar. Testing of two sample proportional intensity assumption for nonhomogeneous Poisson processes. Journal of Statistical Planning and Inference, 81(2):237–251, November 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375899000221.

Dasgupta:1992:CDH

[DMS92] Anirban Dasgupta, Saurabh Mukhopadhyay, and William J. Studden. Compromise designs in heteroscedastic linear models. Journal of Statistical Planning and Inference, 32(3):363– 384, September 1992. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589290017M.

Deheuvels:1994:LLR

[DN94] Paul Deheuvels and Valery B. Nevzorov. Limit laws for K-record times. Journal of Statistical Planning and Inference, 38(3):279– 307, March 1994. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375894900124.

Dette:1996:BOO

[DN96] Holger Dette and H.-M. Neugebauer. Bayesian optimal one point designs for one parameter nonlinear models. Journal of Statistical Planning and Inference, 52(1):17–31, June 1, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895001042.

Dette:1997:BOD

[DN97] Holger Dette and H.-M. Neugebauer. Bayesian D-optimal designs for exponential regression models. Journal of Statistical Planning and Inference, 60(2):331–349, May 15, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896001310.

Dourleijn:1996:CSS

[Dou96] Johan Dourleijn. A case study of subset selection in sugar beet breeding. Journal of Statistical Planning and Inference, 54(3):323–344, September 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837589500176X.

Durham:1991:PSS

[DP91] S. D. Durham and W. J. Padgett. A probabilistic stress-strength model and its application to fatigue failure in gun barrels. Journal of Statistical Planning and Inference, 29(1–2):67–74, September/October 1991. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375892901229.

Denker:1992:ANT

[DP92] Manfred Denker and Madan L. Puri. Asymptotic normality of two sample linear rank statistics under U-statistic structure. Journal of Statistical Planning and Inference, 32(1):89–110, July 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589290154K.

delPino:1997:IGF

[dP97] Guido E. del Pino. Interactions for general factors. Journal of Statistical Planning and Inference, 65(1):145–158, December 1, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375897800018.

Draper:1998:MMB

[DP98] Norman R. Draper and Friedrich Pukelsheim. Mixture models based on homogeneous polynomials. Journal of Statistical Planning and Inference, 71(1–2):303–311, August 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898000123.

Draper:1999:KOS

[DP99] Norman R. Draper and Friedrich Pukelsheim. Kiefer ordering of simplex designs for first- and second-degree mixture models. Journal of Statistical Planning and Inference, 79(2):325–348, July 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375898002638.

Degenhardt:1996:CWC

[DPSvZ96] H. J. A. Degenhardt, Madan L. Puri, Shan Sun, and Martien C. A. van Zuijlen. Characterization of weak convergence for smoothed empirical and quantile processes under φ-mixing. Journal of Statistical Planning and Inference, 53(3):285–295, August 30, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375895001336.

Durairajan:1996:RLM

[DR96] T. M. Durairajan and K. J. Raman. Robustness of locally most powerful invariant test for normal mixture model in control and treatment populations. Journal of Statistical Planning and Inference, 52(1):33–41, June 1, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895001050.

delRioz:1997:NED

[dR97] Alejandro Quintela del Ríoz. Nonparametric estimation of density derivatives of dependent data. Journal of Statistical Planning and Inference, 61(1):155–174, May 30, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896001413.

Draper:1997:BRR

[Dra97] N. R. Draper. Book review: Response surface methodology: Process and product optimization using designed experiments: R. H. Myers and D. C. Montgomery, (Wiley, New York, 1995, $59.95, ISBN: 0-471-58100-3, pp. 714). Journal of Statistical Planning and Inference, 59(1):185–186, March 30, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837589781631X.

Drees:1998:GCE

[Dre98] Holger Drees. A general class of estimators of the extreme value index. Journal of Statistical Planning and Inference, 66(1):95– 112, January 5, 1998. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375897000761.

Druilhet:1999:ONB

[Dru99] Pierre Druilhet. Optimality of neighbour balanced designs. Journal of Statistical Planning and Inference, 81(1):141–152, October 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S037837589900004X.

Datta:1997:MBA

[DS97] Somnath Datta and T. N. Sriram. A modified bootstrap for autoregression without stationarity. Journal of Statistical Planning and Inference, 59(1):19–30, March 30, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896000924.

DeSantis:1998:BFH

[DS98] Fulvio De Santis and Fulvio Spezzaferri. Bayes factors and hierarchical models. Journal of Statistical Planning and Inference, 74(2):323–342, November 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898001098.

DasGupta:1999:NGI

[DS99] Anirban DasGupta and Bimal K. Sinha. A new general interpretation of the Stein estimate and how it adapts: Applications. Journal of Statistical Planning and Inference, 75(2):247– 268, January 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375898001463.

dosSantos:1995:NHV

[dSDF95] Dirley M. dos Santos, Richard B. Davies, and Brian Francis. Nonparametric hazard versus nonparametric frailty distribution in modelling recurrence of breast cancer. Journal of Statistical Planning and Inference, 47(1–2):111–127, October 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589400125F.

Dehling:1991:BSS

[DT91] Herold Dehling and Murad S. Taqqu. Bivariate symmetric statistics of long-range dependent observations. Journal of Statistical Planning and Inference, 28(2):153–165, June 1991. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375891900238.

Du:1996:ODL

[Du96] B. Du. Orthogonal diagonal Latin squares with small subsquares. Journal of Statistical Planning and Inference, 51(2): 163–170, April 1996. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375895000801.

Durairajan:1992:OEF

[Dur92] T. M. Durairajan. Optimal estimating function for nonorthogonal model. Journal of Statistical Planning and Inference, 33(3):381–384, December 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837589290006E.

Dutta:1996:EQF

[Dut96] K. Dutta. Estimation of quadratic functions of the Bernoulli parameter in inverse sampling method. Journal of Statistical Planning and Inference, 52(2):215–224, June 15, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895001131.

DasGupta:1997:SSP

[DV97] Anirban DasGupta and Brani Vidakovic. Sample size problems in ANOVA Bayesian point of view. Journal of Statistical Planning and Inference, 65(2):335–347, December 15, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897000566.

Dwivedi:1995:P

[Dwi95] T. D. Dwivedi. Preface. Journal of Statistical Planning and Inference, 43(1–2):vii–viii, January 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895900217.

Dion:1994:SIB

[DY94a] J. P. Dion and N. M. Yanev. Statistical inference for branching processes with an increasing random number of ancestors. Journal of Statistical Planning and Inference, 39(2):329–351, April 15, 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375894902127.

Draper:1994:NSR

[DY94b] Norman R. Draper and Lisa H. Ying. A note on slope rotatability over all directions. Journal of Statistical Planning and Inference, 41(1):113–119, August 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894901570.

Elliott:1996:EHM

[EA96] Robert J. Elliott and Lakhdar Aggoun. Estimation for hidden Markov random fields. Journal of Statistical Planning and Inference, 50(3):343–351, March 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895000623.

Eaton:1997:AQR

[Eat97] Morris L. Eaton. Admissibility in quadratically regular problems and recurrence of symmetric Markov chains: Why the connection? Journal of Statistical Planning and Inference, 64(2):231–247, November 15, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375897000372.

Ebrahimi:1997:TWL

[Ebr97] Nader Ebrahimi. Testing whether lifetime distribution is decreasing uncertainty. Journal of Statistical Planning and Inference, 64(1):9–19, October 30, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375896002157.

El-Bassiouni:1996:MHM

[EBS96] M. Y. El-Bassiouni and J. F. Seely. A modified harmonic mean test procedure for variance components. Journal of Statistical Planning and Inference, 49(3):319–326, February 1, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895000186.

Ellenberg:1994:SID

[ED94] Susan S. Ellenberg and Dennis O. Dixon. Statistical issues in designing clinical trials of AIDS treatments and vaccines. Journal of Statistical Planning and Inference, 42(1–2):123–135, November 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375894901937.

Eckmann:1997:CDL

[EEK97] Alain Eckmann, Shalom Eliahou, and Michel Kervaire. Compressible difference lists an appendix to ‘A note on the equation θ\ − νgq = n + λΣ’. Journal of Statistical Planning and Inference, 62(1):35–38, July 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896001632.

Einmahl:1992:LTT

[Ein92] John H. J. Einmahl. Limit theorems for tail processes with application to intermediate quantile estimation. Journal of Statistical Planning and Inference, 32(1):137–145, July 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290156M.

Einmahl:1993:CBQ

[Ein93] John H. J. Einmahl. Confidence bands for the quantile function under random censoring. Journal of Statistical Planning and Inference, 36(1):69–75, July 1993. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390102C.

Eisele:1994:DAB

[Eis94] Jeffrey R. Eisele. The doubly adaptive biased coin design for sequential clinical trials. Journal of Statistical Planning and Inference, 38(2):249–261, February 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894900388.

Eccleston:1993:SRC

[EJW93] J. A. Eccleston, J. A. John, and D. Whitaker. Some rowcolumn designs with adjusted orthogonality. Journal of Statistical Planning and Inference, 36(2–3):323–330, August/September 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589390134R.

Eliahou:1997:NE

[EK97] Shalom Eliahou and Michel Kervaire. A note on the equation θ\−gq = n+λΣ. Journal of Statistical Planning and Inference, 62 (1):21–34, July 1997. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375896001644.

Ekstrom:1998:CMS

[Eks98] Magnus Ekström. On the consistency of the maximum spacing method. Journal of Statistical Planning and Inference, 70(2):209– 224, July 15, 1998. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375897001857.

ElBarmi:1996:ELR

[El 96] Hammou El Barmi. Empirical likelihood ratio test for or against a set of inequality constraints. Journal of Statistical Planning and Inference, 55(2):191–204, October 30, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895001883.

Eubank:1993:TNE

[EL93] R. L. Eubank and V. N. LaRiccia. Testing for no effect in nonparametric regression. Journal of Statistical Planning and Inference, 36(1):1–14, July 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390097P.

Elnaggar:1999:IPT

[EM99] Mohamed Elnaggar and Arunava Mukherjea. Identification of the parameters of a trivariate normal vector by the distribution of the minimum. Journal of Statistical Planning and Inference, 78(1– 2):23–37, May 1999. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375898002043.

El-Moalem:1998:NRI

[EMS98] Habib El-Moalem and Pranab K. Sen. Nonparametric recovery of interblock information in clinical trials with a surrogate endpoint. Journal of Statistical Planning and Inference, 72(1–2):185–205, September 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375898000317.

Ennadifi:1995:SAN

[Enn95] Gratiane Ennadifi. Strong approximation of the number of renewal paced record times. Journal of Statistical Planning and Inference, 45(1–2):113–132, ???? 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894000662.

ElBarmi:1999:EWM

[ER99] Hammou El Barmi and Mark D. Rothmann. Estimation of weighted multinomial probabilities under log-convex constraints. Journal of Statistical Planning and Inference, 81(1):1–11, October 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899000105.

Erdos:1998:PMS

[Erd98] Paul Erdős. The probability method: Successes and limitations. Journal of Statistical Planning and Inference, 72(1–2):207–213, September 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375898000329.

Erkanli:1997:BMA

[Erk97] Alaattin Erkanli. Boundary-mode approximations for posterior expectations. Journal of Statistical Planning and Inference, 58(2):217–239, March 15, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375896000894.

Erkanli:1998:OBT

[ESA98] Alaattin Erkanli, Refik Soyer, and Adrian Angold. Optimal Bayesian two-phase designs. Journal of Statistical Planning and Inference, 66(1):175–191, January 5, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837589700075X.

Escobar:1995:NBM

[Esc95] Michael D. Escobar. Nonparametric Bayesian methods in hierarchical models. Journal of Statistical Planning and Inference, 43(1–2):97–106, January 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837589400011J.

Etemadi:1997:CRS

[ESV97] N. Etemadi, T. N. Sriram, and A. N. Vidyashankar. L_{p} convergence of reciprocals of sample means with applications to sequential estimation in linear regression. Journal of Statistical Planning and Inference, 65(1):1–15, December 1, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897000463.

Eubank:1999:SSS

[Eub99] R. L. Eubank. A simple smoothing spline, II. Journal of Statistical Planning and Inference, 81(2):229–235, November 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S037837589900021X.

Evans:1996:MRM

[Eva96] Anthony B. Evans. Magic rectangles and modular magic rectangles. Journal of Statistical Planning and Inference, 51(2):171–180, April 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589500081X.

Falk:1992:BOB

[Fal92] Michael Falk. Bootstrap optimal bandwidth selection for kernel density estimates. Journal of Statistical Planning and Inference, 30(1):13–22, January 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837589290103Y.

Farmakis:1991:COW

[Far91] Nikos Farmakis. Constructions of A-optimal weighing designs when n = 19. Journal of Statistical Planning and Inference, 27(2):249–261, February 1991. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837589190020F.

Farmakis:1992:COW

[Far92] Nikos Farmakis. On constructibility of A-optimal weighing designs (n, k, 5), when n ≡ 3 (mod 4) and k = n − 1, n. Journal of Statistical Planning and Inference, 33(2):275–283, November 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375892900732.

Farebrother:1996:RCP

[Far96] R. W. Farebrother. The role of chaotic processes in econometric models. Journal of Statistical Planning and Inference, 49(2):163– 176, January 1996. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375895000356.

Fong:1994:SNP

[FCA94] Duncan K. H. Fong, Mosuk Chow, and James H. Albert. Selecting the normal population with the best regression value — a Bayesian approach. Journal of Statistical Planning and Inference, 40(1):97–111, June 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894901449.

Fedou:1993:ESF

[Fed93] J. M. Fedou. Enumeration of skew Ferrers diagrams and basic Bessel functions. Journal of Statistical Planning and Inference, 34(1):107–123, January 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375893900377.

Feigin:1994:IME

[Fei94] Paul D. Feigin. Inference and martingale estimating equations for stochastic processes on a semigroup. Journal of Statistical Planning and Inference, 39(2):239–254, April 15, 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894902070.

Ferger:1994:CPE

[Fer94] Dietmar Ferger. Change-point estimators in case of small disorders. Journal of Statistical Planning and Inference, 40(1):33–49, June 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375894901406.

Ferreira:1995:ERN

[Fer95] Helena Ferreira. Extremes of a random number of variables from periodic sequences. Journal of Statistical Planning and Inference, 45(1–2):133–141, ???? 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894000824.

Fernholz:1997:RVS

[Fer97] Luisa Turrin Fernholz. Reducing the variance by smoothing. Journal of Statistical Planning and Inference, 57(1):29–38, January 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S037837589600033X.

Foutz:1997:RET

[FF97] Robert V. Foutz and Patricia B. Foutz. Richard A. Epstein: The theory of gambling and statistical logic, revised edition, (Academic Press, San Diego: 1995). Journal of Statistical Planning and Inference, 60(1):193–194, April 30, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896001565.

Fu:1997:MSC

[FFR97] Chin-Mei Fu, Hung-Lin Fu, and C. A. Rodger. The minimum size of critical sets in latin squares. Journal of Statistical Planning and Inference, 62(2):333–337, August 15, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896001875.

Finner:1994:CSS

[FG94] H. Finner and G. Giani. Closed subset selection procedures for selecting good populations. Journal of Statistical Planning and Inference, 38(2):179–199, February 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894900345.

Finner:1996:DBM

[FG96] H. Finner and G. Giani. Duality between multiple testing and selecting. Journal of Statistical Planning and Inference, 54(2):201– 227, September 1996. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375895001689.

Fisher:1991:BAS

[FH91] Nicholas I. Fisher and Peter Hall. Bootstrap algorithms for small samples. Journal of Statistical Planning and Inference, 27(2):157– 169, February 1991. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375891900135.

Fu:1996:QSD

[FH96] Tung-Shan Fu and Tayuan Huang. Quasi semisymmetric designs with extremal conditions. Journal of Statistical Planning and Inference, 51(3):261–271, May 1, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895000909.

Fuji-Hara:1996:COB

[FHKM96] Ryoh Fuji-Hara, Shinji Kuriki, and Maki Miyake. Cyclic orthogonal and balanced arrays. Journal of Statistical Planning and Inference, 56(2):171–180, December 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837589600016X.

Field:1997:RRS

[Fie97] C. A. Field. Robust regression and small sample confidence intervals. Journal of Statistical Planning and Inference, 57(1):39– 48, January 1997. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375896000341.

Finkelstein:1997:CAD

[Fin97] M. S. Finkelstein. The concealed age of distribution functions and the problem of general repair. Journal of Statistical Planning and Inference, 65(2):315–321, December 15, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897000578.

Fienberg:1998:CRC

[FK98] Stephen E. Fienberg and Sung-Ho Kim. Calibration and refinement for classification trees. Journal of Statistical Planning and Inference, 70(2):241–254, July 15, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837589700195X.

Follmann:1991:IFM

[FL91] Dean A. Follmann and Diane Lambert. Identifiability of finite mixtures of logistic regression models. Journal of Statistical Planning and Inference, 27(3):375–381, March 1991. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589190050O.

Fedorov:1998:ODM

[FMN98] Valery V. Fedorov, Grace Montepiedra, and Christopher J. Nachtsheim. Optimal design and the model validity range. Journal of Statistical Planning and Inference, 72(1–2):215–227, September 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375898000330.

Fedorov:1999:DEL

[FMN99] Valerii V. Fedorov, Grace Montepiedra, and Christopher J. Nachtsheim. Design of experiments for locally weighted regression. Journal of Statistical Planning and Inference, 81 (2):363–382, November 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S037837589900018X.

Foutz:1997:BRI

[Fou97] Robert V. Foutz. Book review: Introduction to statistical time series: Wayne A. Fuller, (2nd edition published by Wiley, New York, 1996, 698 pp. ISBN 0-471-55239-9). Journal of Statistical Planning and Inference, 61(1):191–192, May 30, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897000025.

Foucart:1999:SIC

[Fou99] Thierry Foucart. Stability of the inverse correlation matrix. Partial ridge regression. Journal of Statistical Planning and Inference, 77(1):141–154, February 15, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898001955.

Follmann:1994:EEB

[FP94] Dean Follmann and Michael Proschan. The effect of estimation and biasing strategies on selection bias in clinical trials with permuted blocks. Journal of Statistical Planning and Inference, 39(1):1–17, April 1, 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894901082.

Futschik:1998:LRT

[FP98] Andreas Futschik and Georg Ch. Pflug. The likelihood ratio test for simple tree order: a useful asymptotic expansion. Journal of Statistical Planning and Inference, 70(1):57–68, July 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897001705.

Fortini:1994:CFB

[FR94] Sandra Fortini and Fabrizio Ruggeri. Concentration functions and Bayesian robustness. Journal of Statistical Planning and Inference, 40(2–3):205–220, July 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589490121X.

Finner:1999:ACC

[FR99] H. Finner and M. Roters. Asymptotic comparison of the critical values of step-down and step-up multiple comparison procedures. Journal of Statistical Planning and Inference, 79(1):11–30, June 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S037837589800233X.

Frees:1991:TSE

[Fre91] Edward W. Frees. Trimmed slope estimates for simple linear regression. Journal of Statistical Planning and Inference, 27(2):203–221, February 1991. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375891900168.

From:1990:ACL

[Fro90] Steven G. From. An application of Cheng’s lemma for minimizing the asymptotic variance of best asymptotically normal estimators based on sample quantiles. Journal of Statistical Planning and Inference, 26(3):363–370, November 1990. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589090138K.

Fernandez:1998:SEI

[FRS98] M. A. Fernández, C. Rueda, and B. Salvador. Simultaneous estimation by isotonic regression. Journal of Statistical Planning and Inference, 70(1):111–119, July 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897001833.

Freedman:1994:ESC

[FS94] Benjamin Freedman and Stanley H. Shapiro. Ethics and statistics in clinical research: towards a more comprehensive examination. Journal of Statistical Planning and Inference, 42(1–2):223– 240, November 1994. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375894901988.

Faraggi:1995:MLN

[FS95a] David Faraggi and Richard Simon. The maximum likelihood neural network as a statistical classification model. Journal of Statistical Planning and Inference, 46(1):93–104, July 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895990682.

Fisher:1995:SFS

[FS95b] Gordon Fisher and Ah Boon Sim. Some finite sample theory for bootstrap regression estimates. Journal of Statistical Planning and Inference, 43(1–2):289–300, January 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589400027S.

Floresroux:1996:NME

[FS96] Ernesto M. Floresroux and Michael L. Stein. A new method of edge correction for estimating the nearest neighbor distribution. Journal of Statistical Planning and Inference, 50(3):353– 371, March 1996. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375895000631.

Fosam:1997:ELL

[FS97] E. B. Fosam and D. N. Shanbhag. An extended Laha–Lukacs characterization result based on a regression property. Journal of Statistical Planning and Inference, 63(2):173–186, October 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375897000098.

Fuh:1992:BAC

[Fuh92] C. D. Fuh. Bootstrapping the autocorrelation coefficient of finite Markov chains. Journal of Statistical Planning and Inference, 32(3):291–302, September 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290012H.

Fujikoshi:1993:AES

[Fuj93] Yasunori Fujikoshi. Asymptotic expansions for the standardized and Studentized estimates in the growth curve model. Journal of Statistical Planning and Inference, 36(2–3):165–173, August/ September 1993. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589390121L.

Futschik:1999:CRS

[Fut99] A. Futschik. Confidence regions for the set of global maximizers of nonparametrically estimated curves. Journal of Statistical Planning and Inference, 82(1–2):237–250, December 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375899000464.

Fotopoulos:1993:ANR

[FW93] S. B. Fotopoulos and M. C. Wang. Asymptotic normality for random sums of linear processes. Journal of Statistical Planning and Inference, 34(3):349–360, March 1993. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390145V.

Fang:1998:BCA

[FZ98a] Yue Fang and Bin Zhou. Block-crossed arrays with applications to robust design. Journal of Statistical Planning and Inference, 74(1):169–175, October 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898001013.

Francq:1998:ELR

[FZ98b] Christian Francq and Jean-Michel Zakoïan. Estimating linear representations of nonlinear processes. Journal of Statistical Planning and Inference, 68(1):145–165, May 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897001390.

Gupta:1995:RSW

[GA95] Ramesh C. Gupta and H. Olcay Akman. On the reliability studies of a weighted inverse Gaussian model. Journal of Statistical Planning and Inference, 48(1):69–83, November 1, 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589400148O.

Gao:1999:SRU

[GA99] J. T. Gao and V. V. Anh. Semiparametric regression under long-range dependent errors. Journal of Statistical Planning and Inference, 80(1–2):37–57, August 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898002419.

Gastaldi:1993:IMV

[Gas93] Tommaso Gastaldi. Interpolating missing values in a censored sample. Journal of Statistical Planning and Inference, 35(2): 189–199, May 1993. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375893900447.

Gasparini:1996:NBE

[Gas96a] Mauro Gasparini. Nonparametric Bayes estimation of a distribution function with truncated data. Journal of Statistical Planning and Inference, 55(3):361–369, November 8, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375895001956.

Ghosh:1996:ECD

[GAS96b] Subir Ghosh and Walid S. Al-Sabah. Efficient composite designs with small number of runs. Journal of Statistical Planning and Inference, 53(1):117–132, August 1, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375896001437.

Gan:1995:DOS

[GB95] Gaoxiong Gan and Lee J. Bain. Distribution of order statistics for discrete parents with applications to censored sampling. Journal of Statistical Planning and Inference, 44(1):37–46, March 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375895927815.

Gan:1998:SRT

[GB98] Gaoxiong Gan and Lee J. Bain. Some results for type I censored sampling from geometric distributions. Journal of Statistical Planning and Inference, 67(1):85–97, March 16, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897000967.

Goutis:1995:IIS

[GC95] Constantinos Goutis and George Casella. Improved invariant set estimation for general-scale families. Journal of Statistical Planning and Inference, 44(3):327–340, ???? 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894000573.

Ginebra:1997:TK

[GC97] Josep Ginebra and Murray K. Clayton. On a theorem of Kelley. Journal of Statistical Planning and Inference, 65(2):191–201, December 15, 1997. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837589700061X.

Ginebra:1999:SSP

[GC99] Josep Ginebra and Murray K. Clayton. Small-sample performance of Bernoulli two-armed bandit Bayesian strategies. Journal of Statistical Planning and Inference, 79(1):107–122, June 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375898002304.

Garcia-Escudero:1999:ATM

[GEGM99] L. A. García-Escudero, A. Gordaliza, and C. Matrán. Asymptotics for trimmed k-means and associated tolerance zones. Journal of Statistical Planning and Inference, 77(2):247–262, March 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375898001967.

Gerencser:1994:RPS

[Ger94] László Gerencsér. On Rissanen’s predictive stochastic complexity for stationary ARMA processes. Journal of Statistical Planning and Inference, 41(3):303–325, October 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894900264.

Gessel:1993:CPY

[Ges93] Ira M. Gessel. Counting paths in Young’s lattice. Journal of Statistical Planning and Inference, 34(1):125–134, January 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375893900388.

Giani:1991:SGR

[GF91] Guido Giani and Helmut Finner. Some general results on least favorable parameter configurations with special reference to equivalence testing and the range statistic. Journal of Statistical Planning and Inference, 28(1):33–47, May 1991. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589190057L.

Gil:1992:FEO

[GG92] María Angeles Gil and Pedro Gil. Fuzziness in the experimental outcomes: comparing experiments and removing the loss of information. Journal of Statistical Planning and Inference, 31(1): 93–111, April 1992. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589290043R.

Giles:1996:RHP

[GG96] Judith A. Giles and David E. A. Giles. Risk of a homoscedasticity pre-test estimator of the regression scale under LINEX loss. Journal of Statistical Planning and Inference, 50(1):21–35, February 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375895000437.

Garvan:1999:PPD

[GG99] Cynthia Wilson Garvan and Malay Ghosh. On the property of posteriors for dispersion models fn1. Journal of Statistical Planning and Inference, 78(1–2):229–241, May 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837589800216X.

Gui:1995:CIR

[GGBT95] Rongde Gui, Franklin A. Graybill, Richard K. Burdick, and Naitee Ting. Confidence intervals on ratios of linear combinations for non-disjoint sets of expected mean squares. Journal of Statistical Planning and Inference, 48(2):215–227, November 15, 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589400152L.

Ghysels:1998:KAT

[GGJ98] Eric Ghysels, Christian Gouriéroux, and Joanna Jasiak. Kernel autocorrelogram for time-deformed processes. Journal of Statistical Planning and Inference, 68(1):167–191, May 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897001407.

Ghosal:1999:CSB

[GGR99] Subhashis Ghosal, Jayanta K. Ghosh, and R. V. Ramamoorthi. Consistent semiparametric Bayesian inference about a location parameter. Journal of Statistical Planning and Inference, 77(2):181–193, March 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S037837589800192X.

Gupta:1997:MPD

[GGT97] Pushpa L. Gupta, Ramesh C. Gupta, and Ram C. Tripathi. On the monotonic properties of discrete failure rates. Journal of Statistical Planning and Inference, 65(2):255–268, December 15, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375897000645.

Gupta:1990:SII

[GH90] Shanti S. Gupta and Deng-Yuan Huang. Selecting important independent variables in linear regression models. Journal of Statistical Planning and Inference, 24(2):269, February 1990. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589090047X.

Gupta:1993:SSB

[GH93] Shanti S. Gupta and Sayaji N. Hande. Single-sample Bayes and empirical Bayes rules for ranking and estimating multinomial probabilities. Journal of Statistical Planning and Inference, 35(3): 367–382, June 1993. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589390023Y.

Gaffke:1995:COA

[GH95] Norbert Gaffke and Berthold Heiligers. Computing optimal approximate invariant designs for cubic regression on multidimensional balls and cubes. Journal of Statistical Planning and Inference, 47(3):347–376, October 15, 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589400140Q.

Gombay:1996:ATC

[GH96a] Edit Gombay and Lajos Horváth. Approximations for the time of change and the power function in change-point models. Journal of Statistical Planning and Inference, 52(1):43–66, June 1, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375895000259.

Gupta:1996:DID

[GH96b] Shanti S. Gupta and Deng-Yuan Huang. On detecting influential data and selecting regression variables. Journal of Statistical Planning and Inference, 53(3):421–435, August 30, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895001034.

Gaffke:1998:MSI

[GH98a] Norbert Gaffke and Berthold Heiligers. Minimum support invariant designs for multiple cubic regression. Journal of Statistical Planning and Inference, 72(1–2):229–245, September 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375898000342.

Gombay:1998:RBE

[GH98b] Edit Gombay and Marie Husková. Rank based estimators of the change-point. Journal of Statistical Planning and Inference, 67(1):137–154, March 16, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897000992.

Grechanovsky:1999:CPB

[GH99] Eugene Grechanovsky and Yosef Hochberg. Closed procedures are better and often admit a shortcut. Journal of Statistical Planning and Inference, 76(1–2):79–91, February 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898001256.

Ghorpade:1996:YBL

[Gho96] Sudhir R. Ghorpade. Young bitableaux, lattice paths and Hilbert functions. Journal of Statistical Planning and Inference, 54(1):55–66, September 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375895001565.

Ghosh:1997:NIP

[Gho97] J. K. Ghosh. Non-informative priors do not exist — discussion of a discussion. Journal of Statistical Planning and Inference, 65(1):180–182, December 1, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897900719.

Gardiner:1999:CAC

[GHRG^{+}99] Joseph Gardiner, Margaret Holmes-Rovner, John Goddeeris, David Rovner, and Joel Kupersmith. Covariate-adjusted costeffectiveness ratios. Journal of Statistical Planning and Inference, 75(2):291–304, January 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375898001499.

Giudici:1995:BFZ

[Giu95] Paolo Giudici. Bayes factors for zero partial covariances. Journal of Statistical Planning and Inference, 46(2):161–174, August 1, 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589400101Z.

Golan:1996:RIC

[GJ96] Amos Golan and George Judge. Recovering information in the case of underdetermined problems and incomplete economic data. Journal of Statistical Planning and Inference, 49(1):127–136, January 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589500033X.

Ghosh:1998:SSA

[GJ98a] Kaushik Ghosh and S. Rao Jammalamadaka. Small sample approximations for spacing statistics. Journal of Statistical Planning and Inference, 69(2):245–261, June 15, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837589700164X.

Githinji:1998:DCO

[GJ98b] F. Githinji and M. Jacroux. On the determination and construction of optimal designs for comparing a set of test treatments with a set of controls in the presence of a linear trend. Journal of Statistical Planning and Inference, 66(1):161–174, January 5, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375897000670.

Guerre:1998:GVA

[GJ98c] E. Guerre and F. Jouneau. Geometric versus arithmetic random walk the case of trended variables. Journal of Statistical Planning and Inference, 68(2):203–220, May 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897001420.

Ghosh:1990:DME

[GL90] Subir Ghosh and Eric S. Lagergren. Dispersion models and estimation of dispersion effects in replicated factorial experiments. Journal of Statistical Planning and Inference, 26(3):253– 262, November 1990. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589090129I.

Gupta:1994:EBS

[GL94] Shanti S. Gupta and TaChen Liang. On empirical Bayes selection rules for sampling inspection. Journal of Statistical Planning and Inference, 38(1):43–64, January 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589200154V.

Lei:1996:FRT

[gL96] Jian guo Lei. Further results on transitive Kirkman triple systems. Journal of Statistical Planning and Inference, 51(2):189– 194, April 1996. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375895000836.

Goncalves:1998:SSR

[GL98a] Esmeralda Gonçalves and NazaréMendes Lopes. Some statistical results on autoregressive conditionally heteroscedastic models. Journal of Statistical Planning and Inference, 68(1):193–202, May 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375897001419.

Griffith:1998:QLB

[GL98b] Daniel A. Griffith and Francesco Lagona. On the quality of likelihood-based estimators in spatial autoregressive models when the data dependence structure is misspecified. Journal of Statistical Planning and Inference, 69(1):153–174, June 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897001560.

Gupta:1998:SLC

[GL98c] Shanti S. Gupta and Ta Chen Liang. Simultaneous lower confidence bounds for probabilities of correct selections. Journal of Statistical Planning and Inference, 72(1–2):279–290, September 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375898000378.

Ganju:1999:DRR

[GL99a] Jitendra Ganju and James M. Lucas. Detecting randomization restrictions caused by factors. Journal of Statistical Planning and Inference, 81(1):129–140, October 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375899000026.

Ghosh:1999:OMD

[GL99b] Subir Ghosh and Thomas Liu. Optimal mixture designs for four components in two orthogonal blocks. Journal of Statistical Planning and Inference, 78(1–2):219–228, May 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898002158.

Gupta:1999:EBS

[GL99c] Shanti S. Gupta and TaChen Liang. On empirical Bayes simultaneous selection procedures for comparing normal populations with a standard. Journal of Statistical Planning and Inference, 77(1):73–88, February 15, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898001773.

Gupta:1996:SSH

[GLL96] Shanthi S. Gupta, Lii-Yuh Leu, and TaChen Liang. Simultaneous selection for homogeneous multinomial populations based on entropy function: an empirical Bayes approach. Journal of Statistical Planning and Inference, 54(2):145–157, September 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895001654.

Gerami:1998:EBD

[GLMN98] A. Gerami, S. M. Lewis, D. Majumdar, and W. I. Notz. Efficient block designs for comparing dual with single treatments. Journal of Statistical Planning and Inference, 72(1–2):247–263, September 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375898000354.

Giraitis:1996:CPP

[GLS96] Liudas Giraitis, Remigijus Leipus, and Donatas Surgailis. The change-point problem for dependent observations. Journal of Statistical Planning and Inference, 53(3):297–310, August 30, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375895001484.

Glynn:1997:CRT

[Gly97] David G. Glynn. On cone representations of translation planes. Journal of Statistical Planning and Inference, 58(1):33–41, March 1, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375896000584.

Glynn:1998:SCR

[Gly98] David G. Glynn. A survey of cone representations. Journal of Statistical Planning and Inference, 72(1–2):265–277, September 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375898000366.

George:1993:OIP

[GM93a] Edward I. George and Robert McCulloch. On obtaining invariant prior distributions. Journal of Statistical Planning and Inference, 37(2):169–179, November 1993. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390086L.

Gupta:1993:CSE

[GM93b] Shanti S. Gupta and Klaus J. Miescke. On combining selection and estimation in the search for the largest binomial parameter. Journal of Statistical Planning and Inference, 36(2–3):129–140, August/September 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390118P.

Gupta:1996:BLA

[GM96] Shanti S. Gupta and Klaus J. Miescke. Bayesian look ahead one-stage sampling allocations for selection of the best population. Journal of Statistical Planning and Inference, 54(2):229– 244, September 1996. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375895001697.

Gangopadhyay:1999:ESD

[GMD99] A. K. Gangopadhyay, B. K. Mallick, and D. G. T. Denison. Estimation of spectral density of a stationary time series via an asymptotic representation of the periodogram. Journal of Statistical Planning and Inference, 75(2):281–290, January 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375898001487.

Gourieroux:1996:TSG

[GMR96] Christian Gouriéroux, Alain Monfort, and Eric Renault. Twostage generalized moment method with applications to regressions with heteroscedasticity of unknown form. Journal of Statistical Planning and Inference, 50(1):37–63, February 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895000445.

Ghosh:1999:HBG

[GNWK99] Malay Ghosh, Kannan Natarajan, Lance A. Waller, and Dalho Kim. Hierarchical Bayes GLMs for the analysis of spatial data: an application to disease mapping. Journal of Statistical Planning and Inference, 75(2):305–318, January 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898001505.

Gupta:1997:CMD

[GNZ97] Arjun K. Gupta, Truc T. Nguyen, and Wei-Bin Zeng. Characterization of multivariate distributions through a functional equation of their characteristic functions. Journal of Statistical Planning and Inference, 63(2):187–201, October 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897000104.

Godambe:1990:C

[God90] V. P. Godambe. Comments A. Journal of Statistical Planning and Inference, 25(2):195, June 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375890900664.

Godfrey:1996:MTT

[God96] L. G. Godfrey. Misspecification tests and their uses in econometrics. Journal of Statistical Planning and Inference, 49(2):241– 260, January 1996. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375895000399.

Good:1990:RCG

[Goo90a] I. J. Good. A reply to the comment by G. A. Barnard. Journal of Statistical Planning and Inference, 25(3):439–440, July 1990. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589090087B. See [Bar90b].

Good:1990:SCF

[Goo90b] I. J. Good. Speculations concerning the future of statistics. Journal of Statistical Planning and Inference, 25(3):441–466, July 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589090088C.

Good:1990:SDF

[Goo90c] I. J. Good. Statistical discussion forum. Journal of Statistical Planning and Inference, 26(1):109–125, September 1990. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

0378375890901009. See comment [Jos96c].

Good:1996:FQA

[Goo96a] I. J. Good. F44A. A question about F44. Journal of Statistical Planning and Inference, 55(2):258, October 30, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

S0378375896900671. See [Jos96c] and reply [Jos96d].

Good:1996:FVR

[Goo96b] I. J. Good. F46. Varieties of the regression fallacy. Journal of Statistical Planning and Inference, 55(2):260, October 30, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375896888495.

Good:1996:FCF

[Goo96c] I. J. Good. F49A. A comment on F49. Journal of Statistical Planning and Inference, 55(2):264, October 30, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896900695.

Govindarajulu:1993:BRR

[Gov93] Z. Govindarajulu. A brief review of the role of lattices in rank order statistics. Journal of Statistical Planning and Inference, 34(1):135–149, January 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375893900399.

Gulati:1994:SNE

[GP94] Sneh Gulati and W. J. Padgett. Smooth nonparametric estimation of the hazard and hazard rate functions from record-breaking data. Journal of Statistical Planning and Inference, 42(3):331– 341, December 1994. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589400093X.

Giovagnoli:1998:PFD

[GP98] Alessandra Giovagnoli and Nicolò Pintacuda. Properties of frequency distributions induced by general ‘up-and-down’ methods for estimating quantiles. Journal of Statistical Planning and Inference, 74(1):51–63, October 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898000767.

Grechanovsky:1999:GAS

[GP99] Eugene Grechanovsky and Ilia Pinsker. A general approach to stepup multiple test procedures for free-combinations families. Journal of Statistical Planning and Inference, 82(1–2):35–54, December 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375899000300.

Gutierrez-Pena:1998:ASM

[GPS98] E. Gutierrez-Peña and A. F. M. Smith. Aspects of smoothing and model inadequacy in generalized regression. Journal of Statistical Planning and Inference, 67(2):273–286, April 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897001110.

Gather:1990:OBP

[GR90] U. Gather and B. Rauhut. The outlier behaviour of probability distributions. Journal of Statistical Planning and Inference, 26(2):237–252, October 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837589090128H.

Gross:1993:CTI

[GR93] Kenneth I. Gross and Donald St. P. Richards. Constant term identities and hypergeometric functions on spaces of Hermitian matrices. Journal of Statistical Planning and Inference, 34(1):151–158, January 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837589390040D.

Gupta:1997:IPS

[GR97] Rameshwar D. Gupta and Donald St. P. Richards. Invariance properties of some classical tests for exponentiality. Journal of Statistical Planning and Inference, 63(2):203–213, October 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375897000165.

Gray:1998:SVK

[GR98] Brenton D. Gray and Colin Ramsay. On the spectrum of [ v, k, t ] trades. Journal of Statistical Planning and Inference, 69(1): 1–19, June 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375897001237.

Gross:1998:NCL

[Gro98a] Jürgen Groß. A note on the concepts of linear and quadratic sufficiency. Journal of Statistical Planning and Inference, 70(1): 69–76, July 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375897001778.

Gross:1998:CLR

[Gro98b] Jürgen Gross. On contractions in linear regression. Journal of Statistical Planning and Inference, 74(2):343–351, November 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375898000044.

Gupta:1999:WEB

[GRP99] V. K. Gupta, D. V. V. Ramana, and Rajender Parsad. Weighted A-efficiency of block designs for making treatment-control and treatment-treatment comparisons. Journal of Statistical Planning and Inference, 77(2):301–319, March 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898001876.

Golomb:1997:CEC

[GS97a] Solomon W. Golomb and Hong-Yeop Song. A conjecture on the existence of cyclic Hadamard difference sets. Journal of Statistical Planning and Inference, 62(1):39–41, July 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896001656.

Gupta:1997:NSD

[GS97b] A. K. Gupta and D. Song. L_{p}-norm spherical distribution. Journal of Statistical Planning and Inference, 60(2):241–260, May 15, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375896001292.

Giraitis:1999:CLT

[GS99a] Liudas Giraitis and Donatas Surgailis. Central limit theorem for the empirical process of a linear sequence with long memory. Journal of Statistical Planning and Inference, 80(1–2):81– 93, August 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375898002432.

Glaz:1999:SCI

[GS99b] Joseph Glaz and Cristina P. Sison. Simultaneous confidence intervals for multinomial proportions. Journal of Statistical Planning and Inference, 82(1–2):251–262, December 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375899000476.

Gut:1999:SRR

[GS99c] Allan Gut and Rainer Schwabe. Some remarks on repeated significance tests for linear contrasts. Journal of Statistical Planning and Inference, 82(1–2):25–34, December 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375899000294.

Garcia-Soidan:1997:EEN

[GSGMPS97] Pilar H. García-Soidán, Wenceslao González-Manteiga, and JoséM. Prada-Sánchez. Edgeworth expansions for nonparametric distribution estimation with applications. Journal of Statistical Planning and Inference, 65(2):213–231, December 15, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897000591.

Gnot:1992:BIQ

[GSZ92] Stanislaw Gnot, Jolanta Srzednicka, and Roman Zmyślony. Bayes invariant quadratic estimation in general linear regression models. Journal of Statistical Planning and Inference, 30(2):223– 236, February 1992. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375892900835.

Ghosh:1993:MEP

[GT93] Subir Ghosh and Hooshang Talebi. Main effect plans with an additional search property for 2^{m} factorial experiments. Journal of Statistical Planning and Inference, 36(2–3):367–383, August/September 1993. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589390138V.

Gather:1995:SIO

[GT95] Ursula Gather and Jim Tomkins. On stability of intermediate order statistics. Journal of Statistical Planning and Inference, 45(1–2):175–183, ???? 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375894000891.

Greenshtein:1997:SIE

[GT97] Eitan Greenshtein and Erik Torgersen. Statistical information and expected number of observations for sequential experiments. Journal of Statistical Planning and Inference, 59(2):229–240, April 15, 1997. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375896001140.

Giraitis:1999:CNQ

[GT99] Liudas Giraitis and Murad S. Taqqu. Convergence of normalized quadratic forms. Journal of Statistical Planning and Inference, 80(1–2):15–35, August 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898002407.

Gross:1998:NSC

[GTL98] Jürgen Groß, Götz Trenkler, and Erkki P. Liski. Necessary and sufficient conditions for superiority of misspecified restricted least squares regression estimator. Journal of Statistical Planning and Inference, 71(1–2):109–116, August 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898000779.

Guess:1991:ESC

[GUH91] Frank M. Guess, John S. Usher, and Thom J. Hodgson. Estimating system and component reliabilities under partial information on cause of failure. Journal of Statistical Planning and Inference, 29(1–2):75–85, September/October 1991. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290123A.

Guiard:1996:DDC

[Gui96] V. Guiard. Different definitions of Δ-correct selection for the indifference zone formulation. Journal of Statistical Planning and Inference, 54(2):175–199, September 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895001670.

Guillou:1999:EWB

[Gui99a] A. Guillou. Efficient weighted bootstraps for the mean. Journal of Statistical Planning and Inference, 77(1):11–35, February 15, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375898001694.

Guillou:1999:WBV

[Gui99b] A. Guillou. Weighted bootstraps for the variance. Journal of Statistical Planning and Inference, 81(1):113–120, October 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899000087.

Gupta:1990:ARS

[Gup90] Shanti S. Gupta. Acknowledgement of referees’ services. Journal of Statistical Planning and Inference, 26(1):127–128, September 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589090101Y.

Gupta:1991:ARS

[Gup91] Shanti S. Gupta. Acknowledgement of referees’ services. Journal of Statistical Planning and Inference, 28(1):137, May 1991. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589190066N.

Gupta:1992:ARS

[Gup92] Shanti S. Gupta. Acknowledgement of referees’ services. Journal of Statistical Planning and Inference, 30(1):133–134, January 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375892901137.

Gupta:1998:CMF

[Gup98] Sudhir Gupta. A class of multi-factor designs for test versus control comparisons. Journal of Statistical Planning and Inference, 72(1–2):291–302, September 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837589800038X.

Gutmair:1993:PSM

[Gut93] Susanne Gutmair. Polars and subgradients of mixtures of information functions. Journal of Statistical Planning and Inference, 35(1):93–112, April 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837589390070M.

Goldstein:1994:RMB

[GW94] Michael Goldstein and David A. Wooff. Robustness measures for Bayes linear analyses. Journal of Statistical Planning and Inference, 40(2–3):261–277, July 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894901252.

Hafner:1998:EHF

[Haf98] Christian M. Hafner. Estimating high-frequency foreign exchange rate volatility with nonparametric ARCH models. Journal of Statistical Planning and Inference, 68(2):247–269, May 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897001444.

Hagwood:1992:CPI

[Hag92] Charles Hagwood. The calibration problem as an ill-posed inverse problem. Journal of Statistical Planning and Inference, 31(2): 179–185, May 1992. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589290028Q.

Haiman:1992:SIP

[Hai92] George Haiman. A strong invariance principle for the extremes of multivariate stationary m-dependent sequences. Journal of Statistical Planning and Inference, 32(2):147–163, August 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290046U.

Haiman:1995:P

[Hai95] G. Haiman. Preface. Journal of Statistical Planning and Inference, 45(1–2):5–6, ???? 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375895800018.

Hallin:1998:P

[Hal98] Marc Hallin. Preface. Journal of Statistical Planning and Inference, 68(1):3–4, May 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375897001328.

Hamada:1991:CUI

[Ham91] Michael Hamada. The costs of using incomplete exponential data. Journal of Statistical Planning and Inference, 27(3):317– 324, March 1991. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589190044F.

Hampel:1997:WCF

[Ham97] Frank Hampel. What can the foundations discussion contribute to data analysis? And what may be some of the future directions in robust methods and data analysis? Journal of Statistical Planning and Inference, 57(1):7–19, January 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896000316.

Hamada:1998:NSQ

[Ham98] Noboru Hamada. On the nonexistence of some quaternary linear codes meeting the Griesmer bound. Journal of Statistical Planning and Inference, 72(1–2):303–322, September 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898000391.

Harel:1992:WIM

[Har92] Michel Harel. Weak invariance of the multidimensional rank statistic with unbounded scores for nonstationary absolutely regular processes. Journal of Statistical Planning and Inference, 32(2):197–227, August 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837589290048W.

Harvill:1999:TTS

[Har99] Jane L. Harvill. Testing time series linearity via goodnessof-fit methods. Journal of Statistical Planning and Inference, 75(2):331–341, January 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375898001529.

Hauser:1999:MLE

[Hau99] Michael A. Hauser. Maximum likelihood estimators for ARMA and ARFIMA models: a Monte Carlo study. Journal of Statistical Planning and Inference, 80(1–2):229–255, August 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898002523.

Hawkins:1992:DSF

[Haw92] D. L. Hawkins. Detecting shifts in functions of multivariate location and covariance parameters. Journal of Statistical Planning and Inference, 33(2):233–244, November 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375892900709.

Hayakawa:1994:THM

[Hay94a] Takesi Hayakawa. Test of homogeneity of multiple parameters. Journal of Statistical Planning and Inference, 38(3):351– 357, March 1994. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375894900159.

Hayter:1994:SPT

[Hay94b] Anthony J. Hayter. On the selection probabilities of two-stage decision procedures. Journal of Statistical Planning and Inference, 38(2):223–236, February 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375894900361.

Hayakawa:1998:BPD

[Hay98] Yu Hayakawa. A Bayesian positive dependence of survival times based on the multivariate arrangement increasing property. Journal of Statistical Planning and Inference, 70(2):225–240, July 15, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375897001936.

Hazelton:1999:OLB

[Haz99] Martin L. Hazelton. An optimal local bandwidth selector for kernel density estimation. Journal of Statistical Planning and Inference, 77(1):37–50, February 15, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898001700.

Hwang:1993:PER

[HB93] Sun Young Hwang and I. V. Basawa. Parameter estimation in a regression model with random coefficient autoregressive errors. Journal of Statistical Planning and Inference, 36(1):57–67, July 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589390101B.

Hwang:1994:LSI

[HB94] Sun Young Hwang and I. V. Basawa. Large sample inference for conditional exponential families with applications to nonlinear time series. Journal of Statistical Planning and Inference, 38(2):141–157, February 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894900329.

Hilgers:1995:ODM

[HB95] Ralf-Dieter Hilgers and Peter Bauer. Optimal designs for mixture amount experiments. Journal of Statistical Planning and Inference, 48(2):241–246, November 15, 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589500003R.

Hoef:1998:CFM

[HB98a] Jay M. Ver Hoef and Ronald Paul Barry. Constructing and fitting models for cokriging and multivariable spatial prediction. Journal of Statistical Planning and Inference, 69(2):275–294, June 15, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375897001626.

Hwang:1998:PEG

[HB98b] S. Y. Hwang and I. V. Basawa. Parameter estimation for generalized random coefficient autoregressive processes. Journal of Statistical Planning and Inference, 68(2):323–337, May 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837589700147X.

Hommel:1999:BPL

[HB99] G. Hommel and G. Bernhard. Bonferroni procedures for logically related hypotheses. Journal of Statistical Planning and Inference, 82(1–2):119–128, December 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837589900035X.

Huang:1992:EEM

[HC92] Wen-Tao Huang and Huann-Sheng Chen. Estimation of the exponential mean under type I censored sampling. Journal of Statistical Planning and Inference, 33(2):187–196, November 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375892900662.

Huang:1993:NEC

[HC93] Wen-Tao Huang and Yi-Ping Chang. Nonparametric estimation in change-point models. Journal of Statistical Planning and Inference, 35(3):335–347, June 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390021W.

He:1993:ESM

[He93] Kun He. The estimation of stratum means vector with random sample sizes. Journal of Statistical Planning and Inference, 37(1): 43–49, October 1993. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589390078K.

Heiligers:1991:AED

[Hei91] Berthold Heiligers. Admissibility of experimental designs in linear regression with constant term. Journal of Statistical Planning and Inference, 28(1):107–123, May 1991. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589190064L.

Heiligers:1992:AED

[Hei92] Berthold Heiligers. Admissible experimental designs in multiple polynomial regression. Journal of Statistical Planning and Inference, 31(2):219–233, May 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290031M.

Heiligers:1993:LBM

[Hei93] Berthold Heiligers. Linear Bayes and minimax estimation in linear models with partially restricted parameter space. Journal of Statistical Planning and Inference, 36(2–3):175–183, August/ September 1993. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589390122M.

Heiligers:1996:COP

[Hei96] Berthold Heiligers. Computing E-optimal polynomial regression designs. Journal of Statistical Planning and Inference, 55(2):219– 233, October 30, 1996. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375895001905.

Heiligers:1998:ODP

[Hei98] Berthold Heiligers. E-optimal designs for polynomial spline regression. Journal of Statistical Planning and Inference, 75 (1):159–172, November 15, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375898001335.

Hendry:1996:TLD

[Hen96] David F. Hendry. Typologies of linear dynamic systems and models. Journal of Statistical Planning and Inference, 49(2):177– 201, January 1996. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375895000364.

Heyde:1992:NDI

[Hey92] C. C. Heyde. New developments in inference for temporal stochastic processes. Journal of Statistical Planning and Inference, 33(1):121–129, October 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837589290097C.

Heyde:1996:UQL

[Hey96] C. C. Heyde. On the use of quasi-likelihood for estimation in hidden-Markov random fields. Journal of Statistical Planning and Inference, 50(3):373–378, March 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589500064X.

Hjort:1995:SOA

[HF95] Nils Lid Hjort and Grete Fenstad. Second-order asymptotics for the number of times an estimator is more than ϵ from its target value. Journal of Statistical Planning and Inference, 48(3):261– 275, December 1995. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589500008W.

Hamada:1993:CSM

[HH93a] Noboru Hamada and Tor Helleseth. A characterization of some {v_{2}+2v_{3}, v_{1}+2v_{2}; k−1, 3}-minihypers and some (v_{k−30}, k, 3^{k−1}− 21; 3)-codes meeting the Griesmer bound. Journal of Statistical Planning and Inference, 34(3):387–402, March 1993. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390147X.

Huang:1993:MRL

[HH93b] Mong-Na Lo Huang and Ming-Chun Hsu. Marginally restricted linear-optimal designs. Journal of Statistical Planning and Inference, 35(2):251–266, May 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390048B.

Hamada:1996:CSMa

[HH96a] Noboru Hamada and Tor Helleseth. A characterization of some 3v_{2}+v_{3}, 3v_{1}+v_{2}; 3, 3-minihypers and some [15, 4, 9; 3]-codes with B_{2} = 0. Journal of Statistical Planning and Inference, 56(1):129– 146, December 1996. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375896000146.

Hamada:1996:UCN

[HH96b] Noboru Hamada and Tor Helleseth. The uniqueness of [87,5,57; 3]-codes and the nonexistence of [258,6,171; 3]-codes. Journal of Statistical Planning and Inference, 56(1):105–127, December 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375896000134.

Hettmansperger:1994:MDE

[HHH94] T. P. Hettmansperger, I. Hueter, and J. Hüsler. Minimum distance estimators. Journal of Statistical Planning and Inference, 41(3):291–302, October 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375894900256.

Hinkelmann:1990:P

[Hin90] K. Hinkelmann. Preface. Journal of Statistical Planning and Inference, 25(3):209–210, July 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375890900723.

Hinkelmann:1998:LTF

[HJ98] Klaus Hinkelmann and Jinnam Jo. Linear trend-free Box– Behnken designs. Journal of Statistical Planning and Inference, 72(1–2):347–354, September 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375898000421.

Hadi:1995:URS

[HJL95] Ali S. Hadi, Wendell D. Jones, and Robert F. Ling. A unifying representation of some case-deletion influence measures in univariate and multivariate linear regression. Journal of Statistical Planning and Inference, 46(1):123–135, July 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589400094C.

Hallin:1999:NTI

[HJPZ99] Marc Hallin, Jana Jurecková, Jan Picek, and Toufik Zahaf. Nonparametric tests of independence of two autoregressive time series based on autoregression rank scores. Journal of Statistical Planning and Inference, 75(2):319–330, January 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898001517.

Hyodo:1994:AVB

[HK94] Yoshifumi Hyodo and Masahide Kuwada. Analysis of variance of balanced fractional s^{m} factorial designs of resolution V_{p,q}. Journal of Statistical Planning and Inference, 38(2):263–277, February 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375894900396.

Horvath:1997:ELR

[HK97] Lajos Horváth and Piotr Kokoszka. The effect of long-range dependence on change-point estimators. Journal of Statistical Planning and Inference, 64(1):57–81, October 30, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837589600208X.

Holzmann:1998:EPS

[HKL98] W. H. Holzmann, H. Kharaghani, and M. T. Lavassani. The excess problem and some excess inequivalent matrices of order 32. Journal of Statistical Planning and Inference, 72(1–2):381–391, September 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375898000445.

Haiman:1995:EMS

[HKP95] Goerge Haiman, Maxime Kiki, and Madan L. Puri. Extremes of Markov sequences. Journal of Statistical Planning and Inference, 45(1–2):185–201, ???? 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894000700.

Huang:1994:SVS

[HL94] Deng-Yuan Huang and Kuo-Chuan Liu. Some variable selection procedures in multivariate linear regression models. Journal of Statistical Planning and Inference, 41(2):205–214, September 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375894901643.

Huang:1995:SLC

[HL95] Deng-Yuan Huang and Ren Fen Lee. Selecting the largest capability index from several quality control processes. Journal of Statistical Planning and Inference, 46(3):335–346, August 15, 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589400114B.

Hallin:1990:DFT

[HLM90] Marc Hallin, Annie Laforet, and Guy Melard. Distribution-free tests against serial dependence: signed or unsigned ranks? Journal of Statistical Planning and Inference, 24(2):151–165, February 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589090038V.

Huda:1992:FRO

[HM92] S. Huda and Rahul Mukerjee. Further results on optimal designs with circular string property. Journal of Statistical Planning and Inference, 33(3):403–411, December 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290008G.

Hedayat:1995:GDS

[HM95] A. S. Hedayat and Dibyen Majumdar. Generating desirable sampling plans by the technique of trade-off in experimental design. Journal of Statistical Planning and Inference, 44(2):237– 247, April 1, 1995. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589400044V.

Hamada:1996:CSMb

[HM96] Noboru Hamada and Tomoko Maekawa. A characterization of some 3v_{1} + v_{3}, 3v_{0} + v_{2}; 3, 3-minihypers and its applications to error-correcting codes. Journal of Statistical Planning and Inference, 56(1):147–169, December 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896000158.

He:1997:SNB

[HM97] Kun He and Glen Meeden. Selecting the number of bins in a histogram: a decision theoretic approach. Journal of Statistical Planning and Inference, 61(1):49–59, May 30, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896001425.

Haynes:1997:RRS

[HMM97] Michele A. Haynes, H. L. MacGillivray, and K. L. Mengersen. Robustness of ranking and selection rules using generalised gand-k distributions. Journal of Statistical Planning and Inference, 65(1):45–66, December 1, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375897000505.

Hsu:1999:HTM

[HMT99] Y. S. Hsu, M. H. Metry, and Y. L. Tong. Hypotheses testing for means of dependent and heterogeneous normal random variables. Journal of Statistical Planning and Inference, 78(1–2): 89–99, May 1999. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375898002080.

Hsiao:1991:PSN

[HMW91] Ru-Rong Hsiao, Charles E. McCulloch, and Ruey-Beei Wu. The power of some nonparametric test statistics for the location and mean slippage problems. Journal of Statistical Planning and Inference, 28(3):391–403, July 1991. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589190075P.

He:1999:QSS

[HN99] Xuming He and Pin Ng. Quantile splines with several covariates. Journal of Statistical Planning and Inference, 75(2):343– 352, January 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375898001530.

Hoffmann:1995:AAL

[Hof95] Kurt Hoffmann. All admissible linear estimators of the regression parameter vector in the case of an arbitrary parameter subset. Journal of Statistical Planning and Inference, 48(3):371– 377, December 1995. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375895000163.

Hughes-Oliver:1999:PCM

[HOGF99] Jacqueline M. Hughes-Oliver and Graciela González-Farías. Parametric covariance models for shock-induced stochastic processes. Journal of Statistical Planning and Inference, 77(1):51–72, February 15, 1999. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375898001864.

Holm:1999:CSF

[Hol99] Sture Holm. Confidence sets of fixed size for multinormal means. Journal of Statistical Planning and Inference, 82(1–2):229–236, December 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375899000452.

Hoover:1990:SCA

[Hoo90] Donald R. Hoover. Subset complement addition upper bounds — an improved inclusion-exclusion method. Journal of Statistical Planning and Inference, 24(2):195–202, February 1990. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589090041R.

Hopfner:1997:TPE

[Höp97] Reinhard Höpfner. On tail parameter estimation in certain point process models. Journal of Statistical Planning and Inference, 60(1):169–187, April 30, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896001280.

Horvath:1995:KRA

[Hor95] Lajos Horváth. Kac’s representation from an asymptotic viewpoint. Journal of Statistical Planning and Inference, 46(1):27–45, July 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589400098G.

Horvath:1998:TCU

[Hor98] Lajos Horváth. Tests for changes under random censorship. Journal of Statistical Planning and Inference, 69(2):229–243, June 15, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375897001663.

Haiman:1990:SIP

[HP90a] George Haiman and Madan L. Puri. A strong invariance principle concerning the J-upper order statistics for stationary m-dependent sequences. Journal of Statistical Planning and Inference, 25(1):43–51, May 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589090006G.

Harel:1990:WCS

[HP90b] Michel Harel and Madan L. Puri. Weak convergence of the serial linear rank statistic with unbounded scores and regression constants under mixing conditions. Journal of Statistical Planning and Inference, 25(2):163–186, June 1990. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375890900642.

Harel:1991:LBO

[HP91] Michel Harel and Madan L. Puri. Limiting behavior of one sample rank order statistics with unbounded scores for nonstationary absolutely regular processes. Journal of Statistical Planning and Inference, 27(1):1–23, January 1991. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589190078S.

Hallin:1992:SAR

[HP92] Marc Hallin and Madan L. Puri. Some asymptotic results for a broad class of nonparametric statistics. Journal of Statistical Planning and Inference, 32(2):165–196, August 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290047V.

Hedayat:1997:DTL

[HP97] A. S. Hedayat and H. Pesotan. Designs for two-level factorial experiments with linear models containing main effects and selected two-factor interactions. Journal of Statistical Planning and Inference, 64(1):109–124, October 30, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896002091.

Harel:1998:AQR

[HP98] Michel Harel and Madan L. Puri. Autoregression quantiles and related rank score processes for generalized random coefficient autoregressive processes. Journal of Statistical Planning and Inference, 68(2):271–294, May 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897001456.

Hallin:1992:P

[HR92] Marc Hallin and Jean-Pierre Raoult. Preface. Journal of Statistical Planning and Inference, 32(1):3–4, July 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290148L.

Hochberg:1995:EMT

[HR95] Yosef Hochberg and Dror Rom. Extensions of multiple testing procedures based on Simes’ test. Journal of Statistical Planning and Inference, 48(2):141–152, November 15, 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589500005T.

Hubert:1997:RRB

[HR97] Mia Hubert and Peter J. Rousseeuw. Robust regression with both continuous and binary regressors. Journal of Statistical Planning and Inference, 57(1):153–163, January 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896000419.

Holcroft:1997:EER

[HRR97] Christina A. Holcroft, Andrea Rotnitzky, and James M. Robins. Efficient estimation of regression parameters from multistage studies with validation of outcome and covariates. Journal of Statistical Planning and Inference, 65(2):349–374, December 15, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375897817491.

Heiligers:1992:IAO

[HS92] Berthold Heiligers and Klaus Schneider. Invariant admissible and optimal designs in cubic regression on the v-ball. Journal of Statistical Planning and Inference, 31(1):113–125, April 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290044S.

Hardin:1993:NAC

[HS93] R. H. Hardin and N. J. A. Sloane. A new approach to the construction of optimal designs. Journal of Statistical Planning and Inference, 37(3):339–369, December 1993. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390112J.

Horvath:1995:LTU

[HS95]  Lajos Horváth and Qi-Man Shao.  Limit theorems for

the union-intersection test. Journal of Statistical Planning and Inference, 44(2):133–148, April 1, 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894000492.

Haupt:1997:TTN

[HS97] Elvira Haupt and Hendrik Schäbe. The TTT transformation and a new bathtub distribution model. Journal of Statistical Planning and Inference, 60(2):229–240, May 15, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897897108.

Hedayat:1998:SDC

[HS98a] A. S. Hedayat and John Stufken. Sampling designs to control selection probabilities of contiguous units. Journal of Statistical Planning and Inference, 72(1–2):333–345, September 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S037837589800041X.

Hirschfeld:1998:PPS

[HS98b] J. W. P. Hirschfeld and L. Storme. The packing problem in statistics, coding theory and finite projective spaces. Journal of Statistical Planning and Inference, 72(1–2):355–380, September 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375898000433.

Hedayat:1996:DSO

[HSS96] A. S. Hedayat, John Stufken, and Guoqin Su. On difference schemes and orthogonal arrays of strength t. Journal of Statistical Planning and Inference, 56(2):307–324, December 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896000262.

Hedayat:1997:MNF

[HSS97] A. Hedayat, Esther Seiden, and John Stufken. On the maximal number of factors and the enumeration of 3-symbol orthogonal arrays of strength 3 and index 2. Journal of Statistical Planning and Inference, 58(1):43–63, March 1, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896000596.

Hsu:1992:SMC

[Hsu92] Jason C. Hsu. Stepwise multiple comparisons with the best. Journal of Statistical Planning and Inference, 33(2):197–204, November 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375892900673.

Hayter:1991:SSD

[HT91] Anthony J. Hayter and Ajit C. Tamhane. Sample size determination for step-down multiple test procedures: Orthogonal contrasts and comparisons with a control. Journal of Statistical Planning and Inference, 27(3):271–290, March 1991. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589190042D.

Habib:1992:OES

[HT92a] M. K. Habib and A. Thavaneswaran. Optimal estimation for semimartingale neuronal models. Journal of Statistical Planning and Inference, 33(1):143–156, October 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290099E.

Huang:1992:DPD

[HT92b] Deng-Yuan Huang and Sheng-Tsaing Tseng. A decision procedure for determining the number of components in principal component analysis. Journal of Statistical Planning and Inference, 30(1):63–71, January 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375892901074.

Herrendorfer:1996:SB

[HT96] G. Herrendörfer and A. Tuchscherer. Selection and breeding. Journal of Statistical Planning and Inference, 54(3):307–321, September 1996. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375895001751.

Hardle:1998:NVA

[HTY98] W. Härdle, A. Tsybakov, and L. Yang. Nonparametric vector autoregression. Journal of Statistical Planning and Inference, 68(2): 221–245, May 1998. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375897001432.

Hu:1998:SCB

[Hu98] Inchi Hu. Strong consistency of Bayes estimates in nonlinear stochastic regression models. Journal of Statistical Planning and Inference, 67(1):155–163, March 16, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897000979.

Huang:1995:EBT

[Hua95] Su-Yun Huang. Empirical Bayes testing procedures in some nonexponential families using asymmetric Linex loss function. Journal of Statistical Planning and Inference, 46(3):293–309, August 15, 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375894001129.

Huang:1996:SPL

[Hua96] Deng-Yuan Huang. Selection procedures in linear models. Journal of Statistical Planning and Inference, 54(3):271–277, September 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375895001727.

Husler:1995:NEV

[Hüs95] J. Hüsler. A note on extreme values of locally stationary Gaussian processes. Journal of Statistical Planning and Inference, 45(1–2): 203–213, ???? 1995. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375895800026.

Huskova:1999:GCV

[Hus99] M. Husková. Gradual changes versus abrupt changes. Journal of Statistical Planning and Inference, 76(1–2):109–125, February 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375898001736.

Hutchinson:1996:FNE

[Hut96] T. P. Hutchinson. F49. Nonunique estimates of ability. Journal of Statistical Planning and Inference, 55(2):262–264, October 30, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375896888525.

Hall:1996:ADS

[HW96a] Peter Hall and Andrew T. A. Wood. Approximations to distributions of statistics used for testing hypotheses about the number of modes of a population. Journal of Statistical Planning and Inference, 55(3):299–317, November 8, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896000729.

Hamada:1996:NC

[HW96b] Noboru Hamada and Yoko Watamori. The nonexistence of [71,5,46;3]-codes. Journal of Statistical Planning and Inference, 52(3):379–394, July 1, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375895001263.

Hamada:1998:NTC

[HW98] Noboru Hamada and Yoko Watamori. The nonexistence of ternary [79,6,51] codes. Journal of Statistical Planning and Inference, 72(1–2):323–332, September 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898000408.

Habib:1992:P

[HWD92] Muhammad K. Habib, Edward J. Wegman, and Joel L. Davis. Preface. Journal of Statistical Planning and Inference, 33(1):3– 4, October 1992. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375892900916.

Hyodo:1992:CPI

[Hyo92] Yoshifumi Hyodo. Characteristic polynomials of information matrices of some balanced fractional 2^{m} factorial designs of resolution 2l + 1. Journal of Statistical Planning and Inference, 31(2): 245–252, May 1992. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589290033O.

Hjellvik:1998:LTU

[HYT98] Vidar Hjellvik, Qiwei Yao, and Dag Tjøstheim. Linearity testing using local polynomial approximation. Journal of Statistical Planning and Inference, 68(2):295–321, May 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897001468.

Hwang:1991:KTD

[HZzY91] F. K. Hwang, Lin Zong-zhen, and Y. C. Yao. Knockout tournaments with diluted Bradley-Terry preference schemes. Journal of Statistical Planning and Inference, 28(1):99–106, May 1991. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589190063K.

Iqbal:1994:ERM

[IJ94] I. Iqbal and B. Jones. Efficient repeated measurements designs with equal and unequal period sizes. Journal of Statistical Planning and Inference, 42(1–2):79–88, November 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894901902.

Iliopoulos:1998:IIE

[IK98] George Iliopoulos and Stavros Kourouklis. On improved interval estimation for the generalized variance. Journal of Statistical Planning and Inference, 66(2):305–320, March 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897000906.

Insua:1994:RIU

[IM94] David Ríos Insua and Jacinto Martín. Robustness issues under imprecise beliefs and preferences. Journal of Statistical Planning and Inference, 40(2–3):383–389, July 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894901333.

Imlahi:1995:FLI

[Iml95] A. Imlahi. Functional laws of the iterated logarithm for records. Journal of Statistical Planning and Inference, 45(1–2):215–224, ???? 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375894000727.

Iosifescu:1992:SMS

[Ios92] Marius Iosifescu. On U-statistics and von Mises statistics for a special class of Markov chains. Journal of Statistical Planning and Inference, 30(3):395–400, March 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290165O.

Irony:1993:ISR

[Iro93] Telba Zalkind Irony. Information in sampling rules. Journal of Statistical Planning and Inference, 36(1):27–38, July 1993. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390099R.

Irony:1997:NIP

[IS97] Telba Z. Irony and Nozer D. Singpurwalla. Non-informative priors do not exist: a dialogue with José M. Bernardo. Journal of Statistical Planning and Inference, 65(1):159–177, December 1, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375897000748.

Ionin:1998:RPB

[IS98] Yury J. Ionin and Mohan S. Shrikhande. Resolvable pairwise balanced designs. Journal of Statistical Planning and Inference, 72(1–2):393–405, September 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898000457.

Isogai:1994:BET

[Iso94] Eiichi Isogai. A Berry–Esseen-type bound for recursive estimators of a density and its derivatives. Journal of Statistical Planning and Inference, 40(1):1–14, June 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894901384.

Isogai:1999:SED

[Iso99] Eiichi Isogai. Sequential estimation of a density and its derivatives with bounded MISE. Journal of Statistical Planning and Inference, 79(1):59–78, June 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375898002560.

Izenman:1990:KES

[IT90] Alan Julian Izenman and Lanh Tat Tran. Kernel estimation of the survival function and hazard rate under weak dependence. Journal of Statistical Planning and Inference, 24(2):233–247, February 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589090044U.

Iwashita:1997:ANN

[Iwa97] Toshiya Iwashita. Asymptotic null and nonnull distribution of Hotelling’s T ^{2}-statistic under the elliptical distribution. Journal of Statistical Planning and Inference, 61(1):85–104, May 30, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S037837589600153X.

Iwasa:1998:DMP

[Iwa98] Manabu Iwasa. Directional monotonicity properties of the power functions of likelihood ratio tests for cone-restricted hypotheses of normal means. Journal of Statistical Planning and Inference, 66(2):223–233, March 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897000876.

Iyer:1992:BRI

[Iye92] Hari K. Iyer. Book review: An introduction to generalized linear models: Annette J. Dobson Chapman and Hall, London, 1990 250 pp., $49.95. Journal of Statistical Planning and Inference, 32(3):418–420, September 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290021J.

Jacroux:1990:SOR

[Jac90] Mike Jacroux. Some E-optimal row-column designs having unequally replicated treatments. Journal of Statistical Planning and Inference, 26(1):65–81, September 1990. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589090096D.

Jacroux:1992:OBD

[Jac92] Mike Jacroux. On the A-optimality of block designs. Journal of Statistical Planning and Inference, 32(3):401–415, September 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589290019O.

Jacroux:1998:DCO

[Jac98] Mike Jacroux. On the determination and construction of Eoptimal block designs in the presence of linear trends. Journal of Statistical Planning and Inference, 67(2):331–339, 342, April 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375897001201.

Janardan:1993:SPG

[Jan93a] K. G. Janardan. Some properties of the generalized Eulerian distribution. Journal of Statistical Planning and Inference, 34(2):159–169, February 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837589390001M.

Jandhyala:1993:PPS

[Jan93b] Venkata K. Jandhyala. A property of partial sums of regression least squares residuals and its applications. Journal of Statistical Planning and Inference, 37(3):317–326, December 1993. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390110R.

Janssen:1994:TSG

[Jan94a] Arnold Janssen. Two-sample goodness-of-fit tests when ties are present. Journal of Statistical Planning and Inference, 39(3): 399–424, May 1994. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375894900957.

Janssen:1994:WBS

[Jan94b] Paul Janssen. Weighted bootstrapping of U-statistics. Journal of Statistical Planning and Inference, 38(1):31–41, January 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589200156X.

Janssen:1999:TNS

[Jan99] Arnold Janssen. Testing nonparametric statistical functionals with applications to rank tests. Journal of Statistical Planning and Inference, 81(1):71–93, October 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375899000099.

Jaruskova:1998:TAL

[Jar98] Daniela Jarusková. Testing appearance of linear trend. Journal of Statistical Planning and Inference, 70(2):263–276, July 15, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375897001845.

Joseph:1991:EVO

[JD91] Betty Joseph and T. M. Durairajan. Equivalence of various optimality criteria for estimating functions. Journal of Statistical Planning and Inference, 27(3):355–360, March 1991. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589190048J.

Jensen:1997:BRM

[Jen97a] D. R. Jensen. Book review: Mixture models: Theory, geometry and applications: Bruce G. Lindsay, NSF-CBMS Regional conference series in probability and statics, volume 5 (Institute of Mathematical Statistics and American Statistical Association, 1995 ISBN: 0-940600-32-3, pp. 163). Journal of Statistical Planning and Inference, 59(1):179–181, March 30, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896000973.

Jensen:1997:WAK

[Jen97b] D. R. Jensen. T. W. Anderson, K. T. Fang, I. Olkin (eds.): Review of Multivariate analysis and its applications, lecture notes, monograph series vol. 24 (Institute of Mathematical Statistics Hayward, CA, USA, 1995). Journal of Statistical Planning and Inference, 60(1):189–191, April 30, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896001309.

Jeong:1995:ODE

[Jeo95] Yousceek Jeong. Optimal designs for experiments with a large number of factors. Journal of Statistical Planning and Inference, 46(3):381–393, August 15, 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589400110H.

Jensen:1991:PE

[JF91] D. R. Jensen and R. V. Foutz. Pitman estimation on R^{k}. Journal of Statistical Planning and Inference, 28(2):233–240, June 1991. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589190028D.

John:1995:MDD

[JJMY95] P. W. M. John, M. E. Johnson, L. M. Moore, and D. Ylvisaker. Minimax distance designs in two-level factorial experiments. Journal of Statistical Planning and Inference, 44(2):249–263, April 1, 1995. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589400047Y.

Johnson:1994:FCM

[JK94a] Norman L. Johnson and Samuel Kotz. Further comments on Matveychuk and Petunin’s generalized Bernoulli model, and nonparametric tests of homogeneity. Journal of Statistical Planning and Inference, 41(1):61–72, August 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894901546.

Johnson:1994:BCB

[JK94b] W. O. Johnson and G. E. Kokolakis. Bayesian classification based on multivariate binary data. Journal of Statistical Planning and Inference, 41(1):21–35, August 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589490152X.

Johnson:1997:RTH

[JK97] Norman L. Johnson and Samuel Kotz. Remarks on tests of homogeneity. Journal of Statistical Planning and Inference, 64(1):151– 152, October 30, 1997. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375897000037.

Jones:1992:IMM

[JKW92] B. Jones, J. Kunert, and H. P. Wynn. Information matrices for mixed effects models with applications to the optimality of repeated measurements designs. Journal of Statistical Planning and Inference, 33(2):261–274, November 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290072Z.

Janssen:1995:P

[JL95] Paul Janssen and Emmanuel Lesaffre. Preface. Journal of Statistical Planning and Inference, 47(1–2):3, October 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895900055.

Jandhyala:1991:TPC

[JM91] V. K. Jandhyala and I. B. MacNeill. Tests for parameter changes at unknown times in linear regression models. Journal of Statistical Planning and Inference, 27(3):291–316, March 1991. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589190043E.

Jandhyala:1993:DBT

[JM93] V. K. Jandhyala and C. D. Minogue. Distributions of Bayestype change-point statistics under polynomial regression. Journal of Statistical Planning and Inference, 37(3):271–290, December 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589390107H.

Jacob:1996:ABS

[JM96] Pierre Jacob and Bruno Massé. Asymptotic behavior of samples from general multivariate distributions. Journal of Statistical Planning and Inference, 52(2):183–196, June 15, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895001182.

Jacobson:1998:MMA

[JM98] Sheldon H. Jacobson and Douglas J. Morrice. A mathematical model for assessing the temporal association between health disorders and medical treatments. Journal of Statistical Planning and Inference, 71(1–2):209–228, August 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898000834.

Jureckova:1999:CDI

[JM99] Jana Jurecková and Xavier Milhaud. Characterization of distributions in invariant models. Journal of Statistical Planning and Inference, 75(2):353–362, January 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898001542.

Janardan:1997:SMR

[JMVH97] K. G. Janardan, N. R. Mohan, R. Vasudeva, and H. V. Hebbar. A stochastic model with random change point for the study of oviposition tactics of weevils on beans. Journal of Statistical Planning and Inference, 61(2):193–201, June 16, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896001462.

Johnson:1990:MMD

[JMY90] M. E. Johnson, L. M. Moore, and D. Ylvisaker. Minimax and maximin distance designs. Journal of Statistical Planning and Inference, 26(2):131–148, October 1990. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589090122B.

Janssen:1997:TSR

[JN97] Arnold Janssen and Georg Neuhaus. Two-sample rank tests for censored data with non-predictable weights. Journal of Statistical Planning and Inference, 60(1):45–59, April 30, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896001255.

Johnson:1991:CED

[Joh91] Richard A. Johnson. Confidence estimation with discrete reliability growth models. Journal of Statistical Planning and Inference, 29(1–2):87–97, September/October 1991. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290124B.

Jones:1992:BRA

[Jon92] Richard H. Jones. Book review: Analysis of repeated measures: M. J. Crowder and D. J. Hand Chapman and Hall, London, 1990 256 pp., $55.50. Journal of Statistical Planning and Inference, 32(3):417–418, September 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290020S.

Jost:1991:REM

[Jos91] Steve D. Jost. Relative efficiencies of model robust estimators in two-dimensional situations. Journal of Statistical Planning and Inference, 28(3):369–381, July 1991. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589190073N.

Joshi:1996:FMS

[Jos96a] V. M. Joshi. F42. Minimal sufficiency of a complete sufficient statistic. Journal of Statistical Planning and Inference, 55(2):255–256, October 30, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375896888446.

Joshi:1996:FIL

[Jos96b] V. M. Joshi. F43. Irrelevance of labels in estimation. Journal of Statistical Planning and Inference, 55(2):256–257, October 30, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375896888458.

Joshi:1996:FFP

[Jos96c] V. M. Joshi. F44. Fallacy in the proof of Birnbaum’s theorem: the reply by Berger (F28). Journal of Statistical Planning and Inference, 55(2):257–258, October 30, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

037837589688846X. See [Goo90c, Goo96a].

Joshi:1996:FRF

[Jos96d] V. M. Joshi. F44B. A reply to F44A. Journal of Statistical Planning and Inference, 55(2):258–259, October 30, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375896888471. See [Goo96a].

Joshi:1996:FCI

[Jos96e] V. M. Joshi. F45. A comment on ‘the insufficiency of some definitions of sufficiency’, F25. Journal of Statistical Planning and Inference, 55(2):259–260, October 30, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375896888483.

Joshi:1996:FCS

[Jos96f] V. M. Joshi. F47. Conditionality in structural inference. Journal of Statistical Planning and Inference, 55(2):261, October 30, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375896888501.

Joshi:1996:FLD

[Jos96g] V. M. Joshi. F48. A lacuna in the definition of Pitman-closerness. Journal of Statistical Planning and Inference, 55(2):261–262, October 30, 1996. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375896888513.

Jarrett:1997:ETR

[JPW97] R. G. Jarrett, F. C. Piper, and P. R. Wild. Efficient tworeplicate resolvable row-column designs. Journal of Statistical Planning and Inference, 58(1):65–77, March 1, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896000602.

Jansen:1993:ECL

[JR93] D. R. Jansen and D. E. Ramirez. Efficiency comparisons in linear inference. Journal of Statistical Planning and Inference, 37(1): 51–68, October 1993. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589390079L.

John:1999:RBD

[JR99] J. A. John and Katya Ruggiero. Resolvable block designs for factorial experiments. Journal of Statistical Planning and Inference, 77(2):293–299, March 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898001852.

Jiacheng:1996:CAM

[JRM96] Zhu Jiacheng, Sun Rongguo, and Cheng Murong. A construction of addition-multiplication magic square of order 18. Journal of Statistical Planning and Inference, 51(3):331–337, May 1, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375895000968.

Jacob:1995:EEP

[JS95] Pierre Jacob and Charles Suquet. Estimating the edge of a Poisson process by orthogonal series. Journal of Statistical Planning and Inference, 46(2):215–234, August 1, 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894001033.

Jiang:1999:LRT

[JSH99] Guoyong Jiang, Sanat K. Sarkar, and Francis Hsuan. A likelihood ratio test and its modifications for the homogeneity of the covariance matrices of dependent multivariate normals. Journal of Statistical Planning and Inference, 81(1):95–111, October 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899000245.

Jungnickel:1997:ADS

[JV97] Dieter Jungnickel and Scott A. Vanstone. An application of difference sets to a problem concerning graphical codes. Journal of Statistical Planning and Inference, 62(1):43–46, July 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896001668.

Jones:1992:AES

[JW92] M. C. Jones and M. P. Wand. Asymptotic effectiveness of some higher order kernels. Journal of Statistical Planning and Inference, 31(1):15–21, April 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837589290038T.

Jones:1999:DCT

[JWJB99] B. Jones, J. Wang, P. Jarvis, and W. Byrom. Design of crossover trials for pharmacokinetic studies. Journal of Statistical Planning and Inference, 78(1–2):307–316, May 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898002213.

John:1993:CCD

[JWT93] J. A. John, D. Whitaker, and C. M. Triggs. Construction of cyclic designs using integer programming. Journal of Statistical Planning and Inference, 36(2–3):357–366, August/September 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589390137U.

Kadane:1994:ARB

[Kad94] J. B. Kadane. An application of robust Bayesian analysis to a medical experiment. Journal of Statistical Planning and Inference, 40(2–3):221–232, July 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894901228.

Kageyama:1990:OAA

[Kag90] Sanpei Kageyama. On orthogonal arrays attaining rao’s bounds. Journal of Statistical Planning and Inference, 24(1):135, January 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589090023N.

Kakizawa:1997:HOB

[Kak97] Yoshihide Kakizawa. Higher-order Bartlett-type adjustment. Journal of Statistical Planning and Inference, 65(2):269–280, December 15, 1997. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375897000517.

Kamps:1991:IMO

[Kam91] Udo Kamps. Inequalities for moments of order statistics and characterizations of distributions. Journal of Statistical Planning and Inference, 27(3):397–404, March 1991. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589190052G.

Kamps:1995:CGO

[Kam95a] Udo Kamps. A concept of generalized order statistics. Journal of Statistical Planning and Inference, 48(1):1–23, November 1, 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589400147N.

Kamps:1995:RRM

[Kam95b] Udo Kamps. Recurrence relations for moments of record values. Journal of Statistical Planning and Inference, 45(1–2):225–234, ???? 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375894000735.

Karakostas:1990:EOS

[Kar90] K. X. Karakostas. Exact optimum sampling designs for autocorrelated finite populations. Journal of Statistical Planning and Inference, 24(3):353–361, March 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589090054X.

Karunamuni:1991:NSQ

[Kar91] R. J. Karunamuni. A new smoothness quantification in kernel density estimation. Journal of Statistical Planning and Inference, 27(3):361–373, March 1991. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837589190049K.

Katona:1997:BBS

[Kat97] Gyula Katona. A brief biographical sketch of Paul Erdős. Journal of Statistical Planning and Inference, 58(1):3–4, March 1, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375896000559.

Kaufman:1996:SSS

[Kau96] Gordon M. Kaufman. Successive sampling and software reliability. Journal of Statistical Planning and Inference, 49(3):343–369, February 1, 1996. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375896853274.

Kishino:1998:MLP

[KBH98] Hirohisa Kishino and Avner Bar-Hen. Maximum likelihood procedure adapted to sampling schemes. Journal of Statistical Planning and Inference, 70(2):277–286, July 15, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897001869.

Kim:1993:GAL

[KC93] Peter T. Kim and G. Robert Chapman. Group action on a lattice and an application to time series analysis. Journal of Statistical Planning and Inference, 34(2):183–195, February 1993. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390003O.

Kim:1990:DSO

[KD90] S. H. Kim and H. A. David. On the dependence structure of order statistics and concomitants of order statistics. Journal of Statistical Planning and Inference, 24(3):363–368, March 1990. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589090055Y.

Karia:1999:BHG

[KD99] S. R. Karia and J. V. Deshpande. Bounds for the hazard gradients in the competing risks set up. Journal of Statistical Planning and Inference, 75(2):363–377, January 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898001554.

Kejun:1996:ESS

[Kej96] Chen Kejun. On the existence of super-simple (ν, 4, 4)-BIBDs. Journal of Statistical Planning and Inference, 51(3):339–350, May 1, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375895000976.

Kemp:1997:CDN

[Kem97] Adrienne W. Kemp. Characterizations of a discrete normal distribution. Journal of Statistical Planning and Inference, 63(2):223– 229, October 1997. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375897000207.

Kendall:1990:RDS

[Ken90] David G. Kendall. Random Delaunay simplexes in R^{m}. Journal of Statistical Planning and Inference, 25(3):225–234, July 1990. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375890900745.

Kent:1997:DAS

[Ken97] John T. Kent. Data analysis for shapes and images. Journal of Statistical Planning and Inference, 57(2):181–193, February 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375896000432.

Kestenband:1992:NFB

[Kes92] Barbu C. Kestenband. New families of balanced incomplete block designs with block size q + 1. Journal of Statistical Planning and Inference, 30(1):83–91, January 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375892901096.

Key:1996:ESD

[Key96] J. D. Key. Extendable Steiner designs from finite geometries. Journal of Statistical Planning and Inference, 56(2):181–186, December 1996. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375896000171.

Kumar:1996:SCS

[KGS96] T. Krishna Kumar, Anil P. Gore, and V. Sitaramam. Some conceptual and statistical issues on measurement of poverty. Journal of Statistical Planning and Inference, 49(1):53–71, January 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375895000305.

Kent:1995:CIN

[KH95] John T. Kent and Timothy J. Hainsworth. Confidence intervals for the noncentral chi-squared distribution. Journal of Statistical Planning and Inference, 46(2):147–159, August 1, 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894001044.

Koch:1999:EUD

[KH99a] Hans-Friedrich Koch and Ludwig A. Hothorn. Exact unconditional distributions for dichotomous data in many-to-one comparisons. Journal of Statistical Planning and Inference, 82(1–2):83– 99, December 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375899000336.

Krummenauer:1999:SSG

[KH99b] Frank Krummenauer and Gerhard Hommel. The size of Simes’ global test for discrete test statistics. Journal of Statistical Planning and Inference, 82(1–2):151–162, December 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375899000397.

Khamis:1990:CKS

[Kha90a] H. J. Khamis. The δ-corrected Kolmogorov–Smirnov test for goodness of fit. Journal of Statistical Planning and Inference, 24(3):317–335, March 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837589090051U.

Khattree:1990:SPP

[Kha90b] Ravindra Khattree. Some prediction procedures with special reference to selection. Journal of Statistical Planning and Inference, 24(3):391–401, March 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375890900583.

Khan:1997:MEN

[Kha97] Rasul A. Khan. Monotonic expectations, Neyman–Pearson lemma, and some properties of the SPRT. Journal of Statistical Planning and Inference, 62(2):247–254, August 15, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896001929.

Kambo:1990:SEE

[KHAH90] N. S. Kambo, B. R. Handa, and Z. A. Al-Hemyari. On shrunken estimators for exponential scale parameter. Journal of Statistical Planning and Inference, 24(1):87–94, January 1990. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589090019Q.

Khuri:1990:ETR

[Khu90a] AndréI. Khuri. Exact tests for random models with unequal cell frequencies in the last stage. Journal of Statistical Planning and Inference, 24(2):177–193, February 1990. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375890900402.

Khuri:1990:MR

[Khu90b] AndréI. Khuri. Multiresponse rotatability. Journal of Statistical Planning and Inference, 25(1):1–6, May 1990. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589090002C.

Khuri:1992:TCN

[Khu92] A. I. Khuri. Tests concerning a nested mixed model with heteroscedastic random effects. Journal of Statistical Planning and Inference, 30(1):33–44, January 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375892901052.

Khuri:1996:MDE

[Khu96] AndréI. Khuri. A method for determining the effect of imbalance. Journal of Statistical Planning and Inference, 55(1):115– 129, October 15, 1996. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375895001859.

Khuri:1998:USS

[Khu98] André I. Khuri. On unweighted sums of squares in unbalanced analysis of variance. Journal of Statistical Planning and Inference, 74(1):135–147, October 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898001062.

Kuwada:1998:BFF

[KI98] Masahide Kuwada and Keiko Ikeda. Balanced fractional 3^{m} factorial designs of resolution IV. Journal of Statistical Planning and Inference, 73(1–2):29–45, September 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898000494.

Kim:1997:AOB

[Kim97a] Tae Yoon Kim. Asymptotically optimal bandwidth selection rules for the kernel density estimator with dependent observations. Journal of Statistical Planning and Inference, 59(2):321– 336, April 15, 1997. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375896001176.

Kim:1997:CLT

[Kim97b] Tae Yoon Kim. Central limit theorems for quadratic errors of nonparametric estimators. Journal of Statistical Planning and Inference, 64(2):193–204, November 15, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897000438.

Kim:1998:OEC

[Kim98] Yongdai Kim. Optimal estimation for continuous state branching processes with discrete sampling. Journal of Statistical Planning and Inference, 70(1):77–89, July 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897001821.

King:1996:HTP

[Kin96] Maxwell L. King. Hypothesis testing in the presence of nuisance parameters. Journal of Statistical Planning and Inference, 50(1):103–120, February 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375895000488.

Kyung-Joon:1998:NKR

[KJS98] Cha Kyung-Joon and William R. Schucany. Nonparametric kernel regression estimation near endpoints. Journal of Statistical Planning and Inference, 66(2):289–304, March 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897000827.

Kakiuchi:1995:MMH

[KK95a] Itsuro Kakiuchi and Miyoshi Kimura. Majorization methods on hyperplanes and their applications. Journal of Statistical Planning and Inference, 47(3):217–235, October 15, 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589400137K.

Kochar:1995:SRN

[KK95b] Subhash C. Kochar and S. N. U. A. Kirmani. Some results on normalized spacings from restricted families of distributions. Journal of Statistical Planning and Inference, 46(1):47–57, July 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589400095D.

Kundu:1995:CES

[KK95c] Debasis Kundu and Rameswar Kundu. Consistent estimates of super imposed exponential signals when some observations are missing. Journal of Statistical Planning and Inference, 44(2):205– 218, April 1, 1995. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589400043U.

Kounias:1996:NEC

[KKNK96] S. Kounias, C. Koukouvinos, N. Nikolaou, and A. Kakos. The non-equivalent circulant D-optimal designs for n = 90. Journal of Statistical Planning and Inference, 53(2):253–259, August 15, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375895001026.

Kong:1996:EEC

[KL96] Fanhui Kong and Bruce Levin. Edgeworth expansions for the conditional distributions in logistic regression models. Journal of Statistical Planning and Inference, 52(1):109–129, June 1, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895001069.

Kagan:1999:CND

[KL99] Abram Kagan and Gérard Letac. Characterization of the normal distribution through the power of a one-way ANOVA. Journal of Statistical Planning and Inference, 77(1):1–9, February 15, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375898001943.

Klaassen:1990:RFR

[Kla90] Chris A. J. Klaassen. Random and fixed replacement sampling plans. Journal of Statistical Planning and Inference, 25(2):153– 161, June 1990. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589090063Z.

Klefsjo:1991:TPT

[Kle91] Bengt Klefsjö. TTT-plotting — a tool for both theoretical and practical problems. Journal of Statistical Planning and Inference, 29(1–2):99–110, September/October 1991. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290125C.

Kyriakoussis:1998:CGF

[KLP98] A. Kyriakoussis, Gang Li, and Alex Papadopoulos. On characterization and goodness-of-fit test of some discrete distribution families. Journal of Statistical Planning and Inference, 74(2):215–228, November 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375898001025.

Kent:1996:SCA

[KM96] John T. Kent and Kanti V. Mardia. Spectral and circulant approximations to the likelihood for stationary Gaussian random fields. Journal of Statistical Planning and Inference, 50(3):379– 394, March 1996. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375895000658.

Kageyama:1997:NDB

[KM97] Sanpei Kageyama and Ying Miao. Nested designs with block size five and subblock size two. Journal of Statistical Planning and Inference, 64(1):125–139, October 30, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896002108.

Kageyama:1998:NDS

[KM98a] Sanpei Kageyama and Ying Miao. Nested designs of superblock size four. Journal of Statistical Planning and Inference, 73(1–2):1–5, September 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375898000469.

Kundu:1998:EPL

[KM98b] Debasis Kundu and Amit Mitra. Estimating the parameters of the linear compartment model. Journal of Statistical Planning and Inference, 70(2):317–334, July 15, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897001900.

Kamberova:1999:MRU

[KM99] G. Kamberova and M. Mintz. Minimax rules under zero-one loss for a restricted location parameter. Journal of Statistical Planning and Inference, 79(2):205–221, July 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898002377.

Kubokawa:1993:EVA

[KMMN93] T. Kubokawa, K. Morita, S. Makita, and K. Nagakura. Estimation of the variance and its applications. Journal of Statistical Planning and Inference, 35(3):319–333, June 1993. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375893900207.

Khattree:1994:OTN

[KN94] Ravindra Khattree and Dayanand N. Naik. Optimal tests for nested designs with circular stationary dependence. Journal of Statistical Planning and Inference, 41(2):231–240, September 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589490166X.

Khosrovshahi:1998:CCD

[KN98] G. B. Khosrovshahi and Reza Naserasr. On compressing complete designs. Journal of Statistical Planning and Inference, 74(1):193– 201, October 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375898000895.

Knautz:1999:NUE

[Kna99] Henning Knautz. Nonlinear unbiased estimation in the linear regression model with nonnormal disturbances. Journal of Statistical Planning and Inference, 81(2):293–309, November 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899000233.

Koczkodaj:1998:TAE

[Koc98] Waldemar W. Koczkodaj. Testing the accuracy enhancement of pairwise comparisons by a Monte Carlo experiment. Journal of Statistical Planning and Inference, 69(1):21–31, June 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375897001316.

Kong:1999:ARA

[Kon99] Fan Hui Kong. Adjusting regression attenuation in the Cox proportional hazards model. Journal of Statistical Planning and Inference, 79(1):31–44, June 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898001785.

Koronacki:1992:P

[Kor92] Jacek Koronacki. Preface. Journal of Statistical Planning and Inference, 30(2):137–139, February 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375892900765.

Kott:1990:ECV

[Kot90] Phillip S. Kott. Estimating the conditional variance of a design consistent regression estimator. Journal of Statistical Planning and Inference, 24(3):287–296, March 1990. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589090049Z.

Koul:1991:WCR

[Kou91] Hira L. Koul. A weak convergence result useful in robust autoregression. Journal of Statistical Planning and Inference, 29(3):291–308, November 1991. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837589190005Y.

Koukouvinos:1994:CSS

[Kou94] Christos Koukouvinos. Construction of some SBIBD(4k^{2}, 2k^{2} + k, k^{2} + k) and Hadamard matrices with maximal excess. Journal of Statistical Planning and Inference, 39(1):129–134, April 1, 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375894901171.

Kourouklis:1995:EPS

[Kou95] Stavros Kourouklis. Estimating powers of the scale parameter of an exponential distribution with unknown location under Pitman’s measure of closeness. Journal of Statistical Planning and Inference, 48(2):185–195, November 15, 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589400156P.

Koukouvinos:1998:SNT

[Kou98] Christos Koukouvinos. Some new three and four variable orthogonal designs in order 36. Journal of Statistical Planning and Inference, 73(1–2):21–27, September 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898000482.

Katzenbeisser:1993:SFR

[KP93a] Walter Katzenbeisser and Wolfgang Panny. Some further results on the height of lattice paths. Journal of Statistical Planning and Inference, 34(2):171–181, February 1993. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390002N.

Koreisha:1993:NAD

[KP93b] Sergio G. Koreisha and Tarmo M. Pukkila. New approaches for determining the degree of differencing necessary to induce stationarity in ARIMA models. Journal of Statistical Planning and Inference, 36(2–3):399–412, August/September 1993. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375893901402.

Kirschenhofer:1996:RSS

[KP96] Peter Kirschenhofer and Helmut Prodinger. Return statistics of simple random walks. Journal of Statistical Planning and Inference, 54(1):67–74, September 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895001573.

Kafadar:1997:EDL

[KP97] Karen Kafadar and Philip C. Prorok. Estimating the difference in location parameters of two survival curves, with applications to cancer screening. Journal of Statistical Planning and Inference, 57(2):165–179, February 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896000420.

Kleffe:1993:INO

[KR93] J. Kleffe and J. N. K. Rao. Inadmissibility but near optimality of an estimator of correlated response variance under additive models. Journal of Statistical Planning and Inference, 36(2–3):151–163, August/September 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390120U.

Kaufmann:1995:ARM

[KR95] E. Kaufmann and R.-D. Reiss. Approximation rates for multivariate exceedances. Journal of Statistical Planning and Inference, 45(1–2):235–245, ???? 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375894000743.

Kagan:1997:CRP

[KR97] Abram Kagan and Vijay Rohatgi. A class of random polynomials with an invariance property. Journal of Statistical Planning and Inference, 63(2):215–222, October 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897000177.

Krattenthaler:1993:GFS

[Kra93] C. Krattenthaler. Generating functions for shifted plane partitions. Journal of Statistical Planning and Inference, 34(2):197– 208, February 1993. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589390004P.

Krattenthaler:1996:OTN

[Kra96] C. Krattenthaler. Oscillating tableaux and nonintersecting lattice paths. Journal of Statistical Planning and Inference, 54(1):75– 85, September 1996. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375896001589.

Kreiss:1990:LAN

[Kre90] Jens-Peter Kreiss. Local asymptotic normality for autoregression with infinite order. Journal of Statistical Planning and Inference, 26(2):185–219, October 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589090126F.

Kadane:1990:RBP

[KS90] Joseph B. Kadane and Teddy Seidenfeld. Randomization in a Bayesian perspective. Journal of Statistical Planning and Inference, 25(3):329–345, July 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589090080E.

Kulasekera:1991:ECP

[KS91] K. B. Kulasekera and K. M. Lal Saxena. Estimation of change point in failure rate models. Journal of Statistical Planning and Inference, 29(1–2):111–124, September/October 1991. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290126D.

Kuczmaszewska:1992:CCR

[KS92] Anna Kuczmaszewska and Dominik Szynal. On the complete convergence for randomly indexed partial sums and the limit behaviour of a sequence of quantiles. Journal of Statistical Planning and Inference, 30(3):371–383, March 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290163M.

Kvam:1993:IEA

[KS93] Paul H. Kvam and Francisco J. Samaniego. On the inadmissibility of empirical averages as estimators in ranked set sampling. Journal of Statistical Planning and Inference, 36(1):39–55, July 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589390100K. See [KS11].

Kuchler:1994:EFS

[KS94] Uwe Küchler and Michael Sørensen. Exponential families of stochastic processes and Lévy processes. Journal of Statistical Planning and Inference, 39(2):211–237, April 15, 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894902062.

Koutsoukos:1995:IRH

[KS95] Antonis Koutsoukos and Eric Slud. Inaccuracy rates and Hodges–Lehmann large deviation rates for parametric inferences with nuisance parameters. Journal of Statistical Planning and Inference, 48(1):47–68, November 1, 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589400145L.

Knight:1996:EDM

[KS96] John L. Knight and Stephen E. Satchell. The exact distribution of the maximum likelihood estimators for the linear regression negative exponential model. Journal of Statistical Planning and Inference, 50(1):91–102, February 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589500047X.

Key:1997:STS

[KS97a] J. D. Key and F. D. Shobe. Some transitive Steiner triple systems of Bagchi and Bagchi. Journal of Statistical Planning and Inference, 58(1):79–86, March 1, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896000614.

Khosrovshahi:1997:FCB

[KS97b] G. B. Khosrovshahi and N. M. Singhi. Further characterization of basic trades. Journal of Statistical Planning and Inference, 58(1): 87–92, March 1, 1997. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375896000626.

Koukouvinos:1997:WMT

[KS97c] Christos Koukouvinos and Jennifer Seberry. Weighing matrices and their applications. Journal of Statistical Planning and Inference, 62(1):91–101, July 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375896001723.

Koul:1997:TET

[KS97d] Hira L. Koul and Anton Schick. Testing for the equality of two nonparametric regression curves. Journal of Statistical Planning and Inference, 65(2):293–314, December 15, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897000633.

Krafft:1997:OCB

[KS97e] Olaf Krafft and Martin Schaefer. A-optimal connected block designs with nearly minimal number of observations. Journal of Statistical Planning and Inference, 65(2):375–386, December 15, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375897000542.

Kreher:1997:SGD

[KS97f] D. L. Kreher and D. R. Stinson. Small group-divisible designs with block size four. Journal of Statistical Planning and Inference, 58(1):111–118, March 1, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837589600064X.

Kim:1998:SEU

[KS98a] Yongman Kim and Kesar Singh. Sharpening estimators using resampling. Journal of Statistical Planning and Inference, 66(1):121–146, January 5, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375897000797.

Koul:1998:RMF

[KS98b] Hira L. Koul and Winfried Stute. Regression model fitting with long memory errors. Journal of Statistical Planning and Inference, 71(1–2):35–56, August 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898000160.

Kuchler:1998:EFM

[KS98c] Uwe Küchler and Michael Sørensen. On exponential families of Markov processes. Journal of Statistical Planning and Inference, 66(1):3–19, January 5, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375897000724.

Kulinskaya:1998:TOG

[KS98d] Elena Kulinskaya and Robert G. Staudte. Tests and optimal guarded weights of evidence. Journal of Statistical Planning and Inference, 71(1–2):1–18, August 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898000068.

Koukouvinos:1999:NWM

[KS99] Christos Koukouvinos and Jennifer Seberry. New weighing matrices and orthogonal designs constructed using two sequences with zero autocorrelation function — a review. Journal of Statistical Planning and Inference, 81(1):153–182, October 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375899000063.

Kvam:2011:EIS

[KS11] P. H. Kvam and F. J. Samaniego. Erratum to “On the inadmissibility of standard estimators based on ranked set sampling” [Journal of Statistical Planning and Inference 36 (1993) 39–55]. Journal of Statistical Planning and Inference, 141(5):2001, May 2011. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375810005021. See [KS93].

Krewski:1995:BEB

[KSDS95] Daniel Krewski, Robert T. Smythe, Anup Dewanji, and Mieczyslaw Szyszkowicz. Bootstrapping an empirical Bayes estimator of the distribution of historical controls in carcinogen bioassay. Journal of Statistical Planning and Inference, 43(1–2):215– 228, January 1995. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589400021M.

Kieu:1999:PSS

[KSI99] Kiên Kiêu, Sandie Souchet, and Jacques Istas. Precision of systematic sampling and transitive methods. Journal of Statistical Planning and Inference, 77(2):263–279, March 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898001931.

Kubokawa:1992:IMC

[KSM92] T. Kubokawa, A. K. Md. E. Saleh, and K. Morita. Improving on MLE of coefficient matrix in a growth curve model. Journal of Statistical Planning and Inference, 31(2):169–177, May 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290027P.

Korchmaros:1997:SFS

[KSS97] G. Korchmáros, L. Storme, and T. Szőnyi. Space-filling subsets of a normal rational curve. Journal of Statistical Planning and Inference, 58(1):93–110, March 1, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896000638.

Koukouvinos:1997:ODS

[KSWyX97] C. Koukouvinos, Jennifer Seberry, A. L. Whiteman, and Ming yuan Xia. Optimal designs, supplementary difference sets and multipliers. Journal of Statistical Planning and Inference, 62(1): 81–90, July 1997. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375896001711.

Kariya:1992:BNA

[KT92] Takeaki Kariya and Yasuyuki Toyooka. Bounds for normal approximations to the distributions of generalized least squares predictors and estimators. Journal of Statistical Planning and Inference, 30(2):213–221, February 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

0378375892900824. See errata [UT97].

Katsis:1999:BSS

[KT99a] Athanassios Katsis and Blaza Toman. Bayesian sample size calculations for binomial experiments. Journal of Statistical Planning and Inference, 81(2):349–362, November 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375899000191.

Koziol:1999:BMC

[KT99b] James A. Koziol and Henry C. Tuckwell. A Bayesian method for combining statistical tests fn1. Journal of Statistical Planning and Inference, 78(1–2):317–323, May 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898002225.

Kubokawa:1990:MEC

[Kub90] Tatsuya Kubokawa. Minimax estimation of common coefficients of several regression models under quadratic loss. Journal of Statistical Planning and Inference, 24(3):337–345, March 1990. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589090052V.

Kubacek:1992:EFS

[Kub92] Lubomír Kubácek. Estimation of the first and second order parameters in regression models with special structure. Journal of Statistical Planning and Inference, 30(2):195–211, February 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375892900813.

Kuboki:1998:RPP

[Kub98] Hisataka Kuboki. Reference priors for prediction. Journal of Statistical Planning and Inference, 69(2):295–317, June 15, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375897001638.

Kulperger:1999:CSM

[Kul99] R. J. Kulperger. Countable state Markov process bootstrap. Journal of Statistical Planning and Inference, 76(1–2):19–29, February 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375898001141.

Kunert:1993:NOD

[Kun93] J. Kunert. A note on optimal designs with a non-orthogonal row-column-structure. Journal of Statistical Planning and Inference, 37(2):265–270, November 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837589390095N.

Kunert:1998:ACB

[Kun98] J. Kunert. On the analysis of circular balanced crossover designs. Journal of Statistical Planning and Inference, 69(2):359– 370, June 15, 1998. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375897001596.

Kushner:1999:SOR

[Kus99] H. B. Kushner. H-symmetric optimal repeated measurements designs. Journal of Statistical Planning and Inference, 76(1–2):235– 261, February 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375898001621.

Khuri:1991:CRM

[KV91] A. I. Khuri and G. G. Vining. Conditions required for the mean response to fall within specified bounds. Journal of Statistical Planning and Inference, 28(1):125–136, May 1991. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589190065M.

Khuri:1998:MSM

[KV98] André I. Khuri and Elsie Valeroso. Multiresponse surface models with block effects. Journal of Statistical Planning and Inference, 73(1–2):7–20, September 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898000470.

Khosrovshahi:1990:SSS

[KVAN90] G. B. Khosrovshahi, F. Vatan, and S. Ajoodani-Namini. On some support sizes for BIB designs with υ = 9 and k = 4. Journal of Statistical Planning and Inference, 26(1):83–92, September 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589090097E.

Keener:1992:FWI

[KW92] Robert W. Keener and Michael Woodroofe. Fixed width interval estimation for the reciprocal drift of Brownian motion. Journal of Statistical Planning and Inference, 30(1):1–12, January 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589290102X.

King:1998:OMD

[KW98] Joy King and Weng Kee Wong. Optimal minimax designs for prediction in heteroscedastic models. Journal of Statistical Planning and Inference, 69(2):371–383, June 15, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897001675.

Klonecki:1993:SPE

[KWZ93] W. Klonecki, A. Wiggins, and S. Zontek. Strictly positive estimators for variance components. Journal of Statistical Planning and Inference, 36(2–3):241–252, August/September 1993. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390127R.

Klein:1999:DWS

[KZ99] John P. Klein and Mei-Jie Zhang. Determining when the survival rates of two treatments are the same based on a censored data regression model. Journal of Statistical Planning and Inference, 78(1–2):243–254, May 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898002171.

Lang:1999:NPE

[LA99] Gabriel Lang and Jean-Marc Azaiïs. Non-parametric estimation of the long-range dependence exponent for Gaussian processes. Journal of Statistical Planning and Inference, 80(1–2):59– 80, August 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375898002420.

Labelle:1993:PNC

[Lab93] Jacques Labelle. On pairs of non-crossing generalized Dyck paths. Journal of Statistical Planning and Inference, 34(2):209– 217, February 1993. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589390005Q.

Ladalla:1995:FML

[Lad95] Joseph N. Ladalla. On fitting a modified limiting normal to the posterior distribution. Journal of Statistical Planning and Inference, 43(1–2):229–234, January 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589400022N.

Lahiri:1997:VST

[Lah97] S. N. Lahiri. Variance stabilizing transformations, studentization and the bootstrap. Journal of Statistical Planning and Inference, 61(1):105–123, May 30, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897897157.

Landsman:1996:SQA

[Lan96] Zinoviy Landsman. Sample quantiles and additive statistics: Information, sufficiency, estimation. Journal of Statistical Planning and Inference, 52(1):93–108, June 1, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895001077.

Lavine:1994:AES

[Lav94] Michael Lavine. An approach to evaluating sensitivity in Bayesian regression analyses. Journal of Statistical Planning and Inference, 40(2–3):233–244, July 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894901236.

Lawrence:1997:BRG

[Law97] John Lawrence. Book review: A guide to Chi-squared testing: Priscilla E. Greenwood and Mikhail S. Nikulin, (Wiley, New York, 1996, Price: $39.95, ISBN No. 0-471-55779-X, 280 pp.). Journal of Statistical Planning and Inference, 64(1):157–158, October 30, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375897001018.

Louis:1990:CER

[LB90] Thomas A. Louis and Julia K. Bailey. Controlling error rates using prior information and marginal totals to select tumor sites. Journal of Statistical Planning and Inference, 24(3):297– 316, March 1990. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375890900505.

Lu:1991:IPB

[LB91] Jye-Chyi Lu and Gouri K. Bhattacharyya. Inference procedures for a bivariate exponential model of Gumbel based on life test of component and system. Journal of Statistical Planning and Inference, 27(3):383–396, March 1991. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589190051F.

Lawry:1992:REC

[LB92] K. A. Lawry and D. R. Bellhouse. Relative efficiency of certain randomization procedures in an n × n array when spatial correlation is present. Journal of Statistical Planning and Inference, 32(3):385–399, September 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290018N.

Lee:1994:STS

[LB94] Youngjo Lee and David Birkes. Shrinking toward submodels in regression. Journal of Statistical Planning and Inference, 41(1): 95–111, August 1994. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375894901562.

Lopez-Blazquez:1998:DDM

[LB98] F. López-Blázquez. Discrete distributions with maximum expected value of the maximum. Journal of Statistical Planning and Inference, 70(2):201–207, July 15, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837589700181X.

Levine:1996:CPO

[LC96] Richard A. Levine and George Casella. Convergence of posterior odds. Journal of Statistical Planning and Inference, 55(3):331–344, November 8, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375895001980.

Lu:1997:DFS

[LC97] Zudi Lu and Ping Cheng. Distribution-free strong consistency for nonparametric kernel regression involving nonlinear time series. Journal of Statistical Planning and Inference, 65(1):67–86, December 1, 1997. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375897000451.

Lee:1995:EIS

[LCKP95] Y. H. Lee, S. Cho, W. C. Kim, and B. U. Park. On estimating integrated squared spectral density derivatives. Journal of Statistical Planning and Inference, 48(2):165–184, November 15, 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589400154N.

Liu:1993:CTG

[LD93] Jingyu Liu and H. A. David. Comparing two groups of ranked objects by pairwise matching. Journal of Statistical Planning and Inference, 35(2):157–169, May 1993. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375893900425.

Lopez-Diaz:1998:ROI

[LDG98] Miguel López-Díaz and María Angeles Gil. Reversing the order of integration in iterated expectations of fuzzy random variables, and statistical applications. Journal of Statistical Planning and Inference, 74(1):11–29, October 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898001001.

Leadbetter:1995:HLE

[Lea95] M. R. Leadbetter. On high level exceedance modeling and tail inference. Journal of Statistical Planning and Inference, 45(1–2): 247–260, ???? 1995. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375894000751.

Lecoutre:1999:TUD

[Lec99] Bruno Lecoutre. Two useful distributions for Bayesian predictive procedures under normal models. Journal of Statistical Planning and Inference, 79(1):93–105, June 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898002316.

Lee:1995:TML

[Lee95] Sangyeol Lee. A trimmed mean of location of an AR(∞) stationary process. Journal of Statistical Planning and Inference, 48(2):131–140, November 15, 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589500014Z.

Lee:1998:ABE

[Lee98a] Sangyeol Lee. The asymptotic behavior of the empirical process based on a linear process under some contiguous alternatives. Journal of Statistical Planning and Inference, 67(1):1–15, March 16, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375897000694.

Lee:1998:CCT

[Lee98b] Sangyeol Lee. Coefficient constancy test in a random coefficient autoregressive model. Journal of Statistical Planning and Inference, 74(1):93–101, October 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898000950.

Lee:1998:QPB

[Lee98c] Sangyeol Lee. On the quantile process based on the autoregressive residuals. Journal of Statistical Planning and Inference, 67(1):17– 28, March 16, 1998. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375897000700.

Lee:1999:BCT

[Lee99] Young Jack Lee. Biostatistics and clinical trials: a view. Journal of Statistical Planning and Inference, 78(1–2):349–367, May 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375898002249.

Lesperance:1997:DDP

[Les97] Mary L. Lesperance. Discussion on A. Desmond’s paper: Optimal estimating functions, quasi-likelihood and statistical modelling. Journal of Statistical Planning and Inference, 60(1):106– 108, April 30, 1997. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375897900112. See [Des97a].

Letac:1997:RRI

[Let97] Gérard Letac. The retrieval of a rotation invariant distribution of (X, Y) from E(X^{2}|Y). Journal of Statistical Planning and Inference, 63(2):247–253, October 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897844497.

Levi:1990:CBP

[Lev90] Isaac Levi. Compromising Bayesianism: a plea for indeterminacy. Journal of Statistical Planning and Inference, 25(3):347–361, July 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375890900815.

Lee:1991:LDU

[LG91] Mei-Ling Ting Lee and Alan J. Gross. Lifetime distributions under unknown environment. Journal of Statistical Planning and Inference, 29(1–2):137–143, September/October 1991. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290128F.

Liang:1993:MEB

[LH93] TaChen Liang and Wen-Tao Huang. On monotone empirical Bayes estimators of a binomial parameter. Journal of Statistical Planning and Inference, 34(3):333–348, March 1993. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390144U.

Lau:1997:DSB

[LH97a] Ka-Sing Lau and Alan Ho. On the discount sum of Bernoulli random variables. Journal of Statistical Planning and Inference, 63(2):231–246, October 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897000190.

Lee:1997:BAG

[LH97b] Jack C. Lee and Y. L. Hsu. Bayesian analysis of growth curves with AR(1) dependence. Journal of Statistical Planning and Inference, 64(2):205–229, November 15, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897807501.

Lin:1997:SEF

[LH97c] Y.-X. Lin and C. C. Heyde. On spaces of estimating functions. Journal of Statistical Planning and Inference, 63(2):255–264, October 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375897000219.

Li:1998:OFD

[Li98a] Junfang Li. Orthogonal factorial design of parallel flats type and its application in biomedical research. Journal of Statistical Planning and Inference, 73(1–2):61–75, September 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898000512.

Liao:1998:OPF

[LI98b] C. T. Liao and H. K. Iyer. Orthogonal parallel-flats designs for s_{1}^{n1} × s_{2}^{n2} mixed factorial experiments. Journal of Statistical Planning and Inference, 69(1):175–191, June 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897001298.

Liang:1990:C

[Lia90a] Kung-Yee Liang. Comments B. Journal of Statistical Planning and Inference, 25(2):195–196, June 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375890900675.

Liang:1990:CRM

[Lia90b] TaChen Liang. On the convergence rates of a monotone empirical Bayes test for uniform distributions. Journal of Statistical Planning and Inference, 26(1):25–34, September 1990. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375890900929.

Liang:1997:SSN

[Lia97] TaChen Liang. Simultaneously selecting normal populations close to a control. Journal of Statistical Planning and Inference, 61(2):297–316, June 16, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896001814.

Liero:1992:ANW

[Lie92] Hannelore Liero. Asymptotic normality of a weighted integrated squared error of kernel regression estimates with data-dependent bandwidth. Journal of Statistical Planning and Inference, 30(3): 307–325, March 1992. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589290158O.

Liese:1996:ASB

[Lie96] F. Liese. Adaptive selection of the best population. Journal of Statistical Planning and Inference, 54(3):245–269, September 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375895001719.

Lim:1990:ODC

[Lim90] Yong B. Lim. D-optimal design for cubic polynomial regression on the q-simplex. Journal of Statistical Planning and Inference, 25(2):141–152, June 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589090062Y.

Lindley:1990:GWP

[Lin90] D. V. Lindley. Good’s work in probability, statistics and the philosophy of science. Journal of Statistical Planning and Inference, 25(3):211–223, July 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375890900734.

Lindley:1997:SCN

[Lin97a] Dennis Lindley. Some comments on “non-informative priors do not exist”. Journal of Statistical Planning and Inference, 65(1):182–184, December 1, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375897900732.

Lindley:1997:SCB

[Lin97b] Dennis V. Lindley. Some comments on Bayes factors. Journal of Statistical Planning and Inference, 61(1):181–189, May 30, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375896001899.

Lindsey:1997:SRL

[Lin97c] J. K. Lindsey. Stopping rules and the likelihood function. Journal of Statistical Planning and Inference, 59(1):167–177, March 30, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375896000961.

Lin:1998:MLD

[Lin98] Gwo Dong Lin. On the Mittag-Leffler distributions. Journal of Statistical Planning and Inference, 74(1):1–9, October 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375898000962.

Liu:1994:NCP

[Liu94] W. Liu. A note on the conditional probability of correct selection under the two-element partition. Journal of Statistical Planning and Inference, 39(3):373–383, May 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894900930.

Liu:1995:RSS

[Liu95a] W. Liu. A random-size subset approach to the selection problem. Journal of Statistical Planning and Inference, 48(2):153–164, November 15, 1995. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589400149P.

Liu:1995:ETU

[Liu95b] Wei Liu. On the existence of tests uniformly more powerful than the likelihood ratio test. Journal of Statistical Planning and Inference, 44(1):19–35, March 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589400036U.

Liu:1997:SSD

[Liu97a] W. Liu. On sample size determination of Dunnett’s procedure for comparing several treatments with a control. Journal of Statistical Planning and Inference, 62(2):255–261, August 15, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896001930.

Liu:1997:SSS

[Liu97b] Wei Liu. A k-stage sequential sampling procedure for estimation of normal mean. Journal of Statistical Planning and Inference, 65(1):109–127, December 1, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897000487.

Laskar:1998:ETR

[LK98] Mizan R. Laskar and Maxwell L. King. Estimation and testing of regression disturbances based on modified likelihood functions. Journal of Statistical Planning and Inference, 71(1–2):75– 92, August 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375898000913.

Lindsay:1995:RSM

[LL95a] Bruce G. Lindsay and Mary L. Lesperance. A review of semiparametric mixture models. Journal of Statistical Planning and Inference, 47(1–2):29–39, October 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589400120K.

Lindsey:1995:DGL

[LL95b] J. K. Lindsey and P. Lambert. Dynamic generalized linear models and repeated measurements. Journal of Statistical Planning and Inference, 47(1–2):129–139, October 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589400126G.

Liu:1997:TAC

[LLL97] J. Liu, W. K. Li, and C. W. Li. On a threshold autoregression with conditional heteroscedastic variances. Journal of Statistical Planning and Inference, 62(2):279–300, August 15, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896001966.

Lloyd:1998:IER

[Llo98] Chris J. Lloyd. Improved efficiency for recapture studies from auxiliary experimentation. Journal of Statistical Planning and Inference, 67(1):29–44, March 16, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897000943.

Lavine:1995:NBM

[LM95] M. Lavine and A. Mockus. A nonparametric Bayes method for isotonic regression. Journal of Statistical Planning and Inference, 46(2):235–248, August 1, 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375894001066.

Lin:1999:EAA

[LM99] Yan-Xia Lin and Michellel McCabe. An example of applying the asymptotic quasi-likelihood to dimension estimation for random spatial patterns. Journal of Statistical Planning and Inference, 80(1–2):197–209, August 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898002511.

Liebler:1997:BTR

[LMT97] Robert A. Liebler, Spyros S. Magliveras, and Sergei V. Tsaranov. Block transitive resolutions of t-designs and room rectangles. Journal of Statistical Planning and Inference, 58(1):119– 133, March 1, 1997. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375896000651.

Liski:1995:PIE

[LN95] Erkki P. Liski and Tapio Nummi. Prediction and inverse estimation in repeated-measures models. Journal of Statistical Planning and Inference, 47(1–2):141–151, October 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589400127H.

Lee:1999:UGC

[LN99] Jack C. Lee and Wei-Fang Niu. On an unbalanced growth curve model with random effects and AR(1) errors from a Bayesian and the ML points of view. Journal of Statistical Planning and Inference, 76(1–2):41–55, February 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837589800161X.

Lohr:1993:TSA

[Loh93] Sharon L. Lohr. Three-stage accurate estimation in the general linear model. Journal of Statistical Planning and Inference, 34(3): 317–331, March 1993. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589390143T.

Lecoutre:1995:CCK

[LOS95] Jean-Pierre Lecoutre and Elias Ould-Said. Convergence of the conditional Kaplan–Meier estimate under strong mixing. Journal of Statistical Planning and Inference, 44(3):359–369, ???? 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375894000849.

Lahiri:1991:NBE

[LP91] Parthasarthi Lahiri and Dong Ho Park. Nonparametric Bayes and empirical Bayes estimators of mean residual life at age t. Journal of Statistical Planning and Inference, 29(1–2):125–136, September/October 1991. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589290127E.

Lund:1998:CCG

[LPS98] Robert Lund, W. J. Padgett, and Lynne Seymour. A control chart for a general Gaussian process. Journal of Statistical Planning and Inference, 70(1):19–34, July 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897001808.

Lakshmi:1992:TDU

[LR92] Damaraju V. Lakshmi and Damaraju Raghavarao. A test for detecting untruthful answering in randomized response procedures. Journal of Statistical Planning and Inference, 31(3):387– 390, June 1992. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589290144H.

Liu:1995:ADS

[LR95] Z. J. Liu and C. Radhakrishna Rao. Asymptotic distribution of statistics based on quadratic entropy and bootstrapping. Journal of Statistical Planning and Inference, 43(1–2):1–18, January 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589400005G.

Lee:1996:SEL

[LR96a] David K. C. Lee and Peter M. Robinson. Semiparametric exploration of long memory in stock prices. Journal of Statistical Planning and Inference, 50(2):155–174, March 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895000518.

Listing:1996:RSS

[LR96b] Joachim Listing and Dieter Rasch. Robustness of subset selection procedures. Journal of Statistical Planning and Inference, 54(3):291–305, September 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895001743.

Liang:1997:DFD

[LR97] Kung-Yee Liang and Paul J. Rathouz. Discussion on A. F. Desmond’s paper: Optimal estimating functions, quasilikelihood and statistical modelling. Journal of Statistical Planning and Inference, 60(1):113–116, April 30, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897900148. See [Des97a].

Lee:1995:SCB

[LS95a] Robert Edwin Lee and John D. Spurrier. Successive comparisons between ordered treatments. Journal of Statistical Planning and Inference, 43(3):323–330, February 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589591803B.

Li:1995:TOHa

[LS95b] Teng Li and Bimal K. Sinha. Tests of ordered hypotheses for gamma scale parameters. Journal of Statistical Planning and Inference, 43(3):311–321, February 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589400029U.

Li:1995:TOHb

[LS95c] Teng Li and Bimal K. Sinha. Tests of ordered hypotheses for gamma scale parameters. Journal of Statistical Planning and Inference, 45(3):387–397, June 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589400090I.

Li:1995:ORB

[LS95d] Yuanzhang Li and K. M. Lal Saxena. Optimal robust Bayes estimation. Journal of Statistical Planning and Inference, 46(3):365– 380, August 15, 1995. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375894001099.

Liu:1995:UDB

[LS95e] Regina Y. Liu and Kesar Singh. Using i.i.d. bootstrap inference for general non-i.i.d. models. Journal of Statistical Planning and Inference, 43(1–2):67–75, January 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589400008J.

Lin:1999:FTF

[LS99a] Win-Chin Lin and John Stufken. On finding trend-free block designs. Journal of Statistical Planning and Inference, 78(1–2): 57–70, May 1999. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375898002067.

Lingham:1999:IBF

[LS99b] Rama T. Lingham and S. Sivaganesan. Intrinsic Bayes factor approach to a test for the power law process. Journal of Statistical Planning and Inference, 77(2):195–220, March 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898001815.

Li:1999:RSR

[LSP99] Dayong Li, Bimal K. Sinha, and François Perron. Random selection in ranked set sampling and its applications. Journal of Statistical Planning and Inference, 76(1–2):185–201, February 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375898001360.

Lam:1996:CAR

[LT96] Clement Lam and Vladimir D. Tonchev. Classification of affine resolvable 2 − (27, 9, 4) designs. Journal of Statistical Planning and Inference, 56(2):187–202, December 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

S0378375896000183. See corrigendum [LT00].

Leu:1997:DAS

[LT97] Ching-Ho Leu and Kam-Wah Tsui. Discriminant analysis of survey data. Journal of Statistical Planning and Inference, 60(2):273–290, May 15, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375896001395.

Lam:2000:CCA

[LT00] Clement Lam and Vladimir D. Tonchev. Corrigendum to “Classification of affine resolvable 2 − (27, 9, 4) designs” [Journal of Statistical Planning and Inference 56 (1996) 187–202]. Journal of Statistical Planning and Inference, 86(1):277–278, April 15, 2000. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899000555.

Liski:1993:MMS

[LTT93] Erkki P. Liski, Helge Toutenburg, and Götz Trenkler. Minimum mean square error estimation in linear regression. Journal of Statistical Planning and Inference, 37(2):203–214, November 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589390089O.

Luo:1996:ADM

[Luo96] X. Luo. The asymptotic distribution of MLE of treatment lag threshold. Journal of Statistical Planning and Inference, 53(1):33–61, August 1, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375895001425.

Luschgy:1994:AIS

[Lus94] Harald Luschgy. Asymptotic inference for semimartingale models with singular parameter points. Journal of Statistical Planning and Inference, 39(2):155–186, April 15, 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894902046.

Lutkepohl:1996:TNI

[Lüt96] Helmut Lütkepohl. Testing for nonzero impulse responses in vector autoregressive processes. Journal of Statistical Planning and Inference, 50(1):1–20, February 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895000429.

Lan:1994:DMC

[LW94] K. K. Gordon Lan and Janet Wittes. Data monitoring in complex clinical trials: which treatment is ‘better’? Journal of Statistical Planning and Inference, 42(1–2):241–255, November 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894901996.

Liu:1997:RDA

[LW97] Shawn X. Liu and Douglas P. Wiens. Robust designs for approximately polynomial regression. Journal of Statistical Planning and Inference, 64(2):369–381, November 15, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897000396.

Laskar:1999:CGR

[LW99a] Renu Laskar and Charles Wallis. Chessboard graphs, related designs, and domination parameters. Journal of Statistical Planning and Inference, 76(1–2):285–294, February 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898001323.

Leonenko:1999:PIS

[LW99b] Nikolai N. Leonenko and Wojbor A. Woyczynski. Parameter identification for singular random fields arising in Burgers’ turbulence. Journal of Statistical Planning and Inference, 80(1–2): 1–13, August 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375898002390.

Lin:1995:SIA

[LY95a] D. Y. Lin and Zhiliang Ying. Semiparametric inference for the accelerated life model with time-dependent covariates. Journal of Statistical Planning and Inference, 44(1):47–63, March 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589400039X.

Liu:1995:BRC

[LY95b] Ying-Hsing Liu and Ming-Chung Yang. Bounded risk conditions in simultaneous estimation of independent Poisson means. Journal of Statistical Planning and Inference, 47(3):319–331, October 15, 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589400135I.

Liu:1996:OFG

[LY96] Guizhen Liu and Guiying Yan. Orthogonal [k − 1, k + 1]-factorizations in graphs. Journal of Statistical Planning and Inference, 51(2):195–200, April 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895000844.

Liu:1997:PRS

[LY97] Ying-Hsing Liu and Ming-Chung Yang. Posterior robustness in simultaneous estimation problem with exchangeable contaminated priors. Journal of Statistical Planning and Inference, 65(1):129–143, December 1, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375897000657.

Lyazrhi:1997:BCD

[Lya97] Faouzi Lyazrhi. Bayesian criteria for discriminating among regression models with one possible change point. Journal of Statistical Planning and Inference, 59(2):337–353, April 15, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837589600105X.

Lloyd:1998:CRR

[LYC98] Christopher J. Lloyd, Paul S. F. Yip, and K. S. Chan. A comparison of recapture, removal and resighting designs for population size estimation. Journal of Statistical Planning and Inference, 71(1–2):363–375, August 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898000755.

Lynch:1999:JDC

[Lyn99] J. D. Lynch. On the joint distribution of component failures for monotone load-sharing systems. Journal of Statistical Planning and Inference, 78(1–2):13–21, May 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898002031.

Lamprecht:1998:TAB

[LZ98] E. A. Lamprecht and S. Zacks. Two armed bandits with change point in one arm. Journal of Statistical Planning and Inference, 73(1–2):47–60, September 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898000500.

Ma:1996:RAG

[Ma96] S. L. Ma. Regular automorphism groups on partial geometries. Journal of Statistical Planning and Inference, 51(2):215– 222, April 1996. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375895000860.

Ma:1997:ADM

[Ma97a] Lingyun Ma. The asymptotic distributions of maximum likelihood ratio test and maximally selected χ^{2}-test in binomial observations. Journal of Statistical Planning and Inference, 65(1):17– 43, December 1, 1997. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375897000499.

Ma:1997:SNC

[Ma97b] S. L. Ma. Some necessary conditions on the parameters of partial difference sets. Journal of Statistical Planning and Inference, 62 (1):47–56, July 1997. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837589600167X.

Ma:1998:LRO

[Ma98a] Chunsheng Ma. Likelihood ratio ordering of order statistics. Journal of Statistical Planning and Inference, 70(2):255–261, July 15, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375897001894.

Ma:1998:PDC

[Ma98b] Chunsheng Ma. On peakedness of distributions of convex combinations. Journal of Statistical Planning and Inference, 70(1): 51–56, July 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837589700178X.

Macula:1997:NAS

[Mac97a] Anthony J. Macula. A non-adaptive search algorithm that identifies up to three defects. Journal of Statistical Planning and Inference, 60(2):269–272, May 15, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896001449.

Macula:1997:NVU

[Mac97b] Anthony J. Macula. A nonadaptive version of Ulam’s problem with one lie. Journal of Statistical Planning and Inference, 61(1):175–180, May 30, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375896001450.

Madi:1995:IEN

[Mad95] Mohamed T. Madi. On the invariant estimation of a normal variance ratio. Journal of Statistical Planning and Inference, 44(3): 349–357, ???? 1995. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375894000827.

Maesono:1991:NAS

[Mae91] Yoshihiko Maesono. On the normal approximation of U-statistics of degree two. Journal of Statistical Planning and Inference, 27(1):37–50, January 1991. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837589190080X.

Maesono:1994:NAS

[Mae94] Yoshihiko Maesono. On the normal approximations of V — and L-statistics. Journal of Statistical Planning and Inference, 42(3):291–314, December 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375894901481.

Maesono:1997:EES

[Mae97] Yoshihiko Maesono. Edgeworth expansions of a studentized Ustatistic and a jackknife estimator of variance. Journal of Statistical Planning and Inference, 61(1):61–84, May 30, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896001486.

Maguluri:1992:ESM

[Mag92] Gangaji Maguluri. Estimation in a semiparametric model by the method of minimum distance. Journal of Statistical Planning and Inference, 30(1):23–32, January 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290104Z.

Mahmoodian:1998:DSU

[Mah98] E. S. Mahmoodian. Defining sets and uniqueness in graph colorings: a survey. Journal of Statistical Planning and Inference, 73(1–2):85–89, September 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375898000536.

Maiti:1998:HBE

[Mai98] Tapabrata Maiti. Hierarchical Bayes estimation of mortality rates for disease mapping. Journal of Statistical Planning and Inference, 69(2):339–348, June 15, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897001651.

Makov:1994:SAB

[Mak94] Udi E. Makov. Some aspects of Bayesian loss-robustness. Journal of Statistical Planning and Inference, 38(3):359–370, March 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375894900167.

Malec:1995:SMO

[Mal95] Donald Malec. Selecting multiple-objective fixed-cost sample designs using an admissibility criterion. Journal of Statistical Planning and Inference, 48(2):229–240, November 15, 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589400153M.

Mallick:1998:BCE

[Mal98] Bani. K. Mallick. Bayesian curve estimation by polynomial of random order. Journal of Statistical Planning and Inference, 70(1): 91–109, July 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375897001791.

Mansouri:1990:RTO

[Man90a] H. Mansouri. Rank tests for ordered alternatives in analysis of variance. Journal of Statistical Planning and Inference, 24(1):107–117, January 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837589090021L.

Manteiga:1990:ANG

[Man90b] Wenceslao GonzÁlez Manteiga. Asymptotic normality of generalized functional estimators dependent on covariables. Journal of Statistical Planning and Inference, 24(3):377–390, March 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375890900572.

Mandeli:1995:CAO

[Man95] John P. Mandeli. Construction of asymmetrical orthogonal arrays having factors with a large non-prime power number of levels. Journal of Statistical Planning and Inference, 47(3):377–391, October 15, 1995. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589400143J.

Mansouri:1998:LDA

[Man98] Hossein Mansouri. Limiting distribution of the aligned rank transform tests in balanced incomplete blocks. Journal of Statistical Planning and Inference, 74(2):353–364, November 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375898001086.

Mansouri:1999:ART

[Man99] H. Mansouri. Aligned rank transform tests in linear models. Journal of Statistical Planning and Inference, 79(1):141–155, June 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375898002298.

Martin:1996:SRU

[Mar96] R. J. Martin. Some results on unilateral ARMA lattice processes. Journal of Statistical Planning and Inference, 50(3):395– 411, March 1996. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375895000666.

Markiewicz:1997:PIM

[Mar97a] Augustyn Markiewicz. Properties of information matrices for linear models and universal optimality of experimental designs. Journal of Statistical Planning and Inference, 59(1):127–137, March 30, 1997. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375896001061.

Martin:1997:TDU

[Mar97b] R. J. Martin. A three-dimensional unilateral autoregressive lattice process. Journal of Statistical Planning and Inference, 59(1): 1–18, March 30, 1997. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375896000900.

Matthews:1994:MOD

[Mat94] J. N. S. Matthews. Modelling and optimality in the design of crossover studies for medical applications. Journal of Statistical Planning and Inference, 42(1–2):89–108, November 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894901910.

Maul:1998:DTS

[Mau98] A. Maul. A discrete time sequential process for analyzing censored survival data using the likelihood ratio. Journal of Statistical Planning and Inference, 71(1–2):271–285, August 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S037837589800086X.

Maymin:1990:WSP

[May90] Zakhar Maymin. For what subsets of parameters is the least squares estimator minimax? Journal of Statistical Planning and Inference, 25(2):199–204, June 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375890900697.

Mitra:1994:NFL

[MB94a] Murari Mitra and Sujit K. Basu. On a nonparametric family of life distributions and its dual. Journal of Statistical Planning and Inference, 39(3):385–397, May 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894900949.

Mukhopadhyay:1994:PPT

[MB94b] P. Mukhopadhyay and S. Bhattacharyya. Predicting a population total under balanced samples. Journal of Statistical Planning and Inference, 39(1):85–93, April 1, 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894901147.

MacEachern:1995:AID

[MB95] Steven N. MacEachern and L. Mark Berliner. Asymptotic inference for dynamical systems observed with error. Journal of Statistical Planning and Inference, 46(3):277–292, August 15, 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589400117E.

Mitra:1996:SML

[MB96] Murari Mitra and Sujit K. Basu. Shock models leading to nonmonotonic ageing classes of life distributions. Journal of Statistical Planning and Inference, 55(2):131–138, October 30, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895001921.

Miller:1997:ERP

[MB97a] Gregory K. Miller and U. Narayan Bhat. Estimation for renewal processes with unobservable gamma or Erlang interarrival times. Journal of Statistical Planning and Inference, 61(2):355– 372, June 16, 1997. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375896001760.

Mukhopadhyay:1997:BAI

[MB97b] Chiranjit Mukhopadhyay and Asit P. Basu. Bayesian analysis of incomplete time and cause of failure data. Journal of Statistical Planning and Inference, 59(1):79–100, March 30, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896001036.

Markatou:1997:WLE

[MBL97] Marianthi Markatou, Ayanendranath Basu, and Bruce Lindsay. Weighted likelihood estimating equations: the discrete case with applications to logistic regression. Journal of Statistical Planning and Inference, 57(2):215–232, February 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896000456.

Moreno:1999:DBA

[MBR99] Elías Moreno, Francesco Bertolino, and Walter Racugno. Default Bayesian analysis of the Behrens–Fisher problem. Journal of Statistical Planning and Inference, 81(2):323–333, November 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899000701.

Maul:1999:SPA

[MBS99] A. Maul, A. A. Bartolucci, and K. P. Singh. A stochastic process applied to sequential parametric analysis of censored survival data. Journal of Statistical Planning and Inference, 78(1–2): 191–204, May 1999. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375898002134.

Mukerjee:1990:AOD

[MC90] Rahul Mukerjee and Arijit Chaudhuri. Asymptotic optimality of double sampling plans employing generalized regression estimators. Journal of Statistical Planning and Inference, 26(2):173– 183, October 1990. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589090125E.

Moreno:1995:CBP

[MC95] Elías Moreno and Juan A. Cano. Classes of bidimensional priors specified on a collection of set: Bayesian robustness. Journal of Statistical Planning and Inference, 46(3):325–334, August 15, 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589400134H.

McKenna:1992:PCF

[McK92] Thomas M. McKenna. Physiological constraints on the formal representation of neurons. Journal of Statistical Planning and Inference, 33(1):67–98, October 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375892900949.

McKenzie:1997:SSA

[McK97] Eddie McKenzie. Simulation of a stationary autoregression: a characterization of the normal distribution. Journal of Statistical Planning and Inference, 63(2):265–270, October 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897000220.

McLeish:1999:DFS

[McL99] D. L. McLeish. Designing the future: a simple algorithm for sequential design of a generalized linear model. Journal of Statistical Planning and Inference, 78(1–2):205–218, May 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898002146.

Mukerjee:1994:CRS

[MD94] Rahul Mukerjee and A. M. Dean. Construction of regular, single-replicate, two-factor designs. Journal of Statistical Planning and Inference, 38(1):89–104, January 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375892001503.

Martin:1991:OIB

[ME91] R. J. Martin and J. A. Eccleston. Optimal incomplete block designs for general dependence structures. Journal of Statistical Planning and Inference, 28(1):67–81, May 1991. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589190060R.

Martin:1993:IBD

[ME93] R. J. Martin and J. A. Eccleston. Incomplete block designs with spatial layouts when observations are dependent. Journal of Statistical Planning and Inference, 35(1):77–91, April 1993. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390069I.

Meczarski:1992:FWI

[Mec92] Marek Meczarski. Fixed-width interval estimation of the minimum point of a regression function based on the Kiefer–Wolfowitz procedure. Journal of Statistical Planning and Inference, 30(3): 339–349, March 1992. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589290160T.

Martin:1993:RLB

[MEG93] R. J. Martin, J. A. Eccleston, and A. C. Gleeson. Robust linear block designs for a suspected LV model. Journal of Statistical Planning and Inference, 34(3):433–450, March 1993. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375893901505.

Meier:1997:ANR

[Mei97] Ulrich Meier. On the asymptotic normality of rank tests for independence. Journal of Statistical Planning and Inference, 61(2):279–296, June 16, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375896001796.

Meier:1998:ANT

[Mei98] Ulrich Meier. Asymptotic normality of two-sample linear rank statistics under association. Journal of Statistical Planning and Inference, 66(2):235–252, March 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837589700089X.

Martin:1998:SRM

[MEJ98] R. J. Martin, J. A. Eccleston, and G. Jones. Some results on multi-level factorial designs with dependent observations. Journal of Statistical Planning and Inference, 73(1–2):91–111, September 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375898000548.

Meyer:1999:EMP

[Mey99] Mary C. Meyer. An extension of the mixed primal-dual bases algorithm to the case of more constraints than dimensions. Journal of Statistical Planning and Inference, 81(1):13–31, October 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899000257.

Moser:1996:NSA

[MF96] Barry Kurt Moser and Douglas Faries. A new sequential approximation method. Journal of Statistical Planning and Inference, 55(2):175–189, October 30, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375896888434.

Montepiedra:1997:MBD

[MF97] Grace Montepiedra and Valery V. Fedorov. Minimum bias designs with constraints. Journal of Statistical Planning and Inference, 63(1):97–111, September 30, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896001991.

Mackenzie-Fleming:1998:IFN

[MFS98] Kirsten Mackenzie-Fleming and Ken W. Smith. An infinite family of non-embeddable quasi-residual designs. Journal of Statistical Planning and Inference, 73(1–2):77–83, September 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898000524.

Mukerjee:1995:EDB

[MG95] Rahul Mukerjee and Sudhir Gupta. A-efficient designs for bioassays. Journal of Statistical Planning and Inference, 48 (2):247–259, November 15, 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837589500004S.

Mallick:1996:SEV

[MG96a] Bani K. Mallick and Alan E. Gelfand. Semiparametric errorsin-variables models: a Bayesian approach. Journal of Statistical Planning and Inference, 52(3):307–321, July 1, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895001395.

Muller-Gronbach:1996:ODA

[MG96b] Thomas Müller-Gronbach. Optimal designs for approximating the path of a stochastic process. Journal of Statistical Planning and Inference, 49(3):371–385, February 1, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895000178.

Moreno:1998:EIC

[MG98a] Elías Moreno and Javier Girón. Estimating with incomplete count data: a Bayesian approach. Journal of Statistical Planning and Inference, 66(1):147–159, January 5, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897000736.

Muller-Gronbach:1998:HCD

[MG98b] Thomas Müller-Gronbach. Hyperbolic cross designs for approximation of random fields. Journal of Statistical Planning and Inference, 66(2):321–344, March 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897000888.

Makhuvha:1996:BTB

[MGdW96] V. T. Makhuvha, P. C. N. Groenewald, and D. J. de Waal. Bayesian tests for the balanced two-way analysis of variance model. Journal of Statistical Planning and Inference, 53(3):345– 359, August 30, 1996. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375895001638.

Mandal:1992:MDE

[MH92a] Nripes Kumar Mandal and Berthold Heiligers. Minimax designs for estimating the optimum point in a quadratic response surface.

Journal of Statistical Planning and Inference, 31(2):235–244, May 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589290032N.

Markatou:1992:AAE

[MH92b] Marianthi Markatou and Thomas P. Hettmansperger. Applications of the asymmetric eigenvalue problem techniques to robust testing. Journal of Statistical Planning and Inference, 31(1): 51–65, April 1992. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589290040Y.

Mukhopadhyay:1995:BOD

[MH95] Saurabh Mukhopadhyay and Linda M. Haines. Bayesian Doptimal designs for the exponential growth model. Journal of Statistical Planning and Inference, 44(3):385–397, ???? 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894000562.

Mudholkar:1997:CCA

[MH97] Govind S. Mudholkar and Alan D. Hutson. Continuity corrected approximations for and ‘exact’ inference with Pearson’s χ^{2}. Journal of Statistical Planning and Inference, 59(1):61–78, March 30, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375896000912.

Mudholkar:1998:MAM

[MH98] Govind S. Mudholkar and Alan D. Hutson. LQ-moments: Analogs of L-moments. Journal of Statistical Planning and Inference, 71(1–2):191–208, August 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898000883.

Menon:1990:PAU

[MI90] V. V. Menon and N. K. Indira. A Poisson approximation in an urn model with indistinguishable balls. Journal of Statistical Planning and Inference, 26(1):93–101, September 1990. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589090098F.

Michie:1990:PMR

[Mic90] D. Michie. Personal models of rationality. Journal of Statistical Planning and Inference, 25(3):381–399, July 1990. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375890900837.

Miescke:1990:ORP

[Mie90] Klaus J. Miescke. Optimum replacement policies for using the most reliable components. Journal of Statistical Planning and Inference, 26(3):267–276, November 1990. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589090131D.

Mikaeili:1993:ODF

[Mik93] Fattah Mikaeili. D-optimum design for full cubic on q-simplex. Journal of Statistical Planning and Inference, 35(1):121–130, April 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589390072E.

Mikosch:1994:ASC

[Mik94] T. Mikosch. Almost sure convergence of bootstrapped means and U-statistics. Journal of Statistical Planning and Inference, 41(1): 1–19, August 1994. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375894901511.

Martin:1998:SRT

[MJE98] R. J. Martin, G. Jones, and J. A. Eccleston. Some results on two-level factorial designs with dependent observations. Journal of Statistical Planning and Inference, 66(2):363–384, March 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375897000864.

Maddala:1996:SCU

[MK96a] G. S. Maddala and In-Moo Kim. Structural change and unit roots. Journal of Statistical Planning and Inference, 49(1):73– 103, January 1996. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375895000313.

McKee:1996:MRH

[MK96b] Terry A. McKee and Harry J. Khamis. Multigraph representations of hierarchical loglinear models. Journal of Statistical Planning and Inference, 53(1):63–74, August 1, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895001409.

Mehra:1993:DRS

[MKR93] K. L. Mehra, Y. S. Rama Krishnaiah, and M. Sudhakara Rao. A dominance result for self-dual lattice paths and applications. Journal of Statistical Planning and Inference, 34(2):219–224, February 1993. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589390006R.

Madi:1996:BES

[ML96] Mohamed T. Madi and Tom Leonard. Bayesian estimation for shifted exponential distributions. Journal of Statistical Planning and Inference, 55(3):345–351, November 8, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375895001999.

Mathar:1993:PMF

[MM93] Rudolf Mathar and Renate Meyer. Preorderings, monotone functions, and best rank r approximations with applications to classical MDS. Journal of Statistical Planning and Inference, 37(3):291–305, December 1993. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390108I.

Morris:1995:EDC

[MM95] Max D. Morris and Toby J. Mitchell. Exploratory designs for computational experiments. Journal of Statistical Planning and Inference, 43(3):381–402, February 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589400035T.

Markatou:1996:RTN

[MM96] Marianthi Markatou and George Manos. Robust tests in nonlinear regression models. Journal of Statistical Planning and Inference, 55(2):205–217, October 30, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895001948.

Mathew:1998:BAJ

[MM98a] G. Mathew and W. P. McCormick. A bootstrap approximation to the joint distribution of sum and maximum of a stationary sequence. Journal of Statistical Planning and Inference, 70(2):287– 299, July 15, 1998. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375898000020.

McGilchrist:1998:RRG

[MM98b] C. A. McGilchrist and K. M. Matawie. Recursive residuals in generalised linear models. Journal of Statistical Planning and Inference, 70(2):335–344, July 15, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897001924.

Myers:1994:SAO

[MMC94] William R. Myers, Raymond H. Myers, and Walter H. Carter. Some alphabetic optimal designs for the logistic regression model. Journal of Statistical Planning and Inference, 42(1–2):57–77, November 1994. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375894901899.

Mudholkar:1995:RFI

[MMM95] Govind S. Mudholkar, Michael P. McDermott, and Anil Mudholkar. Robust finite-intersection tests for homogeneity of ordered variances. Journal of Statistical Planning and Inference, 43(1–2):185–195, January 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589400018Q.

Mitchell:1994:AOE

[MMY94] Toby Mitchell, Max Morris, and Donald Ylvisaker. Asymptotically optimum experimental designs for prediction of deterministic functions given derivative information. Journal of Statistical Planning and Inference, 41(3):377–389, October 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894900302.

Morgenthaler:1997:CTM

[MN97] Stephan Morgenthaler and Anna Nicolaou. Composite transformation models: a fiducial perspective. Journal of Statistical Planning and Inference, 57(1):49–57, January 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896000353.

MacEachern:1995:RUD

[MNWZ95] S. N. MacEachern, W. I. Notz, D. C. Whittinghill, and Y. Zhu. Robustness to the unavailability of data in the linear model, with applications. Journal of Statistical Planning and Inference, 48(2):207–213, November 15, 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589500002Q.

Morimune:1996:LIE

[MO96] Kimio Morimune and Kosuke Oya. Limited information estimation and testing subject to linear constraints. Journal of Statistical Planning and Inference, 50(2):223–240, March 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895000550.

Mohanty:1993:P

[Moh93] S. G. Mohanty. Preface. Journal of Statistical Planning and Inference, 34(1):3–4, January 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375893900285.

Mohanty:1996:P

[Moh96] S. G. Mohanty. Preface. Journal of Statistical Planning and Inference, 54(1):1–2, September 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896900075.

Molinski:1996:MCI

[Mol96] Krzysztof Moliński. The modification of confidence intervals for variance components in one-way random model using Stein’s approach. Journal of Statistical Planning and Inference, 52(3):323– 329, July 1, 1996. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375895001344.

Montakhab:1990:EQR

[Mon90] M. S. Montakhab. Embedding of quasi-residual divisible semisymmetric designs with parameters (r, λ, c). Journal of Statistical Planning and Inference, 24(1):55–67, January 1990. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589090016N.

Monod:1991:SBF

[Mon91] Hervé Monod. Serially balanced factorial sequences. Journal of Statistical Planning and Inference, 27(3):325–333, March 1991. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589190045G.

Monticino:1998:CPD

[Mon98] Michael Monticino. Constructing prior distributions with trees of exchangeable processes. Journal of Statistical Planning and Inference, 73(1–2):113–133, September 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837589800055X.

Moorhouse:1996:SRR

[Moo96] G. Eric Moorhouse. Some p-ranks related to Hermitian varieties. Journal of Statistical Planning and Inference, 56(2):229– 241, December 1996. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375896000201.

Morgan:1990:SSC

[Mor90] John P. Morgan. Some series constructions for two-dimensional neighbor designs. Journal of Statistical Planning and Inference, 24(1):37–54, January 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837589090015M.

Morgan:1997:PYD

[Mor97] John P. Morgan. On pairs of Youden designs. Journal of Statistical Planning and Inference, 60(2):367–387, May 15, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897897121.

Morgan:1998:PSB

[Mor98] J. P. Morgan. Properties of superimposed BIBDs. Journal of Statistical Planning and Inference, 73(1–2):135–148, September 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375898000561.

Moreno:1993:BRH

[MP93] Elías Moreno and Luís Raúl Pericchi. Bayesian robustness for hierarchical ϵ-contamination models. Journal of Statistical Planning and Inference, 37(2):159–167, November 1993. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390085K.

Mathon:1996:BTD

[MP96] Rudolf Mathon and Cheryl E. Praeger. On block-transitive designs with affine automorphism group. Journal of Statistical Planning and Inference, 56(2):203–221, December 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896000195.

Mammen:1997:OSA

[MP97] Enno Mammen and Byeong U. Park. Optimal smoothing in adaptive location estimation. Journal of Statistical Planning and Inference, 58(2):333–348, March 15, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

S0378375896000857. See erratum [Ano98q].

Muller:1998:DMA

[MP98] Werner G. Müller and Andrej Pázman. Design measures and approximate information matrices for experiments without replications. Journal of Statistical Planning and Inference, 71(1–2):349– 362, August 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375898000135.

Morales:1994:APD

[MPSM94] D. Morales, L. Pardo, M. Salicrú, and M. L. Menéndez. Asymptotic properties of divergence statistics in a stratified random sampling and its applications to test statistical hypotheses. Journal of Statistical Planning and Inference, 38(2):201–221, February 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375894900353.

Morales:1995:ADE

[MPV95] D. Morales, L. Pardo, and I. Vajda. Asymptotic divergence of estimates of discrete distributions. Journal of Statistical Planning and Inference, 48(3):347–369, December 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589500013Y.

Meisner:1997:CNA

[MPW97] D. B. Meisner, F. C. Piper, and P. R. Wild. A class of nonabelian 2-groups containing Menon difference sets. Journal of Statistical Planning and Inference, 62(1):57–62, July 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896001681.

Morales:1998:IDR

[MPZ98] D. Morales, L. Pardo, and K. Zografos. Informational distances and related statistics in mixed continuous and categorical variables. Journal of Statistical Planning and Inference, 75(1):47–63, November 15, 1998. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375898001207.

Miescke:1996:P

[MR96] K. J. Miescke and D. Rasch. Preface. Journal of Statistical Planning and Inference, 54(2):141–143, September 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375896900312.

Macula:1998:SST

[MR98]  Anthony J. Macula and George R. Reuter.  Simplified

searching for two defects. Journal of Statistical Planning and Inference, 66(1):77–82, January 5, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837589700092X.

Marinucci:1999:AFF

[MR99] D. Marinucci and P. M. Robinson. Alternative forms of fractional Brownian motion. Journal of Statistical Planning and Inference, 80(1–2):111–122, August 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898002456.

Mason:1990:NSC

[MS90a] David M. Mason and Galen R. Shorack. Necessary and sufficient conditions for asymptotic normality of trimmed L-statistics. Journal of Statistical Planning and Inference, 25(2):111–139, June 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589090061X.

Milbrodt:1990:APT

[MS90b] Hartmut Milbrodt and Helmut Strasser. On the asymptotic power of the two-sided Kolmogorov–Smirnov test. Journal of Statistical Planning and Inference, 26(1):1–23, September 1990. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375890900918.

Mukhopadhyay:1992:ASP

[MS92] Nitis Mukhopadhyay and Tumulesh K. S. Solanky. Accelerated sequential procedure for selecting the best exponential population. Journal of Statistical Planning and Inference, 32(3):347– 361, September 1992. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589290016L.

Mi:1993:CBS

[MS93a] Jie Mi and Allan R. Sampson. A comparison of the Bonferroni and Scheffé bounds. Journal of Statistical Planning and Inference, 36(1):101–105, July 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390105F.

Muller:1993:VFE

[MS93b] Hans-Georg Müller and Ulrich Stadtmüller. On variance function estimation with quadratic forms. Journal of Statistical Planning and Inference, 35(2):213–231, May 1993. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375893900469.

Miescke:1995:OMR

[MS95a] K. J. Miescke and Dylan Shi. Optimum Markov replacement policies for using the most reliable components. Journal of Statistical Planning and Inference, 45(3):331–346, June 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589400085A.

Mukerjee:1995:OEF

[MS95b] Rahul Mukerjee and S. Sengupta. Optimal estimation of a finite multidimensional population total. Journal of Statistical Planning and Inference, 48(3):339–346, December 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589500009X.

Mathon:1997:PSD

[MS97] Rudolf Mathon and Anne Penfold Street. Partitions of sets of designs on seven, eight and nine points. Journal of Statistical Planning and Inference, 58(1):135–150, March 1, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896000663.

Meerschaert:1998:SRE

[MS98a] Mark M. Meerschaert and Hans-Peter Scheffler. A simple robust estimation method for the thickness of heavy tails. Journal of Statistical Planning and Inference, 71(1–2):19–34, August 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375898000937.

Mukhopadhyay:1998:MMF

[MS98b] Nitis Mukhopadhyay and Tumulesh K. S. Solanky. Multistage methodologies for fixed-width simultaneous confidence intervals for all pairwise comparisons. Journal of Statistical Planning and Inference, 73(1–2):163–176, September 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898000585.

Mian:1991:MLE

[MST91] I. U. H. Mian, M. M. Shoukri, and D. S. Tracy. Maximum likelihood estimation of sibling correlations in the analysis of family data. Journal of Statistical Planning and Inference, 27(1):125– 141, January 1991. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589190086T.

McDougall:1997:OTS

[MST97] A. J. McDougall, D. S. Stoffer, and D. E. Tyler. Optimal transformations and the spectral envelope for real-valued time series. Journal of Statistical Planning and Inference, 57(2):195– 214, February 1997. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375896000444.

Mardia:1999:BML

[MST99] K. V. Mardia, H. R. Southworth, and C. C. Taylor. On bias in maximum likelihood estimators. Journal of Statistical Planning and Inference, 76(1–2):31–39, February 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898001761.

Monga:1994:CCD

[MT94] Monga and S. Tardif. A class of conditionally distribution-free rank tests for replicated latin square designs. Journal of Statistical Planning and Inference, 41(3):365–376, October 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894900299.

Majumdar:1996:RCD

[MT96] Dibyen Majumdar and Ajit C. Tamhane. Row-column designs for comparing treatments with a control. Journal of Statistical Planning and Inference, 49(3):387–400, February 1, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895000208.

Malamud:1999:SAT

[MT99] Bruce D. Malamud and Donald L. Turcotte. Self-affine time series: measures of weak and strong persistence. Journal of Statistical Planning and Inference, 80(1–2):173–196, August 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375898002493.

Morgan:1993:OCN

[MU93] John P. Morgan and Nizam Uddin. Optimality and construction of nested row and column designs. Journal of Statistical Planning and Inference, 37(1):81–93, October 1993. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390081G.

Morgan:1997:OSB

[MU97] John P. Morgan and Nizam Uddin. Orthogonal sets of balanced incomplete block designs with nested rows and column. Journal of Statistical Planning and Inference, 59(2):369–378, April 15, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375896001188.

Mukhopadhyay:1994:PFP

[Muk94] P. Mukhopadhyay. Prediction in finite population under errorin-variables superpopulation models. Journal of Statistical Planning and Inference, 41(2):151–161, September 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894901600.

Mukerjee:1998:BOM

[Muk98a] Rahul Mukerjee. On balanced orthogonal multi-arrays: Existence, construction and application to design of experiments.

Journal of Statistical Planning and Inference, 73(1–2):149–162, September 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375898000573.

Mukherjee:1998:PTS

[Muk98b] Kanchan Mukherjee. On preliminary test and shrinkage estimation in linear models with long-memory errors. Journal of Statistical Planning and Inference, 69(2):319–328, June 15, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897001602.

Muller:1993:SDO

[Mül93] Marianne Müller. Supersaturated designs for one or two effective factors. Journal of Statistical Planning and Inference, 37(2):237– 244, November 1993. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589390092K.

Muller:1994:ODR

[Mül94] Christine Müller. Optimal designs for robust estimation in conditionally contaminated linear models. Journal of Statistical Planning and Inference, 38(1):125–139, January 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589200155W.

Muller:1995:BPD

[Mül95a] Christine H. Müller. Breakdown points for designed experiments. Journal of Statistical Planning and Inference, 45(3):413– 427, June 1995. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589400086B.

Muller:1995:MED

[Mül95b] Christine H. Müller. Maximin efficient designs for estimating nonlinear aspects in linear models. Journal of Statistical Planning and Inference, 44(1):117–132, March 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589400042T.

Muller:1996:ODL

[Mül96] Werner G. Müller. Optimal design for local fitting. Journal of Statistical Planning and Inference, 55(3):389–397, November 8, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375895001972.

Mullen:1998:OHR

[Mul98] Gary L. Mullen. Orthogonal hypercubes and related designs. Journal of Statistical Planning and Inference, 73(1–2):177–188, September 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375898000597.

Mustafi:1991:MAD

[Mus91] C. K. Mustafi. The median absolute deviation of residual life data and weak convergence of a related stochastic process. Journal of Statistical Planning and Inference, 27(2):231–236, February 1991. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589190018A.

Mukhopadhyay:1996:CSD

[MV96] P. Mukhopadhyay and K. Vijayan. On controlled sampling designs. Journal of Statistical Planning and Inference, 52(3):375– 378, July 1, 1996. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375895001255.

Muliere:1997:BNA

[MW97] Pietro Muliere and Stephen Walker. A Bayesian nonparametric approach to determining a maximum tolerated dose. Journal of Statistical Planning and Inference, 61(2):339–353, June 16, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375896001590.

Mandrell:1996:GCS

[MY96a] R. Mandrell and J. Yucas. A geometric construction of the Steiner system S (5, 8, 24). Journal of Statistical Planning and Inference, 56(2):223–228, December 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896000274.

Miao:1996:CV

[MY96b] Ying Miao and Songlin Yang. Concerning the vector ν(m, t). Journal of Statistical Planning and Inference, 51(2):223–227, April 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375895000879.

Maronna:1997:RES

[MY97] Ricardo A. Maronna and Víctor J. Yohai. Robust estimation in simultaneous equations models. Journal of Statistical Planning and Inference, 57(2):233–244, February 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896000468.

Maekawa:1996:EDR

[MYTH96] K. Maekawa, T. Yamamoto, Y. Takeuchi, and M. Hatanaka. Estimation in dynamic regression with an integrated process. Journal of Statistical Planning and Inference, 49(2):279–303, January 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375895000410.

Nandram:1995:BCP

[Nan95] B. Nandram. Bayesian cuboid prediction intervals: an application to tensile-strength prediction. Journal of Statistical Planning and Inference, 44(2):167–180, April 1, 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589595021N.

Navidi:1995:BMP

[Nav95] W. C. Navidi. Bootstrapping a method of phylogenetic inference. Journal of Statistical Planning and Inference, 43(1–2):169– 184, January 1995. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589400017P.

Nayak:1990:ELS

[Nay90] Tapan K. Nayak. Estimation of location and scale parameters using generalized Pitman nearness criterion. Journal of Statistical Planning and Inference, 24(2):259–268, February 1990. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589090046W.

Nayak:1992:SAS

[Nay92] Tapan K. Nayak. On statistical analysis of a sample from a population of unknown species. Journal of Statistical Planning and Inference, 31(2):187–198, May 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290029R.

Nze:1998:FET

[ND98] P. Ango Nze and P. Doukhan. Functional estimation for time series: Uniform convergence properties. Journal of Statistical Planning and Inference, 68(1):5–29, May 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837589700133X.

Nelder:1997:CDF

[Nel97] John Nelder. Contribution to discussion on A. F. Desmond’s paper: Optimal estimating functions, quasi-likelihood and statistical modelling. Journal of Statistical Planning and Inference, 60(1):108, April 30, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375897900124.

See [Les97].

Neuhaus:1999:TTO

[Neu99] Georg Neuhaus. On testing tumor onset times. Journal of Statistical Planning and Inference, 76(1–2):57–78, February 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375898001724.

Nevzorov:1995:ADR

[Nev95] Valery Nevzorov. Asymptotic distributions of records in nonstationary schemes. Journal of Statistical Planning and Inference, 45(1–2):261–273, ???? 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837589400076X.

Nath:1991:CSC

[NG91] Prem Nath and Kanwar Kumar Gulati. Cyclic symmetry on cyclic and noncyclic sets of probability distributions and the entropies of degree α. Journal of Statistical Planning and Inference, 27(2):237–247, February 1991. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589190019B.

Nieuwenhuis:1992:CLT

[Nie92] Gert Nieuwenhuis. Central limit theorems for sequences with m(n)-dependent main part. Journal of Statistical Planning and Inference, 32(2):229–241, August 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290049X.

Niederhausen:1993:EIS

[Nie93] Heinrich Niederhausen. The enumeration of integer sequences with a given number of colored records. Journal of Statistical Planning and Inference, 34(2):225–237, February 1993. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390007S.

Niederhausen:1996:SSS

[Nie96] Heinrich Niederhausen. Symmetric Sheffer sequences and their applications to lattice path counting. Journal of Statistical Planning and Inference, 54(1):87–100, September 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589500159X.

Nikitin:1991:HLC

[Nik91] Ya. Yu. Nikitin. Hodges-Lehmann and Chernoff efficiencies of linear rank statistics. Journal of Statistical Planning and Inference, 29(3):309–323, November 1991. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589190006Z.

Nishisato:1995:GRQ

[Nis95] Shizuhiko Nishisato. Graphical representation of quantified categorical data: its inherent problems. Journal of Statistical Planning and Inference, 43(1–2):121–132, January 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589400013L.

Niyonsenga:1997:RPE

[Niy97] Théophile Niyonsenga. Response probability estimation. Journal of Statistical Planning and Inference, 59(1):111–126, March 30, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375896000948.

Nanda:1999:SWD

[NJ99] Asok K. Nanda and Kanchan Jain. Some weighted distribution results on univariate and bivariate cases. Journal of Statistical Planning and Inference, 77(2):169–180, March 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898001906.

Naik:1994:TRE

[NK94] Dayanand N. Naik and Ravindra Khattree. Testing for random effects in analysis of covariance model. Journal of Statistical Planning and Inference, 38(3):381–387, March 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894900183.

Nebebe:1995:DPO

[NK95] Fassil Nebebe and Andy C. C. Kwan. Detecting ‘potential’ outliers in time series: Some Monte Carlo evidence on dummy variable plot approaches. Journal of Statistical Planning and Inference, 43(1–2):87–95, January 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589400010S.

Ng:1997:FTS

[NL97] Tze-Man Ng and Wai-Keung Li. F-tests for seasonal differencing with a break-point. Journal of Statistical Planning and Inference, 65(1):87–107, December 1, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897000530.

Nguyen:1997:FFD

[NM97] Nam-Ky Nguyen and Alan J. Miller. 2^{m} fractional factorial designs of resolution V with high A-efficiency, 7 ≤ m ≤ 10. Journal of Statistical Planning and Inference, 59(2):379–384, April 15, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S037837589600122X.

Nagaraja:1997:CBR

[NN97] H. N. Nagaraja and V. B. Nevzorov. On characterizations based on record values and order statistics. Journal of Statistical Planning and Inference, 63(2):271–284, October 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897844503.

Nyblom:1993:ICM

[NP93] Jukka Nyblom and Tarmo M. Pukkila. Information criterion as a multiple testing procedure. Journal of Statistical Planning and Inference, 36(2–3):385–397, August/September 1993. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390139W.

Nieuwenhuis:1990:SSI

[NR90] Gert Nieuwenhuis and Frits Ruymgaart. Some stochastic inequalities and asymptotic normality of serial rank statistics in general linear processes. Journal of Statistical Planning and Inference, 25 (1):53–79, May 1990. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589090007H.

Nasri-Roudsari:1996:EVT

[NR96] Dirk Nasri-Roudsari. Extreme value theory of generalized order statistics. Journal of Statistical Planning and Inference, 55(3):281–297, November 8, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375895002006.

Navarro:1998:UAC

[NRZ98] Jorge Navarro, JoséM. Ruiz, and Noemi Zoroa. A unified approach to characterization problems using conditional expectations. Journal of Statistical Planning and Inference, 69(2):193– 207, June 15, 1998. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375897001614.

Norvaisa:1991:ABL

[NZ91] Rimas Norvaisa and Ricardas Zitikis. Asymptotic behaviour of linear combinations of functions of order statistics. Journal of Statistical Planning and Inference, 28(3):305–317, July 1991. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589190069Q.

Oh:1999:CVP

[OD99] Hyun Sook Oh and Anirban DasGupta. Comparison of the P -value and posterior probability. Journal of Statistical Planning and Inference, 76(1–2):93–107, February 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898001311.

Ogata:1996:ESB

[Oga96] Yosihiko Ogata. Evaluation of spatial Bayesian models — two computational methods. Journal of Statistical Planning and Inference, 51(1):1–18, April 1, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895000674.

Ogawa:1998:OSS

[Oga98] Junjiro Ogawa. Optimal spacing of the selected sample quantiles for the joint estimation of the location and scale parameters of a symmetric distribution. Journal of Statistical Planning and Inference, 70(2):345–360, July 15, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897001948.

OHagan:1991:BHQ

[O’H91] A. O’Hagan. Bayes-Hermite quadrature. Journal of Statistical Planning and Inference, 29(3):245–260, November 1991. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589190002V.

OHagan:1994:RMA

[O’H94] A. O’Hagan. Robust modelling for asset management. Journal of Statistical Planning and Inference, 40(2–3):245–259, July 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375894901244.

Ohtani:1997:MMS

[Oht97] Kazuhiro Ohtani. Minimum mean squared error estimation of each individual coefficient in a linear regression model. Journal of Statistical Planning and Inference, 62(2):301–316, August 15, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375896001802.

Oman:1997:RTC

[Oma97] Samuel D. Oman. A remark on testing covariance structure in a mixed model. Journal of Statistical Planning and Inference, 59(2):355–357, April 15, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896001073.

Omey:1995:SRV

[Ome95] E. Omey. On a subclass of regularly varying functions. Journal of Statistical Planning and Inference, 45(1–2):275–290, ???? 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375894000778.

OConnell:1995:GLC

[ON95] Michael O’Connell and Douglas Nychka. A generalized linear classification model with a smooth link function and predictors obtained from quantile spline fits to high-dimensional data. Journal of Statistical Planning and Inference, 47(1–2):153–164, October 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589400128I.

OBrien:1996:NDP

[OR96] Timothy E. O’Brien and John O. Rawlings. A nonsequential design procedure for parameter estimation and model discrimination in nonlinear regression models. Journal of Statistical Planning and Inference, 55(1):77–93, October 15, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895001824.

Ogunyemi:1998:OMI

[OR98] Theophilus O. Ogunyemi and M. Bhaskara Rao. On s- and (m, s)- optimality of minimal incomplete block designs. Journal of Statistical Planning and Inference, 73(1–2):189–195, September 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375898000603.

Osiewalski:1996:BAE

[OS96] Jacek Osiewalski and Mark F. J. Steel. A Bayesian analysis of exogeneity in models pooling time-series and cross-sectional data. Journal of Statistical Planning and Inference, 50(2):187– 206, March 1996. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375895000534.

Oskrochi:1995:ACC

[Osk95] G. Reza Oskrochi. Analysis of censored correlated observations. Journal of Statistical Planning and Inference, 47(1–2):165– 180, October 1995. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589400129J.

Padmawar:1994:SAN

[Pad94] V. R. Padmawar. Strategies admitting nonnegative unbiased variance estimators. Journal of Statistical Planning and Inference, 40 (1):81–95, June 1994. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375894901430.

Pakes:1997:CIU

[Pak97] Anthony G. Pakes. Characterization by invariance under length-biasing and random scaling. Journal of Statistical Planning and Inference, 63(2):285–310, October 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897000244.

Panchapakesan:1996:RRS

[Pan96] S. Panchapakesan. A review of robustness of selection procedures. Journal of Statistical Planning and Inference, 54(3):279– 290, September 1996. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375896001735.

Park:1994:OLH

[Par94] Jeong-Soo Park. Optimal Latin-hypercube designs for computer experiments. Journal of Statistical Planning and Inference, 39(1): 95–111, April 1, 1994. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375894901155.

Patil:1993:BCN

[Pat93] P. N. Patil. Bandwidth choice for nonparametric hazard rate estimation. Journal of Statistical Planning and Inference, 35(1): 15–30, April 1993. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589390064D.

Patel:1994:RMD

[Pat94] H. I. Patel. A repeated measures design with repeated randomization. Journal of Statistical Planning and Inference, 42(1–2):257– 270, November 1994. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375894902003.

Patil:1997:NHR

[Pat97] Prakash Patil. Nonparametric hazard rate estimation by orthogonal wavelet methods. Journal of Statistical Planning and Inference, 60(1):153–168, April 30, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896001279.

Payne:1998:PGQ

[Pay98] Stanley E. Payne. A problem on generalized quadrangles. Journal of Statistical Planning and Inference, 73(1–2):197–204, September 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375898000615.

Prabakaran:1994:SEE

[PC94] T. Edwin Prabakaran and B. Chandrasekar. Simultaneous equivariant estimation for location-scale models. Journal of Statistical Planning and Inference, 40(1):51–59, June 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894901414.

Park:1999:MCM

[PD99] Hyo-Il Park and M. M. Desu. A multivariate control median test. Journal of Statistical Planning and Inference, 79(1):123–139, June 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375898002328.

Pellikaan:1996:EEC

[Pel96] Ruud Pellikaan. On the existence of error-correcting pairs. Journal of Statistical Planning and Inference, 51(2):229–242, April 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375895000887.

Penev:1991:EES

[Pen91] Spiridon Penev. Efficient estimation of the stationary distribution for exponentially ergodic Markov chains. Journal of Statistical Planning and Inference, 27(1):105–123, January 1991. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589190085S.

Pfanzagl:1991:SCT

[Pfa91] J. Pfanzagl. A supplement to the Convolution Theorem. Journal of Statistical Planning and Inference, 28(1):25–31, May 1991. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589190056K.

Pfanzagl:1993:CAE

[Pfa93] J. Pfanzagl. A case of asymptotic equivalence between conditional and marginal maximum likelihood estimators. Journal of Statistical Planning and Inference, 35(3):301–307, June 1993. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375893900182.

Pfanzagl:1994:ISP

[Pfa94] J. Pfanzagl. On the identifiability of structural parameters in mixtures: applications to psychological tests. Journal of Statistical Planning and Inference, 38(3):309–326, March 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894900132.

Pfanzagl:1998:NCS

[Pfa98] J. Pfanzagl. The nonexistence of confidence sets for discontinuous functionals. Journal of Statistical Planning and Inference, 75(1):9–20, November 15, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375898001219.

Pan:1997:LIA

[PFvR97] Jian-Xin Pan, Kai-Tai Fang, and Dietrich von Rosen. Local influence assessment in the growth curve model with unstructured covariance. Journal of Statistical Planning and Inference, 62(2):263–278, August 15, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896001942.

Phelan:1996:TTB

[Phe96] Michael J. Phelan. Transient tracers and birth and death on flows: Parametric estimation of rates of drift, injection, and decay. Journal of Statistical Planning and Inference, 51(1):19– 40, April 1, 1996. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375895000682.

Pan:1999:ETA

[PK99] Guohua Pan and Ravindra Khattree. On estimation and testing arising from order restricted balanced mixed model. Journal of Statistical Planning and Inference, 77(2):281–292, March 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375898001918.

Pal:1997:NSO

[PL97] Nabendu Pal and Wooi K. Lim. A note on second-order admissibility of the Graybill–Deal estimator of a common mean of several normal populations. Journal of Statistical Planning and Inference, 63(1):71–78, September 30, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896002029.

Priebe:1997:SNA

[PMR97] Carey E. Priebe, David J. Marchette, and George W. Rogers. Semiparametric nonhomogeneity analysis. Journal of Statistical Planning and Inference, 59(1):45–60, March 30, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837589600095X.

Parsian:1996:ESP

[PN96] Ahmad Parsian and Nader Nematollahi. Estimation of scale parameter under entropy loss function. Journal of Statistical Planning and Inference, 52(1):77–91, June 1, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895000267.

Pollock:1993:CFA

[Pol93] D. S. G. Pollock. On the criterion function for ARMA estimation. Journal of Statistical Planning and Inference, 36(2– 3):413–430, August/September 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837589390141R.

Power:1995:MCR

[Pow95] Simon Power. Monte Carlo response surface: a comparative approach. Journal of Statistical Planning and Inference, 43(1–2):301–310, January 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837589400028T.

Padmanabhan:1992:RBE

[PP92a] A. R. Padmanabhan and Madan L. Puri. Rank-based estimation of the ratio of scale parameters and applications. Journal of Statistical Planning and Inference, 31(1):23–49, April 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290039U.

Pazman:1992:NED

[PP92b] Andrej Pázman and Luc Pronzato. Nonlinear experimental design based on the distribution of estimators. Journal of Statistical Planning and Inference, 33(3):385–402, December 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290007F.

Pericchi:1994:PRM

[PP94a] Luis Raúl Pericchi and María Eglée Pérez. Posterior robustness with more than one sampling model. Journal of Statistical Planning and Inference, 40(2–3):279–294, July 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894901260.

Polasek:1994:RBM

[PP94b] Wolfgang Polasek and Klaus Pötzelberger. Robust Bayesian methods in simple ANOVA models. Journal of Statistical Planning and Inference, 40(2–3):295–311, July 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894901279.

Pollock:1996:MDR

[PP96] D. S. G. Pollock and Evangelia Pitta. The misspecification of dynamic regression models. Journal of Statistical Planning and Inference, 49(2):223–239, January 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894000387.

Payne:1995:TTB

[PPO95] Clive Payne, Joan Payne, and Colm O’Muircheartaigh. Testing the tightening bond hypothesis: an application of new models for analysing trends in multiway tables. Journal of Statistical Planning and Inference, 47(1–2):181–201, October 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589400130N.

Pulskamp:1991:GCN

[PR91] Richard J. Pulskamp and Dan A. Ralescu. A general class of nonlinear admissible estimators in the one-parameter exponential case. Journal of Statistical Planning and Inference, 28(3): 383–390, July 1991. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589190074O.

Pena:1993:SSE

[PR93] Edsel A. Peña and Vijay K. Rohatgi. Small sample and efficiency results for the Nelson-aalen estimator. Journal of Statistical Planning and Inference, 37(2):193–202, November 1993. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390088N.

Peszek:1994:ELH

[PR94] Iza Peszek and Andrew L. Rukhin. Estimating lognormal hazard function. Journal of Statistical Planning and Inference, 41(3):281–290, October 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375894900248.

Politis:1996:FTK

[PR96] Dimitris N. Politis and Joseph P. Romano. On flat-top kernel spectral density estimators for homogeneous random fields. Journal of Statistical Planning and Inference, 51(1):41–53, April 1, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375895000690.

Pepe:1994:AOD

[PRF94] Margaret Sullivan Pepe, Marie Reilly, and Thomas R. Fleming. Auxiliary outcome data and the mean score method. Journal of Statistical Planning and Inference, 42(1–2):137–160, November 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375894901945.

Proctor:1993:BPP

[Pro93] Robert A. Proctor. A bideterminant proof of a product identity for plane partitions with symmetries. Journal of Statistical Planning and Inference, 34(2):239–250, February 1993. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390008T.

Provost:1996:CTG

[Pro96] Serge B. Provost. On Craig’s theorem and its generalizations. Journal of Statistical Planning and Inference, 53(3):311–321, August 30, 1996. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375895001506.

Pott:1997:MTP

[PRS97] Alexander Pott, Dirk Reuschling, and Bernhard Schmidt. A multiplier theorem for projections of affine difference sets. Journal of Statistical Planning and Inference, 62(1):63–67, July 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896001693.

Politis:1999:WCD

[PRW99] Dimitris Politis, Joseph P. Romano, and Michael Wolf. Weak convergence of dependent empirical measures with application to subsampling in function spaces. Journal of Statistical Planning and Inference, 79(2):179–190, July 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898001748.

Prasad:1990:EPD

[PS90a] B. Prasad and Radhey S. Singh. Estimation of prior distribution and empirical Bayes estimation in a nonexponential family. Journal of Statistical Planning and Inference, 24(1):81–86, January 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589090018P.

Puri:1990:RFI

[PS90b] Prem S. Puri and Harshinder Singh. On recursive formulas for isotonic regression useful for statistical inference under order restrictions. Journal of Statistical Planning and Inference, 24(1): 1–11, January 1990. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589090012J.

Puri:1992:F

[PS92] Madan L. Puri and Jadgish N. Srivastava. Foreword. Journal of Statistical Planning and Inference, 32(1):1, July 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290147K.

Puri:1995:Fa

[PS95a] Madan L. Puri and Jagdish N. Srivastava. Foreword. Journal of Statistical Planning and Inference, 45(1–2):3, ???? 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589400059X.

Puri:1995:Fb

[PS95b] Madan L. Puri and Jagdish N. Srivastava. Foreword. Journal of Statistical Planning and Inference, 47(1–2):1, October 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895900039.

Pan:1998:SSP

[PS98a] Guohua Pan and Thomas J. Santner. Selection and screening procedures to determine optimal product designs. Journal of Statistical Planning and Inference, 67(2):311–330, April 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897001225.

Park:1998:TPP

[PS98b] Meekyong G. Park and Jiayang Sun. Tests in projection pursuit regression. Journal of Statistical Planning and Inference, 75(1):65–90, November 15, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898001244.

Patil:1997:RSS

[PST97] G. P. Patil, A. K. Sinha, and C. Taillie. Ranked set sampling, coherent rankings and size-biased permutations. Journal of Statistical Planning and Inference, 63(2):311–324, October 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837589700030X.

Pacifico:1998:NGB

[PST98] M. Perone Pacifico, G. Salinetti, and L. Tardella. A note on the geometry of Bayesian global and local robustness. Journal of Statistical Planning and Inference, 69(1):51–64, June 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897001250.

Paul:1992:HTO

[PT92] S. R. Paul and K. Thiagarajah. Hypothesis tests for the one-way layout of type II censored data from Weibull populations. Journal of Statistical Planning and Inference, 33(3):367–380, December 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589290005D.

Pampallona:1994:GSD

[PT94] Sandro Pampallona and Anastasios A. Tsiatis. Group sequential designs for one-sided and two-sided hypothesis testing with provision for early stopping in favor of the null hypothesis. Journal of Statistical Planning and Inference, 42(1–2):19–35, November 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375894901872.

Paul:1997:LBC

[PT97] S. R. Paul and S. Thedchanamoorthy. Likelihood-based confidence limits for the common odds ratio. Journal of Statistical Planning and Inference, 64(1):83–92, October 30, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896002145.

Purkayastha:1995:ASR

[Pur95] Sumitra Purkayastha. An almost sure representation of sample circular median. Journal of Statistical Planning and Inference, 46 (1):77–91, July 1995. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589400092A.

Purohit:1997:RED

[Pur97] Sudha G. Purohit. Recurring event data: a general model and related inference procedures. Journal of Statistical Planning and Inference, 59(1):31–44, March 30, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896001012.

Perlman:1999:LCI

[PW99] Michael D. Perlman and Lang Wu. Lattice conditional independence models for contingency tables with non-monotone missing data patterns. Journal of Statistical Planning and Inference, 79(2):259–287, July 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375899000038.

Poston:1998:ODM

[PWS98] Wendy L. Poston, Edward J. Wegman, and Jeffrey L. Solka. D-optimal design methods for robust estimation of multivariate location and scatter. Journal of Statistical Planning and Inference, 73(1–2):205–213, September 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898000627.

Pan:1999:UOP

[PZ99] Xiao-Rong Pan and Mai Zhou. Using one-parameter sub-family of distributions in empirical likelihood ratio with censored data. Journal of Statistical Planning and Inference, 75(2):379–392, January 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375898001566.

Qian:1996:THS

[QGG96] Guoqi Qian, R. P. Gupta, and George Gabor. Test for homogeneity of several populations by stochastic complexity. Journal of Statistical Planning and Inference, 53(2):133–151, August 15, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375895001301.

Qi:1997:PSO

[Qi97] Yongcheng Qi. A problem on stability of order statistics. Journal of Statistical Planning and Inference, 64(1):21–25, October 30, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375896002066.

Qian:1994:GLS

[Qia94] Shixian Qian. Generalization of least-square isotonic regression. Journal of Statistical Planning and Inference, 38(3):389– 397, March 1994. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375894900191.

Qian:1998:MLE

[Qia98] Lianfen Qian. On maximum likelihood estimators for a threshold autoregression. Journal of Statistical Planning and Inference, 75(1):21–46, November 15, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837589800113X.

Qinglin:1996:EPP

[Qin96] Lu Qinglin. On the existence of pure (ν, 4, 3)-PMD. Journal of Statistical Planning and Inference, 51(2):201–213, April 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375895000852.

Qiu:1996:NPM

[Qiu96] Weisheng Qiu. New progress on the multiplier conjecture. Journal of Statistical Planning and Inference, 51(2):243–259, April 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375895000895.

Qian:1998:MSS

[QK98] Guoqi Qian and Hans R. Künsch. On model selection via stochastic complexity in robust linear regression. Journal of Statistical Planning and Inference, 75(1):91–116, November 15, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898001384.

Qaqish:1999:RFR

[QP99] Bahjat F. Qaqish and John S. Preisser. Resistant fits for regression with correlated outcomes an estimating equations approach. Journal of Statistical Planning and Inference, 75(2):415– 431, January 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837589800158X.

Ragbi:1997:WCW

[Rag97] Bouameur Ragbi. Weak convergence of weighted multivariate empirical U-statistics processes under mixing condition. Journal of Statistical Planning and Inference, 61(1):1–16, May 30, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896001474.

Ralescu:1992:ADB

[Ral92] Stefan S. Ralescu. Asymptotic deviations between perturbed empirical and quantile processes. Journal of Statistical Planning and Inference, 32(2):243–258, August 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375892900503.

Ralescu:1993:ALL

[Ral93a] Stefan S. Ralescu. The asymptotic law of the local oscillation modulus of the empirical process. Journal of Statistical Planning and Inference, 35(2):139–156, May 1993. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375893900414.

Ralescu:1993:NCS

[Ral93b] Stefan S. Ralescu. A note on the Chernoff-Savage theorem. Journal of Statistical Planning and Inference, 35(3):293–299, June 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589390017Z.

Ramamoorthi:1990:SAI

[Ram90] R. V. Ramamoorthi. Sufficiency, ancillarity and independence in invariant models. Journal of Statistical Planning and Inference, 26(1):59–63, September 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589090095C.

Ramachandran:1997:CVC

[Ram97] B. Ramachandran. Complex-valued characteristic functions with (some) real powers — II. Journal of Statistical Planning and Inference, 63(2):337–349, October 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897000323.

Rao:1994:HPI

[Rao94] M. M. Rao. Harmonizable processes and inference: unbiased prediction for stochastic flows. Journal of Statistical Planning and Inference, 39(2):187–209, April 15, 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894902054.

Rasch:1996:SSP

[Ras96] Dieter Rasch. Software for selection procedures. Journal of Statistical Planning and Inference, 54(3):345–358, September 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375895001778.

Rashid:1998:MCC

[Ras98] M. Mushfiqur Rashid. Multiple comparisons with a control in repeated measures incomplete block designs using R-estimators. Journal of Statistical Planning and Inference, 70(2):361–380, July 15, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375897001912.

Rathbun:1996:APM

[Rat96] Stephen L. Rathbun. Asymptotic properties of the maximum likelihood estimator for spatio-temporal point processes. Journal of Statistical Planning and Inference, 51(1):55–74, April 1, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375895000704.

Rom:1995:EPL

[RC95] Meir Rom and Ayala Cohen. Estimation in the polytomous logistic regression model. Journal of Statistical Planning and Inference, 43(3):341–353, February 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589400037V.

Raghavan:1998:APS

[RC98] Nandini Raghavan and Dennis D. Cox. Analysis of the posterior for spline estimators in logistic regression. Journal of Statistical Planning and Inference, 71(1–2):117–136, August 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898000858.

Rodriguez-Campos:1998:THG

[RCGMC98] M. Celia Rodríguez-Campos, Wenceslao González-Manteiga, and Ricardo Cao. Testing the hypothesis of a generalized linear regression model using nonparametric regression estimation. Journal of Statistical Planning and Inference, 67(1):99–122, March 16, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375897000980.

Ray-Chaudhuri:1997:NNC

[RCX97] D. K. Ray-Chaudhuri and Qing Xiang. New necessary conditions for abelian Hadamard difference sets. Journal of Statistical Planning and Inference, 62(1):69–79, July 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837589600170X.

Ray-Chaudhuri:1997:OAO

[RCZ97] D. K. Ray-Chaudhuri and Tianbao Zhu. Orthogonal arrays and ordered designs. Journal of Statistical Planning and Inference, 58(1):177–183, March 1, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896000687.

Reinsel:1999:MLR

[Rei99] Gregory C. Reinsel. On multivariate linear regression shrinkage and reduced-rank procedures. Journal of Statistical Planning and Inference, 81(2):311–321, November 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375899000166.

Ren:1997:SCE

[Ren97] Jian-Jian Ren. On self-consistent estimators and kernel density estimators with doubly censored data. Journal of Statistical Planning and Inference, 64(1):27–43, October 30, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896001887.

Revesz:1995:STE

[Rév95] P. Révész. Strong theorems on the extreme values of stationary Poisson processes. Journal of Statistical Planning and Inference, 45(1–2):291–300, ???? 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894000786.

Rosenberger:1997:ANM

[RFD97] William F. Rosenberger, Nancy Flournoy, and Stephen D. Durham. Asymptotic normality of maximum likelihood estimators from multiparameter response-driven designs. Journal of Statistical Planning and Inference, 60(1):69–76, April 30, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375896001206.

Rom:1995:NCM

[RH95] Dror M. Rom and Burt Holland. A new closed multiple testing procedure for hierarchical families of hypotheses. Journal of Statistical Planning and Inference, 46(3):265–275, August 15, 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589400116D.

Rao:1997:TFN

[RH97] Poduri S. R. S. Rao and Charles E. Heckler. The threefold nested random effects model. Journal of Statistical Planning and Inference, 64(2):341–352, November 15, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897000049.

Rios:1998:TNT

[Rio98] Ricardo Rios. Towards a nonparametric test of linearity for times series. Journal of Statistical Planning and Inference, 68(2):339– 358, May 1998. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375897001481.

Rahme:1998:ESS

[RJ98] Elham Rahme and Lawrence Joseph. Exact sample size determination for binomial experiments. Journal of Statistical Planning and Inference, 66(1):83–93, January 5, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897000682.

Russell:1999:NOR

[RJ99] K. G. Russell and J. A. John. A note on the optimality of resolvable block designs with unequal block sizes. Journal of Statistical Planning and Inference, 81(1):195–199, October 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375899000142.

Reisen:1999:SSA

[RL99] Valdério Anselmo Reisen and Silvia Lopes. Some simulations and applications of forecasting long-memory time-series models. Journal of Statistical Planning and Inference, 80(1–2):269–287, August 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375898002547.

Rao:1995:CBT

[RM95] C. Radhakrishna Rao and Rahul Mukerjee. Comparison of Bartlett-type adjustments for the efficient score statistic. Journal of Statistical Planning and Inference, 46(2):137–146, August 1, 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375894001077.

Ruiz:1993:CCM

[RMZ93] J. M. Ruiz, J. Marín, and P. Zoroa. A characterization of continuous multivariate distributions by conditional expectations.

Journal of Statistical Planning and Inference, 37(1):13–21, October 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589390076I.

Rashid:1995:RBA

[RN95] M. Mushfiqur Rashid and Balgobin Nandram. A rank-based analysis of one-way repeated measures designs with a changing covariate. Journal of Statistical Planning and Inference, 46(2):249– 263, August 1, 1995. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375894001088.

Ronchetti:1997:RII

[Ron97] Elvezio Ronchetti. Robust inference by influence functions. Journal of Statistical Planning and Inference, 57(1):59–72, January 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375896000365.

Ross:1992:GMC

[Ros92] G. J. S. Ross. Graphical methods for constructing data sets for which likelihood functions have multiple optima. Journal of Statistical Planning and Inference, 30(1):107–132, January 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375892901126.

Rosen:1997:ATO

[Ros97a] Bengt Rosén. Asymptotic theory for order sampling. Journal of Statistical Planning and Inference, 62(2):135–158, August 15, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375896001851.

Rosen:1997:SPP

[Ros97b] Bengt Rosén. On sampling with probability proportional to size. Journal of Statistical Planning and Inference, 62(2):159– 191, August 15, 1997. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375896001863.

Roth:1999:MCP

[Rot99] Arthur J. Roth. Multiple comparison procedures for discrete test statistics. Journal of Statistical Planning and Inference, 82(1–2):101–117, December 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375899000348.

Roussas:1990:ANK

[Rou90] George G. Roussas. Asymptotic normality of the kernel estimate under dependence conditions: application to hazard rate. Journal of Statistical Planning and Inference, 25(1):81–104, May 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589090008I.

Rozanski:1992:RSS

[Róz92a] Roman Rózański. Random stopping sets in a sequential analysis of random measures and fields. Journal of Statistical Planning and Inference, 30(3):401–412, March 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290166P.

Rozanski:1992:REI

[Róz92b] Roman Rózański. Recursive estimation of intensity function of a Poisson random field. Journal of Statistical Planning and Inference, 33(2):165–174, November 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290064Y.

Raktoe:1997:NDL

[RP97] B. L. Raktoe and H. Pesotan. Note on Draper and Lin’s article “Capacity considerations for two-level fractional factorial designs”. Journal of Statistical Planning and Inference, 59(1):163– 165, March 30, 1997. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375896001024. See [DL90].

Rao:1997:BSR

[RPK97] C. Radhakrishna Rao, P. K. Pathak, and V. I. Koltchinskii. Bootstrap by sequential resampling. Journal of Statistical Planning and Inference, 64(2):257–281, November 15, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897000414.

Rom:1992:GMA

[RS92] Dror Rom and Sanat K. Sarkar. A generalized model for the analysis of association in ordinal contingency tables. Journal of Statistical Planning and Inference, 33(2):205–212, November 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375892900684.

Ralescu:1993:NSC

[RS93] Stefan S. Ralescu and Shan Sun. Necessary and sufficient conditions for the asymptotic normality of perturbed sample quantiles. Journal of Statistical Planning and Inference, 35(1):55–64, April 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589390067G.

Rahman:1996:UAS

[RS96] Mizan Rahman and Sergei K. Suslov. A unified approach to the summation and integration formulas for q-hypergeometric functions I. Journal of Statistical Planning and Inference, 54(1):101– 118, September 1996. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375895001603.

Rao:1997:CTI

[RS97a] B. L. S. Prakasa Rao and M. Sreehari. Chernoff-type inequality and variance bounds. Journal of Statistical Planning and Inference, 63(2):325–335, October 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897000311.

Rosenberger:1997:EAA

[RS97b] William F. Rosenberger and T. N. Sriram. Estimation for an adaptive allocation design. Journal of Statistical Planning and Inference, 59(2):309–319, April 15, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896001097.

Rissanen:1998:FTP

[RS98] J. Rissanen and G. Shedler. Failure-time prediction. Journal of Statistical Planning and Inference, 66(2):193–210, March 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375897000839.

Rao:1990:SCB

[RT90] B. Raja Rao and Sheela Talwalker. “Setting the clock back to zero” property of a family of life distributions. Journal of Statistical Planning and Inference, 24(3):347–352, March 1990. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589090053W.

Ruszinko:1997:SPM

[RT97] Miklós Ruszinkó and Gábor Tardos. On a search problem in multidimensional grids. Journal of Statistical Planning and Inference, 59(1):101–109, March 30, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896001048.

Rubin:1990:FMS

[Rub90] Donald B. Rubin. Formal mode of statistical inference for causal effects. Journal of Statistical Planning and Inference, 25(3):279– 292, July 1990. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375890900778.

Ruckstuhl:1997:PBT

[Ruc97] Andreas F. Ruckstuhl. Partial breakdown in two-factor models. Journal of Statistical Planning and Inference, 57(2):257– 271, February 1997. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375896000481.

Rukhin:1992:LEL

[Ruk92] Andrew L. Rukhin. Linear estimators of a location parameter and canonical parameter estimators for a Poisson distribution. Journal of Statistical Planning and Inference, 31(1):1–13, April 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589290037S.

Rukhin:1993:ENP

[Ruk93] Andrew L. Rukhin. Estimation of the noncentrality parameter of an F -distribution. Journal of Statistical Planning and Inference, 35(2):201–211, May 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375893900458.

Ruszinko:1993:DSP

[Rus93] Miklós Ruszinkó. On a 2-dimensional search problem. Journal of Statistical Planning and Inference, 37(3):371–383, December 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589390113K.

Ruymgaart:1994:CEQ

[Ruy94] Frits H. Ruymgaart. Conditional empirical, quantile and difference processes for a large class of time series with applications. Journal of Statistical Planning and Inference, 40(1):15–31, June 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375894901392.

Rukhin:1995:ADM

[RV95] Andrew L. Rukhin and Igor Vajda. Adaptive decision making for stochastic processes. Journal of Statistical Planning and Inference, 45(3):313–329, June 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589300087X.

Ruymgaart:1992:ESP

[RvZ92] F. H. Ruymgaart and M. C. A. van Zuijlen. Empirical Ustatistics processes. Journal of Statistical Planning and Inference, 32(2):259–269, August 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837589290051S.

Ruggeri:1995:DBC

[RW95] Fabrizio Ruggeri and Larry Wasserman. Density based classes of priors: Infinitesimal properties and approximations. Journal of Statistical Planning and Inference, 46(3):311–324, August 15, 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375894001118.

Rocke:1997:REM

[RW97] David M. Rocke and David L. Woodruff. Robust estimation of multivariate location and shape. Journal of Statistical Planning and Inference, 57(2):245–255, February 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837589600047X.

Rao:1998:TGL

[RW98a] T. Subba Rao and W. K. Wong. Tests for Gaussianity and linearity of multivariate stationary time series. Journal of Statistical Planning and Inference, 68(2):373–386, May 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837589700150X.

Robertson:1998:IBE

[RW98b] J. M. Robertson and G. R. Wood. Information in Buffon experiments. Journal of Statistical Planning and Inference, 66(1):21– 37, January 5, 1998. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375897000608.

Robinson:1998:NTS

[RZ98] Peter M. Robinson and Paolo Zaffaroni. Nonlinear time series with long memory: a model for stochastic volatility. Journal of Statistical Planning and Inference, 68(2):359–371, May 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897001493.

Saweris:1997:MDC

[SA97] Nashat B. Saweris and M. Masoon Ali. Moments and distributions of certain multivariate test criteria in the canonical correlation case under violation. Journal of Statistical Planning and Inference, 62(2):235–246, August 15, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896001917.

Salehi:1992:BRD

[Sal92] Habib Salehi. Book review: Detection and estimation: Dimitri Kazakos and P. Papantoni-Kazakos, Computer Science Press, New York, 1990, xxi + 428 pp., $51.95. Journal of Statistical Planning and Inference, 32(3):422–423, September 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290023L.

Salinetti:1994:SBD

[Sal94] Gabriella Salinetti. Stability of Bayesian decisions. Journal of Statistical Planning and Inference, 40(2–3):313–329, July 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375894901287.

Sarda:1993:SPS

[Sar93] P. Sarda. Smoothing parameter selection for smooth distribution functions. Journal of Statistical Planning and Inference, 35(1): 65–75, April 1993. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589390068H.

Sarkar:1994:SDS

[Sar94] Sanat K. Sarkar. Shrinkage domination of some usual estimators of the common mean of several multivariate normal populations. Journal of Statistical Planning and Inference, 39(1):43–55, April 1, 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375894901112.

Saran:1995:JDB

[Sar95] Jagdish Saran. Joint distributions based on the number of runs and some other two-sample rank order statistics for arbitrary sample sizes. Journal of Statistical Planning and Inference, 47(3):257–275, October 15, 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589400138L.

Sato:1993:GFN

[Sat93] Masako Sato. Generating functions for the number of lattice paths restricted by two parallel hyperplanes. Journal of Statistical Planning and Inference, 34(2):251–257, February 1993. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390009U.

Sethuraman:1994:LSI

[SB94a] S. Sethuraman and I. V. Basawa. Large sample inference for a multivariate linear model with autocorrelated errors. Journal of Statistical Planning and Inference, 41(2):187–204, September 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375894901635.

Sethuraman:1994:PES

[SB94b] S. Sethuraman and I. V. Basawa. Parameter estimation in a stationary autoregressive process with correlated multiple observations. Journal of Statistical Planning and Inference, 39(2):137– 154, April 15, 1994. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375894902038.

Sen:1995:IPL

[SB95a] Ananda Sen and Gouri K. Bhattacharyya. Inference procedures for the linear failure rate model. Journal of Statistical Planning and Inference, 46(1):59–76, July 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589400099H.

Sethuraman:1995:MLE

[SB95b] S. Sethuraman and I. V. Basawa. Maximum likelihood estimation for a fractionally differenced autoregressive model on a two-dimensional lattice. Journal of Statistical Planning and Inference, 44(2):219–235, April 1, 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589400046X.

Simpson:1997:RAG

[SC97] Douglas G. Simpson and Yuan-Chin Ivan Chang. Reweighting approximate GM estimators: Asymptotics and residualbased graphics. Journal of Statistical Planning and Inference, 57(2):273–293, February 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375896000493.

Samuel-Cahn:1999:NAC

[SC99a] Ester Samuel-Cahn. A note about a curious generalization of Simes’ theorem. Journal of Statistical Planning and Inference, 82(1–2):147–149, December 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375899000385.

Srivastava:1999:MDP

[SC99b] J. N. Srivastava and J. Y. Chu. Multistage design procedures for identifying two-factor interactions, when higher effects are negligible. Journal of Statistical Planning and Inference, 78(1–2): 149–189, May 1999. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375898002122.

Schwabe:1991:ADO

[Sch91] Rainer Schwabe. Adaptive design in the one-way layout. Journal of Statistical Planning and Inference, 28(3):319–338, July 1991. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589190070U.

Schipp:1993:AME

[Sch93] Bernhard Schipp. Approximate minimax estimators in the simultaneous equations model. Journal of Statistical Planning and Inference, 36(2–3):197–213, August/September 1993. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390124O.

Schick:1997:EEL

[Sch97] Anton Schick. Efficient estimates in linear and nonlinear regression with heteroscedastic errors. Journal of Statistical Planning and Inference, 58(2):371–387, March 15, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896000778.

Schick:1998:ESN

[Sch98a] Anton Schick. Estimating a shift in nonparametric regression via U-statistics. Journal of Statistical Planning and Inference, 67(2):259–271, April 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375897001109.

Schwabe:1998:ODH

[Sch98b] Rainer Schwabe. Optimal designs for hierarchical interaction structures. Journal of Statistical Planning and Inference, 70(1): 181–190, July 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375897001742.

Schick:1999:IWL

[Sch99] Anton Schick. Improving weighted least-squares estimates in heteroscedastic linear regression when the variance is a function of the mean response. Journal of Statistical Planning and Inference, 76(1–2):127–144, February 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898001347.

Samuels:1994:EEB

[SCM94] Myra L. Samuels, George Casella, and George P. McCabe. Evaluating the efficiency of blocking without assuming compound symmetry. Journal of Statistical Planning and Inference, 38(2):237– 247, February 1994. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589490037X.

Sobel:1992:QSM

[SE92] Milton Sobel and M. Ebneshahrashoob. Quota sampling for multinomial via Dirichlet. Journal of Statistical Planning and Inference, 33(2):157–164, November 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290063X.

Steinebach:1995:EVA

[SE95] Josef Steinebach and Vera R. Eastwood. On extreme value asymptotics for increments of renewal processes. Journal of Statistical Planning and Inference, 45(1–2):301–312, ???? 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894000794.

Seberry:1998:BMD

[Seb98] Jennifer Seberry. Bose’s method of differences applied to construct Bhaskar Rao designs. Journal of Statistical Planning and Inference, 73(1–2):215–224, September 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898000639.

Seifert:1992:ETU

[Sei92] Burkhardt Seifert. Exact tests in unbalanced mixed analysis of variance. Journal of Statistical Planning and Inference, 30(2):257–266, February 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375892900868.

Seifert:1993:THM

[Sei93] Burkhardt Seifert. Testing hypotheses in mixed linear models. Journal of Statistical Planning and Inference, 36(2–3): 253–268, August/September 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837589390128S.

SenGupta:1991:GCS

[Sen91] Ashis SenGupta. Generalized correlations in the singular case. Journal of Statistical Planning and Inference, 28(2):241–245, June 1991. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589190029E.

Sen:1994:IMD

[Sen94] Pranab K. Sen. Incomplete multiresponse designs and surrogate endpoints in clinical trials. Journal of Statistical Planning and Inference, 42(1–2):161–186, November 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894901953.

Sen:1995:SAS

[Sen95] Pranab Kumar Sen. Statistical analysis of some reliability models: Parametrics, semi-parametrics and nonparametrics. Journal of Statistical Planning and Inference, 43(1–2):41–66, January 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589400007I.

Sen:1999:MCI

[Sen99a] Pranab K. Sen. Multiple comparisons in interim analysis. Journal of Statistical Planning and Inference, 82(1–2):5–23, December 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899000282.

Sen:1999:SRS

[Sen99b] Pranab K. Sen. Some remarks on Simes-type multiple tests of significance. Journal of Statistical Planning and Inference, 82(1–2):139–145, December 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375899000373.

Sen:1999:RNM

[Sen99c]  Pranab Kumar Sen.  Robust nonparametrics in mixed-

MANOCOVA models. Journal of Statistical Planning and Inference, 75(2):433–451, January 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375898001591.

Seoh:1990:EPC

[Seo90] Munsup Seoh. An elementary proof of CLT for signed rank statistics with regression constants under hypothesis of symmetry. Journal of Statistical Planning and Inference, 26(1):103– 108, September 1990. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589090099G.

Seoh:1994:REN

[Seo94] Munsup Seoh. On the relative error in normal approximation for signed rank statistics with a broad range of scores. Journal of Statistical Planning and Inference, 39(1):113–128, April 1, 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375894901163.

Seshadri:1997:PNE

[Ses97] V. Seshadri. A property of natural exponential families in R^{n} with simple quadratic variance functions. Journal of Statistical Planning and Inference, 63(2):351–361, October 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897000335.

Shen:1997:LSP

[SF97] Yu Shen and Thomas R. Fleming. Large sample properties of some survival estimators in heterogeneous samples. Journal of Statistical Planning and Inference, 60(1):123–138, April 30, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375896001243.

Singh:1991:RBD

[SG91] Rajeshwar Singh and V. K. Gupta. Resistance of block designs. Journal of Statistical Planning and Inference, 27(2):263– 269, February 1991. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375891900216.

Suchower:1990:NPA

[SH90] Lisa J. Suchower and Robert A. Hultquist. A new property of the association matrices for EGD and hypercubic association schemes. Journal of Statistical Planning and Inference, 25(1): 105–108, May 1990. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589090009J.

Santner:1992:TSP

[SH92a] Thomas J. Santner and Maika Behaxeteguy Heffernan. A twostage procedure for selecting the largest normal mean whose first stage selects a bounded random number of populations. Journal of Statistical Planning and Inference, 31(2):147–168, May 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589290026O.

Shing:1992:RMD

[SH92b] Chen-Chi Shing and Klaus Hinkelmann. Repeated measurement designs for factorial treatments when high order interactions are negligible. Journal of Statistical Planning and Inference, 31(1): 81–91, April 1992. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589290042Q.

Srivastava:1992:SFP

[SH92c] Jagdish Srivastava and Ranoo Hveberg. Sequential factorial probing designs for identifying and estimating non-negligible factorial effects for the 2^{m} experiment under the tree structure. Journal of Statistical Planning and Inference, 30(2):141–162, February 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375892900776.

Seoh:1997:WDE

[SH97] Munsup Seoh and Marc Hallin. When does Edgeworth beat Berry and Esséen? Numerical evaluations of Edgeworth expansions. Journal of Statistical Planning and Inference, 63(1):19–38, September 30, 1997. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037837589600198X.

Shah:1992:RII

[Sha92] K. R. Shah. Recovery of interblock information: an update. Journal of Statistical Planning and Inference, 30(2):163–172, February 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375892900787.

Shanmugam:1995:GFT

[Sha95] Ramalingam Shanmugam. Goodness-of-fit test for lengthbiased data with discussion on prevalence, sensitivity and specificity of the length bias. Journal of Statistical Planning and Inference, 48(3):277–290, December 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589500007V.

Shanbhag:1997:P

[Sha97] D. N. Shanbhag. Preface. Journal of Statistical Planning and Inference, 63(2):115–116, October 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897001134.

Shaffer:1999:SBS

[Sha99] Juliet Popper Shaffer. A semi-Bayesian study of Duncan’s Bayesian multiple comparison procedure. Journal of Statistical Planning and Inference, 82(1–2):197–213, December 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375899000427.

Sherman:1998:ERS

[She98] Michael Sherman. Efficiency and robustness in subsampling for dependent data. Journal of Statistical Planning and Inference, 75(1):133–146, November 15, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898001232.

Shirakura:1991:MEO

[Shi91] Teruhiro Shirakura. Main effect + one or two plans for 2 ^{m} factorials. Journal of Statistical Planning and Inference, 27(1): 65–74, January 1991. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589190082P.

Shirley:1992:MSL

[Shi92] Angela G. Shirley. Is the minimum of several location parameters positive? Journal of Statistical Planning and Inference, 31(1): 67–79, April 1992. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589290041P.

Shi:1997:TRS

[Shi97] Peide Shi. M-type regression splines involving time series. Journal of Statistical Planning and Inference, 61(1):17–37, May 30, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375897897145.

Shorack:1997:UCW

[Sho97] Galen R. Shorack. Uniform CLT, WLLN, LIL and bootstrapping in a data analytic approach to trimmed L-statistics. Journal of Statistical Planning and Inference, 60(1):1–44, April 30, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375897001213.

Shrikhande:1998:RCB

[Shr98] S. S. Shrikhande. R. C. Bose memorial conference Fort Collins, Colorado. Journal of Statistical Planning and Inference, 73 (1–2):273–276, September 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375898000652.

Shumway:1995:BRL

[Shu95] Robert H. Shumway. Book review: Longitudinal data with serial correlation: a state-space approach: By R. H. Jones, Chapman and Hall, London, 1983. Journal of Statistical Planning and Inference, 47(3):395–397, October 15, 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375895900292.

Silvapulle:1992:MGC

[Sil92] Mervyn J. Silvapulle. On M-methods in growth curve analysis with asymmetric errors. Journal of Statistical Planning and Inference, 32(3):303–309, September 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290013I.

Simon:1994:PMC

[Sim94] Richard Simon. Problems of multiplicity in clinical trials. Journal of Statistical Planning and Inference, 42(1–2):209–221, November 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589490197X.

Simonoff:1995:SCD

[Sim95] Jeffrey S. Simonoff. Smoothing categorical data. Journal of Statistical Planning and Inference, 47(1–2):41–69, October 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589400121B.

Singpurwalla:1991:IUP

[Sin91] Nozer D. Singpurwalla. Inference under planned maintenance, warranties, and other retrospective data. Journal of Statistical Planning and Inference, 29(1–2):171–185, September/October 1991. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589290131B.

Singh:1995:EBL

[Sin95] R. S. Singh. Empirical Bayes linear loss hypothesis testing in a non-regular exponential family. Journal of Statistical Planning and Inference, 43(1–2):107–120, January 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589400012K.

Singhi:1996:P

[Sin96] N. M. Singhi. Preface. Journal of Statistical Planning and Inference, 56(1):1, 4–15, December 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896000031.

Siu:1997:CBA

[Siu97] Man Keung Siu. The combinatorics of binary arrays. Journal of Statistical Planning and Inference, 62(1):103–113, July 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897842085.

Sivaganesan:1991:SSS

[Siv91] S. Sivaganesan. Sensitivity of some standard Bayesian estimates to prior uncertainty: a comparison. Journal of Statistical Planning and Inference, 27(1):85–103, January 1991. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589190084R.

Sivaganesan:1993:RBD

[Siv93] Siva Sivaganesan. Robust Bayesian diagnostics. Journal of Statistical Planning and Inference, 35(2):171–188, May 1993. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375893900436.

Sivaganesan:1994:BPE

[Siv94] Siva Sivaganesan. Bounds on posterior expectations for density bounded class with constant bandwidth. Journal of Statistical Planning and Inference, 40(2–3):331–343, July 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894901295.

Sivaganesan:1997:ORC

[Siv97] Siva Sivaganesan. Optimal robust credible sets for densitybounded priors. Journal of Statistical Planning and Inference, 63(1):9–17, September 30, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375896002005.

Sen:1993:CAM

[SJ93] Kanwar Sen and J. L. Jain. Combinatorial approach to Markovian queueing models. Journal of Statistical Planning and Inference, 34(2):269–279, February 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390011T.

Sen:1996:GMP

[SJ96a] Kanwar Sen and Ritu Jain. Generalized Markov-Polya urn models with predetermined strategies. Journal of Statistical Planning and Inference, 54(1):119–133, September 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895001611.

Seymour:1996:ABM

[SJ96b] Lynne Seymour and Chuanshu Ji. Approximate Bayes model selection procedures for Gibbs-Markov random fields. Journal of Statistical Planning and Inference, 51(1):75–97, April 1, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375895000712.

Sen:1993:LPA

[SJG93] Kanwar Sen, J. L. Jain, and J. M. Gupta. Lattice path approach to transient solution of M/M/1 with (0, k) control policy. Journal of Statistical Planning and Inference, 34(2):259–267, February 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375893900104.

Sinha:1990:STP

[SK90] Kishore Sinha and Sanpei Kageyama. A series of triangular PBIB designs with nested rows and columns. Journal of Statistical Planning and Inference, 24(3):403–404, March 1990. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375890900594.

Sinha:1992:CSO

[SK92] Kishore Sinha and Sanpei Kageyama. Constructions of some Eoptimal 3-concurrence most balanced designs. Journal of Statistical Planning and Inference, 31(1):127–132, April 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290045T.

Sreehari:1997:MSS

[SK97] M. Sreehari and V. A. Kalamkar. Modelling some stationary Markov processes and related characterizations. Journal of Statistical Planning and Inference, 63(2):363–375, October 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375897844515.

Sen:1998:ECI

[SK98] Ananda Sen and Ravindra Khattree. On estimating the current intensity of failure for the power-law process. Journal of Statistical Planning and Inference, 74(2):253–272, November 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375898001165.

Srivastava:1999:SPE

[SK99] Rakesh Srivastava and Zoola Kapasi. Sometimes pool t estimation — via shrinkage. Journal of Statistical Planning and Inference, 79(2):191–204, July 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898002341.

Skouras:1998:ACM

[Sko98] Kostas Skouras. Absolute continuity of Markov chains. Journal of Statistical Planning and Inference, 75(1):1–8, November 15, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375898001177.

Sukhatme:1995:EPW

[SL95] Shashikala Sukhatme and C. H. Lin. Exact powers of the Wilcoxon test and the logrank test under random censoring. Journal of Statistical Planning and Inference, 48(3):291–301, December 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375895000107.

Srivastava:1996:ODP

[SL96] Jagdish N. Srivastava and Junfang Li. Orthogonal designs of parallel flats type. Journal of Statistical Planning and Inference, 53(2):261–283, August 15, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589500145X.

Slud:1992:BPT

[Slu92] Eric V. Slud. Best precedence tests for censored data. Journal of Statistical Planning and Inference, 31(3):283–293, June 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589290138I.

Sengupta:1992:OSS

[SM92] S. Sengupta and Rahul Mukerjee. Optimal spatial sampling plans in the presence of linear trend. Journal of Statistical Planning and Inference, 33(2):285–293, November 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375892900743.

Stute:1996:NGF

[SM96] W. Stute and W. González Manteiga. NN goodness-of-fit tests for linear models. Journal of Statistical Planning and Inference, 53(1):75–92, August 1, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895001441.

Srivastav:1998:ODG

[SM98] Sudesh K. Srivastav and John P. Morgan. Optimality of designs with generalized group divisible structure. Journal of Statistical Planning and Inference, 71(1–2):313–330, August 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898000172.

Sutradhar:1995:STS

[SMD95] B. C. Sutradhar, I. B. MacNeil, and E. B. Dagum. A simple test for stable seasonality. Journal of Statistical Planning and Inference, 43(1–2):157–167, January 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589400016O.

Smith:1990:SDO

[Smi90] Philip J. Smith. Survey design optimization for catch curve analysis. Journal of Statistical Planning and Inference, 26(3):277– 290, November 1990. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589090132E.

Smith:1997:BRA

[Smi97a] Eric P. Smith. Book review: Anant M. Kshirsagar and William Boyce Smith: Growth curves (Marcel Dekker, New York, 1995, Price $135.00, ISBN: 0-8247-9341-2, 392 pp.). Journal of Statistical Planning and Inference, 60(1):195–196, April 30, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837589789708X.

Smith:1997:BRM

[Smi97b] Eric P. Smith. Book review: Methods of multivariate analysis: Alvin Rencher, (John, New York, 1995, $74.95 ISBN 0471571520, xvi + 627 pp.). Journal of Statistical Planning and Inference, 59(1):183–184, March 30, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896000985.

Smith:1997:BRW

[Smi97c] Eric P. Smith. Book review: W. J. Krzanowski and F. H. C. Marriott: Multivariate analysis, part 2, classification, covariance structures and repeated measurements (Arnold, London, 1996, Price $54.95, ISBN Number 0-4702-3413-X, 280 pp.). Journal of Statistical Planning and Inference, 60(1):196–197, April 30, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375897897091.

Swamy:1996:CWD

[SMS96] P. A. V. B. Swamy, J. S. Mehta, and Rao N. Singamsetti. Circumstances in which different criteria of estimation can be applied to estimate policy effects. Journal of Statistical Planning and Inference, 50(1):121–153, February 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589500050X.

Subramanyam:1990:OUS

[SNN90] A. Subramanyam and U. V. Naik-Nimbalkar. Optimal unbiased statistical estimating functions for Hilbert space valued parameters. Journal of Statistical Planning and Inference, 24(1):95– 105, January 1990. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589090020U.

Srivastava:1992:SGE

[SO92] Jagdish Srivastava and Zhao Ouyang. Studies on a general estimator in sampling, utilizing extraneous information through a sample weight function. Journal of Statistical Planning and Inference, 31(2):199–218, May 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290030V.

Scazzero:1993:ODL

[SO93] Joseph A. Scazzero and J. K. Ord. D-optimum designs for the linear logistic model when restrictions exist on p. Journal of Statistical Planning and Inference, 37(2):255–264, November 1993. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390094M.

Somerville:1999:CVM

[Som99] Paul N. Somerville. Critical values for multiple testing and comparisons: one step and step down procedures. Journal of Statistical Planning and Inference, 82(1–2):129–138, December 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899000361.

SenGupta:1991:LOT

[SP91] Ashis SenGupta and ChandraNath Pal. Locally optimal tests for no contamination in standard symmetric multivariate normal mixtures. Journal of Statistical Planning and Inference, 29(1–2):145–155, September/October 1991. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290129G.

Santner:1997:SST

[SP97] Thomas J. Santner and Guohua Pan. Subset selection in twofactor experiments using randomization restricted designs. Journal of Statistical Planning and Inference, 62(2):339–363, August 15, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375896001954.

Spencer:1994:SBC

[Spe94] Bruce D. Spencer. Sensitivity of benefit-cost analysis of data programs to monotone misspecification. Journal of Statistical Planning and Inference, 39(1):19–31, April 1, 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894901090.

Spizzichino:1991:SBP

[Spi91] Fabio Spizzichino. Sequential burn-in procedures. Journal of Statistical Planning and Inference, 29(1–2):187–197, September/ October 1991. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589290132C.

Spiegelman:1994:CES

[Spi94] Donna Spiegelman. Cost-efficient study designs for relative risk modeling with covariate measurement error. Journal of Statistical Planning and Inference, 42(1–2):187–208, November 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894901961.

Spruill:1990:GDP

[Spr90] M. C. Spruill. Good designs for polynomial extrapolation. Journal of Statistical Planning and Inference, 26(2):149–159, October 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589090123C.

Saran:1990:GRW

[SR90] Jagdish Saran and Sarita Rani. Generalized random walk and distributions of some rank order statistics. Journal of Statistical Planning and Inference, 24(1):119–133, January 1990. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589090022M.

Stukel:1999:SAE

[SR99] D. M. Stukel and J. N. K. Rao. On small-area estimation under two-fold nested error regression models. Journal of Statistical Planning and Inference, 78(1–2):131–147, May 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898002110.

Sriram:1990:SER

[Sri90a] T. N. Sriram. Sequential estimation of ratio of normal parameters. Journal of Statistical Planning and Inference, 26(3):305– 324, November 1990. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589090134G.

Srivastava:1990:A

[Sri90b] J. N. Srivastava. Announcement. Journal of Statistical Planning and Inference, 26(1):129, September 1990. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589090102Z.

Srivastava:1990:ACB

[Sri90c] J. N. Srivastava. Announcement concerning book reviews. Journal of Statistical Planning and Inference, 25(1):109, May 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589090010R.

Srivastava:1990:F

[Sri90d] J. N. Srivastava. Foreword. Journal of Statistical Planning and Inference, 25(3):207, July 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375890900712.

Srivastava:1990:SAR

[Sri90e] Jagdish Srivastava. Some aspects of response surface theory, and the philosophy of statistical experimental design. Journal of Statistical Planning and Inference, 25(3):415–435, July 1990. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375890900859.

Srivastava:1991:F

[Sri91] J. N. Srivastava. Foreword. Journal of Statistical Planning and Inference, 29(1–2):1, September/October 1991. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375892901159.

Srivastava:1992:Fa

[Sri92a] J. N. Srivastava. Foreword. Journal of Statistical Planning and Inference, 30(2):135, February 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375892900754.

Srivastava:1992:A

[Sri92b] Jagdish N. Srivastava. Announcement. Journal of Statistical Planning and Inference, 30(1):v, January 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290101W.

Srivastava:1992:Fb

[Sri92c] Jagdish N. Srivastava. Foreword. Journal of Statistical Planning and Inference, 33(1):1, October 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290090F.

Srivastava:1993:F

[Sri93] Jagdish N. Srivastava. Foreword. Journal of Statistical Planning and Inference, 34(1):1, January 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375893900274.

Srivastava:1994:NMS

[Sri94] Jagdish N. Srivastava. A new mathematical space with applications to statistical experimental design — I. Preliminary results. Journal of Statistical Planning and Inference, 40(1):113– 125, June 1994. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375894901457.

Srivastava:1996:P

[Sri96] V. K. Srivastava. Preface. Journal of Statistical Planning and Inference, 49(1):1–2, January 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375896900294.

Srivastava:1997:F

[Sri97] J. N. Srivastava. Foreword. Journal of Statistical Planning and Inference, 63(2):113–114, October 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897001122.

Srivastava:1998:A

[Sri98] J. N. Srivastava. Announcement. Journal of Statistical Planning and Inference, 66(1):1, January 5, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898900052.

Sathe:1990:CMS

[SS90a] Y. S. Sathe and R. G. Shenoy. Construction method for some Aand D-optimal weighing designs when N ≡ 3 (mod 4). Journal of Statistical Planning and Inference, 24(3):369–375, March 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589090056Z.

Schervish:1990:ACC

[SS90b] M. Schervish and T. Seidenfeld. An approach to consensus and certainty with increasing evidence. Journal of Statistical Planning and Inference, 25(3):401–414, July 1990. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375890900848.

Schick:1990:IDC

[SS90c] A. Schick and V. Susarla. An infinite dimensional convolution theorem with applications to random censoring and missing data models. Journal of Statistical Planning and Inference, 24(1):13– 23, January 1990. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589090013K.

Sathe:1991:FRC

[SS91a] Y. S. Sathe and R. G. Shenoy. Further results on construction methods for some A- and D-optimal weighing designs when N ≡ 3 mod 4. Journal of Statistical Planning and Inference, 28(3): 339–352, July 1991. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589190071L.

Shaked:1991:SMM

[SS91b] Moshe Shaked and J. George Shanthikumar. Shock models with MIFRA time to failure distributions. Journal of Statistical Planning and Inference, 29(1–2):157–169, September/October 1991. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589290130K.

Sathe:1992:SMR

[SS92] Y. S. Sathe and Meena R. Satam. Some more robust block designs against the unavailability of data. Journal of Statistical Planning and Inference, 30(1):93–98, January 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290110E.

Sarvate:1993:CBT

[SS93a] Dinesh G. Sarvate and Jennifer Seberry. Constructions of balanced ternary designs based on generalized Bhaskar Rao designs. Journal of Statistical Planning and Inference, 34(3):423– 432, March 1993. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589390149Z.

Shah:1993:OAR

[SS93b] K. R. Shah and Bikas K. Sinha. Optimality aspects of row-column designs with non-orthogonal structure. Journal of Statistical Planning and Inference, 36(2–3):331–346, August/September 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589390135S.

Street:1994:IWD

[SS94] Deborah J. Street and Anne Penfold Street. An investigation of Williams designs with circular structure. Journal of Statistical Planning and Inference, 38(3):399–409, March 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894900205.

Sinha:1995:ABE

[SS95] Bikas K. Sinha and Bimal K. Sinha. An application of bivariate exponential models and related inference. Journal of Statistical Planning and Inference, 44(2):181–191, April 1, 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894000506.

Skeels:1996:SCM

[SS96a] Christopher L. Skeels and Murray D. Smith. Some consequences of model misspecification for t testing in a structural equation. Journal of Statistical Planning and Inference, 50(2):251– 271, March 1996. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375895000577.

Sloane:1996:LPB

[SS96b] N. J. A. Sloane and J. Stufken. A linear programming bound for orthogonal arrays with mixed levels. Journal of Statistical Planning and Inference, 56(2):295–305, December 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896000250.

Srivastava:1996:MEL

[SS96c] Anil K. Srivastava and Praveen Shukla. Minimax estimation in linear regression model. Journal of Statistical Planning and Inference, 50(1):77–89, February 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895000461.

Sebastiani:1997:NOD

[SS97a] Paola Sebastiani and Raffaella Settimi. A note on D-optimal designs for a logistic regression model. Journal of Statistical Planning and Inference, 59(2):359–368, April 15, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896001115.

Shin:1997:RIR

[SS97b] Dong Wan Shin and Sahadeb Sarkar. Regression with integrated regressors. Journal of Statistical Planning and Inference, 64(2):325–340, November 15, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897000426.

Sebastiani:1998:FOO

[SS98a] Paola Sebastiani and Raffaella Settimi. First-order optimal designs for non-linear models. Journal of Statistical Planning and Inference, 74(1):177–192, October 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898000780.

See:1998:ASS

[SS98b] Kyoungah See and Sung Y. Song. Association schemes of small order. Journal of Statistical Planning and Inference, 73 (1–2):225–271, September 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375898000640.

Scarsini:1999:DKI

[SS99] Marco Scarsini and Moshe Shaked. Distributions with known initial hazard rate functions. Journal of Statistical Planning and Inference, 78(1–2):39–55, May 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898002055.

Singh:1998:EAR

[SSE98] Rajesh Singh, Housila P. Singh, and M. Ruiz Espejo. The efficiency of an alternative to ratio estimator under a super population model. Journal of Statistical Planning and Inference, 71(1–2):287–301, August 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375898000093.

Schone:1999:RLE

[SSK99] Alexander Schöne, Wolfgang Schmid, and Sven Knoth. On the run length of the EWMA scheme: a monotonicity result for normal variables. Journal of Statistical Planning and Inference, 79(2):289–297, July 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898002602.

Schmid:1994:IME

[SSR94] Christopher H. Schmid, Mark R. Segal, and Bernard Rosner. Incorporating measurement error in the estimation of autoregressive models for longitudinal data. Journal of Statistical Planning and Inference, 42(1–2):1–18, November 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894901864.

Singh:1994:ERP

[SSS94] Vijay K. Singh, Hari P. Singh, and Housila P. Singh. Estimation of ratio and product of two finite population means in two-phase sampling. Journal of Statistical Planning and Inference, 41(2):163–171, September 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894901619.

Stufken:1999:PDS

[SSSD99] John Stufken, Sung Y. Song, Kyoungah See, and Kenneth R. Driessel. Polygonal designs: some existence and non-existence results. Journal of Statistical Planning and Inference, 77 (1):155–166, February 15, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375898001888.

Stahlecker:1993:MEL

[ST93] Peter Stahlecker and Götz Trenkler. Minimax estimation in linear regression with singular covariance structure and convex polyhedral constraints. Journal of Statistical Planning and Inference, 36(2–3):185–196, August/September 1993. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390123N.

Schipp:1996:FME

[ST96a] Bernhard Schipp and Helge Toutenburg. Feasible minimax estimators in the simultaneous equations model under partial restrictions. Journal of Statistical Planning and Inference, 50(2): 241–250, March 1996. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375895000569.

Shirakura:1996:WOF

[ST96b] Teruhiro Shirakura and Wei-Ping Tong. Weighted A-optimality for fractional 2^{m} factorial designs of resolution V. Journal of Statistical Planning and Inference, 56(2):243–256, December 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375896000213.

Shi:1999:SRM

[ST99] Peide Shi and Chih-Ling Tsai. Semiparametric regression model selections. Journal of Statistical Planning and Inference, 77 (1):119–139, February 15, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S037837589800175X.

Starmer:1992:FDP

[Sta92] C. F. Starmer. Frequency-dependent processes: a model for short-term memory. Journal of Statistical Planning and Inference, 33(1):99–106, October 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837589290095A.

Stanton:1997:RLE

[Sta97] R. G. Stanton. A retrospective look at the Erdős–DeBruijn theorem. Journal of Statistical Planning and Inference, 58(1):185– 191, March 1, 1997. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375896000699.

Stepniak:1991:QUE

[Ste91] Czesław Stepniak. Quadratic unbiased estimation without invariance and its application in the unbalanced one-way random model. Journal of Statistical Planning and Inference, 27(1):25– 35, January 1991. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589190079T.

Steiner:1993:ENO

[Ste93] George Steiner. On estimating the number of order ideals in partial orders, with some applications. Journal of Statistical Planning and Inference, 34(2):281–290, February 1993. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390012U.

Stein:1997:ELP

[Ste97] Michael L. Stein. Efficiency of linear predictors for periodic processes using an incorrect covariance function. Journal of Statistical Planning and Inference, 58(2):321–331, March 15, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896000882.

Sun:1996:BDE

[STL96] Dongchu Sun, Robert K. Tsutakawa, and Wang-Shu Lu. Bayesian design of experiment for quantal responses: What is promised versus what is delivered. Journal of Statistical Planning and Inference, 52(3):289–306, July 1, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895001204.

Street:1992:NSE

[Str92] Deborah J. Street. A note on strongly equineighboured designs. Journal of Statistical Planning and Inference, 30(1):99–105, January 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375892901115.

Stromberg:1997:RCE

[Str97] Arnold J. Stromberg. Robust covariance estimates based on resampling. Journal of Statistical Planning and Inference, 57(2):321–334, February 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375896000511.

Strassler:1998:NCW

[Str98] Yoseph Strassler. New circulant weighing matrices of prime order in cw (31,16), cw (71,25), cw (127,64). Journal of Statistical Planning and Inference, 73(1–2):317–330, September 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898000676.

Stufken:1991:GDT

[Stu91a] John Stufken. On group divisible treatment designs for comparing test treatments with a standard treatment in blocks of size 3. Journal of Statistical Planning and Inference, 28(2):205–221, June 1991. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589190026B.

Stufken:1991:SFO

[Stu91b] John Stufken. Some families of optimal and efficient repeated measurements designs. Journal of Statistical Planning and Inference, 27(1):75–83, January 1991. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589190083Q.

Stute:1992:MCV

[Stu92] W. Stute. Modified cross-validation in density estimation. Journal of Statistical Planning and Inference, 30(3):293–305, March 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589290157N.

Stute:1994:SWR

[Stu94] Winfried Stute. Strong and weak representations of cumulative hazard function and Kaplan–Meier estimators on increasing sets. Journal of Statistical Planning and Inference, 42(3):315– 329, December 1994. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375894000328.

Sulanke:1993:SVD

[Sul93] Robert A. Sulanke. A symmetric variation of a distribution of Kreweras and Poupard. Journal of Statistical Planning and Inference, 34(2):291–303, February 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390013V.

Sun:1997:NNP

[Sun97a] Dongchu Sun. A note on noninformative priors for Weibull distributions. Journal of Statistical Planning and Inference, 61(2):319– 338, June 16, 1997. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375896001553.

Sun:1997:CAD

[Sun97b] Shan Sun. A class of adaptive distribution-free procedures. Journal of Statistical Planning and Inference, 59(2):191–211, April 15, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375896001164.

Suppes:1990:PCQ

[Sup90] Patrick Suppes. Probabilistic causality in quantum mechanics. Journal of Statistical Planning and Inference, 25(3):293–302, July 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375890900789.

Schwabe:1995:ODE

[SW95a] Rainer Schwabe and Werner Wierich. D-optimal designs of experiments with non-interacting factors. Journal of Statistical Planning and Inference, 44(3):371–384, ???? 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589400052W.

Srivastava:1995:IVT

[SW95b] M. S. Srivastava and Y. Wu. An improved version of Taguchi’s on-line control procedure. Journal of Statistical Planning and Inference, 43(1–2):133–145, January 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589400014M.

Soon:1996:NEC

[SW96a] Guoxing Soon and Michael Woodroofe. Nonparametric estimation and consistency for renewal processes. Journal of Statistical Planning and Inference, 53(2):171–195, August 15, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895001271.

Sun:1996:ASP

[SW96b] Jiayang Sun and Michael Woodroofe. Adaptive smoothing for a penalized NPMLE of a non-increasing density. Journal of Statistical Planning and Inference, 52(2):143–159, June 15, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589500114X.

Stahel:1997:ARE

[SW97] Werner A. Stahel and Alan Welsh. Approaches to robust estimation in the simplest variance components model. Journal of Statistical Planning and Inference, 57(2):295–319, February 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S037837589600050X.

Srivastava:1998:RCD

[SW98] J. N. Srivastava and Y. C. Wang. Row-column designs: nonadditivity makes them hazardous to use. Journal of Statistical Planning and Inference, 73(1–2):277–315, September 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375898000664.

Szablowski:1997:IPM

[SWA97] Pawel J. Szablowski, Jacek Wesolowski, and Mohammad Ahsanullah. Identification of probability measures via distribution of quotients. Journal of Statistical Planning and Inference, 63(2):377–385, October 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375897000359.

Szajowski:1992:OLD

[Sza92] K. Szajowski. Optimal on-line detection of outside observations. Journal of Statistical Planning and Inference, 30(3):413– 422, March 1992. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589290167Q.

Thavaneswaran:1991:EMN

[TA91] A. Thavaneswaran and Bovas Abraham. Estimation of multivariate non-linear time series models. Journal of Statistical Planning and Inference, 29(3):351–363, November 1991. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375891900094.

Tahir:1994:SEM

[Tah94] Mohamed Tahir. Sequential estimation of the mean of an exponential family in three stages. Journal of Statistical Planning and Inference, 38(1):1–12, January 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589200148W.

Tahir:1995:ARS

[Tah95] Mohamed Tahir. On the asymptotic regret of a sequential procedure for estimating the mean of a normal distribution. Journal of Statistical Planning and Inference, 48(3):331–338, December 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589500011W.

Takagi:1998:NCC

[Tak98] Yoshiji Takagi. A new criterion of confidence set estimation: Improvement of the Neyman shortness. Journal of Statistical Planning and Inference, 69(2):329–338, June 15, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897001584.

Tallini:1998:CD

[Tal98] Giuseppe Tallini. Composed designs. Journal of Statistical Planning and Inference, 73(1–2):331–351, September 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898000688.

Tang:1999:BBS

[Tan99] Boxin Tang. Balanced bootstrap in sample surveys and its relationship with balanced repeated replication. Journal of Statistical Planning and Inference, 81(1):121–127, October 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375899000130.

Tse:1999:NCT

[TAT99] Y. K. Tse, V. V. Anh, and Q. Tieng. No-cointegration test based on fractional differencing: Some Monte Carlo results. Journal of Statistical Planning and Inference, 80(1–2):257–267, August 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375898002535.

Tamhane:1999:SMT

[TD99] Ajit C. Tamhane and Charles W. Dunnett. Stepwise multiple test procedures with biometric applications. Journal of Statistical Planning and Inference, 82(1–2):55–68, December 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375899000312.

Teirlinck:1997:P

[Tei97] Luc Teirlinck. Preface. Journal of Statistical Planning and Inference, 58(1):1, March 1, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375896000547.

Tendick:1991:ONA

[Ten91] Patrick Tendick. Optimal noise addition for preserving confidentiality in multivariate data. Journal of Statistical Planning and Inference, 27(3):341–353, March 1991. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589190047I.

Tendick:1992:AEN

[Ten92] Patrick Tendick. Assessing the effectiveness of the noise addition method of preserving confidentiality in the multivariate normal case. Journal of Statistical Planning and Inference, 31(3): 273–282, June 1992. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589290137H.

Tenreiro:1998:ADD

[Ten98] Carlos Tenreiro. Asymptotic distribution for a discrete version of integrated square error of multivariate density kernel estimators. Journal of Statistical Planning and Inference, 69(1):133– 151, June 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375897001547.

Terrell:1997:BRS

[Ter97] George R. Terrell. Book review: Statistical theory and computational aspects of smoothing: W. Härdle and M. G. Schimek (eds.), (Physica-Verlag, Heidelberg, 1996, 265 pp., ISBN No. 3-7908-0930-6). Journal of Statistical Planning and Inference, 64(1):167– 169, October 30, 1997. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375897001043.

Toman:1993:RBE

[TG93] Blaza Toman and Joseph L. Gastwirth. Robust Bayesian experimental design and estimation for analysis of variance models using a class of normal mixtures. Journal of Statistical Planning and Inference, 35(3):383–398, June 1993. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390024Z.

Thavaneswaran:1999:PEF

[TH99] A. Thavaneswaran and C. C. Heyde. Prediction via estimating functions. Journal of Statistical Planning and Inference, 77(1):89–101, February 15, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375898001797.

Thompson:1993:RCL

[Tho93] G. L. Thompson. Rates of convergence for linear rank statistics with dependent data. Journal of Statistical Planning and Inference, 37(2):145–157, November 1993. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390084J.

Thoeni:1997:RVV

[Tho97] Hanspeter Thoeni. Rezension: Verfahrensbibliothek: Versuchsplanung und -auswertung, band 1: D. Rasch, G. Herrendörfer, J. Bock, N. Victor, V. Guiard (eds), Lehr- und Handbücher der Statistik. (R. Oldenbourg Verlag München, Wien, 1996. x + 940 pp. ISBN 3-486-23146-4). Journal of Statistical Planning and Inference, 64(1):163–165, October 30, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897001031.

Tiku:1997:NPB

[Tik97] M. L. Tiku. A note on a paper by N. Balakrishnan. Journal of Statistical Planning and Inference, 61(2):317–318, June 16, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375896001528.

Tjur:1993:AOB

[Tju93] Tue Tjur. An algorithm for optimization of block designs. Journal of Statistical Planning and Inference, 36(2–3):277–282, August/September 1993. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589390130X.

Tracy:1991:FEM

[TK91] Derrick S. Tracy and Khushnood A. Khan. Fourth exact moment result for improving MRBP based inferences. Journal of Statistical Planning and Inference, 28(2):263–270, June 1991. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375891900319.

Tripathi:1992:EPU

[TK92] T. P. Tripathi and Ravindra Khattree. Estimation of a parameter using Pitman nearness criterion. Journal of Statistical Planning and Inference, 32(3):281–289, September 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290011G.

Tracy:1996:RJT

[TM96] Derrick S. Tracy and Naurang S. Mangat. On respondent’s jeopardy in two alternate questions randomized response model. Journal of Statistical Planning and Inference, 55(1):107–114, October 15, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375895001832.

Toman:1991:BOE

[TN91] Blaza Toman and William I. Notz. Bayesian optimal experimental design for treatment-control comparisons in the presence of two-way heterogeneity. Journal of Statistical Planning and Inference, 27(1):51–63, January 1991. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589190081O.

Takada:1995:FWS

[TN95] Yoshikazu Takada and Yasushi Nagata. Fixed-width sequential confidence interval for the mean of a gamma distribution. Journal of Statistical Planning and Inference, 44(3):277–289, ???? 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589400053X.

Tiwari:1998:TDO

[TN98]  Neeraj Tiwari and A. K. Nigam.  On two-dimensional

optimal controlled selection. Journal of Statistical Planning and Inference, 69(1):89–100, June 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897001286.

Tomizawa:1993:ODP

[Tom93] Sadao Tomizawa. Orthogonal decomposition of point-symmetry model for two-way contingency tables. Journal of Statistical Planning and Inference, 36(1):91–100, July 1993. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390104E.

Tomkins:1995:RCT

[Tom95] R. J. Tomkins. Rates of convergence for tail series. Journal of Statistical Planning and Inference, 43(1–2):77–85, January 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589400009K.

Thiagarajah:1997:IES

[TP97] K. Thiagarajah and Sudhir R. Paul. Interval estimation for the scale parameter of the two-parameter exponential distribution based on time-censored data. Journal of Statistical Planning and Inference, 59(2):279–289, April 15, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897836403.

Talwalker:1990:MQA

[TR90] Sheela Talwalker and B. Raja Rao. Modified quadratic analysis in prediction with mixed binary and continuous explanatory variables. Journal of Statistical Planning and Inference, 26(1):47– 57, September 1990. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589090094B.

Tran:1990:RSS

[Tra90] Lanh Tat Tran. Rank statistics for serial dependence. Journal of Statistical Planning and Inference, 24(2):215–232, February 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589090043T.

Tran:1994:DET

[Tra94] Lanh Tat Tran. Density estimation for time series by histograms. Journal of Statistical Planning and Inference, 40(1):61–79, June 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375894901422.

Truong:1991:NCE

[Tru91] Young K. Truong. Nonparametric curve estimation with time series errors. Journal of Statistical Planning and Inference, 28(2): 167–183, June 1991. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375891900249.

Tiwari:1991:IHC

[TS91] Ram C. Tiwari and Pranab K. Sen. Incorporating historical controls in testing for a trend in multinomial proportions. Journal of Statistical Planning and Inference, 27(2):143–156, February 1991. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375891900124.

Tiku:1992:NME

[TS92] M. L. Tiku and R. P. Suresh. A new method of estimation for location and scale parameters. Journal of Statistical Planning and Inference, 30(2):281–292, February 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290088A.

Tonghui:1995:CTM

[TS95] Wang Tonghui and Wong Chi Song. Cochran theorems for a multivariate elliptically contoured model. Journal of Statistical Planning and Inference, 43(1–2):257–270, January 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589400025Q.

Trenkler:1996:DVV

[TS96] G. Trenkler and P. Stahlecker. Dropping variables versus use of proxy variables in linear regression. Journal of Statistical Planning and Inference, 50(1):65–75, February 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895000453.

Tazawa:1997:EUB

[TS97] Shinsei Tazawa and Teruhiro Shirakura. Enumeration of unlabelled bicolored graphs by degree parities. Journal of Statistical Planning and Inference, 58(1):193–206, March 1, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896000705.

Tseng:1994:PAL

[Tse94] Sheng-Tsaing Tseng. Planning accelerated life tests for selecting the most reliable product. Journal of Statistical Planning and Inference, 41(2):215–230, September 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894901651.

Tracy:1996:ARP

[TSS96] D. S. Tracy, H. P. Singh, and R. Singh. An alternative to the ratiocum-product estimator in sample surveys. Journal of Statistical Planning and Inference, 53(3):375–387, August 30, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895001360.

Thompson:1992:BNE

[TT92a] M. E. Thompson and A. Thavaneswaran. On Bayesian nonparametric estimation for stochastic processes. Journal of Statistical Planning and Inference, 33(1):131–141, October 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290098D.

Trenkler:1992:PTP

[TT92b] G. Trenkler and H. Toutenburg. Pre-test procedures and forecasting in the regression model under restrictions. Journal of Statistical Planning and Inference, 30(2):249–256, February 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375892900857.

Teverovsky:1999:CLM

[TTW99] Vadim Teverovsky, Murad S. Taqqu, and Walter Willinger. A critical look at Lo’s modified R/S statistic. Journal of Statistical Planning and Inference, 80(1–2):211–227, August 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837589800250X.

Tukey:1997:MHF

[Tuk97] John Tukey. More honest foundations for data analysis. Journal of Statistical Planning and Inference, 57(1):21–28, January 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375896000328.

Tungsarote:1998:SIB

[Tun98] Rarai Tungsarote. Some investigations on the Bose-Srivastava algebra related to the multidimensional partially balanced association scheme. Journal of Statistical Planning and Inference, 73(1–2):363–372, September 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898000706.

Thas:1998:GQW

[TV98] J. A. Thas and H. Van Maldeghem. Generalized quadrangles weakly embedded in finite projective space. Journal of Statistical Planning and Inference, 73(1–2):353–361, September 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S037837589800069X.

Tiessen:1997:MBI

[TvR97] D. Tiessen and G. H. J. van Rees. Many (22, 44, 14, 7, 4) — and (15, 42, 14, 5, 4)-balanced incomplete block designs. Journal of Statistical Planning and Inference, 62(1):115–124, July 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896001747.

Tantiyaswasdikul:1994:ISS

[TW94] Chim Tantiyaswasdikul and Michael B. Woodroofe. Isotonic smoothing splines under sequential designs. Journal of Statistical Planning and Inference, 38(1):75–87, January 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589200151S.

Tonchev:1997:STS

[TW97] Vladimir D. Tonchev and Robert S. Weishaar. Steiner triple systems of order 15 and their codes. Journal of Statistical Planning and Inference, 58(1):207–216, March 1, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896000717.

Tan:1999:SSM

[TX99] Wai-Yuan Tan and Zhihua Xiang. The state-space model of the HIV epidemic with variable infection in the homosexual populations. Journal of Statistical Planning and Inference, 78(1–2): 71–87, May 1999. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375898002079.

Ting:1996:BOD

[TY96] Chao-Ping Ting and Yu-Yun Yang. Bayes A-optimal designs for comparing test treatments with a control. Journal of Statistical Planning and Inference, 53(2):237–251, August 15, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895001468.

Tiwari:1993:NBE

[TZ93] Ram C. Tiwari and Jyoti N. Zalkikar. Nonparametric Bayesian estimation of survival function under random left truncation. Journal of Statistical Planning and Inference, 35(1):31–45, April 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589390065E.

Uddin:1992:CSB

[Udd92] Nizam Uddin. Constructions for some balanced incomplete block designs with nested rows and columns. Journal of Statistical Planning and Inference, 31(2):253–261, May 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290034P.

Uddin:1996:OIB

[Udd96] Nizam Uddin. E-optimal incomplete block designs with two distinct block sizes. Journal of Statistical Planning and Inference, 55(2):235–248, October 30, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589500193X.

Uddin:1997:OPB

[Udd97] Nizam Uddin. Optimal partially balanced nested row-column designs. Journal of Statistical Planning and Inference, 60(2):351– 366, May 15, 1997. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375896001322.

Ullah:1996:EDD

[Ull96] Aman Ullah. Entropy, divergence and distance measures with econometric applications. Journal of Statistical Planning and Inference, 49(1):137–162, January 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895000348.

Uddin:1991:ONO

[UM91] Nizam Uddin and J. P. Morgan. Optimal and near optimal sets of Latin squares for correlated errors. Journal of Statistical Planning and Inference, 29(3):279–290, November 1991. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589190004X.

Usami:1997:EKT

[UT97] Yoshihiro Usami and Yasuyuki Toyooka. Errata of Kariya and Toyooka (1992), “Bounds for normal approximations to the distributions of generalized least squares predictors and estimators”. Journal of Statistical Planning and Inference, 58(2):399– 405, March 15, 1997. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375896000821. See [KT92].

VanAken:1997:SAD

[Van97] Troy D. VanAken. A search for affine difference sets of even order. Journal of Statistical Planning and Inference, 62(1):125–133, July 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375896001759.

Vangel:1998:AEV

[Van98] Mark G. Vangel. ANOVA estimates of variance components for quasi-balanced mixed models. Journal of Statistical Planning and Inference, 70(1):139–148, July 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897001729.

vanBerkum:1995:RAC

[vBLO95] E. E. M. van Berkum, H. N. Linssen, and D. A. Overdijk. The reduction and ancillary classification of observations. Journal of Statistical Planning and Inference, 45(3):347–372, June 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589400088D.

vanBerkum:1996:IRI

[vBLO96] E. E. M. van Berkum, H. N. Linssen, and D. A. Overdijk. Inference rules and inferential distributions. Journal of Statistical Planning and Inference, 49(3):305–317, February 1, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895000151.

vanBerkum:1998:SRS

[vBLO98] E. E. M. van Berkum, H. N. Linssen, and D. A. Overdijk. ‘sampling the reference set’ revisited. Journal of Statistical Planning and Inference, 67(2):167–186, April 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897001171.

vanderLaan:1996:DER

[vdL96] Paul van der Laan. Distributional and efficiency results for subset selection. Journal of Statistical Planning and Inference, 54(2):159–174, September 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0378375896877202.

vanDorp:1997:SID

[vDMS97] J. René van Dorp, Thomas A. Mazzuchi, and Refik Soyer. Sequential inference and decision making for single mission systems development. Journal of Statistical Planning and Inference, 62(2):207–218, August 15, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896001826.

Vellaisamy:1996:NES

[Vel96] P. Vellaisamy. A note on the estimation of the selected scale parameters. Journal of Statistical Planning and Inference, 55(1):39– 46, October 15, 1996. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375895001786.

Voss:1998:MCU

[VH98] Daniel T. Voss and Jason C. Hsu. Multiple comparisons for an unbalanced a×b design under mixed models with interaction. Journal of Statistical Planning and Inference, 67(2):297–309, April 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S037837589700116X.

Vila:1991:LOR

[Vil91] J. P. Vila. Local optimality of replications from a minimal Doptimal design in regression: a sufficient and quasi-necessary condition. Journal of Statistical Planning and Inference, 29(3):261– 277, November 1991. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589190003W.

Vinod:1996:EFS

[Vin96] H. D. Vinod. Empirically feasible solutions and explicit dynamics for rational expectation models. Journal of Statistical Planning and Inference, 49(2):203–222, January 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895000372.

Valeroso:1999:MSM

[VK99] Elsie S. Valeroso and André I. Khuri. Multiresponse surface models with random block effects. Journal of Statistical Planning and Inference, 79(1):157–173, June 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898002353.

Vinod:1995:ECP

[VM95] H. D. Vinod and B. D. McCullough. Estimating cointegration parameters: an application of the double bootstrap. Journal of Statistical Planning and Inference, 43(1–2):147–156, January 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589400015N.

Voss:1993:CCS

[Vos93] Daniel T. Voss. Comparing the classes of single replicate generalized cyclic factorial designs with and without pseudofactors. Journal of Statistical Planning and Inference, 35(1):113– 120, April 1993. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589390071D.

Voss:1999:AOS

[Vos99] Daniel T. Voss. Analysis of orthogonal saturated designs. Journal of Statistical Planning and Inference, 78(1–2):111–130, May 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375898002109.

Vakil:1994:SCS

[VP94] Farok Vakil and Milton Parnes. On the structure of a class of sets useful in nonadaptive group testing. Journal of Statistical Planning and Inference, 39(1):57–69, April 1, 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894901120.

vonRosen:1993:UCM

[vR93] Dietrich von Rosen. Uniqueness conditions for maximum likelihood estimators in a multivariate linear model. Journal of Statistical Planning and Inference, 36(2–3):269–275, August/ September 1993. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589390129T.

Vidakovic:1999:EEB

[VR99] Brani Vidakovic and Fabrizio Ruggeri. Expansion estimation by Bayes rules. Journal of Statistical Planning and Inference, 79(2): 223–235, July 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375899000051.

vanRooij:1996:AMR

[vRR96] Arnoud C. M. van Rooij and Frits H. Ruymgaart. Asymptotic minimax rates for abstract linear estimators. Journal of Statistical Planning and Inference, 53(3):389–402, August 30, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895001212.

Vijayasree:1993:MET

[VS93] G. Vijayasree and Harshinder Singh. Mixed estimators of two ordered exponential means. Journal of Statistical Planning and Inference, 35(1):47–53, April 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390066F.

Vounatsou:1997:SBB

[VS97] P. Vounatsou and A. F. M. Smith. Simulation-based Bayesian inferences for two-variance components linear models. Journal of Statistical Planning and Inference, 59(1):139–161, March 30, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375896000936.

VanLeeuwen:1998:SCO

[VSB98] Dawn M. VanLeeuwen, Justus F. Seely, and David S. Birkes. Sufficient conditions for orthogonal designs in mixed linear models. Journal of Statistical Planning and Inference, 73(1–2):373–389, September 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375898000718.

vanTrung:1996:HOI

[vTrWM96] Tran van Trung, Qiu rong Wu, and Dale M. Mesner. High order intersection numbers of t-designs. Journal of Statistical Planning and Inference, 56(2):257–268, December 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896000225.

VanKeilegom:1998:BQF

[VV98] Ingrid Van Keilegom and Noël Veraverbeke. Bootstrapping quantiles in a fixed design regression model with censored data. Journal of Statistical Planning and Inference, 69(1):115–131, June 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375897001262.

Voss:1999:SCI

[VW99] Daniel T. Voss and Weizhen Wang. Simultaneous confidence intervals in the analysis of orthogonal saturated designs. Journal of Statistical Planning and Inference, 81(2):383–392, November 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899000695.

VanNess:1998:RDA

[VY98] John W. Van Ness and Jim J. Yang. Robust discriminant analysis: Training data breakdown point. Journal of Statistical Planning and Inference, 67(1):67–83, March 16, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897001067.

Wallis:1996:RGD

[Wal96] W. D. Wallis. Regular graph designs. Journal of Statistical Planning and Inference, 51(3):273–281, May 1, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895000917.

Wan:1996:EEV

[Wan96] Alan T. K. Wan. Estimating the error variance after a pre-test for an inequality restriction on the coefficients. Journal of Statistical Planning and Inference, 52(2):197–213, June 15, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895001107.

Wang:1997:VCT

[Wan97a] Tonghui Wang. Versions of Cochran’s theorem for general quadratic expressions in normal matrices. Journal of Statistical Planning and Inference, 58(2):283–297, March 15, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896000845.

Wang:1997:MEW

[Wan97b] Yazhen Wang. Minimax estimation via wavelets for indirect long-memory data. Journal of Statistical Planning and Inference, 64(1):45–55, October 30, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896002054.

Wandji:1998:NGF

[Wan98a] Joseph Ngatchou Wandji. A nonparametric goodness-of-fit test for a class of parametric autoregressive models. Journal of Statistical Planning and Inference, 71(1–2):57–74, August 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837589800010X.

Wang:1998:ECU

[Wan98b] Hsiuying Wang. Estimated confidence under ancillary statistic everywhere-valid constraint. Journal of Statistical Planning and Inference, 67(1):123–135, March 16, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897000931.

Wang:1998:PCI

[Wan98c] Yong-Cheng Wang. Probability of correct identification of nonadditive cells in row-column designs. Journal of Statistical Planning and Inference, 73(1–2):391–408, September 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837589800072X.

Wang:1999:ETV

[Wan99] Weizhen Wang. On equivalence of two variances of a bivariate normal vector. Journal of Statistical Planning and Inference, 81(2):279–292, November 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375899000531.

Wasserman:1996:CBI

[Was96] Larry Wasserman. The conflict between improper priors and robustness. Journal of Statistical Planning and Inference, 52(1): 1–15, June 1, 1996. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375895001093.

Watanabe:1993:PSF

[Wat93] Toshihiro Watanabe. On the product of a Schur function and a skew Schur function. Journal of Statistical Planning and Inference, 34(2):305–315, February 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390014W.

Westfall:1994:EPC

[WB94] Peter H. Westfall and Ronald H. Bremer. Efficiency properties of cell means variance component estimates. Journal of Statistical Planning and Inference, 38(2):159–178, February 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894900337.

Wang:1995:RLC

[WC95] C. Y. Wang and R. J. Carroll. On robust logistic case-control studies with response-dependent weights. Journal of Statistical Planning and Inference, 43(3):331–340, February 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375893E0091T.

Wang:1998:LCT

[WC98a] P. C. Wang and M. H. Chen. Level changes and trend resistance on replacement in asymmetric orthogonal arrays. Journal of Statistical Planning and Inference, 69(2):349–358, June 15, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375897001687.

Weiss:1998:BMI

[WC98b] Robert E. Weiss and Meehyung Cho. Bayesian marginal influence assessment. Journal of Statistical Planning and Inference, 71(1–2):163–177, August 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375898000159.

Wong:1999:CTM

[WC99] Chi Song Wong and Hua Cheng. Cochran theorems for a multivariate vector-elliptically contoured model. Part II. Journal of Statistical Planning and Inference, 79(2):299–324, July 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375898002596.

Whitmore:1991:SEA

[WCP91] G. A. Whitmore, M. Chown, and G. G. Pullum. Some exact algebraic expressions for the reliability of multipath switching networks. Journal of Statistical Planning and Inference, 29(1–2):199–209, September/October 1991. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290133D.

Wu:1998:CRS

[WD98] C. F. J. Wu and Yuan Ding. Construction of response surface designs for qualitative and quantitative factors. Journal of Statistical Planning and Inference, 71(1–2):331–348, August 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375898000032.

William:1999:GQL

[WD99] Martin L. William and T. M. Durairajan. A generalised quasilikelihood estimation. Journal of Statistical Planning and Inference, 79(2):237–246, July 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375898002274.

Wefelmeyer:1997:AEP

[Wef97] Wolfgang Wefelmeyer. Adaptive estimators for parameters of the autoregression function of a Markov chain. Journal of Statistical Planning and Inference, 58(2):389–398, March 15, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896000791.

Weissman:1997:FMD

[Wei97] Ishay Weissman. Fitting a Mejzler distribution to extreme value data. Journal of Statistical Planning and Inference, 61(1):39– 47, May 30, 1997. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375896001371.

Weiss:1999:SUT

[Wei99] Neil A. Weiss. Strategies for using technologies in introductory statistics courses. Journal of Statistical Planning and Inference, 78(1–2):101–110, May 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898002092.

Wertz:1992:QID

[Wer92] W. Wertz. On quasi-invariant density estimation. Journal of Statistical Planning and Inference, 32(2):271–280, August 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290052T.

Wesolowski:1996:SAG

[Wes96] Jacek Wesolowski. Stability for the Arnold-ghosh characterization of the geometric distribution. Journal of Statistical Planning and Inference, 52(3):263–269, July 1, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895001190.

Wetzel:1997:IPT

[Wet97] Nathan R. Wetzel. Improper priors for translation families. Journal of Statistical Planning and Inference, 60(2):215–227, May 15, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S037837589600136X.

Weinberger:1994:PSC

[WF94] Marcelo J. Weinberger and Meir Feder. Predictive stochastic complexity and model estimation for finite-state processes. Journal of Statistical Planning and Inference, 39(3):353–372, May 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894900922.

Wegman:1992:SMN

[WH92] Edward J. Wegman and Muhammad K. Habib. Stochastic methods for neural systems. Journal of Statistical Planning and Inference, 33(1):5–25, October 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375892900927.

Wiens:1992:MDA

[Wie92] Douglas P. Wiens. Minimax designs for approximately linear regression. Journal of Statistical Planning and Inference, 31(3): 353–371, June 1992. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589290142F.

Wiens:1994:RDA

[Wie94] Douglas P. Wiens. Robust designs for approximately linear regression: M-estimated parameters. Journal of Statistical Planning and Inference, 40(1):135–160, June 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894901473.

Wiper:1998:BAM

[Wip98] M. P. Wiper. Bayesian analysis of Er/M/1 and Er/M/c queues. Journal of Statistical Planning and Inference, 69(1):65–79, June 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375897001249.

Wirjanto:1997:CFD

[Wir97] T. S. Wirjanto. Comment on A. F. Desmond’s paper: Optimal estimating functions, quasi-likelihood and statistical modelling. Journal of Statistical Planning and Inference, 60(1):109– 113, April 30, 1997. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375897900136. See [Des97a].

Witkovsky:1998:EVC

[Wit98] Viktor Witkovský. On estimation of variance components with constraints. Journal of Statistical Planning and Inference, 69(1): 81–87, June 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375897001274.

Wang:1996:ASS

[WJ96] Yangxian Wang and Ma Jianmin. Association schemes of symmetric matrices over a finite field of characteristic two. Journal of Statistical Planning and Inference, 51(3):351–371, May 1, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375895000984.

Wang:1993:EOE

[WL93] Song-Gui Wang and Erkki P. Liski. Effects of observations on the eigensystem of a sample covariance matrix. Journal of Statistical Planning and Inference, 36(2–3):215–226, August/September 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589390125P.

Wei:1996:SET

[WL96] Hongzeng Wei and Chong Li. Some enumeration theorems in pseudo-symplectic geometry over a finite field of characteristic two and the construction of association schemes. Journal of Statistical Planning and Inference, 51(3):373–386, May 1, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895000992.

Wasserman:1993:LBR

[WLW93] Larry Wasserman, Michael Lavine, and Robert L. Wolpert. Linearization of Bayesian robustness problems. Journal of Statistical Planning and Inference, 37(3):307–316, December 1993. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589390109J.

Wardrop:1990:SRS

[WM90] Daniel M. Wardrop and Raymond H. Myers. Some response surface designs for finding optimal conditions. Journal of Statistical Planning and Inference, 25(1):7–28, May 1990. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589090003D.

Weissman:1998:BRE

[WN98] I. Weissman and S. Yu. Novak. On blocks and runs estimators of the extremal index. Journal of Statistical Planning and Inference, 66(2):281–288, March 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897000955.

Wong:1994:GOD

[Won94] Weng Kee Wong. G-optimal designs for multi-factor experiments with heteroscedastic errors. Journal of Statistical Planning and Inference, 40(1):127–133, June 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894901465.

Wang:1996:ERP

[WR96] Naisyin Wang and David Ruppert. Estimation of regression parameters in a semiparametric transformation model. Journal of Statistical Planning and Inference, 52(3):331–351, July 1, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589500131X.

Wassmer:1999:PTM

[WRKL99] Gernot Wassmer, Peter Reitmeir, Meinhard Kieser, and Walter Lehmacher. Procedures for testing multiple endpoints in clinical trials: an overview. Journal of Statistical Planning and Inference, 82(1–2):69–81, December 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375899000324.

Wasserman:1994:DPR

[WS94] Larry Wasserman and Teddy Seidenfeld. The dilation phenomenon in robust Bayesian inference. Journal of Statistical Planning and Inference, 40(2–3):345–356, July 1994. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894901309.

Wada:1995:RAT

[WS95] Cicilia Yuko Wada and Pranab Kumar Sen. Restricted alternative test in a parametric model with competing risk data. Journal of Statistical Planning and Inference, 44(2):193–203, April 1, 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589400080F.

Wu:1994:DAC

[Wu94] Margaret C. Wu. Design aspects for clinical trials with repeated measures in the presence of informative censoring. Journal of Statistical Planning and Inference, 42(1–2):109–122, November 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375894901929.

Wu:1997:KDE

[Wu97]  Berlin Wu.  Kernel density estimation under weak de-

pendence with sampled data. Journal of Statistical Planning and Inference, 61(1):141–154, May 30, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896001516.

Wu:1998:ECU

[Wu98] Qi-Guang Wu. Existence conditions of the uniformly minimum risk unbiased estimators in extended growth curve models. Journal of Statistical Planning and Inference, 69(1):101–114, June 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375897001195.

Woodward:1995:MHD

[WWE95] Wayne A. Woodward, Paul Whitney, and Paul W. Eslinger. Minimum Hellinger distance estimation of mixture proportions. Journal of Statistical Planning and Inference, 48(3):303–319, December 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589500006U.

Wei:1996:GCD

[WZ96a] R. Wei and L. Zhu. Generalized H-codes and doubly even self-dual binary codes. Journal of Statistical Planning and Inference, 51(3):387–393, May 1, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589500100X.

Wiens:1996:MRD

[WZ96b] Douglas P. Wiens and Julie Zhou. Minimax regression designs for approximately linear models with autocorrelated errors. Journal of Statistical Planning and Inference, 55(1):95–106, October 15, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375895001875.

Xie:1991:GMC

[XB91] M. Xie and B. Bergman. On a general measure of component importance. Journal of Statistical Planning and Inference, 29(1–2):211–220, September/October 1991. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290134E.

Xiang:1998:RAH

[Xia98] Qing Xiang. On reversible abelian Hadamard difference sets. Journal of Statistical Planning and Inference, 73(1–2):409–416, September 1, 1998. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375898000731.

Yamada:1996:HMS

[Yam96] Mieko Yamada. Hadamard matrices and spin models. Journal of Statistical Planning and Inference, 51(3):309–321, May 1, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375895000941.

Yanagimoto:1990:CME

[Yan90a] Takemi Yanagimoto. Combining moment estimates of a parameter common through strata. Journal of Statistical Planning and Inference, 25(2):187–194, June 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375890900653.

Yanagimoto:1990:R

[Yan90b] Takemi Yanagimoto. Rejoinder. Journal of Statistical Planning and Inference, 25(2):196–198, June 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375890900686.

Yang:1993:CLT

[Yan93] Song Yang. A central limit theorem for the integrated square error of the kernel density estimators with randomly censored data. Journal of Statistical Planning and Inference, 37(2):127– 143, November 1993. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589390083I.

Yang:1995:KDT

[Yan95] Song Yang. Koul-DeWet-type estimation in censored regression using the conditional empirical process. Journal of Statistical Planning and Inference, 46(3):351–363, August 15, 1995. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589400113A.

Yang:1996:SBE

[Yan96a] Hu Yang. On the stability of biased estimates and the regularization method. Journal of Statistical Planning and Inference, 52(1): 67–75, June 1, 1996. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375894000247.

Yanxun:1996:LSD

[Yan96b] Chang Yanxun. On large sets of directed triple systems. Journal of Statistical Planning and Inference, 51(2):137–142, April 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375895000771.

Yatracos:1995:RSP

[Yat95] Yannis G. Yatracos. On the rare species of a population. Journal of Statistical Planning and Inference, 48(3):321–329, December 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375895000196.

Yekutieli:1999:RBF

[YB99] Daniel Yekutieli and Yoav Benjamini. Resampling-based false discovery rate controlling multiple test procedures for correlated test statistics. Journal of Statistical Planning and Inference, 82(1–2):171–196, December 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375899000415.

Yeh:1995:BVS

[YC95] Lam Yeh and S. T. B. Choy. Bayesian variable sampling plans for the exponential distribution with uniformly distributed random censoring. Journal of Statistical Planning and Inference, 47(3):277–293, October 15, 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589400136J.

Yip:1998:SIM

[YC98] Paul S. F. Yip and Qizhi Chen. Statistical inference for a multitype epidemic model. Journal of Statistical Planning and Inference, 71(1–2):229–244, August 1, 1998. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898000871.

Ye:1994:BRP

[Ye94] Keying Ye. Bayesian reference prior analysis on the ratio of variances for the balanced one-way random effect model. Journal of Statistical Planning and Inference, 41(3):267–280, October 1994. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589490023X.

Yamamoto:1996:NLF

[YFHY96] Sumiyasu Yamamoto, Yoshio Fujii, Yoshifumi Hyodo, and Hiromu Yumiba. New light on fractional 2^{m} factorial designs. Journal of Statistical Planning and Inference, 56(2):269–287, December 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375896000237.

Yunqing:1996:ESH

[YGL96] Xu Yunqing, Ge Gennian, and Zhu Lie. Existence of SOLS with holes of type 2^{n}u^{1}. Journal of Statistical Planning and Inference, 51(3):293–307, May 1, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375896837347.

Yao:1990:ONG

[YH90] Y. C. Yao and F. K. Hwang. On optimal nested group testing algorithms. Journal of Statistical Planning and Inference, 24(2):167–175, February 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837589090039W.

Yamada:1999:CTL

[YIHN99] Shu Yamada, Yoshiko T. Ikebe, Hiroki Hashiguchi, and Naoto Niki. Construction of three-level supersaturated design. Journal of Statistical Planning and Inference, 81(1):183–193, October 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375899000075.

Yin:1996:OPE

[Yin96] J. Yin. Optimal packing of even circuits into a DK_{v}. Journal of Statistical Planning and Inference, 51(3):323–330, May 1, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589500095X.

Yakowitz:1992:MLO

[YK92] S. Yakowitz and M. Kollier. Machine learning for optimal blackjack counting strategies. Journal of Statistical Planning and Inference, 33(3):295–309, December 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375892900019.

Yohai:1997:LGR

[Yoh97] Victor J. Yohai. Local and global robustness of regression estimators. Journal of Statistical Planning and Inference, 57(1): 73–92, January 1997. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375896000377.

Yokoyama:1997:TFR

[Yok97] Takahisa Yokoyama. Tests for a family of random-effects covariance structures in a multivariate growth curve model. Journal of Statistical Planning and Inference, 65(2):281–292, December 15, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375897000621.

Yang:1992:EOD

[YS92] Grace L. Yang and Charles E. Swenberg. Estimation of open dwell time and problems of identifiability in channel experiments. Journal of Statistical Planning and Inference, 33(1):107–119, October 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589290096B.

Yu:1999:ECI

[YSS99] Philip L. H. Yu, Yijun Sun, and Bimal K. Sinha. On exact confidence intervals for the common mean of several normal populations. Journal of Statistical Planning and Inference, 81(2):263–277, November 1, 1999. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037837589900052X.

Yao:1998:CVB

[YT98] Qiwei Yao and Howell Tong. Cross-validatory bandwidth selections for regression estimation based on dependent data. Journal of Statistical Planning and Inference, 68(2):387–415, May 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375897001511.

Yu:1996:SCT

[Yu96] Kai Fun Yu. A simple comparison of two sequences of probabilities. Journal of Statistical Planning and Inference, 52(3):271– 287, July 1, 1996. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0378375895001328.

Xia:1996:NFS

[yXL96] Ming yuan Xia and Gang Liu. A new family of supplementary difference sets and Hadamard matrices. Journal of Statistical Planning and Inference, 51(3):283–291, May 1, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895000925.

Yohai:1997:OLR

[YZ97] Víctor J. Yohai and Ruben H. Zamar. Optimal locally robust M-estimates of regression. Journal of Statistical Planning and Inference, 64(2):309–323, November 15, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897000402.

Zahedi:1991:PMR

[Zah91] Hassan Zahedi. Proportional mean remaining life model. Journal of Statistical Planning and Inference, 29(1–2):221–228, September/October 1991. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037837589290135F.

Zhou:1992:TSR

[ZB92] X. H. Zhou and S. Blumenthal. Two stage reliability tests for new series systems. Journal of Statistical Planning and Inference, 33(3):345–366, December 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290004C.

Zhang:1992:RIB

[ZC92] Zhiyi Zhang and Arthur Cohen. Recovery of information in BIBDs with interaction. Journal of Statistical Planning and Inference, 31(3):373–386, June 1992. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290143G.

Zacks:1990:BSE

[ZdBPL90] S. Zacks, C. A. de B. Pereira, and J. G. Leite. Bayes sequential estimation of the size of a finite population. Journal of Statistical Planning and Inference, 25(3):363–380, July 1990. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375890900826.

Zeilberger:1996:SAW

[Zei96] Doron Zeilberger. Self-avoiding walks, the language of science, and Fibonacci numbers. Journal of Statistical Planning and Inference, 54(1):135–138, September 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589500162X.

Zelterman:1990:SNE

[Zel90] Daniel Zelterman. Smooth nonparametric estimation of the quantile function. Journal of Statistical Planning and Inference, 26(3):339–352, November 1990. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837589090136I.

Zellner:1996:PPF

[Zel96] Arnold Zellner. Past, present and future of econometrics. Journal of Statistical Planning and Inference, 49(1):3–8, January 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375895000275.

Zhang:1995:EQE

[Zha95] Biao Zhang. M-estimation and quantile estimation in the presence of auxiliary information. Journal of Statistical Planning and Inference, 44(1):77–94, March 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375894000403.

Zhang:1996:SBP

[Zha96] Jian Zhang. The sample breakdown points of tests. Journal of Statistical Planning and Inference, 52(2):161–181, June 15, 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375895001166.

Zhenfu:1996:FST

[Zhe96] Cao Zhenfu. Finite set theory and its application to cryptology. Journal of Statistical Planning and Inference, 51(2):129–136, April 1996. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375895000763.

Zhu:1996:DGA

[Zhu96] L. Zhu. Duadic group algebra codes. Journal of Statistical Planning and Inference, 51(3):395–401, May 1, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895001018.

Zinoviev:1996:ECC

[Zin96] Victor A. Zinoviev. On the equivalence of certain constant weight codes and combinatorial designs. Journal of Statistical Planning and Inference, 56(2):289–294, December 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896000249.

Zhang:1998:KDE

[ZK98] Shunpu Zhang and Rohana J. Karunamuni. On kernel density estimation near endpoints. Journal of Statistical Planning and Inference, 70(2):301–316, July 15, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897001870.

Zheng:1995:BBC

[ZL95] Xiaodong Zheng and Wei-Yin Loh. Bootstrapping binomial confidence intervals. Journal of Statistical Planning and Inference, 43(3):355–380, February 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037837589400031P.

Zhou:1993:HTV

[ZM93] Leping Zhou and Thomas Mathew. Hypotheses tests for variance components in some multivariate mixed models. Journal of Statistical Planning and Inference, 37(2):215–227, November 1993. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037837589390090S.

Zoubeidi:1992:BDC

[Zou92] Taoufik Zoubeidi. Bounded deficiency in the comparison of the means of two normal populations. Journal of Statistical Planning and Inference, 31(3):295–310, June 1992. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589290139J.

Zou:1997:AEF

[Zou97] Guohua Zou. Admissible estimation for finite population under the Linex loss function. Journal of Statistical Planning and Inference, 61(2):373–384, June 16, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375896001784.

Zhou:1999:SPM

[ZW99] Haibo Zhou and Clarice R. Weinberg. A semi-parametric mixture model in fertility studies. Journal of Statistical Planning and Inference, 75(2):453–462, January 1, 1999. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375898001608.

Zhu:1995:TMN

[ZWF95] Li-Xing Zhu, Hoi Lam Wong, and Kai-Tai Fang. A test for multivariate normality based on sample entropy and projection pursuit. Journal of Statistical Planning and Inference, 45(3):373–385, June 1995. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0378375894000584.

Zehui:1998:INC

[ZX98] Li Zehui and Li Xiaohu. IFRA*t _{0} and NBU*t _{0} classes of life distributions. Journal of Statistical Planning and Inference, 70(2):191–200, July 15, 1998. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375897001882.

Zhigljavsky:1996:ETS

[ZZ96] A. Zhigljavsky and L. Zabalkanskaya. Existence theorems for some group testing strategies. Journal of Statistical Planning and Inference, 55(2):151–173, October 30, 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0378375895001891.